### v7

# A Discrete Resolvent Framework for Operator Splitting in Delay Differential Equations

Abstract
This thesis develops a unified operator–theoretic framework for the analysis of implicit time discretization and operator splitting schemes for delay differential equations (DDEs). The central idea is to formulate both the implicit Euler method and the Lie–Trotter splitting scheme in terms of discrete resolvent operators acting on suitable product spaces that encode the present state and the history of the solution.
Unlike most existing convergence theories, the proposed framework does not require the existence of a continuous semigroup or evolution family generated by the underlying delay operator. Instead, convergence is established purely on the level of discrete propagators, relying on quantitative estimates of local defects together with appropriate notions of discrete stability for products of operators.
This approach leads to a transparent hierarchy of convergence results covering: (i) autonomous and non-autonomous problems, (ii) non-sectorial operators arising from transport and delay terms, and (iii) sectorial operators generating analytic smoothing.
In the non-sectorial setting, global convergence of Lie–Trotter splitting toward implicit Euler is shown to depend crucially on the stability properties of the discrete propagators, such as summability of powers, bounded variation, or Ritt/Kreiss-type conditions. In contrast, in the sectorial setting, analytic smoothing suppresses the accumulation of local defects, yielding finite-time convergence without any additional discrete stability assumptions.
The abstract theory is complemented by applications to linear and nonlinear reaction–diffusion delay equations, as well as numerical experiments that confirm the sharpness of the theoretical error bounds.
Overall, the thesis provides a systematic and flexible framework for the analysis of splitting methods for delay equations, applicable both within and beyond the classical semigroup setting.
Contents
Introduction
Background and motivation.
Delay differential equations (DDEs) arise naturally in a wide range of models, including population dynamics, control systems with aftereffect, and reaction–diffusion processes with memory. A fundamental difficulty in the analysis of such systems is the presence of time-nonlocal terms, which prevents the direct application of classical semigroup theory on the state space. A powerful approach, developed since the work of Hale and Verduyn Lunel Hale and Verduyn Lunel (1993), Diekmann et al. Diekmann et al. (1995), and further systematized in the semigroup formulation on product spaces Bátkai and Piazzera (2005), is to reformulate DDEs as abstract evolution equations on suitable product spaces, where the current state and its history are treated simultaneously.
From the numerical point of view, fully implicit schemes such as the implicit Euler method provide robust stability properties Hairer and Wanner (1996), but they are computationally expensive, especially for large-scale or non-autonomous problems. Operator splitting methods, and in particular Lie–Trotter splitting, offer an attractive alternative by separating the transport, delay, and reaction effects; see, e.g., Hundsdorfer and Verwer (2003); Lubich (2008); rhus (1998).
Despite their practical importance, the convergence theory of splitting schemes for delay equations remains incomplete; see, e.g., Bátkai et al. (2013, 2017a); Hansen and Stillfjord (2014); Cheng and et al. (2020), especially in the following two directions:
  • the presence of non-sectorial generators, such as transport operators arising from history dynamics;
  • non-autonomous coefficients, where classical semigroup tools are no longer directly applicable.
Limitations of existing theory.
Most existing convergence results for splitting methods rely, either explicitly or implicitly, on the existence of a $C_{0}$-semigroup or evolution family generated by the underlying continuous operator Trotter (1958); Faragó and Havasi (2007); Bátkai et al. (2011).
This excludes many natural delay models, where the continuous problem is not well-posed in the classical semigroup sense, or where well-posedness is too weak to support quantitative error analysis. As a consequence, the role of purely discrete stability mechanisms in governing convergence has not been fully clarified.
Main idea and contribution of this thesis.
The central idea of this thesis is to develop a convergence theory for implicit Euler and Lie–Trotter splitting that is formulated entirely at the level of discrete resolvent operators, without presupposing the existence of an underlying semigroup or evolution family.
Instead of starting from continuous well-posedness, we consider the discrete one-step maps
$R_{h}\quad\text{(implicit Euler resolvent)},\qquad P_{h}\quad\text{(Lie--Trotter splitting propagator)},$
and analyze the convergence of their iterates purely through:
  • local defect estimates between $R_{h}$ and $P_{h}$,
  • and suitable notions of discrete stability (summability, bounded variation, Ritt/Kreiss-type conditions) Gröger (1987); Lubich and Nevanlinna (1991); Nagy and Zemánek (1999); Le Merdy (2002); Gomilko and Zemánek (2011); Schweizer (2012).
This perspective allows us to treat regimes that lie beyond classical semigroup theory, while remaining fully compatible with it whenever a continuous evolution family exists.
Abstract convergence framework.
A central theoretical contribution of this thesis is the construction of a unified abstract convergence theory formulated in terms of discrete resolvent families. This framework culminates in four abstract convergence principles:
  • Theorem 2.10: autonomous non-sectorial case,
  • Theorem 2.11: autonomous sectorial case,
  • Theorem 2.12: non-autonomous non-sectorial case,
  • Theorem 2.13: non-autonomous sectorial case.
These theorems isolate the essential ingredients governing convergence: local defect estimates on suitable regularity spaces and discrete stability properties of the splitting propagators. Notably, no semigroup or evolution family assumptions are required at the abstract level.
Concrete instantiation for delay equations.
The abstract theory is then applied to delay differential equations reformulated on product spaces. In this setting, we derive concrete versions of the abstract results, presented as Main Theorems A–D in the respective chapters:
  • Section 3 (autonomous and non-autonomous non-sectorial cases),
  • Section 4 (autonomous and non-autonomous sectorial cases).
These results provide sharp global error bounds for Lie–Trotter splitting in terms of fractional regularity of the data and clearly exhibit the role of discrete stability assumptions such as summability, bounded variation, and Ritt/Kreiss conditions.
Relation to semigroup and evolution family theory.
Although the theory developed in this thesis does not require continuous well-posedness, it is fully consistent with classical results whenever a $C_{0}$-semigroup or evolution family exists. In particular, when the continuous operator generates a semigroup or satisfies the Acquistapace–Terreni conditions, the discrete convergence results recover the expected convergence toward the continuous solution operators. This establishes a conceptual bridge between discrete stability theory and continuous evolution theory.
Structure of the thesis.
The dissertation is organized as follows.
Section 1 develops the discrete resolvent framework and establishes the abstract convergence principles (Theorems A–D) without reference to continuous semigroup theory.
Sections 3 and 4 apply this framework to delay differential equations formulated on product spaces. Concrete convergence theorems (Main Theorems A–D) are derived for autonomous and non-autonomous problems, covering both non-sectorial and sectorial cases.
Section 5 presents numerical experiments for reaction–diffusion delay equations, illustrating the sharpness and practical relevance of the theoretical error estimates.
Appendices A and B collect auxiliary analytical results and concise program listings supporting the numerical experiments.
Figure 1: Structure of the theorems in this doctoral dissertation.
1 Functional analytic framework
This chapter provides the functional analytic framework underlying the continuous and discrete analysis developed in the sequel. Our goal is twofold: first, to place delay differential equations into the classical theory of $C_{0}$-semigroups and evolution families whenever such a framework is available; second, to formulate the operator splitting and implicit time discretization schemes in a manner that remains meaningful even beyond classical well-posedness.
The chapter is organized as follows. We first introduce the product-space formulation of delay equations and the associated operators. We then discuss perturbation theory and semigroup generation. Next, we recall the theory of fractional power spaces and interpolation, which plays a central role in the convergence analysis of numerical schemes. Finally, we explain how this continuous framework connects to the discrete resolvent formulations used throughout the paper.
1.1 Product-space formulation
Let $\tau<0$, let $X$ be a Banach space, and fix $1<p<\infty$. For a function $u:[\tau,T]\to X$, we encode the history segment by
$\rho_{t}(\sigma):=u(t+\sigma),\qquad\sigma\in[\tau,0].$
We work on the product space
$\mathcal{E}_{0}:=X\oplus L^{p}([\tau,0];X),\qquad\|(u,\rho)\|_{\mathcal{E}_{0}}:=\|u\|_{X}+\|\rho\|_{L^{p}([\tau,0];X)}.$
The phase variable is $x(t):=(u(t),\rho_{t})\in\mathcal{E}_{0}$. The compatibility relation $\rho_{t}(0)=u(t)$ is encoded in the domain of the transport operator below.
1.2 Transport operator and semigroup generation
Define the transport (left-shift) operator
$C(u,\rho):=(0,\partial_{\sigma}\rho),$
with domain
$\mathcal{D}(C):=\bigl\{(u,\rho)\in X\oplus W^{1,p}([\tau,0];X):\ \rho(0)=u\bigr\}.$
Then $C$ is the generator of a strongly continuous left-shift semigroup $(T_{C}(t))_{t\geq 0}$ on $\mathcal{E}_{0}$ (see, e.g., (Engel and Nagel, 1999, Chap. VI)). In particular, its resolvent
$S_{h}^{C}:=(I-hC)^{-1}\in\mathcal{L}(\mathcal{E}_{0})$
is well defined for every $h>0$ and admits an explicit representation.
1.3 Delay coupling and perturbation theory
The delay interaction is represented by the evaluation map
$\Phi_{\tau}:W^{1,p}([\tau,0];X)\to X,\qquad\Phi_{\tau}(\rho):=\rho(\tau),$
which is well-defined on $W^{1,p}$ but does not extend to a bounded operator on $L^{p}([\tau,0];X)$.
We therefore consider operator families $D(t):\mathcal{D}(C)\to\mathcal{E}_{0}$ of the form
$D(t)(u,\rho):=\bigl(a(t)u+b(t)\Phi_{\tau}(\rho),\,0\bigr),\qquad t\in[0,T],$
where $a(\cdot)$ and $b(\cdot)$ are bounded scalar functions.
1.4 Relative boundedness and semigroup generation
Definition 1.1 (CC–bounded operators).
Let $C:\mathcal{D}(C)\subset X\to X$ be densely defined. An operator $B:\mathcal{D}(C)\to X$ is called CC–bounded with relative bound $\alpha$ if there exists $c_{\alpha}>0$ such that
$\|Bx\|_{X}\leq\alpha\|Cx\|_{X}+c_{\alpha}\|x\|_{X},\qquad x\in\mathcal{D}(C).$
Lemma 1.2 ($C$-boundedness of the delay/reaction block on $\mathcal{E}_{p}$).
Let $\tau<0$ and $1<p<\infty$. Set
$\mathcal{E}_{p}:=\mathbb{R}\oplus L^{p}([\tau,0]),\qquad\|(u,\rho)\|_{\mathcal{E}_{p}}:=|u|+\|\rho\|_{L^{p}}.$
Define
$C(u,\rho):=(0,\partial_{\sigma}\rho),\qquad\mathcal{D}(C):=\{(u,\rho):\rho\in W^{1,p}([\tau,0]),\ \rho(0)=u\},$
and
$D(u,\rho):=(au+b\,\rho(\tau),0),\qquad(u,\rho)\in\mathcal{D}(C),$
with constants $a,b\in\mathbb{R}$. Then $D$ is $C$-bounded with relative bound 0, i.e., for every $\varepsilon>0$ there exists $c_{\varepsilon}>0$ such that
$\|Dx\|_{\mathcal{E}_{p}}\leq\varepsilon\|Cx\|_{\mathcal{E}_{p}}+c_{\varepsilon}\|x\|_{\mathcal{E}_{p}},\qquad x\in\mathcal{D}(C).$
In particular, $D$ is $C$-bounded with relative bound <1<1.
Proof.
Fix $\varepsilon>0$ and let $x=(u,\rho)\in\mathcal{D}(C)$. Since $1<p<\infty$, the Sobolev space $W^{1,p}([\tau,0])$ embeds continuously into $C([\tau,0])$, and the Gagliardo–Nirenberg inequality yields
$\|\rho\|_{L^{\infty}}\leq K\,\|\rho^{\prime}\|_{L^{p}}^{1/p}\,\|\rho\|_{L^{p}}^{1-1/p}$
for some constant $K=K(p,\tau)>0$. By Young’s inequality, for every $\delta>0$ there exists $C_{\delta}>0$ such that
$\|\rho\|_{L^{\infty}}\leq\delta\|\rho^{\prime}\|_{L^{p}}+C_{\delta}\|\rho\|_{L^{p}}.$
Choosing $\delta>0$ sufficiently small (depending on $\varepsilon,a,b$) and using $\rho(0)=u$ and $|\rho(\tau)|\leq\|\rho\|_{L^{\infty}}$, we obtain
$|a||u|+|b||\rho(\tau)|\leq(|a|+|b|)\|\rho\|_{L^{\infty}}\leq\varepsilon\|\rho^{\prime}\|_{L^{p}}+c_{\varepsilon}\|\rho\|_{L^{p}}.$
Since $Cx=(0,\rho^{\prime})$, we have $\|Cx\|_{\mathcal{E}_{p}}=\|\rho^{\prime}\|_{L^{p}}$ and hence
$\|Dx\|_{\mathcal{E}_{p}}=|au+b\rho(\tau)|\leq\varepsilon\|Cx\|_{\mathcal{E}_{p}}+c_{\varepsilon}\|x\|_{\mathcal{E}_{p}},$
which is the desired estimate. The relative bound is therefore 0. ∎
Corollary 1.3 (Semigroup well-posedness for autonomous DDEs).
For $1<p<\infty$, the operator $\mathcal{L}=C+D$ with domain $\mathcal{D}(C)$ generates a strongly continuous semigroup on $\mathcal{E}_{0}$.
Proof.
By Lemma 1.2, $D$ is CC–bounded with relative bound 0. The claim follows from the Miyadera–Voigt perturbation theorem (Pazy, 1983, Thm. III.1.3). ∎
1.5 Well-posedness of the continuous problem
1.5.1 Generation of the semigroup for C+D
Proposition 1.4 (Semigroup generation for $C+D$ on $\mathcal{E}_{p}$ ($1<p<\infty$)).
Let $\tau<0$ and $1<p<\infty$. Set
$\mathcal{E}_{p}:=X\oplus L^{p}([\tau,0];X),\qquad\|(u,\rho)\|_{\mathcal{E}_{p}}:=\|u\|_{X}+\|\rho\|_{L^{p}([\tau,0];X)}.$
Define the transport operator
$C(u,\rho):=(0,\partial_{\sigma}\rho),\qquad\mathcal{D}(C):=\bigl\{(u,\rho)\in X\oplus W^{1,p}([\tau,0];X):\ \rho(0)=u\bigr\}.$
Let $a,b\in\mathcal{L}(X)$ be bounded operators and define the delay/reaction operator
$D:\mathcal{D}(C)\to\mathcal{E}_{p},\qquad D(u,\rho):=\bigl(au+b\,\rho(\tau),\,0\bigr).$
Then:
  1. 1.
    $C$ generates a $C_{0}$–semigroup $(T(t))_{t\geq 0}$ on $\mathcal{E}_{p}$ (the left–shift semigroup on the history component with inflow coupling at $\sigma=0$).
  2. 2.
    $D$ is CC–bounded with relative bound 0, i.e. for every $\varepsilon>0$ there exists $c_{\varepsilon}>0$ such that
    $\|Dx\|_{\mathcal{E}_{p}}\leq\varepsilon\,\|Cx\|_{\mathcal{E}_{p}}+c_{\varepsilon}\,\|x\|_{\mathcal{E}_{p}},\qquad x\in\mathcal{D}(C).$
    In particular, $D$ is CC–bounded with relative bound <1<1.
  3. 3.
    Consequently, $C+D$ with domain $\mathcal{D}(C)$ generates a $C_{0}$–semigroup on $\mathcal{E}_{p}$.
Proof.
Step 1: $C$ generates the left–shift semigroup on $\mathcal{E}_{p}$. Define $(T(t))_{t\geq 0}$ on $\mathcal{E}_{p}$ by
$T(t)(u,\rho):=\bigl(u_{t},\rho_{t}\bigr),$
where $\rho_{t}$ is the left shift of $\rho$ by $t$ with inflow $u$ at $\sigma=0$:
$\rho_{t}(\sigma):=\begin{cases}\rho(\sigma+t),&\sigma+t\in[\tau,0],\\[5.69054pt] u,&\sigma+t>0,\end{cases}\qquad\sigma\in[\tau,0],$
and $u_{t}:=\rho_{t}(0)=u$. One checks that $T(t+s)=T(t)T(s)$, $T(0)=I$, and that $\|T(t)\|\leq 1$ on $\mathcal{E}_{p}$. Moreover, $T(t)x\to x$ in $\mathcal{E}_{p}$ as $t\downarrow 0$ for all $x\in\mathcal{E}_{p}$, so $(T(t))_{t\geq 0}$ is a $C_{0}$–semigroup. A standard computation shows that its generator is precisely $C$ with domain $\mathcal{D}(C)$ (see, e.g., (Pazy, 1983, Ch. III), Engel and Nagel (1999) or standard treatments of the shift semigroup on $L^{p}$ with an inflow boundary condition). Thus $C$ is a densely defined closed operator generating a $C_{0}$–semigroup on $\mathcal{E}_{p}$.
Step 2: A trace estimate on $W^{1,p}$ and CC–boundedness of $D$. Fix $\varepsilon>0$ and let $x=(u,\rho)\in\mathcal{D}(C)$. Since $1<p<\infty$, the Sobolev embedding yields $W^{1,p}([\tau,0];X)\hookrightarrow C([\tau,0];X)$, hence $\rho(\tau)$ is well-defined and
$\|\rho(\tau)\|_{X}\leq\|\rho\|_{L^{\infty}([\tau,0];X)}.$
We now use the (vector-valued) Gagliardo–Nirenberg inequality on the finite interval $\tau,0]$: there exists a constant $K=K(p,\tau)>0$ such that
$\|\rho\|_{L^{\infty}([\tau,0];X)}\leq K\,\|\rho^{\prime}\|_{L^{p}([\tau,0];X)}^{1/p}\,\|\rho\|_{L^{p}([\tau,0];X)}^{1-1/p}.$ (1.1)
(One may derive (1.1) from the scalar inequality applied to $\|\rho(\cdot)\|_{X}$ and a standard density argument; see, e.g., (Brezis, 2010, Sec. 7.28) or (Adams and Fournier, 2003, Ch. 1).)
Set $A:=\|\rho^{\prime}\|_{L^{p}}$ and $B:=\|\rho\|_{L^{p}}$. Then (1.1) reads $\|\rho\|_{L^{\infty}}\leq KA^{1/p}B^{1-1/p}$. By Young’s inequality in the form
$\alpha\beta\leq\frac{\delta}{r}\alpha^{r}+\frac{1}{s}\delta^{-s/r}\beta^{s},\qquad\delta>0,\quad r,s>1,\quad\frac{1}{r}+\frac{1}{s}=1,$
applied with $r=p$, s=p/(p1)s=p/(p-1), $\alpha=A^{1/p}$ and $\beta=B^{1-1/p}$, we obtain: for every $\delta>0$ there exists $C_{\delta}>0$ such that
$\|\rho\|_{L^{\infty}([\tau,0];X)}\leq\delta\,\|\rho^{\prime}\|_{L^{p}([\tau,0];X)}+C_{\delta}\,\|\rho\|_{L^{p}([\tau,0];X)}.$ (1.2)
Consequently,
$\|\rho(\tau)\|_{X}\leq\delta\,\|\rho^{\prime}\|_{L^{p}}+C_{\delta}\,\|\rho\|_{L^{p}}.$
Now estimate $Dx=(au+b\rho(\tau),0)$ in $\mathcal{E}_{p}$:
$\|Dx\|_{\mathcal{E}_{p}}=\|au+b\rho(\tau)\|_{X}\leq\|a\|\,\|u\|_{X}+\|b\|\,\|\rho(\tau)\|_{X}$
$\leq\|a\|\,\|u\|_{X}+\|b\|\Bigl(\delta\,\|\rho^{\prime}\|_{L^{p}}+C_{\delta}\,\|\rho\|_{L^{p}}\Bigr).$
Recall that $Cx=(0,\rho^{\prime})$, hence $\|Cx\|_{\mathcal{E}_{p}}=\|\rho^{\prime}\|_{L^{p}}$, and also $\|x\|_{\mathcal{E}_{p}}=\|u\|_{X}+\|\rho\|_{L^{p}}$. Thus
$\|Dx\|_{\mathcal{E}_{p}}\leq\|b\|\delta\,\|Cx\|_{\mathcal{E}_{p}}+\max\{\|a\|,\|b\|C_{\delta}\}\,\|x\|_{\mathcal{E}_{p}}.$
Choose $\delta>0$ so small that $\|b\|\delta\leq\varepsilon$ and set $c_{\varepsilon}:=\max\{\|a\|,\|b\|C_{\delta}\}$. Then
$\|Dx\|_{\mathcal{E}_{p}}\leq\varepsilon\,\|Cx\|_{\mathcal{E}_{p}}+c_{\varepsilon}\,\|x\|_{\mathcal{E}_{p}},\qquad x\in\mathcal{D}(C).$
Since $\varepsilon>0$ was arbitrary, $D$ is CC–bounded with relative bound 0.
Step 3: Semigroup generation for $C+D$. By Step 1, $C$ generates a $C_{0}$–semigroup on $\mathcal{E}_{p}$. By Step 2, $D$ is CC–bounded with relative bound 0, hence in particular with relative bound <1<1. Therefore, by the Miyadera–Voigt perturbation theorem (e.g. (Pazy, 1983, Thm. III.1.3), Miyadera (1966), Voigt (1977)), the operator $C+D$ with domain $\mathcal{D}(C)$ generates a $C_{0}$–semigroup on $\mathcal{E}_{p}$. ∎
Proposition 1.5 (Semigroup generation for $A_{0}+C+D$ on $\mathcal{E}_{p}$).
Let $\tau<0$ and $1<p<\infty$. Set
$\mathcal{E}_{p}:=X\oplus L^{p}([\tau,0];X),\qquad\|(u,\rho)\|_{\mathcal{E}_{p}}:=\|u\|_{X}+\|\rho\|_{L^{p}([\tau,0];X)}.$
Let $A:\mathcal{D}(A)\subset X\to X$ generate a $C_{0}$–semigroup $(S(t))_{t\geq 0}$ on $X$ (and in particular, $A$ is densely defined and closed). Define the block operator
$A_{0}:=\mathrm{diag}(A,0),\qquad\mathcal{D}(A_{0}):=\mathcal{D}(A)\oplus L^{p}([\tau,0];X).$
Define the transport operator
$C(u,\rho):=(0,\partial_{\sigma}\rho),\qquad\mathcal{D}(C):=\bigl\{(u,\rho)\in X\oplus W^{1,p}([\tau,0];X):\ \rho(0)=u\bigr\},$
and let $a,b\in\mathcal{L}(X)$ be bounded operators. Define
$D:\mathcal{D}(C)\to\mathcal{E}_{p},\qquad D(u,\rho):=\bigl(au+b\,\rho(\tau),\,0\bigr).$
Set
$G:=A_{0}+C,\qquad\mathcal{D}(G):=\mathcal{D}(A_{0})\cap\mathcal{D}(C),\qquad L:=A_{0}+C+D,\qquad\mathcal{D}(L):=\mathcal{D}(G).$
Then the following hold.
  1. 1.
    $G=A_{0}+C$ generates a $C_{0}$–semigroup $(T(t))_{t\geq 0}$ on $\mathcal{E}_{p}$.
  2. 2.
    $D$ is GG–bounded with relative bound 0, hence in particular with relative bound <1<1.
  3. 3.
    Consequently, $L=A_{0}+C+D$ with domain $\mathcal{D}(G)$ generates a $C_{0}$–semigroup on $\mathcal{E}_{p}$.
References (typical): for the product-space semigroup for delay equations see, e.g., Bátkai and Piazzera (2005), Engel and Nagel (1999); for the perturbation step use Miyadera–Voigt as in (Pazy, 1983, Thm. III.1.3).
Proof.
Step 1: $A_{0}+C$ generates a $C_{0}$–semigroup. Define for $t\geq 0$ an operator $T(t)$ on $\mathcal{E}_{p}$ by
$T(t)(u,\rho):=\bigl(S(t)u,\ \rho_{t}\bigr),$
where $\rho_{t}:[\tau,0]\to X$ is given by the “shift with inflow” rule
$\rho_{t}(\sigma):=\begin{cases}\rho(\sigma+t),&\sigma+t\in[\tau,0],\\[2.84526pt] S(\sigma+t)u,&\sigma+t\in(0,t].\end{cases}$ (1.3)
(So the inflow at $\sigma=0$ is $\rho_{t}(0)=S(t)u$, and the old history is shifted left as long as it stays inside $\tau,0]$.)
We verify that $(T(t))_{t\geq 0}$ is a $C_{0}$–semigroup:
  • Semigroup property. Fix $s,t\geq 0$. Using $S(t+s)=S(t)S(s)$ and the definition (1.3), a direct case distinction on $\sigma+t+s\in[\tau,0]$ or $\sigma+t+s>0$ yields $\rho_{t+s}=(\rho_{s})_{t}$ and hence $T(t+s)=T(t)T(s)$.
  • Strong continuity. Since $S(t)u\to u$ in $X$ as $t\downarrow 0$, it remains to show $\rho_{t}\to\rho$ in $L^{p}$. On the region where $\sigma+t\in[\tau,0]$ we have $\rho_{t}(\sigma)=\rho(\sigma+t)$, which converges to $\rho(\sigma)$ in $L^{p}$ by translation-continuity on a finite interval. The remaining piece where $\sigma+t>0$ has measure $t$ and satisfies $\|\rho_{t}(\sigma)\|_{X}\leq\sup_{0\leq r\leq t}\|S(r)\|\,\|u\|_{X}$; hence its $L^{p}$-norm vanishes as $t\downarrow 0$. Therefore $\rho_{t}\to\rho$ in $L^{p}$ and thus $T(t)x\to x$ in $\mathcal{E}_{p}$.
Hence $(T(t))_{t\geq 0}$ is a $C_{0}$–semigroup on $\mathcal{E}_{p}$.
We now identify its generator. Let $x=(u,\rho)\in\mathcal{D}(G)=\mathcal{D}(A_{0})\cap\mathcal{D}(C)$, so $u\in\mathcal{D}(A)$, $\rho\in W^{1,p}$ and $\rho(0)=u$. A standard differentiation of (1.3) at $t=0$ gives
$\frac{d}{dt}\Big|_{t=0}S(t)u=Au,\qquad\frac{d}{dt}\Big|_{t=0}\rho_{t}(\sigma)=\partial_{\sigma}\rho(\sigma)\ \ \text{in }L^{p},$
and therefore the generator equals $A_{0}+C$ on $\mathcal{D}(G)$. Thus $G=A_{0}+C$ generates $(T(t))_{t\geq 0}$.
Step 2: $D$ is GG–bounded with relative bound 0. Fix $\varepsilon>0$ and let $x=(u,\rho)\in\mathcal{D}(G)$. By the $W^{1,p}\hookrightarrow C$ embedding and the same Gagliardo–Nirenberg estimate as in Lemma 1.2, there exists (for every $\delta>0$) a constant $C_{\delta}>0$ such that
$\|\rho(\tau)\|_{X}\leq\|\rho\|_{L^{\infty}}\leq\delta\,\|\rho^{\prime}\|_{L^{p}}+C_{\delta}\,\|\rho\|_{L^{p}}.$ (1.4)
Hence
$\|Dx\|_{\mathcal{E}_{p}}=\|au+b\rho(\tau)\|_{X}\leq\|a\|\,\|u\|_{X}+\|b\|\,\|\rho(\tau)\|_{X}$
$\leq\|b\|\delta\,\|\rho^{\prime}\|_{L^{p}}+\|a\|\,\|u\|_{X}+\|b\|C_{\delta}\,\|\rho\|_{L^{p}}.$
We next relate $\|\rho^{\prime}\|_{L^{p}}$ to $\|Gx\|_{\mathcal{E}_{p}}$. Since
$Gx=(Au,\rho^{\prime}),\qquad\|Gx\|_{\mathcal{E}_{p}}=\|Au\|_{X}+\|\rho^{\prime}\|_{L^{p}},$
we have the trivial bound $\|\rho^{\prime}\|_{L^{p}}\leq\|Gx\|_{\mathcal{E}_{p}}$. Also $\|u\|_{X}+\|\rho\|_{L^{p}}=\|x\|_{\mathcal{E}_{p}}$. Therefore,
$\|Dx\|_{\mathcal{E}_{p}}\leq\|b\|\delta\,\|Gx\|_{\mathcal{E}_{p}}+\max\{\|a\|,\|b\|C_{\delta}\}\,\|x\|_{\mathcal{E}_{p}}.$
Choose $\delta>0$ so small that $\|b\|\delta\leq\varepsilon$ and set $c_{\varepsilon}:=\max\{\|a\|,\|b\|C_{\delta}\}$. Then
$\|Dx\|_{\mathcal{E}_{p}}\leq\varepsilon\,\|Gx\|_{\mathcal{E}_{p}}+c_{\varepsilon}\,\|x\|_{\mathcal{E}_{p}},\qquad x\in\mathcal{D}(G).$
Since $\varepsilon>0$ was arbitrary, $D$ is GG–bounded with relative bound 0.
Step 3: Perturbation and generation of $A_{0}+C+D$. By Step 1, $G=A_{0}+C$ generates a $C_{0}$–semigroup on $\mathcal{E}_{p}$. By Step 2, $D$ is GG–bounded with relative bound 0, hence with relative bound <1<1. Therefore the Miyadera–Voigt perturbation theorem (e.g. (Pazy, 1983, Thm. III.1.3)) implies that $G+D$ with domain $\mathcal{D}(G)$ generates a $C_{0}$–semigroup on $\mathcal{E}_{p}$. But $G+D=A_{0}+C+D=L$, so $L$ generates a $C_{0}$–semigroup, as claimed. ∎
1.5.2 Non-autonomous case and evolution families
When the coefficients $a(t),b(t)$ depend on time, the associated operator family
$\mathcal{L}(t):=C+D(t)$
defines a non-autonomous evolution equation
$x^{\prime}(t)=\mathcal{L}(t)x(t).$
Lipschitz continuity of $t\mapsto D(t)$ in the CC–graph norm is sufficient for the discrete analysis developed later. However, the existence of an evolution family generally requires stronger time regularity. Accordingly, whenever we appeal to the continuous theory of evolution families, we impose the classical Kato–Tanabe assumption that for each $x\in\mathcal{D}(C)$, the mapping $t\mapsto\mathcal{L}(t)x$ belongs to $C^{1}([0,T];\mathcal{E}_{0})$ (cf. (Pazy, 1983, Thm. 5.3), (Tanabe, 1979, Ch. 7)).
Under this assumption, the family $\mathcal{L}(t)$ generates a unique evolution family $U(t,s)$ satisfying
$\frac{d}{dt}U(t,s)x=\mathcal{L}(t)U(t,s)x,\qquad U(s,s)=I,$
with the growth bound
$\|U(t,s)\|\leq Me^{\omega(t-s)}.$
Remark 1.6.
Time regularity and evolution families. The Lipschitz continuity introduced above is sufficient for the discrete analysis developed in this paper. However, it is in general not sufficient to guarantee the existence of an evolution family for the continuous non-autonomous problem.
Whenever the existence of an evolution family is required, we therefore impose the following additional regularity assumption, in accordance with the classical Kato–Tanabe theory of non-autonomous evolution equations.
For every $x\in\mathcal{D}(C)$, the mapping $t\mapsto\mathcal{L}(t)x$ belongs to $C^{1}([0,T];\mathcal{E}_{0})$.
Under this assumption, the operator family
$\mathcal{L}(t)=C+B(t)\qquad(\text{or more generally }\mathcal{L}(t)=A_{0}+C+B(t))$
with common domain $\mathcal{D}(C)$ generates a unique evolution family $\{U(t,s):0\leq s\leq t\leq T\}\subset\mathcal{L}(\mathcal{E}_{0})$ satisfying
$\frac{d}{dt}U(t,s)x=\mathcal{L}(t)U(t,s)x,\qquad U(s,s)=I.$
Moreover, there exist constants $M,\omega>0$ such that
$\|U(t,s)\|\leq Me^{\omega(t-s)},\qquad 0\leq s\leq t\leq T.$
The combination of sectoriality, perturbation theory, interpolation scales, and sufficient time regularity of $\mathcal{L}(t)$ provides the continuous well-posedness framework when needed. Importantly, the core discrete resolvent analysis developed in this paper does not rely on the existence of such an evolution family, and remains meaningful even beyond the classical well-posedness regime.
1.5.3 Evolution family for the non-autonomous problem
We record a convenient generation result for the non-autonomous operator family
$L(t)=A_{0}+C+D(t),\qquad t\in[0,T],$
under the standard Kato–Tanabe (“hyperbolic”) assumptions. A systematic treatment can be found in Tanabe’s monograph on evolution equations (and also in Pazy’s book); see, e.g., (Pazy, 1983, Ch. 5), (Tanabe, 1979, Ch. 4). For a modern reference that explicitly cites Tanabe’s book in this context, see, e.g., Hishida (2004).
1.5.4 Kato stability and $C^{1}$-regularity in the graph norm
Let $X$ be a Banach space and $D\subset X$ a dense subspace. For each $t\in[0,T]$ let $A(t):D\to X$ be a closed operator. Write $\|\cdot\|_{D}$ for a Banach norm on $D$ equivalent to the graph norm $\|x\|+\|A(t)x\|$ (uniformly in $t$).
Definition 1.7 (Kato stability on a common domain).
A family $\{A(t)\}_{t\in[0,T]}$ on a common domain $D$ is called (uniformly) Kato stable if there exist constants $M\geq 1$ and $\omega\in\mathbb{R}$ such that for every $m\in\mathbb{N}$ and every partition $0\leq t_{1}\leq t_{2}\leq\cdots\leq t_{m}\leq T$ one has
$\bigl\|(\lambda-A(t_{m}))^{-1}\cdots(\lambda-A(t_{1}))^{-1}\bigr\|\leq M(\lambda-\omega)^{-m},\qquad\lambda>\omega.$
Definition 1.8 ($C^{1}$-regularity in the graph norm).
We say that $A(\cdot)$ is $C^{1}$ as a map $0,T]\to\mathcal{L}(D,X)$ (with respect to $\|\cdot\|_{D}$) if for every $x\in D$ the map $t\mapsto A(t)x$ is continuously differentiable in $X$ and
$\sup_{t\in[0,T]}\|A(t)\|_{\mathcal{L}(D,X)}<\infty,\qquad\sup_{t\in[0,T]}\|A^{\prime}(t)\|_{\mathcal{L}(D,X)}<\infty.$
1.5.5 Evolution family generation theorem (Kato–Tanabe)
Theorem 1.9 (Kato–Tanabe evolution family on a common domain).
Let $\{A(t)\}_{t\in[0,T]}$ be a family of closed operators on $X$ with a common dense domain $D$. Assume:
  1. (H1)
    (uniform Kato stability) $\{A(t)\}$ is Kato stable in the sense of Definition 1.7;
  2. (H2)
    (uniform graph norms) there exists a Banach norm $\|\cdot\|_{D}$ on $D$ such that for all $t\in[0,T]$,
    $\|x\|_{D}\simeq\|x\|+\|A(t)x\|\qquad(x\in D),$
    with constants independent of $t$;
  3. (H3)
    ($C^{1}$-regularity) $A(\cdot)\in C^{1}([0,T];\mathcal{L}(D,X))$ in the sense of Definition 1.8.
Then there exists a unique evolution family $\{U(t,s)\}_{0\leq s\leq t\leq T}\subset\mathcal{L}(X)$ such that:
  1. 1.
    $U(s,s)=I$ and $U(t,r)U(r,s)=U(t,s)$ for $s\leq r\leq t$;
  2. 2.
    for each $x\in D$, the map $t\mapsto U(t,s)x$ is differentiable and
    $\partial_{t}U(t,s)x=A(t)U(t,s)x,\qquad t\in[s,T];$
  3. 3.
    there exist $M\geq 1$ and $\omega\in\mathbb{R}$ such that
    $\|U(t,s)\|_{\mathcal{L}(X)}\leq Me^{\omega(t-s)},\qquad 0\leq s\leq t\leq T.$
Remark 1.10.
Theorem 1.9 is the standard “hyperbolic” non-autonomous generation result attributed to Kato and presented in Tanabe’s monograph. See, e.g., (Pazy, 1983, Ch. 5), (Tanabe, 1979, Ch. 4).
Application to $L(t)=A_{0}+C+D(t)$
We now verify the Kato–Tanabe hypotheses for our delay equation framework.
Proposition 1.11 (Evolution family for the non-autonomous delay operator).
Let $1<p<\infty$ and $\tau<0$, and set
$\mathcal{E}_{0}:=X\oplus L^{p}([\tau,0];X),\qquad\mathcal{D}(C):=\{(u,\rho)\in X\oplus W^{1,p}([\tau,0];X):\rho(0)=u\}.$
Let $C(u,\rho):=(0,\partial_{\sigma}\rho)$ on $\mathcal{D}(C)$. Let $A$ be sectorial on $X$ and let $A_{0}:=\mathrm{diag}(A,0)$ on
$\mathcal{D}(A_{0}):=\mathcal{D}(A)\oplus L^{p}([\tau,0];X).$
Assume that the coefficient functions $a(\cdot),b(\cdot)\in C^{1}([0,T])$ are bounded together with their derivatives, and define $D(t):\mathcal{D}(C)\to\mathcal{E}_{0}$ by
$D(t)(u,\rho)=\bigl(a(t)u+b(t)\rho(\tau),\,0\bigr).$
Assume moreover that $D(t)$ is $(A_{0}+C)$–bounded with relative bound <1<1 uniformly in $t$, i.e., there exist $\alpha\in[0,1)$ and $c\geq 0$ such that for all $t\in[0,T]$,
$\|D(t)x\|_{\mathcal{E}_{0}}\leq\alpha\|(A_{0}+C)x\|_{\mathcal{E}_{0}}+c\|x\|_{\mathcal{E}_{0}},\qquad x\in\mathcal{D}(A_{0})\cap\mathcal{D}(C).$
(For the scalar/non-sectorial case take $A_{0}=0$.)
Then the family
$L(t):=A_{0}+C+D(t),\qquad\mathcal{D}(L(t))=\mathcal{D}(A_{0})\cap\mathcal{D}(C),$
generates a unique evolution family $\{U(t,s)\}_{0\leq s\leq t\leq T}$ on $\mathcal{E}_{0}$. Moreover, there exist constants $M\geq 1$ and $\omega\in\mathbb{R}$ such that
$\|U(t,s)\|_{\mathcal{L}(\mathcal{E}_{0})}\leq Me^{\omega(t-s)},\qquad 0\leq s\leq t\leq T.$
Proof.
Set $X_{0}:=\mathcal{E}_{0}$ and $D_{0}:=\mathcal{D}(A_{0})\cap\mathcal{D}(C)$.
Step 1: Each $L(t)$ generates a $C_{0}$-semigroup with uniform bounds. Since $A_{0}$ and $C$ act on different components, $A_{0}+C$ is closed on $D_{0}$. In the sectorial case, $A_{0}$ generates a bounded analytic semigroup on $\mathcal{E}_{0}$ (on the first component and trivially on the second), and $C$ generates the left-shift semigroup on the history component; hence $A_{0}+C$ generates a $C_{0}$-semigroup on $\mathcal{E}_{0}$. (For the non-sectorial case simply take $A_{0}=0$.)
By assumption, $D(t)$ is $(A_{0}+C)$–bounded with relative bound $\alpha<1$ uniformly in $t$. Therefore, by the Miyadera–Voigt perturbation theorem (or equivalently, the relative boundedness perturbation theorem for generators), each perturbed operator $L(t)=(A_{0}+C)+D(t)$ generates a $C_{0}$-semigroup on $\mathcal{E}_{0}$. Moreover, the constants in the corresponding Hille–Yosida estimates can be chosen uniformly in $t$ because the relative bound $\alpha$ and the constant $c$ are uniform. Consequently, there exist $M_{0}\geq 1$ and $\omega_{0}\in\mathbb{R}$ such that
$\|(\lambda-L(t))^{-1}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq\frac{M_{0}}{\lambda-\omega_{0}},\qquad\lambda>\omega_{0},\ \ t\in[0,T].$ (1.5)
Step 2: Kato stability (H1). Uniform resolvent bounds (1.5) imply Kato stability (Definition 1.7) with the same constants $M_{0},\omega_{0}$: for any $0\leq t_{1}\leq\cdots\leq t_{m}\leq T$ and $\lambda>\omega_{0}$,
$\bigl\|(\lambda-L(t_{m}))^{-1}\cdots(\lambda-L(t_{1}))^{-1}\bigr\|\leq\prod_{j=1}^{m}\frac{M_{0}}{\lambda-\omega_{0}}=M_{0}^{m}(\lambda-\omega_{0})^{-m}.$
Absorbing $M_{0}^{m}$ into the standard definition (or using the equivalent formulation adopted in Tanabe (1979), Pazy (1983)) yields (H1). (This is the standard route in Kato–Tanabe theory: uniform Hille–Yosida bounds imply stability.)
Step 3: Uniform equivalence of graph norms (H2). Fix $\lambda>\omega_{0}$. Since $\lambda\in\rho(L(t))$ for all $t$ and the resolvents are uniformly bounded, the norms
$\|x\|_{D_{0},t}:=\|x\|+\|L(t)x\|\quad\text{and}\quad\|x\|_{D_{0}}:=\|x\|+\|(A_{0}+C)x\|$
are uniformly equivalent on $D_{0}$. Indeed, from $L(t)=(A_{0}+C)+D(t)$ and $(A_{0}+C)$–boundedness of $D(t)$ with relative bound $\alpha<1$, we obtain
$\|L(t)x\|\leq\|(A_{0}+C)x\|+\|D(t)x\|\leq(1+\alpha)\|(A_{0}+C)x\|+c\|x\|,$
and conversely,
$\|(A_{0}+C)x\|\leq\|L(t)x\|+\|D(t)x\|\leq\|L(t)x\|+\alpha\|(A_{0}+C)x\|+c\|x\|.$
Hence
$(1-\alpha)\|(A_{0}+C)x\|\leq\|L(t)x\|+c\|x\|,$
which yields the reverse inequality. Collecting both estimates shows that the graph norms are equivalent with constants independent of $t$. Taking $\|\cdot\|_{D_{0}}$ as the reference norm proves (H2).
Step 4: $C^{1}$-regularity in $\mathcal{L}(D_{0},\mathcal{E}_{0})$ (H3). Since $a,b\in C^{1}([0,T])$ and the trace map $\rho\mapsto\rho(\tau)$ is bounded on $W^{1,p}([\tau,0];X)$, the operator $D(t)$ defines a bounded map from $D_{0}$ (with the graph norm of $A_{0}+C$) into $\mathcal{E}_{0}$. Moreover, for each fixed $x=(u,\rho)\in D_{0}$,
$t\mapsto D(t)x=(a(t)u+b(t)\rho(\tau),0)$
is $C^{1}$ in $\mathcal{E}_{0}$, with
$\frac{d}{dt}D(t)x=(a^{\prime}(t)u+b^{\prime}(t)\rho(\tau),0).$
Because $a^{\prime},b^{\prime}$ are bounded, we obtain
$\sup_{t\in[0,T]}\|D(t)\|_{\mathcal{L}(D_{0},\mathcal{E}_{0})}<\infty,\qquad\sup_{t\in[0,T]}\|D^{\prime}(t)\|_{\mathcal{L}(D_{0},\mathcal{E}_{0})}<\infty,$
which yields $D(\cdot)\in C^{1}([0,T];\mathcal{L}(D_{0},\mathcal{E}_{0}))$. Since $A_{0}+C$ is independent of $t$, it follows that $L(\cdot)=A_{0}+C+D(\cdot)\in C^{1}([0,T];\mathcal{L}(D_{0},\mathcal{E}_{0}))$. Thus (H3) holds.
Step 5: Apply the Kato–Tanabe theorem. By Steps 2–4, the family $\{L(t)\}$ satisfies (H1)–(H3) of Theorem 1.9 on $X_{0}=\mathcal{E}_{0}$ with common domain $D_{0}$. Therefore, there exists a unique evolution family $\{U(t,s)\}_{0\leq s\leq t\leq T}$ on $\mathcal{E}_{0}$ solving $\partial_{t}U(t,s)x=L(t)U(t,s)x$ for $x\in D_{0}$, and satisfying the growth bound $\|U(t,s)\|\leq Me^{\omega(t-s)}$. ∎
Remark 1.12 (Why Lipschitz is not enough here).
In Kato–Tanabe hyperbolic theory, Lipschitz continuity of $t\mapsto L(t)$ in the graph norm is often insufficient to ensure differentiability of $t\mapsto U(t,s)x$ for $x$ in the common domain, and the standard convenient criterion is $C^{1}$-regularity in $\mathcal{L}(D,X)$ (as used above); see Tanabe (1979); Pazy (1983). This is the reason why we impose $a,b\in C^{1}([0,T])$ when we need evolution-family generation.
1.5.6 Fractional power and interpolation spaces
In order to measure the regularity of solutions and to formulate sharp convergence estimates for numerical schemes, we work with fractional domain spaces and interpolation spaces associated with sectorial operators. We follow the standard conventions of Lunardi (1995).
Definition 1.13 (Sectorial operator (Lunardi)).
Let $\mathcal{E}$ be a Banach space. A linear operator $A:\mathcal{D}(A)\subset\mathcal{E}\to\mathcal{E}$ is called sectorial if there exist constants $\omega\in\mathbb{R}$, $\vartheta\in(\pi/2,\pi)$, and $M>0$ such that
$\rho(A)\supset S_{\vartheta,\omega}=\{\lambda\in\mathbb{C}:\lambda\neq\omega,|\arg(\lambda-\omega)|<\vartheta\}$
and
$\|R(\lambda,A)\|_{\mathcal{L}(\mathcal{E})}\leq\frac{M}{|\lambda-\omega|},\quad\forall\lambda\in S_{\vartheta,\omega}.$
(See (Lunardi, 1995, Definition 2.0.1).)
Definition 1.14 (Fractional power spaces).
Let $A$ be a sectorial operator on a Banach space $X$ with $0\in\rho(A)$. For $0<\theta<1$, the fractional power $A^{\theta}$ is defined via the Balakrishnan integral. We define the fractional power space by
$\mathcal{D}(A^{\theta}):=\{x\in X:A^{\theta}x\in X\},$
endowed with the graph norm
$\|x\|_{\mathcal{D}(A^{\theta})}:=\|x\|+\|A^{\theta}x\|.$
Since $0\in\rho(A)$, this norm is equivalent to $\|A^{\theta}x\|$. See, e.g., (Lunardi, 1995, Section 1.14), Engel and Nagel (1999); Haase (2006) for details.
Definition 1.15 (Real interpolation spaces).
Let $(X,\mathcal{D}(A))$ be the Banach couple associated with $A$. For $0<\theta<1$ and $1\leq p\leq\infty$, we define the real interpolation space
$(X,\mathcal{D}(A))_{\theta,p}$
via the KK–method. Throughout this paper we fix $p=1$ and define
$X_{\theta}:=(X,\mathcal{D}(A))_{\theta,1}.$
See, e.g., Lunardi (1995) for details.
Theorem 1.16 (Fractional powers and real interpolation).
Let $A$ be a sectorial operator on a Banach space $X$ with $0\in\rho(A)$. Then for every $0<\theta<1$ one has continuous embeddings
$(X,\mathcal{D}(A))_{\theta,1}\hookrightarrow\mathcal{D}(A^{\theta})\hookrightarrow(X,\mathcal{D}(A))_{\theta,\infty}.$
Moreover, there exist constants $C_{\theta}>0$ such that
$\|A^{\theta}x\|\leq C_{\theta}\|x\|_{(X,\mathcal{D}(A))_{\theta,1}},\qquad\|x\|_{(X,\mathcal{D}(A))_{\theta,\infty}}\leq C_{\theta}\|A^{\theta}x\|.$
See, e.g., (Lunardi, 1995, Theorem 1.15.3) for details.
Remark 1.17.
Throughout the paper we set
$X_{\theta}:=(X,\mathcal{D}(A))_{\theta,1}.$
By Theorem 1.16, we have the continuous embedding $X_{\theta}\hookrightarrow\mathcal{D}(A^{\theta})$, and we shall often identify elements of $X_{\theta}$ with elements of $\mathcal{D}(A^{\theta})$ via this embedding. See, e.g., (Lunardi, 1995, Theorem 1.15.3) for details.
1.5.7 Connection to numerical schemes
The abstract formulation developed above serves as the continuous reference framework for the numerical analysis carried out in the subsequent chapters. In the autonomous case, implicit Euler discretization approximates the semigroup generated by $\mathcal{L}=C+D$. In the non-autonomous case, it approximates the evolution family generated by $\mathcal{L}(t)$ when such a family exists.
More importantly, the operator-theoretic formulation allows us to define the implicit Euler and Lie–Trotter splitting schemes directly at the level of resolvents, which remains meaningful even in situations where continuous well-posedness fails. This perspective underlies the entire discrete resolvent framework developed in Chapter 2.
Remark 1.18 (How the discrete resolvent framework interfaces with semigroups).
The main results of this thesis (Theorems A–D) compare two discrete propagators: the Lie–Trotter splitting step $P_{h}(t)$ and the (frozen) implicit Euler step $R_{h}(t)$. No continuous evolution object is required for their formulation.
Whenever the continuous operator family admits a well-posed evolution (either a $C_{0}$-semigroup in the autonomous case, or an evolution family $U(t,s)$ in the non-autonomous case), one can recover error bounds with respect to the continuous solution by a simple triangle argument. For instance, in the autonomous case,
$\|e^{t_{n}L}-P_{h}^{n}\|\leq\|e^{t_{n}L}-R_{h}^{n}\|+\|R_{h}^{n}-P_{h}^{n}\|.$
Here the second term is controlled by Theorem A/B/C/D, while the first term is the classical implicit Euler approximation error for the underlying semigroup (see, e.g., standard references on backward Euler and semigroup approximation).
An analogous decomposition applies in the non-autonomous case:
$\|U(t_{n},0)-P_{h}(t_{n-1})\cdots P_{h}(t_{0})\|\leq\|U(t_{n},0)-R_{h}(t_{n-1})\cdots R_{h}(t_{0})\|+\Bigl\|\prod_{k=0}^{n-1}R_{h}(t_{k})-\prod_{k=0}^{n-1}P_{h}(t_{k})\Bigr\|,$
where the last term is precisely the quantity treated by the discrete telescoping analysis in this thesis, and the first term is covered by classical frozen-step/backward-Euler approximation results for evolution families under Kato/Acquistapace–Terreni type hypotheses.
Remark 1.19 (Link to classical well-posedness).
Assume that $L(t)=A_{0}+C+D(t)$ generates an evolution family $U(t,s)$ on $\mathcal{E}_{0}$ (e.g. under Kato stability / Acquistapace–Terreni type hypotheses Acquistapace and Terreni (1987)). Then the frozen implicit Euler products $R_{h}(t_{n-1})\cdots R_{h}(t_{0})$ converge to $U(t_{n},0)$ as $h\to 0$. See, e.g., Kobayashi (2003, 2005), Tanaka (1994). Combining this with Theorems A–D yields
$\|U(t_{n},0)-P_{h}(t_{n-1})\cdots P_{h}(t_{0})\|\leq\|U(t_{n},0)-R_{h}(t_{n-1})\cdots R_{h}(t_{0})\|+\|R\text{-prod}-P\text{-prod}\|.$
2 Discrete resolvent framework and abstract convergence theory
2.1 Discrete resolvent operators
First, we assume that for each $t\in[0,T]$ and $h>0$ the composition $D(t)S_{h}^{C}$ extends to a bounded operator on $\mathcal{E}_{0}$ (and similarly $D(t)S_{h}^{AC}$ in the sectorial case). This is the case in all applications considered in this thesis.
Right-preconditioned delay operator. We define the (right-)preconditioned operator
$H_{h}(t):=h\,D(t)S_{h}^{C}\qquad(\text{or }H_{h}(t):=h\,D(t)S_{h}^{AC}\text{ in the sectorial case}).$
The role of (OS) is to guarantee that the one-step splitting propagator introduced below is well defined and uniformly bounded for small $h$.
Assumption 2.1 (Operational smallness).
There exist $h_{0}>0$ and $q\in(0,1)$ such that
$\sup_{t\in[0,T]}\ \sup_{0<h\leq h_{0}}\|H_{h}(t)\|_{\mathcal{L}(\mathcal{E}_{0})}\leq q.$
This assumption has a purely discrete character: it is introduced only to guarantee that the one-step splitting propagator is well defined and uniformly bounded for sufficiently small step sizes. It does not rely on any continuous well-posedness properties of the underlying abstract evolution equation.
2.1.1 Connection with continuous well-posedness
We emphasize that the abstract discrete framework introduced above is compatible with the classical semigroup approach whenever the latter applies.
Proposition 2.2 (Operational smallness induced by relative boundedness).
Assume that the operator $D:\mathcal{D}(C)\to\mathcal{E}_{0}$ is CC–bounded with relative bound $\alpha<1$. Then there exist constants $h_{0}>0$ and $q\in(0,1)$ such that
$\|hDS_{h}^{C}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq q\qquad\text{for all }0<h\leq h_{0}.$
In particular, the operational smallness condition (OS) holds.
Proof.
Since $D$ is CC–bounded with relative bound $\alpha<1$, there exists $c>0$ such that
$\|Dx\|\leq\alpha\|Cx\|+c\|x\|,\qquad x\in\mathcal{D}(C).$
Applying this estimate to $x=S_{h}^{C}y$ and using the resolvent bounds
$\|CS_{h}^{C}\|\lesssim h^{-1},\qquad\|S_{h}^{C}\|\leq 1,$
we obtain
$\|hDS_{h}^{C}y\|\leq h\alpha\|CS_{h}^{C}y\|+hc\|S_{h}^{C}y\|\leq\alpha\|y\|+hc\|y\|.$
Choosing $h_{0}>0$ sufficiently small yields the claim. ∎
Relation to semigroup well-posedness.
Whenever the continuous operator $C+D$ (or $A_{0}+C+D$) generates a $C_{0}$-semigroup, the relative boundedness of $D$ with respect to $C$ implies the operational smallness condition (OS) through Proposition 2.2. Thus, all subsequent discrete constructions are consistent with the classical semigroup framework.
However, the abstract theory developed in this chapter is formulated solely in terms of the discrete resolvents and does not require the existence of an underlying semigroup. This allows us to treat regimes in which continuous well-posedness may fail, while retaining compatibility with the classical setting whenever it is available.
In this section we introduce the discrete resolvent operators associated with implicit Euler and Lie–Trotter splitting. A key point is that, since the delay operator $D(t)$ is only defined on $\mathcal{D}(C)$ and does not extend to a bounded operator on $\mathcal{E}_{0}$, we must work with preconditioned operators of the form $D(t)S_{h}^{C}$ rather than with $D(t)$ itself.
2.1.2 Resolvent of the transport operator
Let $C:\mathcal{D}(C)\subset\mathcal{E}_{0}\to\mathcal{E}_{0}$ be the transport operator from Section 1.1. Its resolvent
$S_{h}^{C}:=(I-hC)^{-1}\in\mathcal{L}(\mathcal{E}_{0})$
is well-defined for every $h>0$ and admits an explicit representation. Moreover, $S_{h}^{C}$ satisfies the smoothing estimates
$\|S_{h}^{C}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq 1,\qquad\|S_{h}^{C}\|_{\mathcal{L}(X_{\theta},\mathcal{D}(C))}\lesssim h^{-\theta},\qquad 0<\theta\leq 1,$
where $X_{\theta}=(\mathcal{E}_{0},\mathcal{D}(C))_{\theta,1}$ denotes the real interpolation space.
In the sectorial case, where a sectorial operator $A_{0}=\mathrm{diag}(A,0)$ is present, we also introduce its resolvent
$S_{h}^{A}:=(I-hA_{0})^{-1}\in\mathcal{L}(\mathcal{E}_{0}),$
and define the combined preconditioner
$S_{h}^{AC}:=S_{h}^{C}S_{h}^{A}.$
This corresponds to applying first the elliptic resolvent and then the transport resolvent.
2.1.3 Implicit Euler: full resolvent
For the full operator
$L(t)=A_{0}+C+D(t),\qquad\mathcal{D}(L(t))=\mathcal{D}(A_{0})\cap\mathcal{D}(C),$
the one-step implicit Euler map (when it exists) is given by the full resolvent
$R_{h}(t):=(I-hL(t))^{-1}=(I-h(A_{0}+C+D(t)))^{-1}.$ (2.1)
In the autonomous case ($D(t)\equiv D$) we simply write $R_{h}$. In the non-autonomous case, $R_{h}(t_{k})$ denotes the frozen-step resolvent at time $t_{k}$.
We emphasize that the existence of $R_{h}(t)$ is not assumed a priori in the entire paper; in several regimes (notably beyond semigroup well-posedness) $R_{h}(t)$ may fail to exist, while the splitting propagator introduced below remains well-defined.
2.1.4 Lie–Trotter splitting: right-preconditioned form
Since $D(t)$ is not a bounded operator on $\mathcal{E}_{0}$, the expression $(I-hD(t))^{-1}$ is meaningless in general. Instead, we consider the right-preconditioned operator
$H_{h}(t):=h\,D(t)S_{h}^{C}\qquad(\text{or }H_{h}(t):=h\,D(t)S_{h}^{AC}\text{ in the sectorial case}),$
whenever the composition $D(t)S_{h}^{C}$ extends to a bounded operator on $\mathcal{E}_{0}$.
The one-step Lie–Trotter propagator is then defined by
$P_{h}(t):=(I-H_{h}(t))^{-1}S_{h}^{C}\qquad(\text{or }P_{h}(t):=(I-H_{h}(t))^{-1}S_{h}^{AC}).$ (2.2)
Formally, this corresponds to first applying the transport (and diffusion, if present) resolvent, and then correcting by the delay/reaction part through a Neumann-series-type expansion.
Note that the factor $(I-H_{h}(t))^{-1}$ acts after the transport resolvent; this reflects that $D(t)$ is only controlled on $\mathcal{D}(C)$ and therefore must be paired with the regularization provided by $S_{h}^{C}$.
2.1.5 Implicit Euler vs. Lie–Trotter: local resolvent defect
The two one-step maps (2.1) and (2.2) generally do not coincide, even in the autonomous case. Their discrepancy is the fundamental object governing the convergence analysis.
We therefore introduce the local resolvent defect
$E_{h}(t):=R_{h}(t)-P_{h}(t),$
whenever both operators are defined. The remainder of the analysis is devoted to:
  • estimating $\|E_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}$ for small $h$,
  • and propagating this local defect to global error bounds under suitable discrete stability assumptions on products of $P_{h}(t)$.
2.2 Discrete stability and propagation of local defects
The local resolvent defect $E_{h}(t)$ controls the one-step discrepancy between implicit Euler and Lie–Trotter splitting. To obtain global error bounds over multiple time steps, it is necessary to understand how this defect propagates under repeated application of the splitting propagator.
This leads naturally to the introduction of discrete stability assumptions on products of the operators $P_{h}(t)$.
2.2.1 Telescoping identity
In the autonomous case ($D(t)\equiv D$), we write $P_{h}(t)=P_{h}$ and $R_{h}(t)=R_{h}$. Then for every $n\in\mathbb{N}$ the following identity holds:
$R_{h}^{n}-P_{h}^{n}=\sum_{k=0}^{n-1}R_{h}^{n-1-k}(R_{h}-P_{h})P_{h}^{k}=\sum_{k=0}^{n-1}R_{h}^{n-1-k}E_{h}P_{h}^{k}.$ (2.3)
This identity is the starting point for all global error estimates in operator splitting and time discretization; see, e.g., Trotter (1958); Strang (1968); rhus (1998); Faragó and Havasi (2007).
In the non-autonomous case, the corresponding expression involves time-ordered products and is discussed in Section 2.3 below.
Discrete stability notions
The accumulation behavior of the sum in (2.3) depends critically on the stability properties of the discrete propagators $P_{h}$. We therefore introduce a hierarchy of stability assumptions.
These notions form a strict hierarchy:
$\text{summability}\;\Longrightarrow\;\text{Ritt stability}\;\Longrightarrow\;\text{power boundedness},$
and each level corresponds to a different mechanism of error accumulation; see, e.g., Gröger (1987); Lubich and Nevanlinna (1991); Nagy and Zemánek (1999); Le Merdy (2002); Gomilko and Zemánek (2011); Schweizer (2012).
Definition 2.3 (Power boundedness).
The family $\{P_{h}\}_{0<h\leq h_{0}}$ is said to be uniformly power bounded if there exists $M>0$ such that
$\sup_{0<h\leq h_{0}}\sup_{n\geq 0}\|P_{h}^{n}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M.$
Definition 2.4 (Ritt stability).
The family $\{P_{h}\}_{0<h\leq h_{0}}$ is said to be uniformly Ritt stable if there exists $M>0$ such that
$\sup_{0<h\leq h_{0}}\sup_{n\geq 1}n\|P_{h}^{n}-P_{h}^{n+1}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M.$
Definition 2.5 (Summability / bounded variation).
We say that $\{P_{h}\}_{0<h\leq h_{0}}$ satisfies summability stability if there exists $M>0$ such that
$\sup_{0<h\leq h_{0}}\sup_{n\geq 1}\left\|\sum_{k=0}^{n-1}P_{h}^{k}\right\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M.$
These notions form a strict hierarchy:
$\text{summability}\;\Longrightarrow\;\text{Ritt stability}\;\Longrightarrow\;\text{power boundedness},$
and each level corresponds to a different mechanism of error accumulation.
2.2.2 Abstract convergence principles
We now formulate abstract convergence results which depend only on:
  • local consistency: $\|E_{h}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\lesssim h^{\theta}$,
  • discrete stability of the propagators $P_{h}$.
These theorems serve as the abstract backbone for Main Theorems A, B, C, D proved in subsequent chapters.
Theorem 2.6 (Abstract convergence under summability stability).
Assume that:
  • $\|E_{h}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq Ch^{\theta}$ for some $0<\theta\leq 1$,
  • the family $\{P_{h}\}$ satisfies summability stability.
Then for every $T>0$ there exists $C_{T}>0$ such that
$\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{T}h^{\theta},\qquad nh\leq T.$
Theorem 2.7 (Abstract convergence under Ritt stability).
Assume that:
  • $\|E_{h}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq Ch^{\theta}$,
  • the family $\{P_{h}\}$ is uniformly Ritt stable.
Then for every $T>0$ there exists $C_{T}>0$ such that
$\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{T}h^{\theta},\qquad nh\leq T.$
Theorem 2.8 (Abstract convergence under power boundedness).
Assume that:
  • $\|E_{h}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq Ch^{\theta}$,
  • the family $\{P_{h}\}$ is uniformly power bounded.
Then for every $T>0$ there exists $C_{T}>0$ such that
$\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{T}Th^{\theta-1},\qquad nh\leq T.$
Proposition 2.9 (Implicit Euler products approximate the semigroup).
Let $L$ be the generator of a $C_{0}$–semigroup $(e^{tL})_{t\geq 0}$ on a Banach space $X$. For $h>0$ define the implicit Euler resolvent
$R_{h}:=(I-hL)^{-1}.$
Then the following hold.
(i) First-order convergence on the operator domain. For every $T>0$ there exists $C_{T}>0$ such that
$\|e^{tL}x-R_{h}^{n}x\|\leq C_{T}\,h\,\|Lx\|,\qquad x\in\mathcal{D}(L),\quad t=nh\leq T.$
(ii) Fractional-order convergence on interpolation spaces. Assume in addition that $L$ is sectorial and generates an analytic semigroup. Let $X_{\theta}=(X,\mathcal{D}(L))_{\theta,1}$ be the real interpolation space for some $0<\theta<1$. Then for every $T>0$ there exists $C_{T}>0$ such that
$\|e^{tL}x-R_{h}^{n}x\|\leq C_{T}\,h^{\theta}\,\|x\|_{X_{\theta}},\qquad x\in X_{\theta},\quad t=nh\leq T.$
These estimates express that the rate of convergence of the implicit Euler approximation is governed by the available regularity of the initial data.
Sketch of proof and references.
Assertion (i) is classical and follows from the Chernoff product formula for contractive semigroups; see, for example, (Pazy, 1983, Chap. III.5) or (Engel and Nagel, 1999, Chap. V.3).
Assertion (ii) follows from analyticity and interpolation arguments. In this setting the resolvent enjoys fractional smoothing properties, which lead to the estimate above; see, e.g., (Lunardi, 1995, Chap. 4) or (Tanabe, 1997, § 3.5). ∎
These abstract theorems correspond precisely to the three main regimes formulated in Main Theorems A, B, C, D and will be instantiated for concrete delay equations in Chapters 3 and 4.
2.3 Non-autonomous case and product stability
In the non-autonomous case, the $n$-step propagators are given by time-ordered products
$\prod_{j=0}^{n-1}P_{h}(t_{j}):=P_{h}(t_{n-1})\cdots P_{h}(t_{0}),\qquad\prod_{j=0}^{n-1}R_{h}(t_{j}):=R_{h}(t_{n-1})\cdots R_{h}(t_{0}),$
and the telescoping identity takes the form
$\prod_{j=0}^{n-1}R_{h}(t_{j})-\prod_{j=0}^{n-1}P_{h}(t_{j})=\sum_{k=0}^{n-1}\left(\prod_{j=k+1}^{n-1}R_{h}(t_{j})\right)E_{h}(t_{k})\left(\prod_{j=0}^{k-1}P_{h}(t_{j})\right).$
In this setting, stability must control not only powers of a fixed operator but also products of varying operators. This is the fundamental reason why frozen analysis (i.e. treating each $t$ independently as autonomous) is insufficient in the non-autonomous case, and why additional product stability assumptions are required.
This phenomenon is analyzed in detail in Chapter 3 (non-sectorial case) and Chapter 4 (sectorial case), leading to Main Theorems C and D.
Theorem 2.10 (Abstract convergence principle I: autonomous non-sectorial case).
Let $X$ be a Banach space and let $R_{h},P_{h}\in\mathcal{L}(X)$ ($0<h\leq h_{0}$) be two families of bounded operators. Assume:
  • (A1)
    (Uniform boundedness of implicit scheme)
    $\sup_{0<h\leq h_{0}}\sup_{n\geq 0}\|R_{h}^{n}\|_{\mathcal{L}(X)}<\infty.$
  • (A2)
    (Local defect on a regularity space) There exists a Banach space $X_{\theta}\hookrightarrow X$ and constants $C>0$, $\theta\in(0,1]$ such that
    $\|R_{h}-P_{h}\|_{\mathcal{L}(X_{\theta},X)}\leq Ch^{\theta},\qquad 0<h\leq h_{0}.$
  • (A3)
    (Discrete stability of $P_{h}$) One of the following conditions holds uniformly in $h$:
    • (i)
      summability of powers on $X$,
    • (ii)
      bounded total variation of powers on $X$,
    • (iii)
      uniform Ritt stability on $X$.
Then for every $T>0$ there exists $C_{T}>0$ such that
$\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(X_{\theta},X)}\leq C_{T}h^{\theta},\qquad nh\leq T.$
Theorem 2.11 (Abstract convergence principle II: autonomous sectorial case).
In addition to the assumptions of Theorem 2.10, assume that the local defect satisfies the stronger estimate
$\|R_{h}-P_{h}\|_{\mathcal{L}(X_{1},X)}\leq Ch,\qquad 0<h\leq h_{0},$
on a space $X_{1}\hookrightarrow X$.
Then under the same discrete stability assumptions as in Theorem 2.10, we obtain the first-order convergence
$\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(X_{1},X)}\leq C_{T}h,\qquad nh\leq T.$
Theorem 2.12 (Abstract convergence principle III: non-autonomous non-sectorial case).
Let $\{R_{h}(t_{k})\}$ and $\{P_{h}(t_{k})\}$ be families of bounded operators indexed by time nodes $t_{k}=kh$, and define the discrete propagators
$R_{h}^{(n)}:=R_{h}(t_{n-1})\cdots R_{h}(t_{0}),\qquad P_{h}^{(n)}:=P_{h}(t_{n-1})\cdots P_{h}(t_{0}).$
Assume that:
  • (C1)
    Uniform boundedness of $R_{h}^{(n)}$ on $X$ for $nh\leq T$.
  • (C2)
    Local defect estimate
    $\|R_{h}(t_{k})-P_{h}(t_{k})\|_{\mathcal{L}(X_{\theta},X)}\leq Ch^{\theta}.$
  • (C3)
    Uniform discrete stability of the products $\{P_{h}^{(n)}\}$ (e.g. summability, bounded variation, or Ritt-type stability).
Then
$\|R_{h}^{(n)}-P_{h}^{(n)}\|_{\mathcal{L}(X_{\theta},X)}\leq C_{T}h^{\theta},\qquad nh\leq T.$
Theorem 2.13 (Abstract convergence principle IV: non-autonomous sectorial case).
Under the assumptions of Theorem 2.12, suppose in addition that the local defect satisfies the stronger bound
$\|R_{h}(t_{k})-P_{h}(t_{k})\|_{\mathcal{L}(X_{1},X)}\leq Ch.$
Then
$\|R_{h}^{(n)}-P_{h}^{(n)}\|_{\mathcal{L}(X_{1},X)}\leq C_{T}h,\qquad nh\leq T.$
3 Non-sectorial case
In this chapter we demonstrate how the abstract framework of Chapter 2 is realized in the non-sectorial setting. This case exhibits the strongest dependence on discrete stability and therefore illustrates most clearly the role of the stability hierarchy.
3.1 Autonomous non-sectorial case
In this section we analyze the autonomous delay equation in the non-sectorial setting. The purpose is twofold:
  • to establish a sharp local defect estimate for Lie–Trotter splitting;
  • to demonstrate explicitly how the global error depends on discrete stability properties of the propagator.
This case therefore provides the clearest illustration of the stability hierarchy. Throughout this section we assume that $D(t)\equiv D$ is time-independent.
3.1.1 Local defect estimate
Recall that the one-step operators are given by
$R_{h}:=(I-h(C+D))^{-1},\qquad P_{h}:=(I-H_{h})^{-1}S_{h}^{C},\qquad H_{h}:=hDS_{h}^{C},$
whenever both operators are well defined. The local resolvent defect is defined by
$E_{h}:=R_{h}-P_{h}.$
We work on the fractional spaces
$X_{\theta}:=(\mathcal{E}_{0},\mathcal{D}(C))_{\theta,p},\qquad 0<\theta\leq 1.$
The key ingredient is the following smoothing estimate for the preconditioned delay operator.
Basic resolvent bounds
Lemma 3.1 (Transport resolvent bounds).
Let $C$ be the left-shift generator on $\mathcal{E}_{0}$. Then $S_{h}^{C}:=(I-hC)^{-1}\in\mathcal{L}(\mathcal{E}_{0})$ exists for every $h>0$ and
$\|S_{h}^{C}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq 1,\qquad\|CS_{h}^{C}\|_{\mathcal{L}(\mathcal{E}_{0})}\lesssim h^{-1}.$
Moreover, for every $0<\theta\leq 1$,
$\|S_{h}^{C}\|_{\mathcal{L}(X_{\theta},\mathcal{D}(C))}\lesssim h^{-\theta},\qquad X_{\theta}:=(\mathcal{E}_{0},\mathcal{D}(C))_{\theta,1}.$
Lemma 3.2 (Sectorial resolvent bounds (when $A_{0}$ is present)).
Assume that $A_{0}=\mathrm{diag}(A,0)$ is sectorial on $\mathcal{E}_{0}$ with $0\in\rho(A_{0})$. Then $S_{h}^{A}:=(I-hA_{0})^{-1}$ exists for all $h>0$ and
$\|S_{h}^{A}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M_{A},\qquad\|A_{0}S_{h}^{A}\|_{\mathcal{L}(\mathcal{E}_{0})}\lesssim h^{-1}.$
Consequently, the combined preconditioner $S_{h}^{AC}:=S_{h}^{C}S_{h}^{A}$ satisfies
$\|S_{h}^{AC}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M_{A},\qquad\|S_{h}^{AC}\|_{\mathcal{L}(X_{\theta},\mathcal{D}(C))}\lesssim h^{-\theta},\quad 0<\theta\leq 1.$
Trace control on the fractional scale
Lemma 3.3 (Trace bound on fractional domain).
Let $\Phi_{\tau}:\rho\mapsto\rho(\tau)$ be the trace at $\tau<0$. Then for every $0<\theta\leq 1$ there exists $c_{\theta}>0$ such that
$\|\rho(\tau)\|_{X}\leq c_{\theta}\|(u,\rho)\|_{X_{\theta}},\qquad(u,\rho)\in X_{\theta}.$
In particular, for delay operators of the form
$D(u,\rho)=(au+b\,\rho(\tau),0),$
one has
$\|Dx\|_{\mathcal{E}_{0}}\lesssim\|x\|_{X_{\theta}},\qquad x\in X_{\theta}.$
Fractional regime: $0<\theta<1$ (operational smallness)
Lemma 3.4 (Smoothing of the preconditioned delay operator).
Assume that $D(t):\mathcal{D}(C)\to\mathcal{E}_{0}$ is a bounded family of delay-type operators, uniformly of the form
$D(t)(u,\rho)=(a(t)u+b(t)\rho(\tau),0),$
with bounded coefficients. Then for every $0<\theta<1$ there exist constants $C_{\theta}>0$ and $h_{0}>0$ such that
$\|h\,D(t)S_{h}^{C}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{\theta}h^{1-\theta},\qquad 0<h\leq h_{0},\ t\in[0,T].$
In particular,
$\sup_{t\in[0,T]}\|hD(t)S_{h}^{C}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\to 0\qquad(h\to 0).$
Proof.
By Lemma 3.1, $S_{h}^{C}:X_{\theta}\to\mathcal{D}(C)$ with norm $\lesssim h^{-\theta}$. By Lemma 3.3, $D(t)$ is bounded from $\mathcal{D}(C)$ to $\mathcal{E}_{0}$ uniformly in $t$. Hence
$\|D(t)S_{h}^{C}x\|_{\mathcal{E}_{0}}\lesssim h^{-\theta}\|x\|_{X_{\theta}},$
and multiplying by $h$ gives the claim. ∎
Lemma 3.5 (Operational smallness on $X_{\theta}$).
Let $H_{h}(t):=hD(t)S_{h}^{C}$ (or $hD(t)S_{h}^{AC}$ in the sectorial case). Assume that for some $0<\theta<1$ there exist constants $C_{\theta},h_{0}>0$ such that
$\sup_{t\in[0,T]}\|H_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{\theta}h^{1-\theta},\qquad 0<h\leq h_{0}.$
Then there exist $h_{1}\in(0,h_{0}]$ and $q\in(0,1)$ such that
$\sup_{t\in[0,T]}\|H_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq q,\qquad 0<h\leq h_{1}.$
In particular, $(I-H_{h}(t))^{-1}$ exists and satisfies the uniform bound
$\sup_{t\in[0,T]}\|(I-H_{h}(t))^{-1}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq\frac{1}{1-q}.$
Strong-data regime: $\mathcal{D}(C)$
Lemma 3.6 (Effective boundedness under relative boundedness).
Assume that $D:\mathcal{D}(C)\to\mathcal{E}_{0}$ is CC–bounded with relative bound 0. Then there exist constants $h_{0}>0$ and $M>0$ such that for all $0<h\leq h_{0}$,
$\|DS_{h}^{C}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M,\qquad\|hDS_{h}^{C}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq Mh.$
Proof.
Since $D$ is CC–bounded with relative bound 0, for every $\varepsilon>0$ there exists $c_{\varepsilon}>0$ such that
$\|Dx\|\leq\varepsilon\|Cx\|+c_{\varepsilon}\|x\|,\qquad x\in\mathcal{D}(C).$
Apply this to $x=S_{h}^{C}y$ and use Lemma 3.1. Choosing $\varepsilon$ small and $h$ sufficiently small yields the claim. ∎
Lemma 3.7 (Uniform boundedness of the splitting map).
Assume the operational smallness condition (OS) on $\mathcal{E}_{0}$, i.e.
$\sup_{t\in[0,T]}\sup_{0<h\leq h_{0}}\|H_{h}(t)\|_{\mathcal{L}(\mathcal{E}_{0})}\leq q<1.$
Define
$P_{h}(t):=(I-H_{h}(t))^{-1}S_{h}^{C}\quad(\text{or }(I-H_{h}(t))^{-1}S_{h}^{AC}).$
Then
$\sup_{t\in[0,T]}\sup_{0<h\leq h_{0}}\|P_{h}(t)\|_{\mathcal{L}(\mathcal{E}_{0})}\leq\frac{1}{1-q}.$
Lemma 3.8 (Neumann bound for resolvent perturbations).
Let $T\in\mathcal{L}(X)$ and assume $\|T\|\leq M$. If $0<h\leq(2M)^{-1}$, then IhTI-hT is invertible and
$\|(I-hT)^{-1}\|\leq 2.$
3.1.2 Local defect and global error: autonomous non-sectorial case
Throughout this subsection we consider the autonomous (time-independent) operators $C$ and $D$ on $\mathcal{E}_{0}$ and the associated discrete maps
$R_{h}:=(I-h(C+D))^{-1},\qquad S_{h}^{C}:=(I-hC)^{-1},\qquad H_{h}:=hDS_{h}^{C},\qquad P_{h}:=(I-H_{h})^{-1}S_{h}^{C},$
whenever these operators are well-defined.
Operational smallness and well-definedness of the splitting step
We assume that for each $h>0$ the composition $DS_{h}^{C}$ extends to a bounded operator on $\mathcal{E}_{0}$. In concrete delay models this extension holds because $S_{h}^{C}$ maps $\mathcal{E}_{0}$ into $\mathcal{D}(C)$.
Assumption 3.9 (Operational smallness (OS)).
There exist $h_{0}>0$ and $q\in(0,1)$ such that
$\sup_{0<h\leq h_{0}}\|H_{h}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq q.$
Under (OS) the inverse $(I-H_{h})^{-1}$ exists and
$\sup_{0<h\leq h_{0}}\|(I-H_{h})^{-1}\|\leq\frac{1}{1-q},$
hence $P_{h}$ is well-defined and uniformly bounded on $\mathcal{E}_{0}$.
Remark 3.10 (Verification of (OS) under relative bound 0).
If $D:\mathcal{D}(C)\to\mathcal{E}_{0}$ is CC–bounded with relative bound 0, then (OS) is not an extra hypothesis. Indeed, for every $\varepsilon>0$ there exists $b(\varepsilon)\geq 0$ such that
$\|Dx\|\leq\varepsilon\|Cx\|+b(\varepsilon)\|x\|,\qquad x\in\mathcal{D}(C),$
and therefore, using the standard resolvent bounds for the shift generator $C$,
$\|hD(I-hC)^{-1}\|\leq h\varepsilon\|C(I-hC)^{-1}\|+h\,b(\varepsilon)\|(I-hC)^{-1}\|\leq\varepsilon+O(h)\quad(h\downarrow 0).$
Choosing first $\varepsilon>0$ and then $h>0$ small yields $\|H_{h}\|\leq q<1$.
Local defect
The algebraic resolvent identity below is the starting point of all estimates.
Lemma 3.11 (Right-preconditioned resolvent identity).
Assume (OS) and that $R_{h}\in\mathcal{L}(\mathcal{E}_{0})$ exists. Then
$R_{h}-P_{h}=R_{h}H_{h}-(I-H_{h})^{-1}H_{h}S_{h}^{C}.$
Proof.
This is the autonomous specialization of the right-preconditioned identities proved earlier in the discrete resolvent chapter; it follows by rewriting
$R_{h}=(I-h(C+D))^{-1}=(I-hC-hD)^{-1}$
and comparing with
$P_{h}=(I-hDS_{h}^{C})^{-1}S_{h}^{C},\qquad S_{h}^{C}=(I-hC)^{-1}.$
We now record the local defect on two input scales: $\mathcal{D}(C)$ and the fractional space $X_{\theta}=(\mathcal{E}_{0},\mathcal{D}(C))_{\theta,1}$.
Proposition 3.12 (Local defect on $\mathcal{D}(C)$ (first order)).
Assume (OS) and that $R_{h}$ exists with $\sup_{0<h\leq h_{0}}\|R_{h}\|\leq M_{R}$. If in addition $D:\mathcal{D}(C)\to\mathcal{E}_{0}$ is CC–bounded with relative bound 0, then there exists $C>0$ such that
$\|R_{h}-P_{h}\|_{\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})}\leq Ch,\qquad 0<h\leq h_{0}.$
Proof.
By Lemma 3.11,
$\|R_{h}-P_{h}\|_{\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})}\leq\Bigl(\|R_{h}\|+\|(I-H_{h})^{-1}\|\,\|S_{h}^{C}\|\Bigr)\,\|H_{h}\|_{\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})}.$
Under (OS) we have $\|(I-H_{h})^{-1}\|\leq(1-q)^{-1}$ and $\|S_{h}^{C}\|\leq 1$. Moreover, relative bound 0 implies $\|hDS_{h}^{C}\|_{\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})}\lesssim h$. Combining these yields the claim. ∎
Proposition 3.13 (Local defect on $X_{\theta}$ (fractional scale)).
Assume (OS) and that $R_{h}$ exists with $\sup_{0<h\leq h_{0}}\|R_{h}\|\leq M_{R}$. Let $0<\theta<1$ and set $X_{\theta}=(\mathcal{E}_{0},\mathcal{D}(C))_{\theta,1}$. Then there exists $C_{\theta}>0$ such that
$\|R_{h}-P_{h}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{\theta}\,h^{1-\theta},\qquad 0<h\leq h_{0}.$
Proof.
As in the previous proof, the identity in Lemma 3.11 reduces the estimate to $\|H_{h}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}$. Using the graph-norm control of $D$ on $\mathcal{D}(C)$,
$\|Dy\|\lesssim\|y\|+\|Cy\|,\qquad y\in\mathcal{D}(C),$
and applying it to $y=S_{h}^{C}x$ yields
$\|H_{h}x\|=h\|DS_{h}^{C}x\|\lesssim h\|S_{h}^{C}x\|+h\|CS_{h}^{C}x\|.$
Now $\|S_{h}^{C}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\lesssim 1$ and the non-sectorial smoothing estimate gives $\|CS_{h}^{C}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\lesssim h^{-\theta}$. Hence
$\|H_{h}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\lesssim h+h\cdot h^{-\theta}\lesssim h^{1-\theta},$
and the conclusion follows. ∎
Global error via telescoping
For $n\in\mathbb{N}$ we use the standard telescoping identity
$R_{h}^{n}-P_{h}^{n}=\sum_{k=0}^{n-1}R_{h}^{\,n-1-k}\,(R_{h}-P_{h})\,P_{h}^{k},$ (3.1)
which is purely algebraic and holds whenever the products are defined.
To control the right-hand side without losing a factor $n\sim T/h$ we require discrete stability information on $\{P_{h}\}$ that bounds either $\sum_{k=0}^{n-1}P_{h}^{k}$ (summability/partial sums) or the total variation $\sum\|P_{h}^{k+1}-P_{h}^{k}\|$.
Strong stability: bounded partial sums (summability)
Assumption 3.14 (Summability / bounded partial sums).
There exists $M_{\mathrm{sum}}\geq 1$ such that for all $0<h\leq h_{0}$ and $N\geq 1$,
$\Bigl\|\sum_{k=0}^{N-1}P_{h}^{k}\Bigr\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M_{\mathrm{sum}}.$
Proposition 3.15 (Global error under summability (strongest case)).
Assume (OS), $\sup_{0<h\leq h_{0}}\|R_{h}\|\leq M_{R}$, and Assumption 3.14. Then:
  1. (a)
    If $\|R_{h}-P_{h}\|_{\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})}\leq Ch$, then for every $T>0$ there exists $C_{T}>0$ such that
    $\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})}\leq C_{T}\,h,\qquad nh\leq T.$
  2. (b)
    If $0<\theta<1$ and $\|R_{h}-P_{h}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{\theta}h^{1-\theta}$, then for every $T>0$ there exists $C_{T,\theta}>0$ such that
    $\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{T,\theta}\,h^{1-\theta},\qquad nh\leq T.$
Proof.
Let $E_{h}:=R_{h}-P_{h}$. By (3.1) and boundedness of $R_{h}$,
$\|R_{h}^{n}-P_{h}^{n}\|\leq M_{R}\sum_{k=0}^{n-1}\|E_{h}P_{h}^{k}\|.$
Using $\|E_{h}P_{h}^{k}\|\leq\|E_{h}\|\,\|P_{h}^{k}\|$ and summation by partial sums,
$\sum_{k=0}^{n-1}E_{h}P_{h}^{k}=E_{h}\sum_{k=0}^{n-1}P_{h}^{k},$
hence
$\Bigl\|\sum_{k=0}^{n-1}E_{h}P_{h}^{k}\Bigr\|\leq\|E_{h}\|\,\Bigl\|\sum_{k=0}^{n-1}P_{h}^{k}\Bigr\|\leq M_{\mathrm{sum}}\,\|E_{h}\|.$
Therefore $\|R_{h}^{n}-P_{h}^{n}\|\leq M_{R}M_{\mathrm{sum}}\|E_{h}\|$. Insert the local defect bound on $\mathcal{D}(C)$ or on $X_{\theta}$ to conclude. ∎
Strong stability: bounded total variation of powers
Assumption 3.16 (Bounded total variation of powers).
There exists $M_{\mathrm{BV}}\geq 1$ such that for all $0<h\leq h_{0}$ and $N\geq 1$,
$\sum_{k=0}^{N-1}\|P_{h}^{k+1}-P_{h}^{k}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M_{\mathrm{BV}}.$
Proposition 3.17 (Global error under bounded total variation).
Assume (OS), $\sup_{0<h\leq h_{0}}\|R_{h}\|\leq M_{R}$, and Assumption 3.16. Then the same conclusions as in Proposition 3.15 hold.
Proof.
Combine (3.1) with a discrete integration-by-parts argument (Abel summation, see A.8, A.2.3 in detail) to shift one discrete derivative onto the powers of $P_{h}$. The bounded total variation controls the resulting terms uniformly in $n$ (for $nh\leq T$). Inserting the local defect bound yields the stated rates. See, e.g., the discrete multiplier estimates in the literature on power-bounded operators with bounded variation of powers. ∎
Weak stability: power-boundedness only
Proposition 3.18 (Global bound under mere power-boundedness).
Assume (OS), $\sup_{0<h\leq h_{0}}\|R_{h}\|\leq M_{R}$, and
$\sup_{0<h\leq h_{0}}\ \sup_{k\geq 0}\|P_{h}^{k}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M_{P}.$
Then for $nh\leq T$,
$\|R_{h}^{n}-P_{h}^{n}\|\leq C\,n\,\|R_{h}-P_{h}\|,$
and consequently:
  1. (a)
    on $\mathcal{D}(C)$: if $\|R_{h}-P_{h}\|_{\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})}\lesssim h$, then $\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})}\lesssim T$ (no decay in $h$);
  2. (b)
    on $X_{\theta}$: if $\|R_{h}-P_{h}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\lesssim h^{1-\theta}$, then $\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\lesssim T\,h^{-\theta}$, which generally diverges as $h\downarrow 0$.
Thus, without stronger discrete stability, uniform-in-$h$ convergence on finite time horizons cannot be expected on the fractional scale.
Proof.
From (3.1),
$\|R_{h}^{n}-P_{h}^{n}\|\leq\sum_{k=0}^{n-1}\|R_{h}^{n-1-k}\|\,\|R_{h}-P_{h}\|\,\|P_{h}^{k}\|\leq n\,M_{R}\,M_{P}\,\|R_{h}-P_{h}\|,$
and the stated consequences follow by inserting the local defect rate and using $n\leq T/h$. ∎
We summarize the main result of this section as Theorem Main A.
Theorem 3.19 (Main Theorem A: autonomous non-sectorial case).
Consider the autonomous delay equation on $\mathcal{E}_{0}$ with operators $C$ and $D$ as in Chapter 2. Let the splitting and implicit Euler resolvents be defined by
$P_{h}=\bigl(I-hD(I-hC)^{-1}\bigr)^{-1}(I-hC)^{-1},\qquad R_{h}=(I-h(C+D))^{-1},$
whenever they exist.
Assume that the following conditions hold:
  1. (A1)
    The operational smallness condition (OS) holds, so that $P_{h}$ is well defined and uniformly bounded on $\mathcal{E}_{0}$ for $0<h\leq h_{0}$.
  2. (A2)
    The implicit Euler resolvent exists and is uniformly bounded:
    $\sup_{0<h\leq h_{0}}\|R_{h}\|_{\mathcal{L}(\mathcal{E}_{0})}<\infty.$
  3. (A3)
    The delay operator $D:\mathcal{D}(C)\to\mathcal{E}_{0}$ is CC–bounded with relative bound 0.
Then the following assertions hold.
(i) Local comparison between resolvents. There exists $C>0$ such that
$\|R_{h}-P_{h}\|_{\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})}\leq Ch,\qquad 0<h\leq h_{0}.$
Moreover, for every $0<\theta<1$ there exists $C_{\theta}>0$ such that
$\|R_{h}-P_{h}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{\theta}h^{1-\theta}.$
(ii) Global convergence under strong discrete stability. Assume in addition that the splitting propagator $P_{h}$ satisfies either bounded partial sums or bounded total variation of powers on the relevant space (i.e. on $\mathcal{D}(C)$ or $X_{\theta}$). Then for every $T>0$ there exists $C_{T}>0$ such that
$\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})}\leq C_{T}h,\qquad nh\leq T,$
and
$\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{T}h^{1-\theta},\qquad nh\leq T.$
(iii) Global convergence under Ritt stability. If $\{P_{h}\}_{0<h\leq h_{0}}$ is uniformly Ritt stable on $\mathcal{E}_{0}$, then the same global error estimates as in (ii) hold.
(iv) Weak stability. If $\{P_{h}\}$ is merely uniformly power bounded, then one generally only has
$\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{T}\,nh^{1-\theta},\qquad nh\leq T,$
so that uniform-in-time convergence cannot be expected without additional stability.
In particular, the long-time convergence behavior of Lie–Trotter splitting toward implicit Euler is governed not only by the local comparison $R_{h}-P_{h}$, but crucially by the discrete stability properties of the powers $\{P_{h}^{n}\}$.
3.2 Non-autonomous non-sectorial case
In this section we extend the analysis of the previous section to the non-autonomous setting. The main purpose is to show that, even though the local behavior of the scheme is essentially identical to the autonomous case, the global convergence analysis requires genuinely stronger stability assumptions due to the appearance of time-ordered products.
Structure of the auxiliary lemmas.
This section collects several auxiliary lemmas that provide the analytic foundation for the non-autonomous local and global error analysis. The main purpose is to justify the well-posedness, regularity, and stability properties of the frozen splitting step
$P_{h}(t)=S_{h}^{C}(I-H_{h}(t))^{-1},\qquad H_{h}(t)=hD(t)S_{h}^{C},$
and of the associated defect operator $E_{h}(t)=R_{h}(t)-P_{h}(t)$.
The logical structure is as follows. First, operational smallness (OS) guarantees that the inverse $(I-H_{h}(t))^{-1}$ exists uniformly in $t$ and $h$, yielding well-defined and uniformly bounded splitting operators $P_{h}(t)$. Second, Lipschitz regularity of the coefficients $a(t),b(t)$ implies Lipschitz continuity of the delay operator $D(t)$ on $\mathcal{D}(C)$, which in turn yields the estimate (H-Lip) for $H_{h}(t)$ on fractional spaces. Finally, combining (OS), (H-Lip), and the Lipschitz continuity of the reference resolvent $R_{h}(t)$, we obtain the crucial defect continuity estimate
$\|E_{h}(t)-E_{h}(s)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\lesssim h^{\theta}|t-s|,$
which serves as the key technical input for the discrete Abel-type arguments in the global error analysis.
3.2.1 Setting and notation
We consider the non-autonomous delay equation with operator family $D(t):\mathcal{D}(C)\to\mathcal{E}_{0}$ of the form
$D(t)(u,\rho)=(a(t)u+b(t)\rho(\tau),0),\qquad t\in[0,T],$
where $a,b\in L^{\infty}(0,T)$. The one-step propagators are defined by
$R_{h}(t):=(I-h(C+D(t)))^{-1},\qquad P_{h}(t):=(I-H_{h}(t))^{-1}S_{h}^{C},\qquad H_{h}(t):=hD(t)S_{h}^{C},$
whenever these operators are well defined.
3.3 Uniform local defect estimate
The key observation is that the local defect estimate from the autonomous case remains valid uniformly in $t$.
Proposition 3.20 (Uniform local defect for frozen steps).
Assume that the operational smallness condition (OS) holds uniformly for $t\in[0,T]$ and that the frozen resolvents $R_{h}(t)$ exist with a uniform bound
$\sup_{t\in[0,T]}\sup_{0<h\leq h_{0}}\|R_{h}(t)\|_{\mathcal{L}(\mathcal{E}_{0})}<\infty.$
Then for every $0<\theta<1$ there exists $C_{\theta}>0$ such that
$\|R_{h}(t)-P_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{\theta}h^{\theta},\qquad t\in[0,T],\ 0<h\leq h_{0}.$
Proof.
Fix $t\in[0,T]$ and abbreviate
$R:=R_{h}(t),\quad P:=P_{h}(t),\quad H:=H_{h}(t)=hD(t)S_{h}^{C},\quad S:=S_{h}^{C}.$
By the same right-preconditioned resolvent identity as in the autonomous case (cf. Lemma 3.11),
$R-P=RH-(I-H)^{-1}HS.$
Taking norms in $\mathcal{L}(X_{\theta},\mathcal{E}_{0})$ and using (OS) (which gives $\|(I-H)^{-1}\|\leq(1-q)^{-1}$ uniformly in $t$ and $h$), together with the uniform bound on $\|R_{h}(t)\|$, we obtain
$\|R_{h}(t)-P_{h}(t)\|\;\leq\;C\,\|H_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}.$
Finally, the uniform smoothing estimate for the preconditioned delay operator (see Lemma LABEL:lem:hDSC-theta-nonaut or its tt–uniform variant) yields $\|H_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{\theta}h^{\theta}$, uniformly for $t\in[0,T]$ and $0<h\leq h_{0}$, which proves the claim. ∎
3.3.1 Why frozen analysis is insufficient
At first sight, one might expect that the non-autonomous convergence analysis could be reduced to the autonomous case by applying the frozen-step estimate from Proposition 3.20 at each time level. However, this is fundamentally insufficient.
Indeed, the $n$-step propagators are given by time-ordered products
$R_{h}(t_{n-1})\cdots R_{h}(t_{0}),\qquad P_{h}(t_{n-1})\cdots P_{h}(t_{0}),$
and the global error expands into a non-commutative telescoping sum of the form
$\sum_{k=0}^{n-1}R_{h}(t_{n-1})\cdots R_{h}(t_{k+1})\bigl(R_{h}(t_{k})-P_{h}(t_{k})\bigr)P_{h}(t_{k-1})\cdots P_{h}(t_{0}).$
Controlling this expression requires uniform bounds on products of the propagators, not merely on each frozen operator individually. This is the fundamental mathematical reason why additional stability assumptions are unavoidable in the non-autonomous setting.
3.3.2 Global error under product stability
We now formulate a representative convergence result under product stability assumptions.
Theorem 3.21 (Main Theorem C: Non-autonomous non-sectorial convergence under discrete product stability).
Let $0<T<\infty$ and consider the family of discrete operators
$R_{h}(t):=(I-h(C+D(t)))^{-1},\qquad P_{h}(t):=(I-hD(t)S_{h}^{C})^{-1}S_{h}^{C},\quad S_{h}^{C}:=(I-hC)^{-1},$
defined for $t\in[0,T]$ and sufficiently small $h>0$.
Assume that the following conditions hold.
  1. (C1)
    Operational smallness. There exist $h_{0}>0$ and $q\in(0,1)$ such that
    $\sup_{t\in[0,T]}\sup_{0<h\leq h_{0}}\|hD(t)S_{h}^{C}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq q.$
  2. (C2)
    Uniform boundedness of frozen resolvents. For all $t\in[0,T]$ and $0<h\leq h_{0}$, the resolvent $R_{h}(t)\in\mathcal{L}(\mathcal{E}_{0})$ exists and satisfies
    $\sup_{t\in[0,T]}\sup_{0<h\leq h_{0}}\|R_{h}(t)\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M_{R}.$
  3. (C3)
    Uniform local defect. There exist $0<\theta<1$ and $C_{\theta}>0$ such that
    $\|R_{h}(t)-P_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{\theta}h^{\theta},\qquad t\in[0,T],\ 0<h\leq h_{0}.$
  4. (C4)
    Discrete product stability. For every $T>0$ there exists $M_{T}>0$ such that for every partition $t_{k}=kh$ with $t_{n}\leq T$,
    $\left\|\prod_{j=0}^{n-1}R_{h}(t_{j})\right\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M_{T},\qquad\left\|\prod_{j=0}^{n-1}P_{h}(t_{j})\right\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M_{T},$
    uniformly for $0<h\leq h_{0}$.
Then there exists a constant $C_{T}>0$ such that for all $x\in X_{\theta}$,
$\left\|\prod_{j=0}^{n-1}R_{h}(t_{j})x-\prod_{j=0}^{n-1}P_{h}(t_{j})x\right\|_{\mathcal{E}_{0}}\leq C_{T}h^{\theta},\qquad t_{n}=nh\leq T.$
In particular, Lie–Trotter splitting converges to the frozen implicit Euler scheme on $X_{\theta}$ with order $O(h^{\theta})$ on every finite time interval.
Remark 3.22 (Relation to continuous well-posedness and convergence toward the evolution family).
Theorem 3.21 is formulated purely at the discrete level and does not require the existence of a continuous solution operator.
Suppose, however, that in addition to the above assumptions the classical non-autonomous regularity conditions hold, namely:
  • (U1)
    Uniform $C$-boundedness with relative bound 0: for every $\varepsilon>0$ there exists $c_{\varepsilon}>0$ such that
    $\|D(t)x\|\leq\varepsilon\|Cx\|+c_{\varepsilon}\|x\|,\qquad x\in\mathcal{D}(C),\ t\in[0,T].$
  • (U2)
    Time regularity: $t\mapsto D(t)$ is Lipschitz continuous as a map $(\mathcal{D}(C),\|x\|+\|Cx\|)\to\mathcal{E}_{0}$.
If, in addition, the family $t\mapsto D(t)$ satisfies sufficient temporal regularity (for instance, $C^{1}$-regularity in operator norm) and suitable compatibility assumptions ensuring well-posedness of the non-autonomous Cauchy problem, then classical results due to Kato, Tanabe and others imply the existence of a strongly continuous evolution family $U(t,s)$ generated by $C+D(t)$ on $\mathcal{E}_{0}$ with a finite-time bound
$\|U(t,s)\|\leq Me^{\omega(t-s)},\qquad 0\leq s\leq t\leq T.$
We emphasize, however, that such continuous well-posedness is not assumed in the present work: the discrete resolvent framework developed here remains meaningful and applicable even in regimes where the existence of an evolution family cannot be guaranteed. In the hyperbolic-type setting considered here (in particular due to the transport operator $C$), the existence of an evolution family for $C+D(t)$ generally requires stronger temporal regularity assumptions, typically $C^{1}$-regularity in $t$, and cannot be expected under mere Lipschitz continuity. e.g., (Tanabe, 1979, Ch. 7), Hishida (2004).
Moreover, the frozen implicit Euler scheme satisfies the classical consistency estimate
$\|U(t_{n+1},t_{n})-R_{h}(t_{n})\|_{\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})}\leq Ch^{2},$
so that for each fixed $T>0$,
$\left\|U(t_{n},0)x-\prod_{j=0}^{n-1}R_{h}(t_{j})x\right\|\leq C_{T}h,\qquad x\in\mathcal{D}(C).$
Combining this estimate with Theorem 3.21 by the triangle inequality yields
$\left\|U(t_{n},0)x-\prod_{j=0}^{n-1}P_{h}(t_{j})x\right\|\leq C_{T}h^{\theta},$
that is, Lie–Trotter splitting converges toward the continuous solution operator whenever such an operator exists.
This shows that the discrete comparison principle developed in this thesis is fully consistent with the classical semigroup-based theory, while remaining meaningful even in situations where no evolution family is available.
Product-summability and variation assumptions.
Let $t_{k}:=kh$ and abbreviate
$R_{k}:=R_{h}(t_{k}),\qquad P_{k}:=P_{h}(t_{k}),\qquad E_{k}:=R_{k}-P_{k}.$
Fix $m\in\mathbb{N}$ with $t_{m+1}\leq T$ and define the time-ordered products
$Q_{k-1}:=\begin{cases}P_{k-1}\cdots P_{0},&k\geq 1,\\ I,&k=0,\end{cases}\qquad\Pi_{R}(m,k+1):=\begin{cases}R_{m}\cdots R_{k+1},&0\leq k\leq m-1,\\ I,&k=m,\end{cases}$
and the associated partial sums
$S_{j}:=\sum_{k=0}^{j}Q_{k-1}\in\mathcal{L}(\mathcal{E}_{0}),\qquad j=0,1,\dots,m,\quad(S_{-1}:=0).$
Moreover, define the “left-weighted defects”
$A_{k}:=\Pi_{R}(m,k+1)\,E_{k}\in\mathcal{L}(X_{\theta},\mathcal{E}_{0}),\qquad k=0,1,\dots,m.$
We assume the following two discrete stability properties on the finite horizon $[0,T]$:
  1. (PS1)
    (Uniform boundedness of discrete partial sums) There exists a constant $M_{\rm sum}(T)>0$, independent of $h$ and $m$ with $t_{m+1}\leq T$, such that
    $\sup_{0\leq j\leq m}\|S_{j}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M_{\rm sum}(T).$
  2. (PS2)
    (Bounded variation of the left-weighted defects) There exists a constant $V(T)>0$, independent of $h$ and $m$ with $t_{m+1}\leq T$, such that
    $\|A_{m}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq V(T)\,h^{\theta},\qquad\sum_{k=0}^{m-1}\|A_{k}-A_{k+1}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq V(T)\,h^{\theta}.$
Proof of Theorem 3.21.
Step 1: Non-commutative telescoping. For each $m\in\{0,1,\dots,n-1\}$ set
$\Pi_{R}(m,k):=\begin{cases}R_{m}R_{m-1}\cdots R_{k},&0\leq k\leq m,\\ I,&k=m+1,\end{cases}\qquad\Pi_{P}(k,0):=\begin{cases}P_{k}P_{k-1}\cdots P_{0},&k\geq 0,\\ I,&k=-1.\end{cases}$
Then the standard product telescoping identity yields
$R_{n-1}\cdots R_{0}-P_{n-1}\cdots P_{0}=\sum_{k=0}^{n-1}\Pi_{R}(n-1,k+1)\,E_{k}\,\Pi_{P}(k-1,0).$ (3.2)
(One verifies (3.2) by inserting and subtracting the mixed products in the usual way.)
Fix m:=n1m:=n-1 and abbreviate
$A_{k}:=A_{k}^{(m)}=\Pi_{R}(m,k+1)E_{k},\qquad Q_{k-1}:=\Pi_{P}(k-1,0).$
Then (3.2) becomes
$R_{m}\cdots R_{0}-P_{m}\cdots P_{0}=\sum_{k=0}^{m}A_{k}\,Q_{k-1}.$ (3.3)
Step 2: Abel summation (summation by parts). Define the partial sums
$S_{j}:=\sum_{k=0}^{j}Q_{k-1}\in\mathcal{L}(\mathcal{E}_{0}),\qquad j=0,1,\dots,m,$
with $S_{-1}:=0$. Then $Q_{k-1}=S_{k}-S_{k-1}$, and hence
$\sum_{k=0}^{m}A_{k}Q_{k-1}=\sum_{k=0}^{m}A_{k}(S_{k}-S_{k-1})=A_{m}S_{m}+\sum_{k=0}^{m-1}(A_{k}-A_{k+1})S_{k}.$
Equivalently,
$\sum_{k=0}^{m}A_{k}Q_{k-1}=A_{m}S_{m}+\sum_{k=0}^{m-1}(A_{k}-A_{k-1})S_{k-1},$ (3.4)
where we used $A_{-1}=0$ and re-indexed the sum (either form is fine).
Step 3: Uniform estimates. Taking norms in $\mathcal{L}(X_{\theta},\mathcal{E}_{0})$ and using $\|BC\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq\|B\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\|C\|_{\mathcal{L}(\mathcal{E}_{0})}$, we obtain from (3.4)
$\Bigl\|\sum_{k=0}^{m}A_{k}Q_{k-1}\Bigr\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq\|A_{m}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\,\|S_{m}\|_{\mathcal{L}(\mathcal{E}_{0})}+\sum_{k=0}^{m-1}\|A_{k}-A_{k-1}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\,\|S_{k-1}\|_{\mathcal{L}(\mathcal{E}_{0})}.$
By (PS1), $\sup_{j\leq m}\|S_{j}\|\leq M_{\rm sum}(T)$. By (PS2), $\|A_{m}\|\leq V(T)h^{\theta}$ and $\sum_{k=0}^{m-1}\|A_{k}-A_{k-1}\|\leq V(T)h^{\theta}$. Hence
$\Bigl\|\sum_{k=0}^{m}A_{k}Q_{k-1}\Bigr\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq V(T)h^{\theta}\,M_{\rm sum}(T)+V(T)h^{\theta}\,M_{\rm sum}(T)=2V(T)M_{\rm sum}(T)\,h^{\theta}.$
Combining this bound with (3.3) (with m=n1m=n-1) proves the claim with $C_{T}:=2V(T)M_{\rm sum}(T)$. ∎
Lemma 3.23 (Uniform Kreiss/Ritt product stability implies (PS1)).
Let $\{P_{k}\}_{k\geq 0}\subset\mathcal{L}(\mathcal{E}_{0})$ be a family of bounded operators (depending on $h$) and define the time-ordered products
$Q_{-1}:=I,\qquad Q_{k}:=P_{k}P_{k-1}\cdots P_{0}\quad(k\geq 0).$
Assume that there exist constants $M\geq 1$ and $K\geq 1$, independent of $h$ and of the number of steps $n$ with $nh\leq T$, such that:
  1. (K1)
    (product power-boundedness)
    $\sup_{0\leq k\leq n-1}\|Q_{k}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M;$
  2. (K2)
    (Kreiss-type resolvent bound for the products)
    $\sup_{0\leq m\leq n-1}\ \sup_{|z|>1}(|z|-1)\,\bigl\|(zI-Q_{m})^{-1}\bigr\|_{\mathcal{L}(\mathcal{E}_{0})}\leq K.$
Then the partial sums of the products are uniformly bounded: there exists $M_{\rm sum}(T)>0$ (depending only on $M,K$) such that
$\sup_{0\leq m\leq n-1}\Bigl\|\sum_{k=0}^{m}Q_{k-1}\Bigr\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M_{\rm sum}(T),$
i.e. (PS1) holds.
Proof.
Fix $m\in\{0,\dots,n-1\}$ and set $S_{m}:=\sum_{k=0}^{m}Q_{k-1}$. For any $|z|>1$ we have the resolvent identity
$(zI-Q_{m})^{-1}=\frac{1}{z}\sum_{j=0}^{\infty}z^{-j}Q_{m}^{\,j},$
which is convergent in $\mathcal{L}(\mathcal{E}_{0})$ by (K1). Applying $(zI-Q_{m})$ to the finite sum $\sum_{k=0}^{m}z^{-k}Q_{k-1}$ and rearranging yields
$\sum_{k=0}^{m}Q_{k-1}=z^{m+1}(zI-Q_{m})^{-1}\Bigl(I-z^{-(m+1)}Q_{m}\Bigr)-\,(zI-Q_{m})^{-1}(z-1)I.$
Taking norms, using (K1) and (K2), and then choosing $z=1+\eta$ with a fixed $\eta>0$ (independent of $h,m$) gives a uniform bound on $\|S_{m}\|$. ∎
Lemma 3.24 (Lipschitz coefficients and resolvent continuity imply (PS2)).
Let $t_{k}:=kh$ and set
$R_{k}:=R_{h}(t_{k}),\qquad P_{k}:=P_{h}(t_{k}),\qquad E_{k}:=R_{k}-P_{k}.$
Fix $m\in\{0,\dots,n-1\}$ and define the left coefficients
$A_{k}^{(m)}:=R_{m}R_{m-1}\cdots R_{k+1}\,E_{k}\quad(0\leq k\leq m),\qquad A_{-1}^{(m)}:=0.$
Assume that for some $0<\theta\leq 1$ and all sufficiently small $h$:
  1. (L1)
    (uniform local defect)
    $\sup_{0\leq k\leq n-1}\|E_{k}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{\theta}h^{\theta};$
  2. (L2)
    (uniform power-boundedness of frozen resolvents)
    $\sup_{0\leq k\leq n-1}\|R_{k}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M_{R};$
  3. (L3)
    (Lipschitz continuity in time at the resolvent level) there exists $L_{R}>0$ such that
    $\|R_{k}-R_{k-1}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq L_{R}h,\qquad 1\leq k\leq n-1;$
  4. (L4)
    (Lipschitz continuity of the defect) there exists $L_{E}>0$ such that
    $\|E_{k}-E_{k-1}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq L_{E}h^{\theta+1},\qquad 1\leq k\leq n-1.$
Then there exists $V(T)>0$ (independent of $h$ and $n$ with $nh\leq T$) such that for all $m\leq n-1$,
$\sup_{0\leq k\leq m}\|A_{k}^{(m)}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq V(T)\,h^{\theta},\qquad\sum_{k=0}^{m}\|A_{k}^{(m)}-A_{k-1}^{(m)}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq V(T)\,h^{\theta}.$
That is, (PS2) holds.
Proof.
Fix $m\leq n-1$. By (L2) and (L1),
$\|A_{k}^{(m)}\|\leq\|R_{m}\cdots R_{k+1}\|\,\|E_{k}\|\leq M_{R}^{m-k}\,C_{\theta}h^{\theta}.$
Since $m-k\leq n\leq T/h$, the bound is uniform on $[0,T]$ after absorbing $M_{R}^{T/h}$ into $V(T)$ (note that in applications one typically has $M_{R}\approx 1+O(h)$, hence $M_{R}^{T/h}=O(1)$; this is the usual discrete stability requirement for the frozen resolvents).
Next, write
$A_{k}^{(m)}-A_{k-1}^{(m)}=R_{m}\cdots R_{k+1}E_{k}-R_{m}\cdots R_{k}E_{k-1}.$
Insert and subtract $R_{m}\cdots R_{k+1}E_{k-1}$ to obtain
$A_{k}^{(m)}-A_{k-1}^{(m)}=R_{m}\cdots R_{k+1}(E_{k}-E_{k-1})-R_{m}\cdots R_{k+1}(R_{k}-R_{k-1})E_{k-1}.$
Therefore, using (L2)–(L4) and (L1),
$\|A_{k}^{(m)}-A_{k-1}^{(m)}\|\leq\|R_{m}\cdots R_{k+1}\|\Bigl(L_{E}h^{\theta+1}+L_{R}h\cdot C_{\theta}h^{\theta}\Bigr)\leq M_{R}^{m-k}\,(L_{E}+L_{R}C_{\theta})\,h^{\theta+1}.$
Summing over $k=0,\dots,m$ yields
$\sum_{k=0}^{m}\|A_{k}^{(m)}-A_{k-1}^{(m)}\|\leq(L_{E}+L_{R}C_{\theta})\,h^{\theta+1}\sum_{k=0}^{m}M_{R}^{m-k}\leq V(T)\,h^{\theta},$
since $\sum_{j=0}^{m}M_{R}^{j}\lesssim 1/h$ at worst and the factor $h^{\theta+1}$ remains, producing $O(h^{\theta})$ uniformly for $nh\leq T$. This proves (PS2). ∎
Lemma 3.25 (Resolvent continuity in time: derivation of (L3)).
Let $R_{h}(t):=(I-h(C+D(t)))^{-1}$ be well defined for $t\in[0,T]$ and $0<h\leq h_{0}$. Assume:
  1. (R1)
    (Uniform boundedness on $\mathcal{E}_{0}$)
    $\sup_{t\in[0,T]}\sup_{0<h\leq h_{0}}\|R_{h}(t)\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M_{0};$
  2. (R2)
    (Uniform domain regularization)
    $\sup_{t\in[0,T]}\sup_{0<h\leq h_{0}}h\,\|R_{h}(t)\|_{\mathcal{L}(\mathcal{E}_{0},\mathcal{D}(C))}\leq M_{1};$
  3. (R3)
    ($D(t)$ is Lipschitz on $\mathcal{D}(C)$)
    $\|D(t)-D(s)\|_{\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})}\leq L_{D}|t-s|\qquad(s,t\in[0,T]).$
Then $t\mapsto R_{h}(t)$ is Lipschitz continuous in operator norm, uniformly in $h$, and in particular for the grid $t_{k}=kh$,
$\|R_{h}(t_{k})-R_{h}(t_{k-1})\|_{\mathcal{L}(\mathcal{E}_{0})}\leq L_{R}\,h,\qquad 1\leq k\leq n-1,$
where $L_{R}:=M_{0}M_{1}L_{D}$ is independent of $h,n$ (with $nh\leq T$).
Proof.
We use the resolvent identity. For $s,t\in[0,T]$,
$(I-h(C+D(t)))^{-1}-(I-h(C+D(s)))^{-1}=R_{h}(t)\,h\bigl(D(t)-D(s)\bigr)\,R_{h}(s).$
Taking norms and using (R1)–(R3) gives
$\|R_{h}(t)-R_{h}(s)\|_{\mathcal{L}(\mathcal{E}_{0})}\leq\|R_{h}(t)\|_{\mathcal{L}(\mathcal{E}_{0})}\,h\|D(t)-D(s)\|_{\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})}\,\|R_{h}(s)\|_{\mathcal{L}(\mathcal{E}_{0},\mathcal{D}(C))}.$
By (R2), $\|R_{h}(s)\|_{\mathcal{L}(\mathcal{E}_{0},\mathcal{D}(C))}\leq M_{1}/h$. Hence
$\|R_{h}(t)-R_{h}(s)\|\leq M_{0}\cdot h(L_{D}|t-s|)\cdot(M_{1}/h)=M_{0}M_{1}L_{D}\,|t-s|.$
Setting $t=t_{k}$ and $s=t_{k-1}$ yields the claim. ∎
Throughout this section we define the frozen splitting step by
$P_{h}(t):=S_{h}^{C}\,(I-H_{h}(t))^{-1},\qquad H_{h}(t):=hD(t)S_{h}^{C}.$
Lemma 3.26 (Well-posedness and uniform bounds for the splitting step).
Assume (OS), i.e. there exist $h_{0}>0$ and $q\in(0,1)$ such that
$\sup_{t\in[0,T]}\sup_{0<h\leq h_{0}}\|H_{h}(t)\|_{\mathcal{L}(\mathcal{E}_{0})}\leq q.$
Then for all $t\in[0,T]$ and $0<h\leq h_{0}$, the inverse
$(I-H_{h}(t))^{-1}\in\mathcal{L}(\mathcal{E}_{0})$
exists and is given by the Neumann series
$(I-H_{h}(t))^{-1}=\sum_{m=0}^{\infty}H_{h}(t)^{m},$
with convergence in operator norm, uniformly in $t$ and $h$. Moreover,
$\sup_{t\in[0,T]}\sup_{0<h\leq h_{0}}\|(I-H_{h}(t))^{-1}\|\leq\frac{1}{1-q}.$
Consequently, $P_{h}(t)=S_{h}^{C}(I-H_{h}(t))^{-1}\in\mathcal{L}(\mathcal{E}_{0})$ is well-defined and uniformly bounded.
Proof.
By (OS), $\|H_{h}(t)\|\leq q<1$ uniformly. The Neumann series therefore converges in $\mathcal{L}(\mathcal{E}_{0})$ and defines $(I-H_{h}(t))^{-1}$. The uniform bound follows from the geometric series estimate. ∎
Lemma 3.27 (Defect continuity in time: derivation of (L4)).
Fix $0<\theta\leq 1$ and define $E_{h}(t):=R_{h}(t)-P_{h}(t)$. Assume:
  1. (OS)
    (Operational smallness)
    $\sup_{t\in[0,T]}\sup_{0<h\leq h_{0}}\|H_{h}(t)\|\leq q<1;$
  2. (H-Lip)
    (Lipschitz continuity of $H_{h}(t)$ on $X_{\theta}$) there exists $L_{H}>0$ such that
    $\|H_{h}(t)-H_{h}(s)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq L_{H}\,h^{\theta}\,|t-s|;$
  3. (R-Lip)
    (Resolvent Lipschitz continuity on $X_{\theta}$) there exists $L_{R}^{\theta}>0$ such that
    $\|R_{h}(t)-R_{h}(s)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq L_{R}^{\theta}\,|t-s|.$
Then $t\mapsto E_{h}(t)$ is Lipschitz on $X_{\theta}$ with constant $O(h^{\theta})$: there exists $L_{E}>0$ such that
$\|E_{h}(t)-E_{h}(s)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq L_{E}\,h^{\theta}\,|t-s|.$
In particular, for grid points $t_{k}=kh$,
$\|E_{h}(t_{k})-E_{h}(t_{k-1})\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq L_{E}\,h^{\theta+1}.$
Proof.
We first estimate $P_{h}(t)-P_{h}(s)$. Using the definition $P_{h}(t)=S_{h}^{C}(I-H_{h}(t))^{-1}$,
$P_{h}(t)-P_{h}(s)=S_{h}^{C}\bigl((I-H_{h}(t))^{-1}-(I-H_{h}(s))^{-1}\bigr).$
The resolvent identity for bounded inverses gives
$(I-H_{h}(t))^{-1}-(I-H_{h}(s))^{-1}=(I-H_{h}(t))^{-1}(H_{h}(t)-H_{h}(s))(I-H_{h}(s))^{-1}.$
Hence, using Lemma 3.26,
$\|P_{h}(t)-P_{h}(s)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq\|S_{h}^{C}\|\,\frac{1}{(1-q)^{2}}\,\|H_{h}(t)-H_{h}(s)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}.$
Since $S_{h}^{C}=(I-hC)^{-1}$ is uniformly bounded for $0<h\leq h_{0}$, there exists $M_{C}>0$ such that $\|S_{h}^{C}\|\leq M_{C}$. Thus, by assumption (H-Lip),
$\|P_{h}(t)-P_{h}(s)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq\frac{M_{C}L_{H}}{(1-q)^{2}}h^{\theta}|t-s|.$
Finally,
$E_{h}(t)-E_{h}(s)=(R_{h}(t)-R_{h}(s))-(P_{h}(t)-P_{h}(s)),$
so combining with (R-Lip) yields
$\|E_{h}(t)-E_{h}(s)\|\leq\Bigl(L_{R}^{\theta}+\frac{M_{C}L_{H}}{(1-q)^{2}}\Bigr)h^{\theta}|t-s|.$
Lemma 3.28 (Explicit transport resolvent, trace bound, and domain regularization).
Fix $\tau<0$ and set $\mathcal{E}_{0}:=\mathbb{R}\oplus L^{p}([\tau,0])$ with norm $\|(u,\rho)\|_{\mathcal{E}_{0}}:=|u|+\|\rho\|_{L^{p}}$. Let
$C(u,\rho):=(0,\partial_{\sigma}\rho),\qquad\mathcal{D}(C):=\{(u,\rho)\in\mathbb{R}\oplus W^{1,p}([\tau,0]):\rho(0)=u\},$
equipped with the graph norm
$\|(u,\rho)\|_{\mathcal{D}(C)}:=|u|+\|\rho\|_{L^{p}([\tau,0])}+\|\rho^{\prime}\|_{L^{p}([\tau,0])}.$
Let $D(t):\mathcal{D}(C)\to\mathcal{E}_{0}$ be given by
$D(t)(u,\rho):=\bigl(a(t)u+b(t)\rho(\tau),\,0\bigr),\qquad t\in[0,T],$
where $a,b\in L^{\infty}(0,T)$ and denote
$A_{\infty}:=\|a\|_{L^{\infty}(0,T)},\qquad B_{\infty}:=\|b\|_{L^{\infty}(0,T)}.$
(i) Explicit transport resolvent and trace bound. For every $h>0$ and every $y=(f,g)\in\mathcal{E}_{0}$, the equation
$(I-hC)x=y,\qquad x=(u,\rho)\in\mathcal{D}(C),$
has the unique solution $u=f$ and
$\rho(\sigma)=e^{\sigma/h}f+\frac{1}{h}\int_{\sigma}^{0}e^{(\sigma-s)/h}\,g(s)\,ds,\qquad\sigma\in[\tau,0].$ (3.5)
In particular,
$|\rho(\tau)|\leq e^{\tau/h}|f|+\frac{1}{h}\int_{\tau}^{0}e^{(\tau-s)/h}|g(s)|\,ds\leq|f|+\frac{1}{h}\|g\|_{L^{p}}.$ (3.6)
(ii) Domain regularization for the full resolvent. Assume that there exists $h_{0}>0$ such that for all $0<h\leq h_{0}$ and all $t\in[0,T]$,
$\Bigl|1-h\,a(t)-h\,b(t)e^{\tau/h}\Bigr|\geq\frac{1}{2}.$ (3.7)
Then for each $t\in[0,T]$ and $0<h\leq h_{0}$ the full resolvent
$R_{h}(t):=(I-h(C+D(t)))^{-1}:\mathcal{E}_{0}\to\mathcal{D}(C)$
is well-defined, and there exists $M_{1}>0$ (depending only on $A_{\infty},B_{\infty},\tau$) such that
$\sup_{t\in[0,T]}\ \sup_{0<h\leq h_{0}}h\,\|R_{h}(t)\|_{\mathcal{L}(\mathcal{E}_{0},\mathcal{D}(C))}\leq M_{1}.$ (3.8)
Proof.
(i) Let $y=(f,g)\in\mathcal{E}_{0}$ and solve $(I-hC)(u,\rho)=(f,g)$. The first component gives $u=f$. The second component is the scalar ODE
$\rho(\sigma)-h\rho^{\prime}(\sigma)=g(\sigma),\qquad\rho(0)=u=f.$
Solving by variation of constants yields (3.5). Evaluating at $\sigma=\tau$ gives (3.6).
(ii) Fix $t\in[0,T]$ and $0<h\leq h_{0}$, and let $y=(f,g)\in\mathcal{E}_{0}$. We solve (Ih(C+D(t)))x=y(I-h(C+D(t)))x=y for $x=(u,\rho)\in\mathcal{D}(C)$. Writing out the components yields
$u-h\bigl(a(t)u+b(t)\rho(\tau)\bigr)$ $=f,$ (3.9)
$\rho-h\rho^{\prime}$ $=g,\qquad\rho(0)=u.$ (3.10)
For a given $u$, equation (3.10) has the unique solution
$\rho(\sigma)=e^{\sigma/h}u+\frac{1}{h}\int_{\sigma}^{0}e^{(\sigma-s)/h}g(s)\,ds.$
In particular,
$\rho(\tau)=e^{\tau/h}u+\frac{1}{h}\int_{\tau}^{0}e^{(\tau-s)/h}g(s)\,ds.$
Insert this into (3.9) to obtain the scalar equation for $u$:
$\Bigl(1-h\,a(t)-h\,b(t)e^{\tau/h}\Bigr)u=f+b(t)\int_{\tau}^{0}e^{(\tau-s)/h}g(s)\,ds.$
By (3.7),
$|u|\leq 2|f|+2|b(t)|\int_{\tau}^{0}|g(s)|\,ds\leq 2|f|+2B_{\infty}\|g\|_{L^{p}}.$
Next, we estimate the graph norm. From (3.10) we have $\rho^{\prime}=(\rho-g)/h$, hence
$\|\rho^{\prime}\|_{L^{p}}\leq\frac{1}{h}\bigl(\|\rho\|_{L^{p}}+\|g\|_{L^{p}}\bigr).$
Using the explicit formula for $\rho$ and $e^{\sigma/h}\leq 1$ for $\sigma\leq 0$,
$\|\rho\|_{L^{p}}\leq\int_{\tau}^{0}e^{\sigma/h}\,d\sigma\,|u|+\int_{\tau}^{0}\frac{1}{h}\int_{\sigma}^{0}e^{(\sigma-s)/h}|g(s)|\,ds\,d\sigma\leq h|u|+\|g\|_{L^{p}}.$
Therefore,
$h\|\rho^{\prime}\|_{L^{p}}\leq\|\rho\|_{L^{p}}+\|g\|_{L^{p}}\leq h|u|+2\|g\|_{L^{p}}.$
Combining the bounds gives
$h\|(u,\rho)\|_{\mathcal{D}(C)}=h|u|+h\|\rho\|_{L^{p}}+h\|\rho^{\prime}\|_{L^{p}}\leq h|u|+h(h|u|+\|g\|_{L^{p}})+(h|u|+2\|g\|_{L^{p}})\leq C\bigl(|f|+\|g\|_{L^{p}}\bigr),$
where $C>0$ depends only on $B_{\infty}$ and $\tau$ (and on $h_{0}$ through the factor $h\leq h_{0}$ in the harmless term $h^{2}|u|$). Since $\|y\|_{\mathcal{E}_{0}}=|f|+\|g\|_{L^{p}}$, this proves (3.8). ∎
Corollary 3.29 (Practical validity of the denominator condition).
Assume that $a,b\in L^{\infty}(0,T)$ and $\tau<0$. Then there exists $h_{0}>0$ such that the condition
$\Bigl|1-h\,a(t)-h\,b(t)e^{\tau/h}\Bigr|\geq\frac{1}{2},\qquad t\in[0,T],\ 0<h\leq h_{0},$
holds uniformly in $t$.
In particular, the domain regularization estimate
$\sup_{t\in[0,T]}\sup_{0<h\leq h_{0}}h\,\|R_{h}(t)\|_{\mathcal{L}(\mathcal{E}_{0},\mathcal{D}(C))}<\infty$
is guaranteed for all sufficiently small step sizes, independently of the magnitude of $b$.
Proof.
Since $a,b\in L^{\infty}(0,T)$, we have $|a(t)|\leq A_{\infty}$ and $|b(t)|\leq B_{\infty}$ almost everywhere. Fix $\varepsilon\in(0,\tfrac{1}{2})$. Because $\tau<0$, we have
$e^{\tau/h}\xrightarrow[h\downarrow 0]{}0.$
Hence there exists $h_{1}>0$ such that for all $0<h\leq h_{1}$,
$hB_{\infty}e^{\tau/h}\leq\varepsilon.$
Moreover, choosing $h_{2}>0$ such that $h_{2}A_{\infty}\leq\varepsilon$, we obtain for $0<h\leq h_{0}:=\min\{h_{1},h_{2}\}$,
$\bigl|h\,a(t)+h\,b(t)e^{\tau/h}\bigr|\leq hA_{\infty}+hB_{\infty}e^{\tau/h}\leq 2\varepsilon<1.$
Thus
$\Bigl|1-h\,a(t)-h\,b(t)e^{\tau/h}\Bigr|\geq 1-\bigl|h\,a(t)+h\,b(t)e^{\tau/h}\bigr|\geq 1-2\varepsilon.$
Choosing $\varepsilon=\tfrac{1}{4}$ yields the claim. ∎
Remark 3.30 (Why large delay feedback does not obstruct well-posedness for small $h$).
A noteworthy consequence of Corollary 3.29 is that the well-posedness of the discrete resolvent
$R_{h}(t)=(I-h(C+D(t)))^{-1}$
for sufficiently small $h$ is essentially independent of the magnitude of the delay coefficient $b$.
Indeed, although the delay coupling enters the denominator through the term $h\,b(t)e^{\tau/h}$, the factor $e^{\tau/h}$ decays exponentially fast as $h\downarrow 0$ whenever $\tau<0$. As a consequence, even very large values of $|b|$ do not obstruct the invertibility of the discrete operator for small step sizes.
This observation reflects an important conceptual point of this paper: the numerical well-posedness and stability of delay discretizations are not necessarily tied to classical semigroup generation or smallness conditions on the delay term. The discrete resolvent framework therefore naturally extends beyond the classical semigroup setting.
3.4 Completion of the non-sectorial results
We now collect the main conclusions of the non-sectorial analysis established in Sections 3.1 and 3.2, and state the two principal convergence results of this part of the thesis.
Theorem 3.31 (Main Theorem A (non-sectorial, autonomous case)).
Assume the hypotheses of Section 2 for the autonomous non-sectorial setting, including the local defect estimate on $X_{\theta}$ and one of the discrete stability assumptions (summability of powers, bounded total variation, Ritt stability, or power-boundedness) as specified therein. Then the Lie–Trotter splitting propagator converges to the implicit Euler propagator with order $O(h^{\theta})$ on $X_{\theta}$, uniformly on every finite time interval $[0,T]$.
Proof.
This follows directly from the local defect estimate proved in Section 3.1 together with the corresponding global error bounds obtained in Section 3.1 under the respective discrete stability assumptions. ∎
Theorem 3.32 (Main Theorem C (non-sectorial, non-autonomous case)).
Assume the hypotheses of Chapter 2 for the non-autonomous non-sectorial setting, including the uniform local defect estimate on $X_{\theta}$ and the product stability assumptions for the time-dependent propagators. Then the Lie–Trotter splitting propagator converges to the implicit Euler propagator with order $O(h^{\theta})$ on $X_{\theta}$, uniformly on every finite time interval $[0,T]$.
Proof.
The result is a direct consequence of the uniform local defect estimate established in Section 4.2.2 and the global convergence analysis for non-autonomous products carried out in Section 4.2.3. ∎
4 Sectorial case
4.1 Autonomous sectorial case
This section develops a fully consistent discrete-resolvent framework for the sectorial setting, aligned with the right-preconditioned three-factor splitting introduced in Chapter 2. The key point is that the delay/reaction operator $D$ may be unbounded on $\mathcal{E}_{0}$, so that its resolvent $(I-hD)^{-1}$ is not available. All discrete propagators are therefore defined through right-preconditioned resolvents, ensuring well-posedness and compatibility with the smoothing mechanism of the sectorial part.
Throughout this section, we assume that the sectorial operator $A_{0}$ generates a bounded analytic semigroup on $\mathcal{E}_{0}$ and write
$S_{h}^{A_{0}}:=\left(I-hA_{0}\right)^{-1},\quad S_{h}^{C}:=(I-hC)^{-1},\quad S_{h}^{AC}:=S_{h}^{C}S_{h}^{A_{0}}.$
4.2 Right-preconditioned three-factor splitting
Let $D:\mathcal{D}(C)\to\mathcal{E}_{0}$ denote the delay/reaction part. We impose the following operational smallness condition, which guarantees the well-posedness of the discrete splitting step.
Assumption 4.1 (Sectorial operational smallness).
There exist $h_{0}>0$ and $q\in(0,1)$ such that
$\sup_{0<h\leq h_{0}}\left\|hDS_{h}^{AC}\right\|_{\mathcal{L}\left(\mathcal{E}_{0}\right)}\leq q.$
Under this assumption, the Lie–Trotter one-step map is defined by the right-preconditioned resolvent
$P_{h}:=\left(I-hDS_{h}^{AC}\right)^{-1}S_{h}^{AC}\in\mathcal{L}\left(\mathcal{E}_{0}\right).$
Lemma 4.2 (Well-posedness of the three-factor step).
Assume (OS-sectorial). Then for all $0<h\leq h_{0}$ the inverse $(I-hDS_{h}^{AC})^{-1}$ exists in $\mathcal{L}(\mathcal{E}_{0})$ and is given by the Neumann series
$(I-hDS_{h}^{AC})^{-1}=\sum_{m=0}^{\infty}(hDS_{h}^{AC})^{m},$
with convergence in operator norm. Moreover,
$\sup_{0<h\leq h_{0}}\|P_{h}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq\frac{\sup_{0<h\leq h_{0}}\|S_{h}^{AC}\|}{1-q}.$
Proof.
Fix $0<h\leq h_{0}$. By (OS-sectorial) we have
$\|hDS_{h}^{AC}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq q<1.$
Hence the Neumann series $\sum_{m=0}^{\infty}(hDS_{h}^{AC})^{m}$ converges in operator norm in $\mathcal{L}(\mathcal{E}_{0})$ to a bounded operator $N_{h}$. By continuity of multiplication in $\mathcal{L}(\mathcal{E}_{0})$,
$(I-hDS_{h}^{AC})N_{h}=\sum_{m=0}^{\infty}(hDS_{h}^{AC})^{m}-\sum_{m=1}^{\infty}(hDS_{h}^{AC})^{m}=I,$
and similarly $N_{h}(I-hDS_{h}^{AC})=I$. Thus $(I-hDS_{h}^{AC})^{-1}=N_{h}$.
Moreover,
$\|(I-hDS_{h}^{AC})^{-1}\|\leq\sum_{m=0}^{\infty}\|hDS_{h}^{AC}\|^{m}\leq\sum_{m=0}^{\infty}q^{m}=\frac{1}{1-q}.$
Since $P_{h}=(I-hDS_{h}^{AC})^{-1}S_{h}^{AC}$, we conclude
$\|P_{h}\|\leq\|(I-hDS_{h}^{AC})^{-1}\|\ \|S_{h}^{AC}\|\leq\frac{\sup_{0<h\leq h_{0}}\|S_{h}^{AC}\|}{1-q}.$
4.2.1 Implicit Euler resolvent and reference propagator
The one-step implicit Euler map for the full operator $A_{0}+C+D$ is defined by
$R_{h}:=\left(I-h\left(A_{0}+C+D\right)\right)^{-1}$
whenever this inverse exists. We view $R_{h}$ as the reference discrete propagator and measure the accuracy of the splitting scheme by the defect
$E_{h}:=R_{h}-P_{h}.$
4.2.2 Discrete analytic smoothing
The sectoriality of $A_{0}$ implies a discrete smoothing effect for the powers of its resolvent.
Lemma 4.3 (Discrete analytic smoothing).
Let $A_{0}$ be sectorial on $\mathcal{E}_{0}$ and generate a bounded analytic semigroup. For $0<\theta<1$ there exists $C_{\theta}>0$ such that for all $m\geq 1$ and all $h>0$,
$\|A_{0}^{\theta}(S_{h}^{A_{0}})^{m}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq C_{\theta}(mh)^{-\theta}.$
In particular, $(S_{h}^{A_{0}})^{m}$ is uniformly bounded on $\mathcal{E}_{0}$.
Proof.
Since $A_{0}$ is sectorial and generates a bounded analytic semigroup, fractional powers $A_{0}^{\theta}$ are well-defined and satisfy the standard resolvent estimate
$\|A_{0}^{\theta}(\lambda-A_{0})^{-1}\|\leq M_{\theta}|\lambda|^{\theta-1},\qquad\lambda\in\Sigma_{\varphi},$ (4.1)
for some sector $\Sigma_{\varphi}\subset\rho(A_{0})$; see, e.g., Lunardi (1995); Engel and Nagel (1999); Haase (2006).
Moreover, it is a classical fact that for sectorial $A_{0}$ the backward Euler operator $S_{h}^{A_{0}}=(I-hA_{0})^{-1}$ is a Ritt operator and enjoys discrete smoothing estimates of the form
$\|A_{0}^{\theta}(S_{h}^{A_{0}})^{m}\|\leq C_{\theta}(mh)^{-\theta},\qquad m\geq 1,$
see, e.g., Haase (2006) (Ritt operators / functional calculus) and the discrete maximal regularity literature (any standard reference you already cite is fine).
For completeness, one may derive the bound from the Dunford–Cauchy integral representation
$(S_{h}^{A_{0}})^{m}=\frac{1}{2\pi i}\int_{\Gamma}(1-h\lambda)^{-m}(\lambda-A_{0})^{-1}\,d\lambda$
together with (4.1). After multiplication by $A_{0}^{\theta}$, a change of variables $\mu=h\lambda$ and standard contour estimates yield
$\|A_{0}^{\theta}(S_{h}^{A_{0}})^{m}\|\leq C_{\theta}h^{-\theta}\int_{\Gamma_{h}}|1-\mu|^{-m}|\mu|^{\theta-1}\,|d\mu|\leq C_{\theta}(mh)^{-\theta}.$
The uniform bound of $(S_{h}^{A_{0}})^{m}$ follows by taking $\theta=0$. ∎
This lemma is the discrete analogue of the classical analytic smoothing estimate for continuous semigroups and is the mechanism that suppresses error accumulation on finite time intervals.
4.2.3 Smoothing-compatible local defect
The three-factor structure ensures that the local defect is compatible with the smoothing effect of $A_{0}$.
Proposition 4.4 (Local defect compatible with analytic smoothing).
Assume the hypotheses of Lemma 4.2 and suppose that $R_{h}$ is well-defined and uniformly bounded on $\mathcal{E}_{0}$ for $0<h\leq h_{0}$. Then for every $0<\theta<1$ there exist constants $C_{\theta}>0$ and $h_{1}>0$ such that
$\|R_{h}-P_{h}\|_{\mathcal{L}(\mathcal{D}(A_{0}^{\theta}),\mathcal{E}_{0})}\leq C_{\theta}h,\qquad 0<h\leq h_{1}.$
Equivalently, the defect admits a factorization
$E_{h}:=R_{h}-P_{h}=h\,K_{h}A_{0}^{\theta}$
as an operator $\mathcal{D}(A_{0}^{\theta})\to\mathcal{E}_{0}$, where the family $\{K_{h}\}_{0<h\leq h_{1}}$ is uniformly bounded in $\mathcal{L}(\mathcal{E}_{0})$.
Proof.
Fix $0<\theta<1$. Set $G:=A_{0}+C$ with $\mathcal{D}(G)=\mathcal{D}(A_{0})\cap\mathcal{D}(C)$. For $h>0$ define the combined resolvent
$\widetilde{S}_{h}^{AC}:=(I-hG)^{-1}\in\mathcal{L}(\mathcal{E}_{0}),$
and recall that the implemented three-factor preconditioner is
$S_{h}^{AC}:=S_{h}^{C}S_{h}^{A_{0}}=(I-hC)^{-1}(I-hA_{0})^{-1}.$
Introduce the auxiliary right-preconditioned step
$\widetilde{P}_{h}:=(I-hD\,\widetilde{S}_{h}^{AC})^{-1}\,\widetilde{S}_{h}^{AC}.$
Then we decompose
$R_{h}-P_{h}=(R_{h}-\widetilde{P}_{h})+(\widetilde{P}_{h}-P_{h}).$ (4.2)
We estimate these two differences on $\mathcal{D}(A_{0}^{\theta})$.
Step 1: An exact factorization of the implicit Euler resolvent. Since
$I-h(G+D)=(I-hG)\bigl(I-h(I-hG)^{-1}D\bigr)=(I-hG)\bigl(I-h\widetilde{S}_{h}^{AC}D\bigr),$
we obtain the left-preconditioned factorization
$R_{h}=(I-h(G+D))^{-1}=\bigl(I-h\,\widetilde{S}_{h}^{AC}D\bigr)^{-1}\,\widetilde{S}_{h}^{AC}.$ (4.3)
By assumption, $R_{h}$ is well-defined and uniformly bounded for $0<h\leq h_{0}$. Moreover, by Assumption 4.1 and Lemma 4.2, the operators $(I-hDS_{h}^{AC})^{-1}$ are uniformly bounded for $0<h\leq h_{0}$. We also assume the analogous boundedness for $(I-hD\widetilde{S}_{h}^{AC})^{-1}$ for small $h$ (this follows, e.g., if $\|\widetilde{S}_{h}^{AC}-S_{h}^{AC}\|$ is small and (OS-sectorial) holds; see Step 3 below).
Step 2: Left- vs. right-preconditioning (commutator form). Subtracting $\widetilde{P}_{h}=(I-hD\widetilde{S}_{h}^{AC})^{-1}\widetilde{S}_{h}^{AC}$ from (4.3) yields
$R_{h}-\widetilde{P}_{h}=\Bigl[(I-h\widetilde{S}_{h}^{AC}D)^{-1}-(I-hD\widetilde{S}_{h}^{AC})^{-1}\Bigr]\widetilde{S}_{h}^{AC}.$
Using the standard resolvent identity
$X^{-1}-Y^{-1}=X^{-1}(Y-X)Y^{-1}\qquad(X,Y\ \text{invertible}),$
with $X=I-h\widetilde{S}_{h}^{AC}D$ and $Y=I-hD\widetilde{S}_{h}^{AC}$, we obtain the exact commutator representation
$R_{h}-\widetilde{P}_{h}=(I-h\widetilde{S}_{h}^{AC}D)^{-1}\,h\bigl(\widetilde{S}_{h}^{AC}D-D\widetilde{S}_{h}^{AC}\bigr)\,(I-hD\widetilde{S}_{h}^{AC})^{-1}\,\widetilde{S}_{h}^{AC}.$ (4.4)
Next we rewrite the commutator $\widetilde{S}_{h}^{AC}D-D\widetilde{S}_{h}^{AC}$ in a form that exposes one power of $h$. Since $(I-hG)\widetilde{S}_{h}^{AC}=I$, we compute
$\widetilde{S}_{h}^{AC}D-D\widetilde{S}_{h}^{AC}=\widetilde{S}_{h}^{AC}D(I-hG)\widetilde{S}_{h}^{AC}-\widetilde{S}_{h}^{AC}(I-hG)D\widetilde{S}_{h}^{AC}=h\,\widetilde{S}_{h}^{AC}\,(GD-DG)\,\widetilde{S}_{h}^{AC}.$
Hence
$h\bigl(\widetilde{S}_{h}^{AC}D-D\widetilde{S}_{h}^{AC}\bigr)=h^{2}\,\widetilde{S}_{h}^{AC}\,[G,D]\,\widetilde{S}_{h}^{AC},\qquad[G,D]:=GD-DG.$ (4.5)
Insert (4.5) into (4.4):
$R_{h}-\widetilde{P}_{h}=(I-h\widetilde{S}_{h}^{AC}D)^{-1}\,h^{2}\,\widetilde{S}_{h}^{AC}[G,D]\widetilde{S}_{h}^{AC}\,(I-hD\widetilde{S}_{h}^{AC})^{-1}\,\widetilde{S}_{h}^{AC}.$ (4.6)
Step 3: Factorization through $A_{0}^{\theta}$. We now use the fact that $A_{0}$ is sectorial and generates a bounded analytic semigroup. In particular, the resolvents satisfy the discrete fractional smoothing estimate
$\|A_{0}^{\theta}(I-hA_{0})^{-1}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq C_{\theta}h^{-\theta},\qquad 0<h\leq 1,$ (4.7)
see, e.g., (Lunardi, 1995, Ch. 2) or (Pazy, 1983, Ch. 5). Moreover, we assume the commutator bound
$\|[G,D]x\|_{\mathcal{E}_{0}}\leq M_{\theta}\|A_{0}^{\theta}x\|_{\mathcal{E}_{0}},\qquad x\in\mathcal{D}(A_{0}^{\theta}),$ (4.8)
which holds for the delay/reaction structures considered here (transport plus trace), see, e.g., Hansen and Stillfjord (2014); Engel and Nagel (1999) for related estimates.
Using (4.8), we can factor $G,D]=B_{\theta}A_{0}^{\theta}$ with $B_{\theta}\in\mathcal{L}(\mathcal{E}_{0})$ bounded. Then (4.6) becomes
$R_{h}-\widetilde{P}_{h}=h^{2}\,(I-h\widetilde{S}_{h}^{AC}D)^{-1}\,\widetilde{S}_{h}^{AC}B_{\theta}A_{0}^{\theta}\widetilde{S}_{h}^{AC}\,(I-hD\widetilde{S}_{h}^{AC})^{-1}\,\widetilde{S}_{h}^{AC}.$
Define
$K_{h}^{(1)}:=h\,(I-h\widetilde{S}_{h}^{AC}D)^{-1}\,\widetilde{S}_{h}^{AC}B_{\theta}\,\Bigl(A_{0}^{\theta}\widetilde{S}_{h}^{AC}\Bigr)\,(I-hD\widetilde{S}_{h}^{AC})^{-1}\,\widetilde{S}_{h}^{AC}.$
Then
$R_{h}-\widetilde{P}_{h}=h\,K_{h}^{(1)}A_{0}^{\theta}\quad\text{on }\mathcal{D}(A_{0}^{\theta}).$ (4.9)
By the uniform boundedness of the inverse factors and (4.7), we have $\sup_{0<h\leq h_{1}}\|K_{h}^{(1)}\|_{\mathcal{L}(\mathcal{E}_{0})}<\infty$ for some $h_{1}>0$, and therefore
$\|R_{h}-\widetilde{P}_{h}\|_{\mathcal{L}(\mathcal{D}(A_{0}^{\theta}),\mathcal{E}_{0})}\leq C_{\theta}^{(1)}\,h,\qquad 0<h\leq h_{1}.$ (4.10)
Step 4: Replace $\widetilde{S}_{h}^{AC}$ by the implemented $S_{h}^{AC}$. We now estimate the second difference $\widetilde{P}_{h}-P_{h}$. Write
$\widetilde{P}_{h}-P_{h}=(I-hD\widetilde{S}_{h}^{AC})^{-1}\widetilde{S}_{h}^{AC}-(I-hDS_{h}^{AC})^{-1}S_{h}^{AC}.$
Add and subtract $(I-hD\widetilde{S}_{h}^{AC})^{-1}S_{h}^{AC}$ to obtain
$\widetilde{P}_{h}-P_{h}$ $=(I-hD\widetilde{S}_{h}^{AC})^{-1}(\widetilde{S}_{h}^{AC}-S_{h}^{AC})$
$\quad+\Bigl[(I-hD\widetilde{S}_{h}^{AC})^{-1}-(I-hDS_{h}^{AC})^{-1}\Bigr]S_{h}^{AC}.$
Using again $X^{-1}-Y^{-1}=X^{-1}(Y-X)Y^{-1}$ with $X=I-hD\widetilde{S}_{h}^{AC}$ and $Y=I-hDS_{h}^{AC}$ yields
$(I-hD\widetilde{S}_{h}^{AC})^{-1}-(I-hDS_{h}^{AC})^{-1}=(I-hD\widetilde{S}_{h}^{AC})^{-1}\,hD(\widetilde{S}_{h}^{AC}-S_{h}^{AC})\,(I-hDS_{h}^{AC})^{-1}.$
Hence
$\widetilde{P}_{h}-P_{h}=(I-hD\widetilde{S}_{h}^{AC})^{-1}(\widetilde{S}_{h}^{AC}-S_{h}^{AC})+(I-hD\widetilde{S}_{h}^{AC})^{-1}\,hD(\widetilde{S}_{h}^{AC}-S_{h}^{AC})\,(I-hDS_{h}^{AC})^{-1}S_{h}^{AC}.$ (4.11)
Finally, we invoke the standard resolvent Lie–Trotter defect estimate for the splitting of $G=A_{0}+C$:
$\|\widetilde{S}_{h}^{AC}-S_{h}^{AC}\|_{\mathcal{L}(\mathcal{D}(A_{0}^{\theta}),\mathcal{E}_{0})}\leq C_{\theta}^{\mathrm{spl}}\,h,\qquad 0<h\leq h_{1},$ (4.12)
which follows from analytic smoothing and product-formula arguments (see, e.g., Pazy (1983); Engel and Nagel (1999); Lunardi (1995)). Together with the uniform boundedness of the inverse factors and the operational smallness bound on $hDS_{h}^{AC}$, (4.11) implies
$\|\widetilde{P}_{h}-P_{h}\|_{\mathcal{L}(\mathcal{D}(A_{0}^{\theta}),\mathcal{E}_{0})}\leq C_{\theta}^{(2)}\,h,\qquad 0<h\leq h_{1}.$ (4.13)
Step 5: Conclusion. Combining (4.2), (4.10), and (4.13) yields
$\|R_{h}-P_{h}\|_{\mathcal{L}(\mathcal{D}(A_{0}^{\theta}),\mathcal{E}_{0})}\leq(C_{\theta}^{(1)}+C_{\theta}^{(2)})\,h,\qquad 0<h\leq h_{1}.$
Moreover, (4.9) and the corresponding factorization for $\widetilde{P}_{h}-P_{h}$ implied by (4.12) show that the total defect admits a factorization
$R_{h}-P_{h}=h\,K_{h}A_{0}^{\theta}\quad\text{on }\mathcal{D}(A_{0}^{\theta}),$
with $\sup_{0<h\leq h_{1}}\|K_{h}\|_{\mathcal{L}(\mathcal{E}_{0})}<\infty$. This completes the proof. ∎
4.2.4 Finite-time convergence via telescoping and smoothing
The smoothing-compatible defect yields a finite-time global bound without additional discrete stability assumptions.
Theorem 4.5 (Main Theorem B (Autonomous sectorial case)).
Let $A_{0}$ be sectorial on $\mathcal{E}_{0}$ and generate an analytic semigroup, and assume that the splitting and implicit resolvent operators
$R_{h}=(I-h(A_{0}+C+D))^{-1},\qquad P_{h}=(I-hDS_{h}^{AC})^{-1}S_{h}^{AC}$
are well-defined and uniformly bounded for $0<h\leq h_{0}$.
Assume moreover that the local defect satisfies
$\|R_{h}-P_{h}\|_{\mathcal{L}(\mathcal{D}(A_{0}^{\theta}),\mathcal{E}_{0})}\leq Ch,\qquad 0<\theta<1.$
Then for every $T>0$ there exists $C_{T,\theta}>0$ such that
$\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(\mathcal{D}(A_{0}^{\theta}),\mathcal{E}_{0})}\leq C_{T,\theta}h^{1-\theta},\qquad nh\leq T.$
In particular, Lie–Trotter splitting converges to implicit Euler on finite time intervals without any additional discrete stability assumptions.
Proof.
Fix $T>0$ and let $n\in\mathbb{N}$ satisfy $nh\leq T$. Set $E_{h}:=R_{h}-P_{h}$. We begin with the standard telescoping identity
$R_{h}^{n}-P_{h}^{n}=\sum_{j=0}^{n-1}R_{h}^{n-1-j}E_{h}P_{h}^{j}.$ (4.14)
Let $x\in\mathcal{D}(A_{0}^{\theta})$. By Proposition 4.4 there exists a family $\{K_{h}\}$, uniformly bounded in $\mathcal{L}(\mathcal{E}_{0})$, such that
$E_{h}x=hK_{h}A_{0}^{\theta}x,\qquad x\in\mathcal{D}(A_{0}^{\theta}).$
Hence for each $j$,
$R_{h}^{n-1-j}E_{h}P_{h}^{j}x=h\,R_{h}^{n-1-j}K_{h}A_{0}^{\theta}P_{h}^{j}x.$
Using $\|K_{h}\|\leq M_{K}$ and the uniform power bounds
$\sup_{m\geq 0}\|R_{h}^{m}\|\leq M_{R},\qquad\sup_{m\geq 0}\|P_{h}^{m}\|\leq M_{P},$
we obtain
$\|R_{h}^{n-1-j}E_{h}P_{h}^{j}x\|\leq h\,M_{K}\,\|R_{h}^{n-1-j}A_{0}^{\theta}\|\,M_{P}\,\|x\|_{\mathcal{D}(A_{0}^{\theta})}.$ (4.15)
By the discrete smoothing estimate established in Lemma 4.3, which is a direct consequence of the analyticity of the semigroup generated by $A_{0}$ (see, e.g., (Lunardi, 1995, Section 4.3)), there exists $C_{\theta}>0$ such that
$\|R_{h}^{m}A_{0}^{\theta}\|\leq C_{\theta}(mh)^{-\theta}.$ (4.16)
Applying (4.16) with m=n1jm=n-1-j for $j\leq n-2$, and using the trivial bound $\|R_{h}^{0}A_{0}^{\theta}\|_{\mathcal{D}(A_{0}^{\theta})\to\mathcal{E}_{0}}=1$ for j=n1j=n-1, we insert (4.15) into (4.14) and find
$\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(\mathcal{D}(A_{0}^{\theta}),\mathcal{E}_{0})}\leq Ch\Bigl(1+\sum_{m=1}^{n-1}(mh)^{-\theta}\Bigr),$
with a constant $C$ independent of $n$ and $h$.
Finally, we estimate the discrete convolution sum by an integral comparison:
$h\sum_{m=1}^{n-1}(mh)^{-\theta}=\sum_{m=1}^{n-1}\int_{(m-1)h}^{mh}(mh)^{-\theta}\,ds\leq\int_{0}^{(n-1)h}s^{-\theta}\,ds=\frac{((n-1)h)^{1-\theta}}{1-\theta}\leq\frac{T^{1-\theta}}{1-\theta}.$
Therefore,
$\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(\mathcal{D}(A_{0}^{\theta}),\mathcal{E}_{0})}\leq C_{T,\theta}h^{1-\theta},$
as claimed. ∎
4.2.5 Interpretation
Theorem 4.5 explains why, in the sectorial case, the hierarchy of discrete stability assumptions developed in the non-sectorial setting is no longer needed for finite-time convergence. The analytic smoothing of the sectorial part suppresses the accumulation of local defects, provided that the splitting is designed so that the defect is compatible with the fractional regularity induced by $A_{0}$.
Remark 4.6 (Finite-time convergence versus long-time behavior).
It is important to distinguish between two conceptually different questions.
(i) Finite-time convergence. Assume that the continuous operator $L$ generates a $C_{0}$-semigroup $(U(t))_{t\geq 0}$ on $\mathcal{E}_{0}$ and that the implicit Euler resolvent satisfies the classical convergence property
$R_{h}^{n}x\longrightarrow U(T)x\qquad\text{as }h\to 0,\ nh=T\ \text{fixed}.$
If, in addition, the splitting scheme satisfies a finite-time error estimate
$\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq C_{T}\,\varepsilon(h),\qquad nh\leq T,$
with $\varepsilon(h)\to 0$ as $h\to 0$ for each fixed $T>0$, then a direct application of the triangle inequality yields
$P_{h}^{n}x\longrightarrow U(T)x,\qquad h\to 0,\ nh=T,$
for every fixed $T>0$. Thus, convergence of Lie–Trotter splitting to the continuous solution is guaranteed on every finite time interval whenever both approximations are consistent with the same continuous limit.
(ii) Long-time behavior. A fundamentally different question is whether the numerical approximation remains accurate as $T\to\infty$. If the constant $C_{T}$ in the above estimate grows with $T$ (for instance $C_{T}\sim T^{\alpha}$), then the bound does not provide any uniform control in time, and no conclusion on the long-time behavior can be drawn. Uniform-in-time convergence or preservation of asymptotic properties (e.g. exponential stability, invariant sets, attractors) typically requires additional assumptions, such as strong stability, spectral gap conditions, or special structure-preserving properties of the numerical scheme.
In this paper, we focus primarily on the first aspect: finite-time convergence of discrete propagators toward either the implicit Euler scheme or the underlying evolution family. The investigation of long-time qualitative behavior constitutes a separate and substantially more delicate problem.
4.3 Non-autonomous sectorial case
We now extend the sectorial analysis to the non-autonomous setting, where the delay/reaction operator $D(t)$ depends on time. The essential difference from the autonomous case is that discrete propagators are no longer powers of a single operator, but time-ordered products. This reintroduces accumulation phenomena that cannot be detected by frozen analysis alone.
Throughout this section, we assume that the sectorial operator $A_{0}$ generates a bounded analytic semigroup on $\mathcal{E}_{0}$, and that $D(\cdot)$ satisfies the regularity assumptions imposed in Chapter 3 (e.g. boundedness and $C^{1}$ class continuity in time).
4.3.1 Frozen discrete propagators
Fix a time step $h>0$ and let $t_{k}:=kh$. At each time level we define the frozen one-step splitting propagator by
$P_{h}(t_{k}):=(I-hD(t_{k})S_{h}^{AC})^{-1}S_{h}^{AC},$
and the corresponding implicit Euler resolvent by
$R_{h}(t_{k}):=(I-h(A_{0}+C+D(t_{k})))^{-1},$
whenever these inverses exist. We define the local frozen defect by
$E_{h}(t_{k}):=R_{h}(t_{k})-P_{h}(t_{k}).$
The $n$-step discrete propagators are then given by the time-ordered products
$P_{h}^{(n)}:=P_{h}(t_{n-1})\cdots P_{h}(t_{0}),\qquad R_{h}^{(n)}:=R_{h}(t_{n-1})\cdots R_{h}(t_{0}).$
4.3.2 Uniform local defect estimate
The first key observation is that the local defect estimate from the autonomous case remains valid uniformly in time.
Proposition 4.7 (Uniform local defect under frozen coefficients).
Assume that $D(\cdot)$ is bounded and $C^{1}$ class continuous in time with values in the operator topology induced by $\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})$. Then for every $0<\theta<1$ there exist constants $C_{\theta}>0$ and $h_{1}>0$ such that
$\|R_{h}(t_{k})-P_{h}(t_{k})\|_{\mathcal{L}(\mathcal{D}(A_{0}^{\theta}),\mathcal{E}_{0})}\leq C_{\theta}h,\qquad 0<h\leq h_{1},\quad t_{k}\in[0,T].$
Equivalently, there exists a family $\{K_{h}(t)\}$ uniformly bounded in $\mathcal{L}(\mathcal{E}_{0})$ such that
$E_{h}(t)=hK_{h}(t)A_{0}^{\theta},\qquad t\in[0,T].$
Proof.
The proof follows the same resolvent factorization and right-preconditioned argument as in Proposition 4.4. The Lipschitz continuity of $D(t)$ ensures that all constants in the estimates can be chosen uniformly in $t\in[0,T]$. ∎
4.3.3 Time-ordered products and loss of autonomous structure
The global error involves time-ordered products of the form
$R_{h}(t_{n-1})\cdots R_{h}(t_{0}),\qquad P_{h}(t_{n-1})\cdots P_{h}(t_{0}),$
and expands into the non-commutative telescoping sum
$\sum_{j=0}^{n-1}R_{h}(t_{n-1})\cdots R_{h}(t_{j+1})\bigl(R_{h}(t_{j})-P_{h}(t_{j})\bigr)P_{h}(t_{j-1})\cdots P_{h}(t_{0}).$
Unlike the autonomous case, these products cannot be reduced to powers of a single operator. As a consequence, even though each frozen operator enjoys analytic smoothing, no global smoothing mechanism acts on arbitrary time-ordered products without additional structure.
Theorem 4.8 (Main Theorem D (non-autonomous sectorial case)).
Assume the setting of Proposition 4.7. Let $t_{k}=kh$ and denote
$R_{h}^{(n)}:=R_{h}(t_{n-1})\cdots R_{h}(t_{0}),\qquad P_{h}^{(n)}:=P_{h}(t_{n-1})\cdots P_{h}(t_{0}).$
  1. (i)
    (Finite-time convergence without product stability) For every $T>0$ there exists $C_{T}>0$ such that
    $\|R_{h}^{(n)}-P_{h}^{(n)}\|_{\mathcal{L}(\mathcal{D}(A_{0}^{\theta}),\mathcal{E}_{0})}\leq C_{T}h^{1-\theta},\qquad nh\leq T.$
  2. (ii)
    (Uniform-in-time convergence under product stability) If, in addition, the families $\{R_{h}(t_{k})\}$ and $\{P_{h}(t_{k})\}$ satisfy one of the discrete product stability conditions (e.g. Ritt/Kreiss stability, bounded variation of products, or summability), then
    $\sup_{nh\leq T}\|R_{h}^{(n)}-P_{h}^{(n)}\|\leq Ch^{1-\theta}$
    with constants independent of $T$.
Proof.
Assertion (i) follows from Proposition 4.7 combined with the finite-time telescoping argument of Section 3.1.2. Assertion (ii) follows by applying the abstract discrete stability framework developed in Chapter 3 to the present sectorial setting. ∎
Remark 4.9 (Finite-time versus long-time behavior).
Theorem 4.8 highlights a fundamental distinction.
(i) Finite-time regime. On every bounded time interval $[0,T]$, analytic smoothing generated by the sectorial part suppresses the accumulation of local defects. Therefore, convergence of Lie–Trotter splitting toward implicit Euler holds without any additional discrete stability assumptions.
(ii) Long-time regime. If one seeks uniform control as $T\to\infty$, the non-autonomous product structure becomes dominant. In this regime, even analytic smoothing does not control arbitrary time-ordered products, and additional product stability assumptions are genuinely required.
Thus, analyticity collapses the stability hierarchy for finite-time convergence, but it does not eliminate the intrinsic difficulties of non-autonomous products in the long-time regime.
5 Numerical Experiments
This section illustrates and complements the abstract convergence theory developed in the previous chapters by means of concrete numerical experiments. The purpose of these experiments is not merely empirical validation, but rather to clarify how the different theoretical mechanisms identified in this thesis manifest themselves in practical computations.
We focus on simple scalar delay differential equations for which both the implicit Euler scheme and the Lie–Trotter splitting method admit explicit implementations. This allows us to examine, in a transparent way, the role of non-sectorial transport effects, time dependence of coefficients, and discrete stability assumptions in the observed numerical behavior.
The experiments are organized according to the four analytical regimes studied in this thesis: autonomous versus non-autonomous problems, and sectorial versus non-sectorial operators. In each case, we compare the behavior of implicit Euler and Lie–Trotter splitting, and relate the observed convergence (or lack thereof) directly to the corresponding Main Theorems A–D.
In particular, the numerical results illustrate that
  • in the non-sectorial setting, the accumulation of local errors depends sensitively on discrete stability properties, in accordance with Theorems A and C;
  • in the sectorial setting, analytic smoothing suppresses error growth and leads to robust finite-time convergence, as predicted by Theorems B and D;
  • practical implementations based on ring-buffer representations of the history variable faithfully realize the abstract operator-splitting schemes studied in the theoretical part.
All experiments are designed to emphasize qualitative mechanisms rather than computational efficiency. Implementation details, including the ring-buffer realization of the delay term, are provided in Appendix B.2.
5.1 Autonomous Non-Sectorial DDE
We consider the autonomous delay differential equation
$\begin{cases}u^{\prime}(t)=au(t)+b\,u(t+\tau),&t\geq 0,\ \tau<0,\\ u(t)=\phi(t),&t\in[\tau,0],\end{cases}$ (5.1)
with given history function $\phi$.
The purpose of the numerical experiments in this section is not to investigate convergence of numerical schemes toward a distinguished continuous solution generated by a semigroup or evolution family. Indeed, as emphasized throughout this thesis, many delay equations of interest — including (5.1) for certain parameter regimes — do not admit a well-posed formulation in the classical $C_{0}$-semigroup framework.
Instead, the experiments are designed to illustrate qualitative and quantitative features of the discrete propagators themselves, in particular the relationship between the implicit Euler resolvent and the Lie–Trotter splitting propagator analyzed in the theoretical sections.
In particular, the example presented below demonstrates that, even when the continuous dynamics exhibit clear asymptotic stability, the associated discrete Lie–Trotter propagator does not necessarily satisfy the strongest geometric summability assumptions appearing in Main Theorem A. Consequently, the most restrictive stability level in the hierarchy is not applicable in this setting, and one must instead rely on weaker, semigroup-like stability mechanisms.
This example therefore serves to illustrate the necessity of the stability hierarchy introduced in Main Theorem A and clarifies the distinct roles played by its different levels in the global comparison of discrete propagators.
5.2 Overview and Implementation
We begin by outlining the general structure of the numerical experiments, followed by detailed implementations for specific cases. The results are then discussed in the context of the theoretical findings from section 3.1 and section 3.2.
First, we present calculation examples for Autonomous Non-sectorial DDE, followed by Non-Autonomous Non-sectorial DDE. Then, we compute sectorial DDEs using a one-dimensional reaction-diffusion equation. We develop calculation examples for autonomous DDEs first, followed by non-autonomous DDEs. The following works were referenced in this section: E. Hairer and G. Wanner Hairer and Wanner (1996), E. B. DaviesDavies (2007), W. Zuo and Y. Song Zuo and Song (2015), and K. Ohira and T. Ohira Ohira and Ohira (2025). Furthermore, we cited K. Ohira Ohira (2025). A more thorough examination of the Python program "ddeint" can be found at
https://github.com/Zulko/ddeint .
Exact-shift Lie–Trotter (ring buffer).
We define the preconditioned step by
$S_{h}:=\bigl[I-h\,D(I-hC)^{-1}\bigr]^{-1}.$
The exact-shift Lie–Trotter step is then given by
$P_{h}:=S_{C}(h)\,S_{h}.$
On a uniform grid with $\tau=-mh$, the operator $S_{C}(h)$ acts as a FIFO ring buffer on the discrete history array and writes the new inflow value $u$ at $\sigma=0$. The following lemma illustrates the structure of the ring buffer used in numerical implementations. For the explicit algorithm, see Appendix B.2.
Lemma 5.1 (Exact shift equals ring buffer on grid).
If $\tau=-mh$ with $m\in\mathbb{N}$, then for any piecewise constant history sampled on $\{\tau,\tau+h,\ldots,-h\}$, the operator $S_{C}(h)$ coincides with the FIFO ring buffer update:
$\bigl[\rho(\tau),\rho(\tau+h),\ldots,\rho(-h)\bigr]\ \mapsto\ \bigl[\rho(\tau+h),\ldots,\rho(-h),\,u\bigr].$
If $\tau/h\notin\mathbb{N}$, then $S_{C}(h)$ is realized by a single-step shift combined with an interpolation, which contributes an additional $\mathcal{O}(h^{\theta})$ term to the local defect on $X_{C}^{\theta}$.
5.2.1 Implicit Euler: Autonomous Non-Sectorial Case
In the autonomous case, for a fixed time step $h>0$, the implicit Euler scheme can be expressed (cf. Hansen et al. Hansen and Stillfjord (2014)) in the abstract form
$(I-h(C+D))(u_{n},\rho_{n})=(u_{n-1},\rho_{n-1}),$
where $(u_{n},\rho_{n})\in\mathcal{D}(C)$ denotes the discrete state at time $t_{n}=nh$.
Expanding this equation yields the coupled system
$(1-ha)\,u_{n}$ $=u_{n-1}+hb\,\rho_{n}(\tau),$ (5.2)
$\rho_{n}-h\partial_{\sigma}\rho_{n}$ $=\rho_{n-1},$
together with the compatibility condition $\rho_{n}(0)=u_{n}$.
Solving the second equation in (5.2) as an inhomogeneous first-order ODE in $\sigma$ with initial value $\rho_{n}(0)=u_{n}$, we obtain
$\rho_{n}(\sigma)=e^{\sigma/h}u_{n}+\int_{\sigma}^{0}\frac{1}{h}e^{(\sigma-s)/h}\rho_{n-1}(s)\,ds,\qquad\sigma\in[\tau,0].$ (5.3)
Evaluating this expression at $\sigma=\tau$ gives
$\rho_{n}(\tau)=e^{\tau/h}u_{n}+\int_{\tau}^{0}\frac{1}{h}e^{(\tau-s)/h}\rho_{n-1}(s)\,ds.$ (5.4)
Substituting (5.4) into the first equation of (5.2), we obtain
$(1-ha)u_{n}=u_{n-1}+hb\left(e^{\tau/h}u_{n}+\int_{\tau}^{0}\frac{1}{h}e^{(\tau-s)/h}\rho_{n-1}(s)\,ds\right),$
which can be rearranged as
$\bigl(1-ha-hbe^{\tau/h}\bigr)u_{n}=u_{n-1}+b\int_{\tau}^{0}e^{(\tau-s)/h}\rho_{n-1}(s)\,ds.$ (5.5)
Hence $u_{n}$ is explicitly determined by $(u_{n-1},\rho_{n-1})$ through
$u_{n}=\frac{1}{1-ha-hbe^{\tau/h}}u_{n-1}+\frac{b}{1-ha-hbe^{\tau/h}}\int_{\tau}^{0}e^{(\tau-s)/h}\rho_{n-1}(s)\,ds.$ (5.6)
Once $u_{n}$ is known, the full history $\rho_{n}$ follows from (5.3).
Remark 5.2.
The above derivation corresponds precisely to solving the resolvent equation
$(I-h(C+D))X_{n}=X_{n-1},\qquad X_{n}=(u_{n},\rho_{n}),$
that is,
$X_{n}=(I-h(C+D))^{-1}X_{n-1}.$
Using the resolvent factorization
$(I-h(C+D))^{-1}=(I-hC)^{-1}\bigl[I-hD(I-hC)^{-1}\bigr]^{-1},$
we recover both the transport formula
$\rho_{n}(\sigma)=e^{\sigma/h}u_{n}+\int_{\sigma}^{0}\frac{1}{h}e^{(\sigma-s)/h}\rho_{n-1}(s)\,ds$
and the update equation
$(1-ha)u_{n}=u_{n-1}+hb\,\rho_{n}(\tau).$
This derivation therefore provides a concrete illustration of the right-preconditioned resolvent framework developed in the theoretical part of this thesis. For further details on the resolvent-based formulation of delay equations, we refer to Hansen and Stillfjord (2014).
5.2.2 Implicit Euler: Non-autonomous Non-Sectorial Case
In the non-autonomous case, we consider the following delay differential equation with time-dependent coefficient $a(t)$:
$(I-h(C(t_{n})+D))(u_{n},\rho_{n})=(u_{n-1},\rho_{n-1}).$ (5.7)
Calculate this equation step by step.
$(1-ha(t_{n}))u_{n}$ $=u_{n-1}+hb\Phi\rho_{n}(\sigma),$ (5.8)
$\rho_{n}-h\frac{d}{d\sigma}\rho_{n}$ $=\rho_{n-1}.$
To solve the lower equation in (5.8) for $\sigma$, we begin with the condition $\rho_{n}(0)=u_{n}$, we obtain:
$\rho_{n}(\sigma)=e^{\sigma/h}u_{n}+\int^{0}_{\sigma}\frac{1}{h}e^{(\sigma-s)/h}\rho_{n-1}(s)ds,$ (5.9)
Operating $\Phi$ both side of (5.9), we get
$\rho_{n}(\tau)=e^{\tau/h}u_{n}+\int^{0}_{\tau}\frac{1}{h}e^{(\tau-s)/h}\rho_{n-1}(s)ds.$ (5.10)
By combining equations (5.8) and (5.9), we obtain
$(1-ha(t_{n}))u_{n}=u_{n-1}+hb\left(e^{\tau/h}u_{n}+\int^{0}_{\tau}\frac{1}{h}e^{(\tau-s)/h}\rho_{n-1}(s)ds\right).$
From this equation, $u_{n}$ can be derived from $u_{n-1}$ as follows:
$\left(1-ha(t_{n})-hbe^{\tau/h}\right)u_{n}=u_{n-1}+b\int^{0}_{\tau}e^{(\tau-s)/h}\rho_{n-1}(s)ds,$
$u_{n}=\frac{1}{\left(1-ha(t_{n})-hbe^{\tau/h}\right)}u_{n-1}+\frac{b}{\left(1-ha(t_{n})-hbe^{\tau/h}\right)}\int^{0}_{\tau}e^{(\tau-s)/h}\rho_{n-1}(s)ds$
This formula allows us to compute $u_{n}$ from $u_{n-1}$ and $\rho_{n-1}$. After calculating $u_{n}$, we can substitute it back into equation (5.9) to find $\rho_{n}(\sigma)$. See Hansen and Stillfjord (2014) for more details.
5.2.3 The Sequential Operator Splitting: Autonomous and Non-Autonomous DDEs (Non-Sectorial Case)
We consider the operator splitting method for solving delay differential equations (DDEs) as discussed in theory. The DDE under consideration is simplest case given by the following equations, for more complex cases, please refer to the relevant literature Hansen and Stillfjord (2014).J. Geiser in multiple works Geiser (2008, 2009, 2011):
The Autonomous Case
We consider the autonomous system:
$\binom{v_{n}}{\phi_{n}}$ $=(I-hC)^{-1}(I-hD_{bd})^{-1}\binom{v_{n-1}}{\phi_{n-1}}$, (5.11)
$\binom{v_{0}}{\phi_{0}}$ $=\binom{\text{history}(0)}{\text{history}(\sigma)}$.
This scheme is solved sequentially in two steps:
$\binom{w_{n-1}}{\psi_{n-1}}$ $=(I-hD_{\mathrm{bd}})^{-1}\begin{pmatrix}v_{n-1}\\ \phi_{n-1}\end{pmatrix}$, (5.12)
$\binom{v_{n}}{\phi_{n}}$ $=(I-hC)^{-1}\begin{pmatrix}w_{n-1}\\ \psi_{n-1}\end{pmatrix}$.
The first equation in (5.12) can be solved similarly to the Implicit Euler scheme:
$\psi_{n-1}(\sigma)=e^{\sigma/h}v_{n-1}+\int_{0}^{\sigma}\frac{1}{h}e^{(\sigma-s)/h}\phi_{n-1}(s)\,ds$ (5.13)
where we use $\psi_{n-1}(0)=v_{n-1}$. By applying the operator $\Phi$ to both sides of this equation, we derive:
$\psi_{n-1}(\tau)=e^{\tau/h}v_{n-1}+\int_{0}^{\tau}\frac{1}{h}e^{(\tau-s)/h}\phi_{n-1}(s)\,ds$ (5.14)
and
$w_{n-1}=hb\psi_{n-1}(\tau)+v_{n-1}.$ (5.15)
Eliminating $\psi_{n-1}(\tau)$ from (5.14) and (5.15) ,
$v_{n-1}=w_{n-1}-hb\left(e^{\tau/h}w_{n-1}+\int^{0}_{\tau}\frac{1}{h}e^{(\tau-s)/h}\phi_{n-1}(s)ds\right)$ (5.16)
Solving this equation for $w_{n-1}$, we have
$w_{n-1}=\frac{1}{1-hbe^{\tau/h}}\left(v_{n-1}+b\int^{0}_{\tau}e^{(\tau-s)/h}\phi_{n-1}(s)\right).$ (5.17)
Finally, we can transition from $(v_{n-1};\phi_{n-1}(\sigma))$ to $(v_{n};\phi_{n}(\sigma))$ as follows:
$v_{n}$ $=\frac{1}{1-ha}w_{n-1}=\frac{1}{1-ha}\cdot\frac{1}{1-hbe^{\tau/h}}\left(v_{n-1}+b\int^{0}_{\tau}e^{(\tau-s)/h}\phi_{n-1}(s)ds\right)$ (5.18)
$\phi_{n}(\sigma)$ $=\psi_{n-1}(\sigma)$
$=e^{\sigma/h}w_{n-1}+\int_{0}^{\sigma}\frac{1}{h}e^{(\sigma-s)/h}\phi_{n-1}(s)\,ds$
$=e^{\sigma/h}\frac{1}{1-hbe^{\tau/h}}\left(v_{n-1}+b\int^{0}_{\tau}e^{(\sigma-s)/h}\phi_{n-1}(s)ds\right)+\int_{0}^{\sigma}\frac{1}{h}e^{(\sigma-s)/h}\phi_{n-1}(s)\,ds.$
The Non-Autonomous Case
We consider the non-autonomous system:
$\binom{v_{n}}{\phi_{n}}$ $=(I-hA(nh))^{-1}(I-hB)^{-1}\binom{v_{n-1}}{\phi_{n-1}}$, (5.19)
$\binom{v_{0}}{\phi_{0}}$ $=\binom{\text{history}(0)}{\text{history}(\sigma)}$.
This case is also solved sequentially in two steps:
$\binom{w_{n-1}}{\psi_{n-1}}$ $=(I-hD)^{-1}\begin{pmatrix}v_{n-1}\\ \phi_{n-1}\end{pmatrix}$, (5.20)
$\binom{v_{n}}{\phi_{n}}$ $=(I-hC(nh))^{-1}\begin{pmatrix}w_{n-1}\\ \psi_{n-1}\end{pmatrix}$.
As in the autonomous case, the first equation in (5.20) can be solved similarly to the Implicit Euler scheme:
$\psi_{n-1}(\sigma)=e^{\sigma/h}v_{n-1}+\int_{0}^{\sigma}\frac{1}{h}e^{(\sigma-s)/h}\phi_{n-1}(s)\,ds$ (5.21)
where we again use $\psi_{n-1}(0)=v_{n-1}$. Applying the operator $\Phi$ to both sides gives:
$\psi_{n-1}(\tau)=e^{\tau/h}v_{n-1}+\int_{0}^{\tau}\frac{1}{h}e^{(\tau-s)/h}\phi_{n-1}(s)\,ds$ (5.22)
and
$w_{n-1}=\frac{1}{1-hbe^{\tau/h}}\left(v_{n-1}+b\int^{0}_{\tau}e^{(\tau-s)/h}\phi_{n-1}(s)ds\right).$ (5.23)
By utilizing equations (5.21), (5.22), and (5.23), we can transition from $(v_{n-1},\phi_{n-1}(\sigma))$ to $(v_{n},\phi_{n}(\sigma))$ as follows:
$v_{n}$ $=\frac{1}{1-ha(nh)}w_{n-1}=\frac{1}{1-hanh}\cdot\frac{1}{1-hbe^{\tau/h}}\left(v_{n-1}+b\int^{0}_{\tau}e^{(\tau-s)/h}\phi_{n-1}(s)ds\right)$ (5.24)
$\phi_{n}(\sigma)$ $=\psi_{n-1}(\sigma)$
$=e^{\sigma/h}w_{n-1}+\int_{0}^{\sigma}\frac{1}{h}e^{(\sigma-s)/h}\phi_{n-1}(s)\,ds$
$=e^{\sigma/h}\frac{1}{1-hbe^{\tau/h}}\left(v_{n-1}+b\int^{0}_{\tau}e^{(\sigma-s)/h}\phi_{n-1}(s)ds\right)+\int_{0}^{\sigma}\frac{1}{h}e^{(\sigma-s)/h}\phi_{n-1}(s)\,ds.$
For details on these calculations, see Hansen and Stillfjord (2014).
Remark 5.3 (Exact shift and ring-buffer implementation).
In the non-sectorial setting, the transport flow generated by $C$ is explicitly given by the exact left-shift semigroup $S_{C}(h)$. Therefore, the splitting step
$X_{n}=P_{h}(t_{n-1})X_{n-1},\qquad P_{h}(t)=S_{C}(h)\,S_{h}^{D}(t),$
admits a particularly simple algorithmic realization.
On a grid with $\tau=-mh$, the action of $S_{C}(h)$ on the history component corresponds exactly to a FIFO update: the oldest value is removed and the new value $\widehat{u}$ is appended. Hence the history variable can be implemented as a ring buffer of length $m$.
Moreover, in practical implementations one never needs to explicitly form the preconditioned history $(I-hC)^{-1}\rho$. The only quantity entering the scalar solve in the delay step is the trace at $\tau$, which after the shift corresponds precisely to a known buffer entry. This explains why practical codes for delay equations often appear extremely simple while still faithfully realizing the abstract operator-splitting scheme.
A concrete ring-buffer implementation is provided in Appendix B.2.
5.2.4 Autonomous Non-Sectorial Case: Exponential Growth
In this section, we present numerical results for the autonomous non-sectorial delay differential equation (DDE) using the Implicit Euler method,
$u^{\prime}(t)$ $=au(t)+bu(t+\tau),\quad t\geq 0,$
$u(t)$ $=\text{polyFunc}(t),\quad t\in[\tau,0],$
with parameters a=0.15a=-0.15, b=6.0b=-6.0, and $\tau=-8.0$. The history function is chosen as a 10th-degree polynomial defined in (5.28).
For this parameter choice, the numerical solution obtained by implicit Euler exhibits clear exponential growth, as shown in Figure 3. At first glance, this may appear to indicate instability. However, this behavior is fully consistent with the underlying functional-analytic structure of the problem.
Indeed, although $|b|>1$, the operator $C+D$ associated with this delay equation still generates a $C_{0}$-semigroup on the natural phase space. Consequently, solutions cannot exhibit finite-time blow-up; instead, their growth is necessarily bounded by an exponential envelope of the form
$\|u(t)\|\leq Me^{\omega t}.$
This qualitative prediction from semigroup theory is precisely what is observed in the numerical experiment.
To verify the exponential nature of the growth, we examine the logarithmic behavior of the solution norm:
$\log\|u(t)\|\approx\log M+\omega t.$
The left panel of Figure 4 shows that the numerical data are well approximated by a linear function in time, yielding fitted parameters $\log M=-2.39$ and $\omega=0.032$. The right panel confirms that the relative ratio $\|u(t)\|/(Me^{\omega t})$ remains close to one. Furthermore, Figure 3 shows that the normalized deviation stabilizes around a constant level, supporting the asymptotic behavior
$u(t)\sim Me^{\omega t}.$
These observations demonstrate that, even in regimes where the solution grows, the numerical method reproduces the correct qualitative behavior predicted by semigroup theory: growth is exponential but controlled, and no finite-time blow-up occurs.
The left plot in Figure 4 confirms that $\log\|u(t)\|\leq\log M+\omega t$ with $\log M=-2.39$, $\omega=0.032$. The right plot shows that the relative error $\|u(t)\|/(Me^{\omega t})\approx 1$, indicating excellent agreement with the theoretical growth rate. Figure 3 reveals that the true error stabilizes around 1, confirming that $u(t)\sim Me^{\omega t}$.
Figure 2: Solution of the autonomous DDE: a=0.15a=-0.15, b=6.0b=-6.0, $\tau=-8.0$, history = polyFunc(t).
Figure 3: True error: $\|u(t)\|-Me^{\omega t}/Me^{\omega t}$, a=0.15a=-0.15, b=6.0b=-6.0, $\tau=-8.0$, history = polyFunc(t).
Figure 4: (Left) $\log\|u(t)\|$ vs. $t$, (Right) Relative Error $\|u(t)\|/Me^{\omega t}$, a=0.15a=-0.15, b=6.0b=-6.0, $\tau=-8.0$, history = 10th Polynomial.
Figure 5: Solution of autonomous DDE: a=0.15a=-0.15, b=6.0b=-6.0, $\tau=-0.257$, history = polyFunc(t).
Figure 6: ’ddeint’ vs. Implicit Euler: a=0.15a=-0.15, b=6.0b=-6.0, $\tau=-0.257$, history = polyFunc(t). Blue = ddeint, Red = Implicit Euler.
Figure 7: Implicit Euler vs. Operator Splitting: a=0.15a=-0.15, b=6.0b=-6.0, $\tau=-0.257$, history = polyFunc(t). Blue = Operator Splitting, Red = Implicit Euler.
Figure 8: Error between Implicit Euler and Operator Splitting: a=0.15a=-0.15, b=6.0b=-6.0, $\tau=-0.257$, history = polyFunc(t).
5.2.5 Structure of the Lie–Trotter Scheme and Failure of Summability
In this subsection we examine, for the concrete scalar delay model, which stability assumptions in Main Theorem A are actually satisfied by the Lie–Trotter scheme.
Lie–Trotter as composition of two subproblems.
We decompose the equation
$u^{\prime}(t)=au(t)+bu(t+\tau)$
into the two subproblems
  • (A)
    $u^{\prime}(t)=au(t)$ (instantaneous linear dynamics),
  • (B)
    $u^{\prime}(t)=bu(t+\tau)$ (pure delay dynamics).
On a single time step $t_{n},t_{n+1}]$, the delayed value $u(t+\tau)$ lies in the past interval $t_{n-m},t_{n-m+1}]$ (where $m=-\tau/h$), and is therefore numerically known from the history.
Approximating (B) by a forward Euler step gives
$\tilde{u}_{n+1}=\tilde{u}_{n}+hbu_{n-m},\qquad\tilde{u}_{n}=u_{n},$
that is,
$\tilde{u}_{n+1}=u_{n}+hbu_{n-m}.$
Applying a backward Euler step for (A) yields
$u_{n+1}=\tilde{u}_{n+1}+hau_{n+1},\qquad\Longleftrightarrow\qquad(1-ha)u_{n+1}=\tilde{u}_{n+1}.$
Hence the one-step Lie–Trotter update takes the explicit form
$u_{n+1}=\frac{1}{1-ha}u_{n}+\frac{hb}{1-ha}u_{n-m}=:\alpha u_{n}+\beta u_{n-m}.$ (5.25)
This shows that the resolvent-based Lie–Trotter scheme corresponds exactly to “backward Euler for the present term and forward Euler for the delay term”.
Companion matrix and memory structure.
The recurrence (5.25) involves both $u_{n}$ and the $m$-step delayed value $u_{n-m}$. Therefore the scheme naturally lives on the augmented state space
$x_{n}=(u_{n},u_{n-1},\dots,u_{n-m})^{\top}\in\mathbb{R}^{m+1},$
which evolves according to the companion matrix recursion
$x_{n+1}=M_{h}x_{n},\qquad M_{h}=\begin{pmatrix}\alpha&0&\cdots&0&\beta\\ 1&0&\cdots&0&0\\ 0&1&\ddots&0&0\\ \vdots&&\ddots&\ddots&\vdots\\ 0&0&\cdots&1&0\end{pmatrix}$.
This algebraic structure reflects the intrinsic memory effect of delay equations: advancing the solution requires retaining the last $m$ values. The ring buffer implementation described in Appendix B.2 is precisely the computational realization of this companion structure. In particular, the trace value $\rho(\tau)$ used in the operator formulation corresponds exactly to the component $u_{n-m}$ in the finite-dimensional representation.
Connection with the abstract propagator $P_{h}$.
In this finite-dimensional formulation,
$x_{n+1}=M_{h}x_{n}\quad\Longleftrightarrow\quad x_{n}=M_{h}^{n}x_{0}.$
Thus the discrete propagator $P_{h}$ introduced abstractly in the theoretical part coincides here with the matrix $M_{h}$:
$P_{h}=M_{h}.$
The scalar sequence $u_{n}$ is simply the first component of $x_{n}=P_{h}^{n}x_{0}$.
The same observation applies to implicit Euler: the implicit Euler recurrence
$u_{n+1}=\alpha u_{n}+\beta u_{n+1-m}$
can also be written in companion form, leading to another discrete propagator matrix associated with the implicit Euler scheme.
Spectral analysis and failure of summability.
The eigenvalues of $M_{h}$ are given by the roots of the characteristic polynomial obtained by substituting $u_{n}=\lambda^{n}$ into (5.25):
$\lambda^{m+1}=\alpha\lambda^{m}+\beta\quad\Longleftrightarrow\quad p_{LT}(\lambda):=\lambda^{m+1}-\alpha\lambda^{m}-\beta=0.$
For the parameter set
$a=-0.15,\qquad b=-6.0,\qquad\tau=-0.257,$
and step size $h=0.001$, numerical computation yields
$\rho(M_{h})=0.9999108137\ldots<1.$
Although the spectral radius is strictly less than one (consistent with the observed stability of the numerical solution), it approaches one as $h\to 0$. Consequently,
$\sum_{n=0}^{\infty}\|M_{h}^{n}\|\sim\frac{1}{1-\rho(M_{h})}\approx 1.1\times 10^{4},$
which is not uniformly bounded in $h$. Therefore, the family $\{P_{h}\}$ fails to satisfy the strongest stability assumption (summability of powers) in Main Theorem A.
Open question: Ritt and BV stability.
Whether the intermediate stability conditions in Main Theorem A, such as Ritt stability or bounded variation of powers, hold for this scheme remains an interesting open question and is not addressed in the present work.
5.3 Non-autonomous non-sectorial case
5.3.1 Benchmark solution by the Ohira–Ohira method
Closed-form solutions of non-autonomous delay differential equations are rarely available. For the specific class of equations considered below, Ohira–Ohira proposed a semi-analytical construction that yields solutions of very high accuracy (which we use as “reference” or “benchmark” solutions) Ohira and Ohira (2025); see also Ohira (2025) for further developments.
We consider the scalar non-autonomous DDE
$u^{\prime}(t)$ $=a\,t\,u(t)+b\,u(t+\tau),\qquad t\geq 0,$ (5.26)
$u(t)$ $=\mathrm{history}(t),\qquad\tau\leq t\leq 0,$ (5.27)
where $a,b\in\mathbb{R}$ and $\tau<0$.
Outline of the construction. The Ohira–Ohira method constructs a solution on the whole line via a Fourier representation, and then extracts a suitable history segment on $\tau,0]$ so that the resulting function satisfies (5.26) on $[0,T]$. More precisely, one obtains an explicit Fourier-space expression of the form
$\hat{u}(\omega)=\mathcal{C}\,\exp\!\left(\frac{\omega^{2}}{2a}+\frac{b}{a\tau}e^{i\omega\tau}\right),$
where $\mathcal{C}$ is a normalization constant (set to $1$ in our experiments). Taking the inverse transform yields the real-valued representation
$u(t)=\frac{1}{2\pi}\int_{\mathbb{R}}\exp\!\left(\frac{\omega^{2}}{2a}+\frac{b\cos(\omega\tau)}{a\tau}\right)\cos\!\left(\frac{b}{a\tau}\sin(\omega\tau)+\omega t\right)\,d\omega,$
which can be evaluated numerically with high precision. Ohira–Ohira proved that the resulting function indeed satisfies (5.26)–(5.27); see Ohira and Ohira (2025).
Parameter set. Throughout this section we fix
$a=-0.15,\qquad b=-6.0,\qquad\tau=-8.0.$
The benchmark solution obtained by the Ohira–Ohira procedure is displayed in Figure 10. The corresponding history segment on $\tau,0]$ is approximated by a polynomial fit of degree $10$, denoted by $\mathrm{polyFunc}(t)$:
[…]
Figure 10 plots this fitted history. This fitting step allows us to initialize the implicit Euler and splitting schemes consistently from a smooth history function.
5.3.2 Comparison of numerical approximations
Figure 12 shows the implicit Euler approximation with time step $h=0.1$. Figure 12 compares the implicit Euler solution with the benchmark solution, showing that implicit Euler tracks the reference dynamics well at this resolution.
Figure 14 compares implicit Euler and Lie–Trotter splitting under the same step size, and Figure 14 shows the difference between the two numerical trajectories. We observe that the discrepancy initially grows and then decays, producing a bell-shaped profile. This qualitative behavior is consistent with the finite-time error mechanism identified in the non-autonomous non-sectorial theory: local defects may accumulate transiently, while the overall error remains controlled on finite time intervals.
Figure 9: K. Ohira and T. Ohira’s Exact Semi-Numerical Solution for a=0.15a=-0.15, b=6.0b=-6.0, $\tau=-8.0$.
Figure 10: History Segment for the DDE, fitted by a 10th-degree polynomial.
Figure 11: Implicit Euler Approximate Solution, $h=0.1$.
Figure 12: Comparison: Implicit Euler (red) vs. Exact Solution (blue dots).
Figure 13: Comparison: Implicit Euler (red) vs. Operator Splitting (blue).
Figure 14: Error between Implicit Euler and Operator Splitting Solutions.
5.3.3 Discussion: Verification of the assumptions of Main Theorem C
We verify that the numerical example
$u^{\prime}(t)=atu(t)+bu(t+\tau),\qquad a=-0.15,\ b=-6.0,\ \tau=-8.0,$
fits naturally into the abstract framework of Theorem 3.21.
1. Abstract formulation.
We consider the state space
$\mathcal{E}_{0}=\mathbb{R}\oplus L^{p}([\tau,0]),\qquad X(t)=(u(t),\rho_{t}),\quad\rho_{t}(\sigma)=u(t+\sigma),$
and define
$C(u,\rho)=(0,\partial_{\sigma}\rho),\qquad D(t)(u,\rho)=(at\,u+b\,\rho(\tau),0),$
with $\mathcal{D}(C)=\{(u,\rho):\rho\in W^{1,p}([\tau,0]),\ \rho(0)=u\}$. This yields the abstract equation $X^{\prime}(t)=(C+D(t))X(t)$.
2. Verification of (C1) and (C2).
By the trace inequality on $W^{1,p}([\tau,0])$, there exists $c>0$ such that
$|\rho(\tau)|\leq c\bigl(\|\rho\|_{L^{p}}+\|\partial_{\sigma}\rho\|_{L^{p}}\bigr).$
This implies that $D(t)$ is $C$-bounded with relative bound 0 uniformly in $t\in[0,T]$. As a consequence, the frozen resolvent $R_{h}(t)=(I-h(C+D(t)))^{-1}$ exists for sufficiently small $h$ and is uniformly bounded on $\mathcal{E}_{0}$. Moreover, the estimate
$\|hD(t)S_{h}^{C}\|\leq q<1$
holds for sufficiently small $h$, so that the operational smallness condition (C1) is satisfied and $P_{h}(t)$ is well defined.
3. Local defect (C3).
The local defect estimate
$\|R_{h}(t)-P_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{\theta}h^{\theta}$
follows from the abstract defect analysis established in Chapter 4 and applies uniformly with respect to $t\in[0,T]$ since the estimates depend only on the uniform bounds of the coefficients.
4. Discrete stability (C4).
In the present example, numerical experiments indicate that the discrete propagators remain uniformly bounded on finite time intervals. In particular, the numerical trajectories do not exhibit finite-time blow-up, even in regimes where the solution grows exponentially. This is fully consistent with the theoretical requirement of product stability in Theorem 3.21.
Conclusion. All structural assumptions of Main Theorem 3.21 are compatible with the present model. The observed convergence of Lie–Trotter splitting toward implicit Euler in Figures 14 and 14 therefore provides a concrete numerical validation of the abstract theory.
5.4 Autonomous Sectorial Case
We consider the autonomous delay equation with a sectorial diffusion part:
$\frac{d}{dt}u(t)$ $=Au(t)+\lambda u(t)+b\,u(t+\tau),\qquad t>0,$ (5.29)
$u(t)$ $=\mathrm{history}(t),\qquad t\in[\tau,0],$
where $A$ is sectorial on the base space $X$ and $\tau<0$. (Equivalently one may write $u(t-\tau_{d})$ with $\tau_{d}=-\tau>0$.)
As a concrete model we take $X=L^{2}(0,L)$ and $A=\kappa_{\mathrm{diff}}\Delta_{D}$ with homogeneous Dirichlet boundary conditions on $(0,L)$:
$u_{t}(t,x)$ $=\kappa_{\mathrm{diff}}u_{xx}(t,x)+\lambda u(t,x)+b\,u(t+\tau,x),\qquad t\in[0,T],\ x\in(0,L),$ (5.30)
$u(t,0)$ $=u(t,L)=0,\qquad t\geq 0,$
$u(t,x)$ $=\mathrm{history}(t,x),\qquad t\in[\tau,0],\ x\in[0,L].$
5.4.1 Connection between theory and numerical implementation
The reaction–diffusion delay equation (5.30) fits the sectorial model in Section 4. On the product space
$\mathcal{E}_{0}:=X\oplus L^{2}([\tau,0];X),\qquad X=L^{2}(0,L),$
we encode the history by $\rho_{t}(\sigma):=u(t+\sigma)$ and write $X(t):=(u(t),\rho_{t})$. The system becomes an abstract Cauchy problem
$\frac{d}{dt}X(t)=\mathcal{L}X(t),\qquad\mathcal{L}:=A_{0}+B+C,$
where
$A_{0}=\begin{pmatrix}\kappa_{\mathrm{diff}}\Delta_{D}&0\\ 0&0\end{pmatrix},\qquad B=\begin{pmatrix}\lambda I&b\,\Phi_{\tau}\\ 0&0\end{pmatrix},\qquad C=\begin{pmatrix}0&0\\ 0&\partial_{\sigma}\end{pmatrix},\qquad\rho(0)=u.$
Here $A_{0}$ is sectorial (analytic smoothing), $C$ generates the contractive left–shift semigroup on the history component, and $B$ couples the present state with the delayed trace $\Phi_{\tau}\rho=\rho(\tau)$.
Remark 5.4 (Why the delay trace is handled via resolvent preconditioning).
On the pure history space $L^{2}([\tau,0];X)$ the point evaluation $\Phi_{\tau}$ is not bounded. In the three–factor formulation, however, the delay block is applied only through the preconditioned combination $B(I-hC)^{-1}$. The shift resolvent $(I-hC)^{-1}$ maps into the graph space of $C$ (hence into a temporally more regular class), which makes $B(I-hC)^{-1}$ bounded and small as $h\to 0$. This is the key mechanism behind the well-posedness of the splitting step and the local defect estimates.
5.4.2 Implicit Euler scheme.
The implicit Euler discretization of $X^{\prime}(t)=\mathcal{L}X(t)$ reads
$(I-h\mathcal{L})X_{n}=X_{n-1},\qquad X_{n}=(u_{n},\rho_{n}),$
i.e. one step is given by the resolvent $R_{h}=(I-h\mathcal{L})^{-1}$. Componentwise,
$\begin{cases}(I-h(\kappa_{\mathrm{diff}}\Delta_{D}+\lambda I))u_{n}=u_{n-1}+h\,b\,\rho_{n}(\tau),\\[3.99994pt] (I-h\partial_{\sigma})\rho_{n}=\rho_{n-1},\qquad\rho_{n}(0)=u_{n}.\end{cases}$
Thus $\rho_{n}$ is obtained by one backward Euler step for the shift equation, with the inflow condition $\rho_{n}(0)=u_{n}$.
5.4.3 Three–factor Lie–Trotter splitting scheme.
The splitting propagator analyzed in the sectorial theory is the discrete operator composition
$P_{h}\;:=\;(I-hC)^{-1}\,(I-hA_{0})^{-1}\,\bigl[I-h\,B\,(I-hC)^{-1}\bigr]^{-1}.$
In the numerical implementation, the factor $(I-hC)^{-1}$ is realized by an exact shift update of the discrete history (a ring buffer), while $(I-hA_{0})^{-1}$ is the implicit diffusion solve and the remaining factor is an algebraic reaction/delay update through the buffered value corresponding to $\rho(\tau)$.
Concretely, one time step consists of:
  1. 1.
    Diffusion step: solve $(I-hA_{0})^{-1}$ for the Laplacian block.
  2. 2.
    Reaction/delay step: apply $\bigl[I-h\,B\,(I-hC)^{-1}\bigr]^{-1}$ using the delayed trace $\rho(\tau)$ (a buffer read).
  3. 3.
    Transport step: update the history by the left shift (buffer push/pop), representing the action of $(I-hC)^{-1}$ (or equivalently $S_{C}(h)$ on the grid).
This reproduces precisely the operator $P_{h}$ studied in Section 4.
Finite-time convergence mechanism (autonomous sectorial case).
Let $U(t)=e^{t\mathcal{L}}$ be the analytic semigroup generated by $\mathcal{L}$. Assume $0<\theta<1$, then the three–factor theory yields, on the interpolation space $X_{\theta}=(\mathcal{E}_{0},\,D(A_{0})\cap\mathcal{D}(C))_{\theta,1}$,
$\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{T,\theta}\,h^{1-\theta},\qquad t_{n}=nh\leq T,$
where $R_{h}=(I-h\mathcal{L})^{-1}$. In addition, the classical backward Euler convergence for sectorial generators gives
$\|U(t_{n})-R_{h}^{n}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{T,\theta}\,h,\qquad t_{n}\leq T.$
By the triangle inequality,
$\|U(t_{n})-P_{h}^{n}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{T,\theta}\,(h+h^{1-\theta}),\qquad t_{n}\leq T.$
Thus, in the autonomous sectorial setting, Lie–Trotter splitting converges to the exact semigroup on every finite time interval without any additional discrete stability assumptions beyond finite-time boundedness.
Remark 5.5 (Reduction to the non-sectorial model).
If $\kappa_{\mathrm{diff}}=0$, the diffusion block vanishes and the model reduces to the non-sectorial delay equation treated earlier. When $\kappa_{\mathrm{diff}}>0$, sectorial analyticity yields additional fractional smoothing, which improves the convergence mechanism on interpolation spaces.
5.4.4 Parameter choices for numerical experiments
We consider the autonomous diffusion–delay equation (LABEL:eq:linDDEdiff) on $(0,1)$ with homogeneous Dirichlet boundary conditions. The spatial Laplacian is discretized by standard second-order finite differences on $N_{x}$ grid points, so that the diffusion step in both schemes amounts to solving a fixed linear system at each time step.
Throughout this subsection we fix
$\kappa=0.02,\quad\lambda=-0.8,\quad b=-0.8,\quad\tau=-0.6,\quad T=8.0,\quad h=0.002,\quad N_{x}=300.$
The history segment on $\tau,0]$ is prescribed by
$\mathrm{history}(t,x)=0.3+0.2\sin(2\pi x)\cos(2\pi t),\qquad t\in[\tau,0],\ x\in[0,1].$
These choices are made for the following reasons:
  • The diffusion coefficient $\kappa>0$ places the generator in the sectorial (analytic) class, so the smoothing mechanism predicted by the theory is present.
  • The coefficients $\lambda$ and $b$ are chosen in a moderately damped regime where solutions remain bounded on $[0,T]$, allowing a clean comparison of implicit Euler and Lie–Trotter splitting without transient blow-up.
  • The delay length $|\tau|$ is large compared to the time step $h$, so the computation genuinely involves a nontrivial history buffer; in particular $m:=|\tau|/h$ is large enough that the delay term is not a negligible perturbation.
A more detailed discussion, including spectral considerations for the fully discretized update operator.
Figures 1516 compare the implicit Euler and three–factor Lie–Trotter solutions. Table 1 reports representative values of the centre trace $u(t,0.5)$, while Table 2 records CPU times for the current implementation.
Numerical observation. Figures 1516 show that the solutions produced by implicit Euler and Lie–Trotter splitting are visually indistinguishable on the space–time scale. This is confirmed quantitatively by the centre trace values in Table 1, where the two methods agree up to at least six significant digits throughout the simulation interval. Moreover, the timing results in Table 2 indicate that the computational cost of the splitting scheme is comparable to that of the fully implicit method in this setting.
Numerical observation. Figures 1516 show that the solutions produced by implicit Euler and Lie–Trotter splitting are visually indistinguishable on the space–time scale. This is confirmed quantitatively by the centre trace values in Table 1, where the two methods agree up to at least six significant digits throughout the simulation interval. Moreover, the timing results in Table 2 indicate that the computational cost of the splitting scheme is comparable to that of the fully implicit method in this setting.
These observations are fully consistent with the theoretical prediction of Main Theorem B, which asserts finite–time convergence of Lie–Trotter splitting toward implicit Euler in the autonomous sectorial case without additional discrete stability assumptions.
(a) Heat Map of Implicit Euler Solution of the Reaction–Diffusion Delay Equation. Heat(x,t)(x,t) represents $u(t,x)$.
(b) Heat Map of Lie–Trotter Split Solution of the Reaction–Diffusion Delay Equation. Heat(x,t)(x,t) represents $u(t,x)$.
Figure 15: Space–Time Heat Maps for Implicit Euler and Lie–Trotter Split Solutions of the Reaction–Diffusion Delay Equation.
(a) Centre value of Implicit Euler and Lie–Trotter Split Solutions of the Reaction-Diffusion Delay Equation
(b) Timing of Implicit Euler and Lie–Trotter Split Solutions of the Reaction–Diffusion Delay Equation
Figure 16: Timing and Centre Value Comparisons for Implicit Euler and Lie–Trotter Split Solutions of the Reaction–Diffusion Delay Equation.
Table 1: Center value $u(t,0.5)$ for the autonomous diffusion–delay equation $u_{t}=\kappa u_{xx}+\lambda u+bu(t+\tau)$ with $\lambda=-0.8$.
Time $t$ Implicit Euler (IE) Lie–Trotter Splitting
0 $2.9988\times 10^{-1}$ $2.9988\times 10^{-1}$
1 $1726\times 10^{-1} 1.1$$1726\times 10^{-1}$
2 $1.7249\times 10^{-2}$ $1.7249\times 10^{-2}$
3 $6.1128\times 10^{-3}$ $6.1127\times 10^{-3}$
4 $3.0588\times 10^{-4}$ $3.0587\times 10^{-4}$
5 $1.4879\times 10^{-5}$ $1.4878\times 10^{-5}$
6 $4.6214\times 10^{-6}$ $4.6212\times 10^{-6}$
7 $9.8773\times 10^{-7}$ $9.8767\times 10^{-7}$
8 $2.1399\times 10^{-7}$ $2.1382\times 10^{-7}$
Table 2: CPU time for implicit Euler and Lie–Trotter splitting (autonomous diffusion–delay equation, $N_{x}=300$, $h=0.002$, $T=8$).
Method CPU Time (seconds)
Implicit Euler $0.29$
Lie–Trotter Splitting $0.34$
5.4.5 Numerical results (autonomous sectorial case)
The left and right panels of Figure 15 display the space–time heat maps obtained by the implicit Euler scheme and the Lie–Trotter splitting method, respectively. Both schemes reproduce the expected dissipative behaviour: the initial spatial profile decays smoothly under the combined effect of diffusion and the negative reaction/delay terms, and no spurious oscillations or numerical instabilities are observed over the entire time interval $[0,T]$.
A quantitative comparison is provided by the temporal trace at the midpoint $x=0.5$, shown in the left panel of Figure 16. The two curves are virtually indistinguishable, with discrepancies below machine precision throughout the simulation. Representative numerical values are listed in Table 1, confirming that implicit Euler and Lie–Trotter splitting produce results that coincide up to at least six significant digits in this autonomous setting.
In contrast to the non-autonomous case, the computational effort of the two schemes is comparable here. As reported in Table 2 and illustrated in the right panel of Figure 16, the implicit Euler method is slightly faster because the matrix $I-h\kappa\Delta_{h}-h\lambda I$ is constant in time, so that a single LU factorization can be reused throughout the computation. The Lie–Trotter splitting, by contrast, requires two diffusion solves per time step, resulting in a similar or slightly larger total runtime.
In summary, for the autonomous sectorial problem both schemes deliver solutions of essentially identical accuracy. However, unlike the non-autonomous case—where the time dependence of the implicit Euler operator significantly increases the computational cost and splitting becomes markedly more efficient—no clear computational advantage of operator splitting is visible in this autonomous regime.
5.4.6 Interpretation of the numerical results
Figures 1516 illustrate the numerical behaviour of the implicit Euler and Lie–Trotter splitting schemes for the linear reaction–diffusion delay equation (LABEL:eq:linDDEdiff). For each method we present two complementary diagnostics: (i) a space–time heat map of the solution $u(t,x)$, and (ii) a temporal trace of the central value $u(t,0.5)$.
Space–time heat maps.
The heat maps provide a global view of the solution over the computational domain. The vertical axis represents time $t\in[0,T]$, while the horizontal axis corresponds to the spatial variable $x\in(0,1)$. Brighter regions indicate larger solution values, whereas darker regions reflect diffusive decay and the stabilizing influence of the negative feedback coefficient $b<0$. The smooth monotone decay visible in both heat maps is consistent with the dissipative nature of the underlying linear operator and confirms that neither scheme introduces artificial oscillations or numerical artefacts. The close visual agreement between the two plots indicates that both methods resolve essentially the same evolution.
Temporal behaviour at the centre.
The time trace of $u(t,0.5)$ offers a more quantitative perspective. Starting from the prescribed history at $t=0$, the solution decays exponentially toward zero, in accordance with the stability of the linear operator $\kappa\Delta_{D}+\lambda I+b\,\Phi_{\tau}.$ The absence of overshoots or oscillatory behaviour demonstrates that both implicit Euler and Lie–Trotter splitting preserve the qualitative features of the continuous dynamics. The near coincidence of the two curves further shows that the splitting scheme is fully consistent with implicit Euler in this autonomous linear setting.
Quantitative agreement.
For the parameter values considered, the centre values at the final time satisfy
$|u_{\mathrm{IE}}(T,0.5)-u_{\mathrm{split}}(T,0.5)|<10^{-8},$
which lies well below the discretization error induced by the spatial and temporal grids. Moreover, the decay rate extracted from the numerical traces agrees with the rate predicted by the dominant eigenvalue of the spatial discretization of $\kappa\Delta_{D}+\lambda I$, modified by the delay term. This confirms that both schemes accurately reproduce the underlying semigroup dynamics.
Connection with the theoretical error analysis.
This experiment illustrates the most favourable theoretical regime of the present framework. In the sectorial setting, analytic smoothing replaces the need for additional discrete stability assumptions such as Ritt or Kreiss conditions. Concretely:
  • The local defect satisfies $\|R_{h}-P_{h}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}=\mathcal{O}(h^{\theta})$ for data in the fractional space $X_{\theta}$.
  • Sectoriality and the associated smoothing mechanism yield a global finite-time error bound, typically of order $\mathcal{O}(h)$ for sufficiently regular data, and $\mathcal{O}(h^{\theta})$ on $X_{\theta}$.
Thus, convergence is guaranteed without requiring any additional discrete stability hypotheses.
Overall, the autonomous sectorial example exhibits excellent agreement between implicit Euler and Lie–Trotter splitting, both qualitatively and quantitatively, and therefore serves as a baseline for interpreting the more delicate behaviour observed in the non-autonomous case. The theoretical background for these observations can be found, for instance, in Pazy (1983), Lunardi (1995), Bátkai and Piazzera (2005), and Bellen and Zennaro (2003).
5.5 Non-Autonomous Sectorial Case
We finally consider the non-autonomous sectorial case, corresponding to Main Theorem D. Here the delay differential equation takes the form
$\frac{d}{dt}u(t)=Au(t)+B(t)u(t)+Cu(t+\tau),$
where $A$ is a sectorial operator and $B(t)$ is a time-dependent perturbation.
As a concrete model we study the heat equation with time-dependent reaction and delayed feedback:
$u_{t}(t,x)$ $=\kappa_{\mathrm{diff}}u_{xx}(t,x)+\lambda(t)u(t,x)+b\,u(t+\tau,x),\quad t\in[0,T],\ x\in(0,L),$ (5.31)
$u(t,0)$ $=u(t,L)=0,\quad t\geq 0,$
$u(t,x)$ $=\mathrm{history}(t,x),\quad t\in[\tau,0],\ x\in[0,L],$
where $\lambda(t)$ is a bounded Lipschitz function. This model coincides with the autonomous sectorial case studied in the previous section, except that the constant coefficient $\lambda$ is replaced by a time-dependent coefficient $\lambda(t)$.
5.5.1 Connection between theory and numerical implementation
The above non-autonomous reaction–diffusion delay equation fits into the abstract framework developed in Section LABEL:sec:nonauto-sectorial. On the product space $\mathcal{E}_{0}=X\oplus L^{1}([\tau,0];X)$ with $X=L^{2}(0,L)$, the equation can be written as
$\frac{d}{dt}X(t)=\bigl(A+B(t)+C\bigr)X(t),\qquad X(t)=\bigl(u(t),\rho(t,\cdot)\bigr),$
where
$A=\begin{pmatrix}\kappa_{\mathrm{diff}}\Delta_{D}&0\\ 0&0\end{pmatrix},\quad B(t)=\begin{pmatrix}\lambda(t)I&b\,\Phi_{\tau}\\ 0&0\end{pmatrix},\quad C=\begin{pmatrix}0&0\\ 0&\partial_{\sigma}\end{pmatrix}$.
Here $A$ is a sectorial operator generating an analytic semigroup, $C$ is the generator of the contractive left–shift semigroup on the history space, and $B(t)$ is a Lipschitz family of CC–bounded perturbations. By standard non-autonomous perturbation theory (Kato–Tanabe framework), the operator family $\mathcal{L}(t)=A+B(t)+C$ generates a unique evolution family $U(t,s)$ on $\mathcal{E}_{0}$.
5.5.2 Verification of the hypotheses of Main Theorem D
We briefly verify the assumptions (S1)–(S4) of Main Theorem D for the present model.
  • (S1)
    $A=\kappa_{\mathrm{diff}}\Delta_{D}$ with homogeneous Dirichlet boundary conditions is sectorial on $X=L^{2}(0,L)$ and generates an analytic semigroup.
  • (S2)
    $C(u,\rho)=(0,\partial_{\sigma}\rho)$ with $\mathcal{D}(C)=\{(u,\rho):\rho\in W^{1,1}([\tau,0];X),\ \rho(0)=u\}$ is the standard left–shift generator on the history space.
  • (S3)
    The trace operator $\Phi_{\tau}:\rho\mapsto\rho(\tau)$ is bounded $\mathcal{D}(C)\to X$ with respect to the graph norm $\|x\|+\|Cx\|$. Hence each $B(t)$ is CC–bounded with relative bound 0, uniformly in $t$.
  • (S4)
    If $\lambda(\cdot)$ is bounded and Lipschitz on $[0,T]$, then $t\mapsto B(t)$ is Lipschitz as a map $0,T]\to\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})$ endowed with the graph norm.
All hypotheses of Main Theorem D are therefore satisfied.
5.5.3 Frozen-time discretization
At each time step $t_{n}=nh$, we approximate the operator by freezing the coefficients:
$\mathcal{L}(t)\approx\mathcal{L}(t_{n})=A+B(t_{n})+C.$
The associated implicit Euler step is then given by the resolvent
$R_{h}(t_{n})=(I-h(A+B(t_{n})+C))^{-1},\qquad X_{n+1}=R_{h}(t_{n})X_{n}.$
This is exactly the frozen-time backward Euler scheme used in the local defect analysis of Proposition 4.7.
5.5.4 Frozen-time Lie–Trotter splitting
The frozen-time Lie–Trotter splitting step at time $t_{n}$ is defined by
$S_{h}^{B}(t_{n})=\bigl[I-h\,B(t_{n})(I-hC)^{-1}\bigr]^{-1},\qquad P_{h}(t_{n})=S_{C}(h)\,(I-hA)^{-1}\,S_{h}^{B}(t_{n}).$
In computation, each factor corresponds to a familiar substep:
  1. 1.
    $(I-hA)^{-1}$ — implicit diffusion solve;
  2. 2.
    $S_{h}^{B}(t_{n})$ — algebraic update involving the delayed value;
  3. 3.
    $S_{C}(h)$ — exact shift of the history profile (implemented by a ring buffer).
In practical implementations, the algebraic step is often written as $\bigl(I-hB(t_{n})\bigr)^{-1}$. At the discrete level (after spatial and temporal discretization of the history), this coincides with the preconditioned step up to the same order of accuracy, because the delay is evaluated on the already shifted buffer. Thus, the implemented numerical scheme is consistent with the abstract frozen splitting operator analyzed in the theory.
Local defect measured against implicit Euler.
In this dissertation we quantify the splitting defect at the level of the frozen-time backward Euler resolvent
$R_{h}(t_{n}):=(I-h(A+B(t_{n})+C))^{-1},$
rather than against the evolution family. Accordingly, the local consistency estimate takes the form
$\|R_{h}(t_{n})-P_{h}(t_{n})\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{\theta}\,h^{\theta}\quad(0<h\leq h_{0}),$
with the corresponding first-order variant on $\mathcal{D}(C)$ when additional regularity is available. This is exactly the defect bound used throughout the sectorial analysis of the present thesis.
Remark 5.6 (Relation to evolution-family based error estimates).
The analysis developed in this thesis is formulated entirely at the level of the backward Euler resolvent $R_{h}(t_{n})$ and the splitting operator $P_{h}(t_{n})$, and does not rely on the existence of an evolution family.
If, however, one additionally assumes the standard non-autonomous hypotheses used in the Tanabe–Kato framework (e.g. sectoriality of $A$, uniform $C$-boundedness of $B(t)$ with relative bound zero, and Lipschitz continuity of $t\mapsto B(t)$ in the graph norm), then the family $\mathcal{L}(t)=A+B(t)+C$ generates an evolution family $U(t,s)$. In that case, the one-step error against $U$ can be decomposed as
$U(t_{n+1},t_{n})-P_{h}(t_{n})=\bigl(U(t_{n+1},t_{n})-R_{h}(t_{n})\bigr)+\bigl(R_{h}(t_{n})-P_{h}(t_{n})\bigr),$
and the first term admits the classical estimate $\|U(t_{n+1},t_{n})-R_{h}(t_{n})\|=\mathcal{O}(h^{2})$ under the above regularity assumptions. Combining both contributions yields a local error of order $\mathcal{O}(h)+\mathcal{O}(h^{2})$ and, under appropriate discrete stability, a global $\mathcal{O}(h)$ bound with respect to $U$.
Combined with the stability mechanism available in the sectorial framework, this yields a global finite-time error of order $\mathcal{O}(h)$, in full agreement with Main Theorem D.
Remark 5.7.
In the autonomous limit $\lambda(t)\equiv\lambda$, the evolution family reduces to a semigroup discussed in Section 5.4.1.
Remark 5.8 (Boundedness, preconditioning, and the role of ring buffers).
1. Boundedness is a property of the function space, not of the data. Even for smooth history functions, the evaluation map $\Phi_{\tau}:\rho\mapsto\rho(\tau)$ is not bounded on the infinite-dimensional spaces $L^{p}([\tau,0];X)$. Thus, the delay operator is not bounded on $\mathcal{E}_{0}$, and naive resolvent steps of the form $(I-hB(t))^{-1}$ are not well-defined at the abstract level. This difficulty is intrinsic to the infinite-dimensional setting.
2. Why the theory uses the preconditioner $(I-hC)^{-1}$.
The transport resolvent $(I-hC)^{-1}$ maps $\mathcal{E}_{0}$ into $\mathcal{D}(C)$, where point evaluation is bounded. As a result, the composition $B(t)(I-hC)^{-1}$ becomes a bounded operator, and the preconditioned step
$S_{h}^{B}(t)=\bigl[I-h\,B(t)(I-hC)^{-1}\bigr]^{-1}$
is well-defined and uniformly bounded for sufficiently small $h$. This is the analytic mechanism that makes delay evaluation controllable in the infinite-dimensional framework.
3. Why numerical implementations can ignore this difficulty.
After discretization (e.g. via a ring buffer), the history space becomes finite dimensional, and evaluation at $\tau$ reduces to a coordinate projection. All operators are then automatically bounded, and algebraic inverses $(I-hB_{h}(t))^{-1}$ are perfectly well-defined.
4. Bridge between abstract theory and numerics.
Although the abstract analysis explicitly relies on the preconditioner $(I-hC)^{-1}$, this regularization is implicitly built into the discrete scheme through the history shift. Consequently, the numerical Lie–Trotter step coincides, up to order $\mathcal{O}(h)$, with the preconditioned frozen splitting operator analyzed in the theory.
5. Interpretation.
Ring buffers are therefore not merely an implementation detail; they provide a finite-dimensional surrogate for the transport regularization mechanism that is essential in the infinite-dimensional analysis.
5.5.5 Parameter justification
We now present a numerical experiment for the non-autonomous sectorial delay equation, comparing the Lie–Trotter splitting scheme with the implicit Euler method. The parameter setting is identical to that of the autonomous sectorial case (Section 5.4), except that the constant reaction coefficient $\lambda$ is replaced by a smooth time-dependent function $\lambda(t)$, while the diffusion coefficient $\kappa$ and the delay feedback parameter $b$ remain unchanged. Specifically, we consider
$\lambda(t)=\lambda_{0}+\lambda_{1}\sin\!\Bigl(\frac{2\pi t}{T_{\lambda}}\Bigr),\qquad\lambda_{0}=-0.8,\quad\lambda_{1}=0.2,\quad T_{\lambda}=4.0,$
so that $\lambda(t)$ oscillates slowly in time within the interval [1.0,0.6][-1.0,-0.6]. This choice ensures that the operator family
$A(t)=\kappa\Delta_{D}+\lambda(t)I$
remains uniformly sectorial on the spatial domain $(0,1)$ with homogeneous Dirichlet boundary conditions.
The remaining parameters are fixed as
$\kappa=0.02,\qquad b=-0.8,\qquad\tau=-0.6,\qquad h=0.002,\qquad T=8.0,\qquad N_{x}=300.$
The history function is prescribed as
$u(t,x)=0.3+0.2\sin(2\pi x)\cos(2\pi t),\qquad t\in[\tau,0],$
which provides a smooth, non-constant initial profile compatible with the boundary conditions.
5.5.6 Numerical results
Table 3: Center value $u(t,0.5)$ for the non-autonomous diffusion–delay equation with $\lambda(t)=-0.8+0.2\sin(\tfrac{2\pi t}{4})$.
Time $t$ Implicit Euler (IE) Lie–Trotter Splitting
0 $2.9988\times 10^{-1}$ $2.9988\times 10^{-1}$
1 $1.2276\times 10^{-1}$ $1.2276\times 10^{-1}$
2 $814\times 10^{-2} 1.881$$814\times 10^{-2}$
3 $5.6897\times 10^{-3}$ $5.6898\times 10^{-3}$
4 $3.9143\times 10^{-4}$ $3.9141\times 10^{-4}$
5 $2.0877\times 10^{-5}$ $2.0878\times 10^{-5}$
6 $6.3167\times 10^{-6}$ $6.3166\times 10^{-6}$
7 $1.4738\times 10^{-6}$ $1.4738\times 10^{-6}$
8 $3.9015\times 10^{-7}$ $3.8738\times 10^{-7}$
Table 4: CPU time for implicit Euler and Lie–Trotter splitting (non-autonomous diffusion–delay equation, $N_{x}=300$, $h=0.002$, $T=8$).
Method CPU Time (seconds)
Implicit Euler $3.1516$
Lie–Trotter Splitting $0.1100$
Figures 17, 18, and 19 display the numerical solutions of the non-autonomous diffusion–delay equation with $\lambda(t)=-0.8+0.2\sin\!\left(\tfrac{2\pi t}{4}\right)$. The two heat maps (implicit Euler and Lie–Trotter splitting) are visually indistinguishable and exhibit the same qualitative space–time behaviour: the initial profile decays smoothly under the influence of diffusion, while a mild transient modulation is induced by the time-dependent reaction and the delay term. The temporal trace at the spatial midpoint $x=0.5$ (Figure 19) confirms this agreement quantitatively. The values produced by both schemes remain virtually identical over the entire time interval $[0,T]$, as also reflected in Table 3.
A substantial difference is observed in computational cost. Figure 20 and Table 4 show that the Lie–Trotter splitting method is more than an order of magnitude faster than the implicit Euler scheme under the same spatial and temporal discretisation. This performance gain is explained by the structure of the two methods. In the non-autonomous case, the implicit Euler method requires solving at each time step a linear system of the form
$\left(I-h\kappa\Delta_{h}-h\lambda(t_{n+1})I\right)u^{n+1}=u^{n}+hb\,u^{n-m},$
where the coefficient matrix changes with $t$ and therefore cannot be factorised once and reused. By contrast, the Lie–Trotter splitting method involves only a single diffusion solve with a time-independent matrix and an explicit algebraic update for the reaction–delay substep. As a result, the splitting approach achieves comparable accuracy at significantly reduced computational cost in the non-autonomous regime. For further discussion of efficiency aspects in non-autonomous problems, see Bellen and Zennaro (2003); Acquistapace and Terreni (1987); Bátkai et al. (2017).
Figure 17: Heat Map of Lie–Trotter Split Solution of the Non-Autonomous Reaction–Diffusion Delay Equation. Heat(x,t)(x,t) represents $u(t,x)$.
Figure 18: Heat Map of Implicit Euler Solution of the Non-Autonomous Reaction–Diffusion Delay Equation. Heat(x,t)(x,t) represents $u(t,x)$.
Figure 19: Center Value $u(t,L/2)$ over Time by Implicit Euler and Lie–Trotter Split Methods. Blue = Implicit Euler, Orange = Lie–Trotter Split.
Figure 20: Computational Cost between Implicit Euler and Lie–Trotter Split Solutions.
5.5.7 Discussion of numerical results
The numerical experiment in this subsection examines the non-autonomous delay–diffusion equation
$u_{t}(t,x)=\kappa\Delta_{D}u(t,x)+\lambda(t)\,u(t,x)+b\,u(t+\tau,x),\qquad\tau<0,$
where the reaction coefficient is prescribed as $\lambda(t)=\lambda_{0}+\lambda_{\mathrm{amp}}\sin\!\left(2\pi t/T_{a}\right).$ Both the Lie–Trotter splitting and the implicit Euler scheme were applied using the same discretization parameters and the same history data on the interval $\tau,0]$.
The results demonstrate that the two numerical schemes produce virtually identical space–time dynamics and temporal traces at the midpoint $x=0.5$. The solution remains smooth for all times, and its amplitude decays monotonically, reflecting the dissipative character and analytic smoothing property of the uniformly sectorial operator $\kappa\Delta_{D}+\lambda(t)I.$ This observation is consistent with the theoretical result that, in the sectorial non-autonomous setting, analytic smoothing suppresses error accumulation and leads to the same convergence order for Lie–Trotter splitting and implicit Euler.
The time-modulated coefficient $\lambda(t)$ introduces a mild variation in the decay rate but does not compromise stability, and both methods remain robust throughout the simulation. The close agreement in Table 3 further confirms that the Lie–Trotter scheme achieves the same accuracy as the implicit Euler method in tracking the temporal behaviour of the solution.
In contrast, the computational cost differs markedly. As shown in Table 4, the Lie–Trotter method is more than an order of magnitude faster than the implicit Euler method. The reason is structural: in the non-autonomous setting, implicit Euler must solve, at every time step, the linear system
$\left(I-h\kappa\Delta_{h}-h\lambda(t_{n+1})I\right)u^{n+1}=u^{n}+hbu^{n-m},$
whose coefficient matrix changes with $t_{n+1}$. No LU factorization can therefore be reused, resulting in significant computational overhead. In contrast, the Lie–Trotter splitting reuses a single LU factorization of the diffusion operator and treats the reaction and delay components analytically or pointwise. This yields a substantial reduction in runtime while maintaining the same level of accuracy.
Such a computational advantage of splitting methods in comparison with fully implicit schemes is well known in the numerical analysis of time-dependent PDEs, especially in large-scale or non-autonomous settings.
Overall, the experiment illustrates that operator splitting provides an effective and computationally efficient approximation of the non-autonomous evolution system governed by $A+B(t)+C$. It preserves the qualitative behavior of the exact solution and achieves accuracy comparable to implicit Euler, but at dramatically reduced computational expense.
For references to theoretical results related to the sectorial non-autonomous setting, see, e.g., Tanabe (1979); Pazy (1983); Lunardi (1995); Bátkai and Piazzera (2005). For numerical analysis of splitting methods and their computational advantages, see, e.g., Hundsdorfer and Verwer (2003); Lubich (2008); Hansen and Stillfjord (2014); Cheng and et al. (2020); Blanes and Casas (2016).
Conclusion
This thesis developed a unified operator-theoretic framework for the analysis of implicit Euler and Lie–Trotter splitting schemes applied to delay differential equations. The central contribution is a convergence theory formulated at the level of discrete resolvent operators, which makes it possible to separate systematically the roles of local defect structure and global discrete stability.
Summary of contributions.
  • A discrete resolvent framework was introduced that allows the analysis of splitting schemes without assuming the existence of an underlying $C_{0}$-semigroup or evolution family.
  • For non-sectorial problems dominated by transport and delay, it was shown that convergence is not governed by local defect alone but depends critically on discrete stability properties of the splitting propagator, leading to the hierarchy described in Main Theorem A and Main Theorem C.
  • In the sectorial setting, both in the autonomous and non-autonomous cases, analytic smoothing was shown to suppress error accumulation. This yields convergence results for Lie–Trotter splitting of the same order as implicit Euler without the need for additional discrete stability assumptions (Main Theorem B and Main Theorem D).
  • The abstract results were illustrated by numerical experiments for reaction–diffusion delay equations, where the observed behaviour closely matched the theoretical predictions.
Conceptual impact. Beyond individual convergence results, the thesis clarifies the conceptual roles of three mechanisms in the numerical analysis of delay equations:
  1. 1.
    local defect structure of splitting methods,
  2. 2.
    discrete stability of numerical propagators,
  3. 3.
    analytic smoothing generated by sectorial operators.
This perspective explains why classical splitting theory works well in parabolic-type problems while fundamentally different phenomena appear in transport-dominated or non-autonomous settings.
Perspectives. The discrete-resolvent viewpoint developed here suggests several directions for future work, including higher-order splitting schemes, adaptive time stepping under discrete stability constraints, and the extension to nonlinear delay equations and stochastic delay systems.
It is hoped that the framework introduced in this thesis will provide a useful basis for further systematic studies of numerical methods for delay equations from an operator-theoretic perspective.
Appendix A Auxiliary Functional–Analytic Material
This appendix collects basic functional-analytic properties of the transport operator used throughout the paper. These results ensure that the abstract framework introduced in Section 1 is mathematically well posed and compatible with classical semigroup theory.
A.1 Functional-analytic properties of the transport operator
This appendix collects basic functional-analytic properties of the transport operator used throughout the paper. These results ensure that the abstract framework introduced in Chapter 2 is mathematically well posed and compatible with classical semigroup theory.
A.1.1 Hille–Yosida Estimate and Graph Norms
Lemma A.1.
Let $\mathcal{E}_{0}$ is a Banach space, and let $C:\mathcal{D}(C)\subset\mathcal{E}_{0}\to\mathcal{E}_{0}$ be a densely defined closed linear operator that generates a strongly continuous semigroup $(e^{tC})_{t\geq 0}$ on $\mathcal{E}_{0}$. Then, by the Hille–Yosida theorem (cf. (Engel and Nagel, 1999, Theorem II.3.8)), for all $h>0$ sufficiently small, the following estimates hold.
$\left\|C(I-hC)^{-1}\right\|\sim 1/h.$ (A.1)
That is, there exists constant $M>0$ such that
$\left\|C(I-hC)^{-1}\right\|\leq M/h,$
Proof.
Using the resolvent identity
$CR(\lambda,C)=\lambda R(\lambda,C)-I,$
we obtain
$\|CR(\lambda,C)\|\leq\lambda\|R(\lambda,C)\|+1.$
Setting $\lambda=\frac{1}{h}$ and using the Hille–Yosida estimate
$\|R(\lambda,C)\|\leq\frac{M}{\lambda-\omega},$
we obtain
$\|CR(\tfrac{1}{h},C)\|\leq\frac{M}{1-h\omega}+1.$
Since $CR(\tfrac{1}{h},C)=h\,C(I-hC)^{-1}$, we obtain
$\|h\,C(I-hC)^{-1}\|=\|CR(\tfrac{1}{h},C)\|\leq\frac{M}{1-h\omega}+1.$
Hence
$\|C(I-hC)^{-1}\|\leq\frac{1}{h}\left(\frac{M}{1-h\omega}+1\right)\lesssim\frac{1}{h},$
for $0<h\leq h_{0}$. ∎
A.1.2 Closedness and density of the transport operator
Lemma A.2 (closedness and density of the transport/shift generator).
Let $I=[\tau,0]$ (finite interval). Define
$C:\mathcal{D}(C)\subset L^{p}(I)\rightarrow L^{p}(I),\quad C\rho=\rho^{\prime},$
with the domain
$\mathcal{D}(C):=\left\{\rho\in W^{1,p}(I):\rho(\tau)=0\right\}.$
Then $C$ is closed and dense in $L^{p}(I)$.
Proof.
  • 1.
    Density of $\mathcal{D}(C)$ in $\mathbb{R}\oplus L^{p}(I)$
    Take any $\left(u_{0},\rho_{0}\right)\in\mathbb{R}\oplus L^{p}(I)$. Choose a standard mollification $\tilde{\rho}_{n}\in C^{\infty}(I)\subset W^{1,p}(I)$ with $\tilde{\rho}_{n}\rightarrow\rho_{0}$ in $L^{p}(I)$. Now adjust the trace at 0 without changing $L^{p}$ much: pick a smooth bump $\psi_{n}\in C_{c}^{\infty}([-1/n,0])$ with $\psi_{n}(0)=1$ and $\left\|\psi_{n}\right\|_{L^{p}}\leq 1/n$. Set
    $\rho_{n}:=\tilde{\rho}_{n}+\left(u_{0}-\tilde{\rho}_{n}(0)\right)\psi_{n}.$
    Then $\rho_{n}\in W^{1,p}(I),\rho_{n}(0)=u_{0}$, and $\left\|\rho_{n}-\rho_{0}\right\|_{L^{p}}\leq\left\|\tilde{\rho}_{n}-\rho_{0}\right\|_{L^{p}}+\left|u_{0}-\tilde{\rho}_{n}(0)\right|\left\|\psi_{n}\right\|_{L^{p}}\rightarrow 0$. Hence $\left(u_{0},\rho_{n}\right)\in\mathcal{D}(C)$ and $\left(u_{0},\rho_{n}\right)\rightarrow\left(u_{0},\rho_{0}\right)$ in $\mathcal{E}_{0}$. So $\mathcal{D}(C)$ is dense.
  • 2.
    Closedness of $C$ on $\mathcal{E}_{0}$:
    Suppose $\left(u_{n},\rho_{n}\right)\in\mathcal{D}(C)$ with
    $\left(u_{n},\rho_{n}\right)\rightarrow(u,\rho)\text{ in }\mathcal{E}_{0},\quad A\left(u_{n},\rho_{n}\right)=\left(au_{n},\rho_{n}^{\prime}\right)\rightarrow(v,\eta)\text{ in }\mathcal{E}_{0}.$
    Then $u_{n}\rightarrow u$ in $\mathbb{R}$ and $au_{n}\rightarrow v$ in $\mathbb{R}$, hence $v=au$. For the second component: $\rho_{n}\rightarrow\rho$ in $L^{p}(I)$ and $\rho_{n}^{\prime}\rightarrow\eta$ in $L^{p}(I)$. Therefore $\left(\rho_{n},\rho_{n}^{\prime}\right)\rightarrow(\rho,\eta)$ in $L^{p}\times L^{p}$. The graph of the derivative operator with domain $W^{1,p}(I)$ is closed in $L^{p}\times L^{p}$, so $\rho\in W^{1,p}(I)$ and $\rho^{\prime}=\eta$.
    Finally, the trace map $T:W^{1,p}(I)\rightarrow\mathbb{R},T(\varphi)=\varphi(0)$, is continuous. Since $\rho_{n}\rightarrow\rho$ in $W^{1,p}$ (we have both $\rho_{n}\rightarrow\rho$ and $\rho_{n}^{\prime}\rightarrow\rho^{\prime}$ in $L^{p}$ ), we get $\rho_{n}(0)\rightarrow\rho(0)$. But $\rho_{n}(0)=u_{n}\rightarrow u$. Hence $\rho(0)=u$.
    Thus $(u,\rho)\in\mathcal{D}(C)$ and $C(u,\rho)=\left(au,\rho^{\prime}\right)=(v,\eta)$. So $C$ is closed.
A.1.3 Generator property of $C$ on $L^{p}$
Lemma A.3 (generator property of the transport/shift operator).
Let $I=[\tau,0]$ (finite interval). Define
$C:\mathcal{D}(C)\subset L^{p}(I)\rightarrow L^{p}(I),\quad C\rho=\rho^{\prime},$
with the domain
$\mathcal{D}(C):=\left\{\rho\in W^{1,p}(I):\rho(\tau)=0\right\}.$
Then $C$ generates the left-shift semigroup $(S(t))_{t\geq 0}$ on $L^{p}(I)$ defined by
$(S(t)\rho)(\sigma):=\begin{cases}\rho(\sigma+t),&\sigma+t\in I\\ 0,&\sigma+t\notin I\end{cases}$
for $\rho\in L^{p}(I),t\geq 0,\sigma\in I$.
Proof.
We verify the Hille–Yosida estimate. For $\lambda>0$ and $\rho\in L^{p}(I)$, the resolvent equation $(\lambda I-C)\varphi=\rho$ reads
$\lambda\varphi(\sigma)-\varphi^{\prime}(\sigma)=\rho(\sigma),\quad\sigma\in I,\quad\varphi(\tau)=0.$
The unique solution is given by
$\varphi(\sigma)=\int_{\tau}^{\sigma}e^{-\lambda(\sigma-s)}\rho(s)ds,\quad\sigma\in I.$
Using Minkowski’s integral inequality, we estimate
$\|\varphi\|_{L^{p}(I)}$ $=\left(\int_{\tau}^{0}\left|\int_{\tau}^{\sigma}e^{-\lambda(\sigma-s)}\rho(s)ds\right|^{p}d\sigma\right)^{1/p}$
$\leq\int_{\tau}^{0}\left(\int_{s}^{0}e^{-\lambda p(\sigma-s)}d\sigma\right)^{1/p}|\rho(s)|ds$
$=\left(\frac{1-e^{-\lambda p|\tau|}}{\lambda p}\right)^{1/p}\int_{\tau}^{0}e^{-\lambda s}|\rho(s)|ds$
$\leq\left(\frac{1-e^{-\lambda p|\tau|}}{\lambda p}\right)^{1/p}\left(\int_{\tau}^{0}e^{-\lambda ps}ds\right)^{1/p}\|\rho\|_{L^{p}(I)}$
$=\frac{1-e^{-\lambda|\tau|}}{\lambda}\|\rho\|_{L^{p}(I)}\leq\frac{|\tau|}{\lambda}\|\rho\|_{L^{p}(I)}.$
Thus the Hille–Yosida estimate holds:
$\|R(\lambda,C)\|_{\mathcal{L}(L^{p}(I))}\leq\frac{1}{\lambda},\quad\lambda>0.$
By the Hille–Yosida theorem, $C$ generates a $C_{0}$-semigroup on $L^{p}(I)$. It is straightforward to verify that this semigroup coincides with the left-shift semigroup defined above. ∎
A.1.4 Relative boundedness of $D$
Lemma A.4 (relative boundedness of the delay operator).
Let $I=[\tau,0]$ (finite interval) and $1<p<\infty$. Define
$D:L^{p}(I)\rightarrow\mathbb{R},\quad D\rho=\int_{\tau}^{0}\rho(\sigma)\mu(d\sigma),$
where $\mu$ is a finite Borel measure on $I$. Then $D$ is relatively bounded with respect to $C$ with relative bound 0, i.e., for every $\varepsilon>0$ there exists $b(\varepsilon)\geq 0$ such that
$\|D\rho\|_{\mathbb{R}}\leq\varepsilon\|C\rho\|_{L^{p}(I)}+b(\varepsilon)\|\rho\|_{L^{p}(I)}$
for all $\rho\in\mathcal{D}(C)$
Proof.
Since $1<p<\infty$, the Sobolev embedding theorem and the Gagliardo–Nirenberg inequality yields
$\|\rho\|_{L^{\infty}}\leq K\,\|\rho^{\prime}\|_{L^{p}}^{1/p}\,\|\rho\|_{L^{p}}^{1-1/p}$
for some constant $K=K(p,\tau)>0$. By Young’s inequality, for every $\delta>0$ there exists $C_{\delta}>0$ such that
$\|\rho\|_{L^{\infty}}\leq\delta\|\rho^{\prime}\|_{L^{p}}+C_{\delta}\|\rho\|_{L^{p}}.$
for all $\rho\in W^{1,p}(I)$. Therefore
$|D\rho|$ $=\left|\int_{\tau}^{0}\rho(\sigma)\mu(d\sigma)\right|\leq\|\rho\|_{\infty}|\mu|(I)$
$\leq(\delta\|\rho^{\prime}\|_{L^{p}}+C_{\delta}\|\rho\|_{L^{p}})|\mu|(I)$
$=|\mu|(I)\delta\|\rho^{\prime}\|_{L^{p}}+|\mu|(I)C_{\delta}\|\rho\|_{L^{p}}$
$=\varepsilon\|\rho^{\prime}\|_{L^{p}}+b(\varepsilon)\|\rho\|_{L^{p}}$
for all $\rho\in\mathcal{D}(C)$. Given $\varepsilon>0$, we can choose $\delta=\varepsilon/|\mu|(I)$ and set $b(\varepsilon)=|\mu|(I)C_{\delta}$. ∎
A.1.5 Generation of $C+D$ via Miyadera-Voigt
Proposition A.5 (generation of $C+D$).
Let $I=[\tau,0]$ (finite interval) and $1<p<\infty$. Define
$C:\mathcal{D}(C)\subset L^{p}(I)\rightarrow L^{p}(I),\quad C\rho=\rho^{\prime},$
with the domain
$\mathcal{D}(C):=\left\{\rho\in W^{1,p}(I):\rho(\tau)=0\right\},$
and
$D:L^{p}(I)\rightarrow\mathbb{R},\quad D\rho=\int_{\tau}^{0}\rho(\sigma)\mu(d\sigma),$
where $\mu$ is a finite Borel measure on $I$. Then the operator $C+D$ with domain $\mathcal{D}(C)$ generates a $C_{0}$-semigroup on $L^{p}(I)$.
Proof.
By Lemma A.3, $C$ generates a $C_{0}$-semigroup $(S(t))_{t\geq 0}$ on $L^{p}(I)$. Moreover, by Lemma A.4, $D$ is relatively bounded with respect to $C$ with relative bound 0. Therefore, by the Miyadera-Voigt perturbation theorem (see, e.g., (Engel and Nagel, 1999, Theorem III.3.9)), $C+D$ with domain $\mathcal{D}(C)$ generates a $C_{0}$-semigroup on $L^{p}(I)$. ∎
A.2 Technical lemmas for discrete error analysis
This section collects purely discrete identities and estimates used to control error propagation for (possibly non-autonomous) splitting propagators. No commutativity is assumed.
A.2.1 Non-autonomous telescoping identity (discrete Duhamel formula)
Lemma A.6 (Non-autonomous telescoping identity).
Let $\mathcal{E}_{0}$ be a Banach space and let $R_{k},P_{k}\in\mathcal{L}(\mathcal{E}_{0})$ $(k=0,1,\dots,n-1)$ be bounded operators. Set the (time-ordered) products
$\mathcal{R}_{j:k}:=\begin{cases}R_{j}R_{j-1}\cdots R_{k},&j\geq k,\\ I,&j<k,\end{cases}\qquad\mathcal{P}_{j:k}:=\begin{cases}P_{j}P_{j-1}\cdots P_{k},&j\geq k,\\ I,&j<k.\end{cases}$
Then with $E_{k}:=R_{k}-P_{k}$ one has the exact identity
$\mathcal{R}_{n-1:0}-\mathcal{P}_{n-1:0}=\sum_{k=0}^{n-1}\mathcal{R}_{n-1:k+1}\,E_{k}\,\mathcal{P}_{k-1:0}.$ (A.2)
Proof.
Expand
$\mathcal{R}_{n-1:0}-\mathcal{P}_{n-1:0}=(R_{n-1}-P_{n-1})\mathcal{P}_{n-2:0}+R_{n-1}\bigl(\mathcal{R}_{n-2:0}-\mathcal{P}_{n-2:0}\bigr),$
and iterate. This yields (A.2). ∎
Remark A.7 (How the identity is used).
In applications $R_{k}$ is the frozen implicit Euler resolvent and $P_{k}$ is the frozen splitting step, so $E_{k}$ is the local resolvent defect. The identity (A.2) reduces global error bounds to estimating sums of the form $\sum\mathcal{R}_{n-1:k+1}E_{k}\mathcal{P}_{k-1:0}$ under suitable discrete stability assumptions.
A.2.2 Abel transformation (summation-by-parts) in the non-commutative setting
Lemma A.8 (Abel (summation-by-parts) identity).
Let $X$ be a Banach space, let $T\in\mathcal{L}(X)$, and let $(\tau_{k})_{k\geq 0}\subset X$. Define partial sums $A_{k}:=\sum_{j=0}^{k}\tau_{j}$ (with $A_{-1}:=0$). Then for every $n\geq 1$ one has the exact identity
$\sum_{k=0}^{n-1}T^{\,n-1-k}\tau_{k}=T^{\,n-1}A_{0}+\sum_{m=1}^{n-1}\bigl(T^{\,m-1}-T^{\,m}\bigr)\,A_{n-1-m}+A_{n-1}.$ (A.3)
Proof.
Write $\tau_{k}=A_{k}-A_{k-1}$ and expand:
$\sum_{k=0}^{n-1}T^{n-1-k}(A_{k}-A_{k-1})=\sum_{k=0}^{n-1}T^{n-1-k}A_{k}-\sum_{k=0}^{n-1}T^{n-1-k}A_{k-1}.$
Shift indices in the second sum and simplify to obtain (A.3). ∎
Remark A.9 (Where Abel enters).
In the autonomous case, applying (A.3) to the standard error representation $e_{n}=\sum_{k=0}^{n-1}T^{n-1-k}\tau_{k}$ moves derivatives (finite differences) from the partial sums $A_{k}$ to the powers $T^{m}-T^{m-1}$. This is precisely where Ritt/Kreiss-type assumptions yield improved bounds.
A.2.3 Abel summation gains under discrete stability assumptions
Proposition A.10 (Abel estimate under power-boundedness / Ritt / bounded variation).
Let $T\in\mathcal{L}(X)$ and assume $\sup_{m\geq 0}\|T^{m}\|\leq M$.
(i) Power-bounded case. For any sequence $(\tau_{k})$ with partial sums $A_{k}=\sum_{j=0}^{k}\tau_{j}$,
$\left\|\sum_{k=0}^{n-1}T^{n-1-k}\tau_{k}\right\|\leq(M+1)\,\max_{0\leq k\leq n-1}\|A_{k}\|+\Bigl(\sum_{m=1}^{n-1}\|T^{m-1}-T^{m}\|\Bigr)\max_{0\leq k\leq n-1}\|A_{k}\|.$ (A.4)
(ii) Ritt case (logarithmic accumulation). If in addition
$\sup_{m\geq 1}m\|T^{m}-T^{m-1}\|\leq K,$
then
$\left\|\sum_{k=0}^{n-1}T^{n-1-k}\tau_{k}\right\|\leq C(M,K)\,(1+\log n)\,\max_{0\leq k\leq n-1}\|A_{k}\|.$ (A.5)
(iii) Bounded-variation case (uniform accumulation). If instead
$\sum_{m=0}^{\infty}\|T^{m+1}-T^{m}\|<\infty,$
then
$\left\|\sum_{k=0}^{n-1}T^{n-1-k}\tau_{k}\right\|\leq C_{\mathrm{BV}}\,\max_{0\leq k\leq n-1}\|A_{k}\|,\qquad n\geq 1.$ (A.6)
Proof.
(i) is obtained by taking norms in the Abel identity (A.3). (ii) follows since $\|T^{m-1}-T^{m}\|\leq K/m$ and $\sum_{m=1}^{n-1}1/m\leq 1+\log n$. (iii) follows by replacing the harmonic bound with the summable series. ∎
Remark A.11 (Why this is exactly the “stability hierarchy” in the main text).
The three regimes above correspond to increasingly strong control of $\sum\|T^{m}-T^{m-1}\|$. In the error propagation formula, this quantity measures how local defects accumulate over time. This is the structural reason why the main theorems distinguish power-boundedness, Ritt stability, and bounded-variation assumptions.
Remark A.12.
The identity (A.3) is a discrete analogue of summation-by-parts (Abel transformation) and is a standard tool in the analysis of error propagation for operator iterations and time-discretization schemes; see, e.g., Lubich and Nevanlinna (1991); Gröger (1987). Its role in connection with Ritt-type stability conditions is discussed in Nagy and Zemánek (1999); Schweizer (2012).
A.3 Transport Resolvents
This section collects several local and global error estimates for operator splitting schemes and resolvent approximations that are used throughout the main text Chapter 2. We begin with the fundamental local error estimate for the implicit Euler method applied to a $C_{0}$-semigroup, and then proceed to splitting-specific bounds.
Remark A.13 (Notation).
Throughout this section, $\mathcal{E}_{0}$ denotes a Banach space, $L:\mathcal{D}(L)\subset\mathcal{E}_{0}\to\mathcal{E}_{0}$ is a densely defined closed operator generating a $C_{0}$-semigroup $(U(t))_{t\geq 0}$, and $R_{h}:=(I-hL)^{-1}$ denotes the resolvent at $\lambda=1/h$. The graph norm on $\mathcal{D}(L)$ is denoted by $\|x\|_{\mathcal{D}(L)}:=\|x\|_{\mathcal{E}_{0}}+\|Lx\|_{\mathcal{E}_{0}}$.
The following lemma estimates the local error of the implicit Euler method for a $C_{0}$-semigroup.
Lemma A.14 (Local Error Per Step of the Implicit Euler Method).
Let $L:\mathcal{D}(L)\subset\mathcal{E}_{0}\to\mathcal{E}_{0}$ be a densely defined closed linear operator such that $(U(t))_{t\geq 0}$ is a $C_{0}$-semigroup defined by $C+D_{\text{bd}}$ on a Banach space $\mathcal{E}_{0}$, and define the implicit Euler step by $R_{h}:=(I-hL)^{-1}$. Then for every $x\in\mathcal{E}_{0}$, the local error satisfies
$\|U(h)x-R_{h}x\|_{\mathcal{E}_{0}}\leq Ch\|Lx\|_{\mathcal{E}_{0}},$
where $C>0$ is a constant independent of $x$ and $h$, for all sufficiently small $h>0$. In particular,
$\|U(h)-R_{h}\|_{\mathcal{L}(\mathcal{D}(L)\to\mathcal{E}_{0})}\leq Ch.$
Proof.
Let $x\in\mathcal{D}(L)$ be arbitrary. First, observe that
$R_{h}x-x=(I-hL)^{-1}x-x=h(I-hL)^{-1}Lx=hR_{h}Lx,$
where we used that $L$ and $(I-hL)^{-1}$ commute on $\mathcal{D}(L)$. On the other hand, by the variation-of-constants formula, the semigroup satisfies
$U(h)x-x=\int_{0}^{h}\frac{d}{dt}U(t)x\,dt=\int_{0}^{h}LU(t)x\,dt.$
Since $LU(t)x=U(t)Lx$ for $x\in\mathcal{D}(L)$, thus:
$U(h)x-x=\int_{0}^{h}U(t)Lx\,dt.$
Subtracting the two expressions gives
$U(h)x-R_{h}x=\int_{0}^{h}U(t)Lx\,dt-hR_{h}Lx.$
Taking norms and applying the triangle inequality yields
$\|U(h)x-R_{h}x\|_{\mathcal{E}_{0}}\leq\int_{0}^{h}\|U(t)Lx\|_{\mathcal{E}_{0}}dt+h\|R_{h}Lx\|_{\mathcal{E}_{0}}.$
Using $\|U(t)\|\leq Me^{\omega t}$ and $\|R_{h}\|\leq M^{\prime}$ for small $h$, we estimate
$\int_{0}^{h}\|U(t)Lx\|_{\mathcal{E}_{0}}dt\leq Me^{\omega h}h\|Lx\|_{\mathcal{E}_{0}},\quad h\|R_{h}Lx\|_{\mathcal{E}_{0}}\leq hM^{\prime}\|Lx\|_{\mathcal{E}_{0}}.$
Thus,
$\|U(h)x-R_{h}x\|_{\mathcal{E}_{0}}\leq\left(Me^{\omega h}+M^{\prime}\right)h\|Lx\|_{\mathcal{E}_{0}}\leq Ch\|Lx\|_{\mathcal{E}_{0}},$
where $C:=Me^{\omega h_{0}}+M^{\prime}$ is independent of $h\in(0,h_{0}]$. ∎
This estimate is classical; see, e.g., (Pazy, 1983, Chap. 9), (Tanabe, 1979, Chap. 5), or Kobayashi (2003).
Lemma A.15 (Transport resolvent).
Let $C:\mathcal{D}(C)\subset\mathcal{E_{0}}\to\mathcal{E_{0}}$ be the generator of a $C_{0}$-semigroup and let $D:\mathcal{D}(C)\to\mathcal{E}_{0}$ be $C$-bounded with relative bound 0 . Fix $h>0$ small and set
$L:=C+D\quad S_{h}:=\left[I-hD(I-hC)^{-1}\right]^{-1}$
Then on $\mathcal{E}_{0}$ the following identities hold:
$(I-hL)^{-1}=(I-hC)^{-1}S_{h}\quad\text{ and }\quad S_{h}=(I-hC)(I-hL)^{-1}$
Consequently, the exact resolvent step for $L$ can be written as the transport resolvent $(I-hC)^{-1}$ followed by the preconditioned inverse $S_{h}$. Moreover, $S_{h}$ is well-defined and uniformly bounded for small $h$ (by a Neumann-series argument from the relative bound 0 assumption).
Proof.
Using the definition of $S_{h}$, we have
$S_{h}^{-1}=I-hD(I-hC)^{-1}.$
Multiplying both sides of the equation by (IhC)(I-hC) from the right gives
$S_{h}^{-1}(I-hC)=(I-hC)-hD=I-hL.$
Taking inverses yields
$(I-hC)^{-1}S_{h}=(I-hL)^{-1},$
which is the first identity. Rearranging gives the second identity:
$S_{h}=(I-hC)(I-hL)^{-1}.$
This yields the desired factorization of the resolvent:
$(I-hL)^{-1}=(I-hC)^{-1}S_{h}.$
Finally, the relative bound 0 implies $\left\|hD(I-hC)^{-1}\right\|\leq q<1$ for $h$ small, hence $S_{h}=\sum_{m\geq 0}\left(hD(I-hC)^{-1}\right)^{m}$ exists and $\left\|S_{h}\right\|\leq(1-q)^{-1}$. ∎
Similar resolvent factorizations appear implicitly in perturbation arguments for delay equations and operator splitting; see, e.g., Bátkai and Piazzera (2005); Bátkai et al. (2017b).
Lemma A.16 (Boundedness of several composite operators).
Let $C:\mathcal{D}(C)\subset\mathcal{E}_{0}\to\mathcal{E}_{0}$ be the generator of the left–shift (transport) semigroup acting only on the history component; in block form
$C\begin{pmatrix}u\\ \rho\end{pmatrix}=\begin{pmatrix}0\\ \partial_{\sigma}\rho\end{pmatrix}$.
Let
$A_{0}:=\begin{pmatrix}A&0\\ 0&0\end{pmatrix}$,
where $A$ is the diffusion operator on the uu–component, and define the resolvent
$X:=(I-hA_{0})^{-1}=\operatorname{diag}\bigl((I-hA)^{-1},\,I\bigr).$
Then, for sufficiently small $h>0$, the following composite operators are bounded:
$\left\|(I-hCX)^{-1}\right\|,\quad\left\|C(X-I)(I-hC)^{-1}\right\|$
Proof.
  • I
    Note that $A_{0}$ acts only on $u$-components and that $C$ acts only on $\rho$-components. Since
    $X:=\left(I-hA_{0}\right)^{-1}=\operatorname{diag}\left((I-hA)^{-1},I\right),$
    we have
    $CX=C\cdot\operatorname{diag}\left((I-hA)^{-1},I\right)=C.$
    Because $C$ ignores the first block. Consequently,
    $I-hCX=I-hC\quad\text{implies }\quad(I-hCX)^{-1}=(I-hC)^{-1}$
    which is bounded for small $h>0$ by the Hille–Yosida theorem, so we have
    $\left\|(I-hCX)^{-1}\right\|=\left\|(I-hC)^{-1}\right\|\leq M<\infty.$
  • II
    Next, we consider the second operator:
    $C(X-I)(I-hC)^{-1}=C\left(\left(I-hA_{0}\right)^{-1}-I\right)(I-hC)^{-1}.$
    Again in our block setting,
    $X-I=\operatorname{diag}\left((I-hA)^{-1}-I,0\right)$
    so ( XIX-I ) is zero on the history block. Since $C$ acts only on the history block,
    $C(X-I)=0,\quad C(X-I)(I-hC)^{-1}=0$
    Thus any expression where we worry about $C(X-I)(I-hC)^{-1}$ simply vanishes in the block model.
Lemma A.17 (Boundedness of the preconditioned composition).
Let $C$ generate the left-shift semigroup $S_{C}(t)$ on $\mathcal{E}_{0}$, and let $D$ be $C$-bounded with relative bound 0 . For $0<h\leq h_{0}$, define
$S_{h}:=\left[I-hD(I-hC)^{-1}\right]^{-1},\quad P_{h}:=S_{C}(h)S_{h}$
Assume that $S_{h}$ exists and
$\sup_{0<h\leq h_{0}}\left\|S_{h}\right\|\leq M_{S}$
Then there exists $h_{1}\in\left(0,h_{0}\right]$ and a constant $M_{R}\geq 1$ such that
$I-hP_{h}\text{ is invertible and }\left\|\left(I-hP_{h}\right)^{-1}\right\|\leq M_{R}$
for all $0<h\leq h_{1}$. In particular, one may take $M_{R}=2$.
Proof.
Since $\left\|S_{C}(h)\right\|\leq 1$ and $\left\|S_{h}\right\|\leq M_{S}$, we have $\left\|P_{h}\right\|\leq M_{P}:=M_{S}$ for all $0<h\leq h_{0}$. Choose $h_{1}\in\left(0,h_{0}\right]$ so that $hM_{P}\leq 1/2$ whenever $0<h\leq h_{1}$. Then the Neumann series $\left(I-hP_{h}\right)^{-1}=\sum_{k=0}^{\infty}h^{k}P_{h}^{k}$ converges and satisfies
$\left\|\left(I-hP_{h}\right)^{-1}\right\|\leq\sum_{k=0}^{\infty}\left(hM_{P}\right)^{k}\leq\frac{1}{1-hM_{P}}\leq 2.$
Remark A.18 (Why $B$ is always paired with $(I-hC)^{-1}$).
In our product-space setting the delay operator $B$ is typically not bounded on $\mathcal{E}_{0}=X\oplus L^{p}([\tau,0];X)$, because point evaluation $\Phi_{\tau}:\rho\mapsto\rho(\tau)$ is not continuous on $L^{p}$. However, $B$ is well-defined on $\mathcal{D}(C)$ and is CC–bounded with relative bound 0: for every $\varepsilon>0$ there exists $c_{\varepsilon}$ such that
$\|Bx\|\leq\varepsilon\|Cx\|+c_{\varepsilon}\|x\|,\qquad x\in\mathcal{D}(C).$
Since the transport resolvent satisfies $(I-hC)^{-1}:\mathcal{E}_{0}\to\mathcal{D}(C)$ and $\|(I-hC)^{-1}\|\leq 1$, we obtain
$\|h\,B(I-hC)^{-1}\|\leq h\Bigl(\varepsilon\|C(I-hC)^{-1}\|+c_{\varepsilon}\|(I-hC)^{-1}\|\Bigr)\lesssim\varepsilon+c_{\varepsilon}h.$
Choosing $\varepsilon$ and then $h$ sufficiently small yields $\|h\,B(I-hC)^{-1}\|<1$, so the preconditioned step
$\bigl[I-h\,B(I-hC)^{-1}\bigr]^{-1}$
is well-defined by a Neumann series. This is the only bounded combination of $B$ that appears in our resolvent identities and splitting schemes.
This mechanism is closely related to Miyadera-type perturbations and the regularizing effect of resolvents for unbounded delay operators; see, e.g., Miyadera (1966); Engel and Nagel (1999); Bátkai and Piazzera (2005).
A.4 Non–autonomous telescoping identity
Remark A.19 (Heuristic interpretation of the preconditioned BB–step).
The preconditioned one–step map
$S_{h}^{B}(t_{n}):=\bigl[I-h\,B(t_{n})(I-hC)^{-1}\bigr]^{-1}$
is introduced at the resolvent level. It is a bounded operator on $\mathcal{E}_{0}$ for sufficiently small $h>0$.
Formally, recalling that $(I-hK)^{-1}=I+hK+\mathcal{O}(h^{2})$ as $h\to 0$ for bounded $K$, one may view $S_{h}^{B}(t_{n})$ as the resolvent analogue of a short-time $B(t_{n})$–flow. In this sense, the frozen three–factor step
$P_{h}(t_{n})=S_{C}(h)\,(I-hA_{0})^{-1}\,S_{h}^{B}(t_{n})$
can be interpreted as a Lie–Trotter–type factorization associated with the frozen operator $A_{0}+C+B(t_{n})$.
We emphasize that this viewpoint is used only as intuition. All rigorous estimates in this thesis are derived from resolvent identities, preconditioning by $(I-hC)^{-1}$, and interpolation–space bounds, without invoking operator exponentials of the non–sectorial part.
A.4.1 From Ritt/Kreiss stability to variation bounds
Definition A.20 (Ritt and Kreiss stability).
Let $X$ be a Banach space and $T\in\mathcal{L}(X)$.
  • (i)
    $T$ is a Ritt operator if
    $\sup_{n\geq 0}\|T^{n}\|\leq M,\qquad\sup_{n\geq 1}n\,\|T^{n}-T^{n-1}\|\leq M_{1}$
    for some constants $M,M_{1}>0$.
  • (ii)
    $T$ satisfies a Kreiss resolvent condition if
    $\sup_{|z|>1}(|z|-1)\,\|(zI-T)^{-1}\|\leq K$
    for some constant $K>0$.
Lemma A.21 (Ritt $\Rightarrow$ logarithmic variation bound).
Let $T\in\mathcal{L}(X)$ be a Ritt operator. Then for every $N\in\mathbb{N}$,
$\sum_{n=0}^{N-1}\|T^{n+1}-T^{n}\|\;\leq\;M_{1}\sum_{n=0}^{N-1}\frac{1}{n+1}\;\leq\;M_{1}\,(1+\log N).$
In particular, Ritt stability yields at most logarithmic growth of the total variation of the powers.
Proof.
By the Ritt bound $n\|T^{n}-T^{n-1}\|\leq M_{1}$ we have $\|T^{n+1}-T^{n}\|\leq M_{1}/(n+1)$ for $n\geq 0$. Summing this estimate gives the claim. ∎
Remark A.22 (On bounded variation assumptions).
The uniform bounded-variation property $\sup_{N}\sum_{n=0}^{N-1}\|T^{n+1}-T^{n}\|<\infty$ is strictly stronger than the Ritt condition in general. In the main text we therefore treat bounded-variation assumptions as separate discrete stability hypotheses when needed.
A.4.2 Non–autonomous telescoping identity
Lemma A.23 (Non–autonomous telescoping identity).
Let $\{U(t,s)\}_{0\leq s\leq t\leq T}$ be an evolution family on a Banach space $\mathcal{E}_{0}$. Let $\{P_{h}(t_{k})\}_{k\geq 0}$ be bounded one–step operators with step size $h>0$ and $t_{k}:=kh$. Define the local defect
$E_{h}(t_{k}):=U(t_{k+1},t_{k})-P_{h}(t_{k}).$
Then for every $n\in\mathbb{N}$ with $t_{n}\leq T$,
$U(t_{n},0)-P_{h}(t_{n-1})\cdots P_{h}(t_{0})=\sum_{k=0}^{n-1}U(t_{n},t_{k+1})\,E_{h}(t_{k})\,P_{h}(t_{k-1})\cdots P_{h}(t_{0}),$ (A.7)
where $P_{h}(t_{-1})\cdots P_{h}(t_{0})$ is interpreted as the identity.
Proof.
Set
$Q_{n}:=P_{h}(t_{n-1})\cdots P_{h}(t_{0}),\qquad Q_{0}:=I.$
For each $k=0,\dots,n-1$ define
$W_{k}:=U(t_{n},t_{k})\,Q_{k}.$
Then $W_{0}=U(t_{n},0)$ and $W_{n}=Q_{n}$. Moreover,
$W_{k+1}-W_{k}$ $=U(t_{n},t_{k+1})Q_{k+1}-U(t_{n},t_{k})Q_{k}$
$=U(t_{n},t_{k+1})\bigl(P_{h}(t_{k})-U(t_{k+1},t_{k})\bigr)Q_{k}$
$=-\,U(t_{n},t_{k+1})E_{h}(t_{k})Q_{k}.$
Summing $W_{k+1}-W_{k}$ from $k=0$ to n1n-1 yields
$Q_{n}-U(t_{n},0)=\sum_{k=0}^{n-1}(W_{k+1}-W_{k})=-\sum_{k=0}^{n-1}U(t_{n},t_{k+1})E_{h}(t_{k})Q_{k},$
which is equivalent to (A.7). ∎
See also Nevanlinna (1993); Borovykh et al. (2000); Le Merdy (2002) for related identities.
Appendix B Program Listings: Core Numerical Algorithms
This appendix summarizes the essential Python implementations used for the numerical experiments in Section 5. The purpose is not to provide exhaustive code, but to document the algorithmic structure underlying the implementations and to clarify the computational distinction between the fully implicit Euler method and the Lie–Trotter splitting scheme.
B.1 Algorithmic comparison: Implicit Euler vs. Lie–Trotter
Both methods approximate the same delayed reaction–diffusion equation under identical discretization settings. The difference lies in how the nonlinear and history-coupled system is treated.
Implicit Euler (fully coupled solve)
At each time step, implicit Euler requires solving a nonlinear system
$u^{n+1}=u^{n}+hF(t_{n+1},u^{n+1},u^{n+1-m}),$
which is implemented via Newton iteration. A simplified schematic is:
for n in range(steps):
    u_iter = u_prev.copy()
    for k in range(max_iter):
        F = residual(u_iter, u_prev, u_delay)
        J = jacobian(u_iter)              # depends on current iterate
        du = solve(J, -F)
        u_iter += du
    u = u_iter
The Jacobian depends on the iterate $u^{n+1,(k)}$ and must be recomputed and factorized at each Newton step, resulting in substantial computational cost.
Lie–Trotter splitting (decoupled steps)
The Lie–Trotter method separates the diffusion and reaction–delay components. The linear diffusion step is solved implicitly using a constant matrix, while the nonlinear reaction–delay step is computed explicitly or in closed form.
# Pre-factorization (once)
LU = factorize(I - h * kappa * Laplacian)
for n in range(steps):
    # A-step: diffusion
    uA = LU.solve(u_prev)
    # B-step: reaction + delay
    u = reaction_delay_flow(uA, u_delay, h)
    update_history_buffer(u)
Since the diffusion matrix is constant, its factorization is reused throughout the computation, leading to a substantial reduction in computational cost. This structural difference explains the performance gap observed in Chapter 5.
The algorithmic distinction between fully implicit schemes and splitting methods follows the classical numerical analysis literature; see, e.g., Hundsdorfer and Verwer (2003); Blanes and Casas (2016).
B.2 Ring buffer for delay handling
The discrete delay term $u(t+\tau)$ is handled using a ring buffer (circular queue). When $|\tau|=mh$ for some integer $m$, the delayed value corresponds exactly to the solution from $m$ time steps earlier.
Let $\{u^{n}\}_{n\geq 0}$ denote the discrete solution at $t_{n}=nh$. The ring buffer stores the most recent $m+1$ values:
$\mathcal{B}_{n}=(u^{n-m},\dots,u^{n}),$
so that
$u^{n-m}=\mathcal{B}_{n}[0],\qquad u^{n}=\mathcal{B}_{n}[m].$
Each update discards the oldest value and appends the newest one, requiring only constant memory and constant-time updates.
Algorithmic structure
Algorithm 1 Ring buffer update for discrete delay
1:step size $h$, delay $\tau=mh$, initial history $\{u^{-m},\dots,u^{0}\}$
2:Initialize buffer $\mathcal{B}\leftarrow[u^{-m},\dots,u^{0}]$
3:for $n=0,1,\dots$ do
4:  $u^{n}\leftarrow\mathcal{B}[m]$
5:  $u^{n-m}\leftarrow\mathcal{B}[0]$
6:  Compute $u^{n+1}=\Psi_{h}(u^{n},u^{n-m})$
7:  Update buffer: $\mathcal{B}\leftarrow[u^{n-m+1},\dots,u^{n+1}]$
8:end for
Minimal Python implementation
The following short routine implements one time step of a scalar delayed equation using a ring buffer:
def step(u_n, hist, h, a, b):
    u_delay = hist[0]
    u_next  = (u_n + h*b*u_delay) / (1 - h*a)
    hist = np.roll(hist, -1)
    hist[-1] = u_next
    return u_next, hist
This implementation mirrors exactly the discrete history transport structure used throughout the numerical experiments.
The ring buffer thus provides a computational realization of the abstract history formulation
$u_{t}(\sigma)=u(t+\sigma),\qquad\partial_{t}u_{t}=\partial_{\sigma}u_{t},$
used in the operator-theoretic analysis of delay equations Hale and Verduyn Lunel (1993); Bátkai and Piazzera (2005). From a numerical viewpoint, this corresponds to the standard discrete history management in DDE solvers; see, e.g., Bellen and Zennaro (2003); Lubich (1984).
Acknowledgments
First, I would like to express my sincere gratitude to my supervisor, Professor Toru Ohira of the Graduate School of Mathematics at Nagoya University, for his invaluable guidance, encouragement, and insightful discussions throughout this research project. I would also like to thank Professor Toshiaki Hishida of the Graduate School of Mathematics at Nagoya University for enthusiastically answering my questions and for his invaluable advice on generating semigroups and evolution families while preparing this paper. Thanks also go to: Professor Mitsuru Sugimoto, who allowed me to attend his seminar and provided guidance. Thanks also go to Yukihiko Nakata of Aoyama Gakuin University and Junya Nishiguchi of Tohoku University. Lastly, I would like to thank Professor Jong Son Shin at Shizuoka University for his continuous support.