### v9

# A Discrete Resolvent Framework for Operator Splitting in Delay Differential Equations

Abstract
This thesis develops a unified operator–theoretic framework for the analysis of implicit time discretization and operator splitting schemes for delay differential equations (DDEs). The central idea is to formulate both the implicit Euler method and the Lie–Trotter splitting scheme in terms of discrete resolvent operators acting on suitable product spaces that encode the present state and the history of the solution.
The theory is constructed entirely at the level of discrete propagators and does not presuppose the existence of a continuous semigroup or evolution family generated by the underlying delay operator. Nevertheless, whenever standard well-posedness assumptions hold and the continuous operator generates a $C_{0}$-semigroup (in the autonomous case) or a strongly continuous evolution family (in the non-autonomous case), the discrete framework is compatible with the classical continuous theory. In particular, convergence of Lie–Trotter splitting toward implicit Euler implies convergence toward the continuous solution on finite time intervals.
The approach yields a transparent hierarchy of convergence results covering autonomous and non-autonomous problems, non-sectorial operators arising from transport and delay terms, and block-structured models with sectorial operators acting on the present-state component.
A central finding is that, for delay equations formulated on product spaces, the dominant contribution to the splitting defect originates from the non-sectorial transport dynamics on the history component. As a consequence, the local defect typically exhibits order $O(h^{1-\theta})$ on fractional spaces, both in non-sectorial models and in sectorial block-structured models. This shows that analytic smoothing on the present-state component alone does not automatically improve the convergence order as long as the history dynamics remain non-sectorial.
Accordingly, global convergence of Lie–Trotter splitting toward implicit Euler is governed by discrete stability mechanisms such as summability of powers, bounded variation, or Ritt/Kreiss-type conditions, in both autonomous and non-autonomous settings. The abstract theory is complemented by applications to reaction–diffusion delay equations and numerical experiments confirming the sharpness of the theoretical error bounds.
Overall, the thesis provides a systematic discrete-resolvent framework for the analysis of splitting methods for delay equations and clarifies the structural role of discrete stability in relation to classical semigroup and evolution family theory.
Contents
Introduction
Background and motivation.
Delay differential equations (DDEs) arise naturally in a wide range of models, including population dynamics, control systems with aftereffect, and reaction–diffusion processes with memory. A fundamental difficulty in the analysis of such systems is the presence of time-nonlocal terms, which prevents the direct application of classical semigroup theory on the state space. A powerful approach, developed since the seminal works of Hale and Verduyn Lunel (1993); Diekmann et al. (1995) and further systematized through the product-space formulation in Bátkai and Piazzera (2005), is to reformulate DDEs as abstract evolution equations on suitable product spaces, in which the present state and its history are treated simultaneously.
From the numerical point of view, fully implicit schemes such as the implicit Euler method provide robust stability properties but are computationally expensive, especially for large-scale or non-autonomous problems. Operator splitting methods, and in particular Lie–Trotter splitting, offer an efficient alternative by separating transport, delay, and reaction effects. Despite their practical relevance, the convergence theory of splitting schemes for delay equations remains incomplete, especially in the presence of
  • non-sectorial generators, such as transport operators arising from history dynamics;
  • non-autonomous coefficients, for which classical semigroup tools are no longer directly applicable.
Limitations of existing theory.
Most existing convergence results for splitting methods rely, either explicitly or implicitly, on the existence of a $C_{0}$-semigroup or an evolution family generated by the underlying continuous operator. This excludes many natural delay models, in which continuous well-posedness is unavailable or too weak to support quantitative error analysis. As a result, the role of purely discrete stability mechanisms in governing convergence has not been fully clarified.
Main idea and contribution of this thesis.
The central idea of this thesis is to develop a convergence theory for implicit Euler and Lie–Trotter splitting formulated entirely at the level of discrete resolvent operators, without presupposing the existence of an underlying semigroup or evolution family.
Instead of starting from continuous well-posedness, we analyze the discrete one-step maps
$R_{h}\quad\text{(implicit Euler resolvent)},\qquad P_{h}\quad\text{(Lie--Trotter splitting propagator)},$
and study convergence of their iterates in terms of
  • local defect estimates between $R_{h}$ and $P_{h}$, and
  • suitable notions of discrete stability (summability, bounded variation, Ritt/Kreiss-type conditions).
This viewpoint allows us to treat regimes beyond classical semigroup theory, while remaining fully compatible with it whenever a continuous semigroup or evolution family does exist.
Abstract convergence framework.
A central theoretical contribution of this thesis is the construction of a unified abstract convergence framework formulated in terms of discrete resolvent families. This framework culminates in two abstract convergence principles:
  • Theorem 2.14: autonomous non-sectorial case,
  • Theorem 2.16: non-autonomous non-sectorial case.
These theorems isolate the essential ingredients governing convergence: local defect estimates on suitable regularity spaces and discrete stability properties of the splitting propagators. Notably, no semigroup or evolution family assumptions are required at the abstract level.
Main theoretical findings.
The analysis developed in this thesis reveals a structural limitation in the convergence of operator splitting schemes for delay equations formulated on product spaces.
Even in the presence of a sectorial diffusion operator acting on the present-state component, the dominant contribution to the local defect is governed by the transport operator on the history component. As a consequence, the local error typically satisfies
$\|R_{h}(t)-P_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}=O(h^{1-\theta}),$
both in non-sectorial models and in sectorial block-structured models. This shows that analytic smoothing on the present-state component alone does not automatically improve the convergence order when the history dynamics remain non-sectorial.
Concrete instantiation for delay equations.
The abstract theory is applied to delay differential equations reformulated on product spaces. Concrete convergence results are derived as Main Theorems A and C in the respective chapters:
  • Section 3: autonomous and non-autonomous non-sectorial cases,
  • Section 4: autonomous and non-autonomous sectorial block-structured cases.
These results provide sharp global error bounds for Lie–Trotter splitting in terms of fractional regularity of the data and clearly exhibit the role of discrete stability assumptions.
Relation to continuous theory.
Although the theory developed in this thesis does not require continuous well-posedness, it is fully consistent with classical results whenever a $C_{0}$-semigroup or evolution family exists. In particular, under additional regularity assumptions such as the Acquistapace–Terreni or Kato–Tanabe conditions, the discrete convergence results imply convergence of Lie–Trotter splitting not only to implicit Euler but also to the continuous solution on finite time intervals. In this sense, the discrete resolvent framework provides both an extension of and a conceptual bridge to the traditional semigroup-based approach.
Central message of the thesis.
The main contribution of this thesis is a convergence theory for operator splitting in delay differential equations that is formulated entirely at the level of discrete resolvent operators, without presupposing continuous well-posedness.
The discrete framework is designed to remain meaningful even in regimes where classical well-posedness is unavailable.
This discrete viewpoint reveals a structural limitation intrinsic to delay equations on product spaces: the dominant contribution to the splitting defect is governed by the non-sectorial transport dynamics of the history component. As a consequence, analytic smoothing acting only on the present-state component cannot, by itself, improve the convergence order of splitting schemes. The characteristic fractional rate $h^{1-\theta}$ persists even in block-structured models containing sectorial operators.
The thesis therefore shifts the focus of convergence analysis from classical semigroup generation toward discrete stability mechanisms. Global convergence is shown to depend on summability, bounded variation, or Ritt/Kreiss-type conditions of the discrete propagators, rather than on continuous smoothing properties alone.
This perspective provides both an extension of and a bridge to traditional semigroup theory: whenever a continuous semigroup or evolution family exists, the discrete results recover convergence toward the continuous solution. However, the theory remains meaningful even in regimes where classical well-posedness is unavailable.
Unlike classical numerical analysis, the main convergence results of this thesis are formulated entirely at the discrete level and do not require the prior existence of a continuous solution operator. Whenever a continuous evolution family does exist, the discrete results are consistent with it, but such existence is not assumed.
Structure of the thesis.
Section 1 sets the stage by introducing delay differential equations, reviewing the product-space formulation, and presenting the motivation, scope, and principal contributions of the thesis.
Section 2 constructs the discrete resolvent framework and formulates the abstract convergence theory at the level of discrete propagators.
Sections 3 and 4 provide concrete instantiations for delay equations, treating autonomous and non-autonomous problems in both non-sectorial and sectorial block settings.
Section 5 presents numerical experiments for reaction–diffusion delay equations, illustrating the sharpness and practical relevance of the theoretical error estimates.
Section Conclusion concludes the thesis with a summary of findings and an outlook on future research directions.
Appendices A and B collect auxiliary analytical results and concise program listings supporting the numerical experiments.
Figure 1: Structure of this paper.
1 Functional analytic framework
This section provides the functional analytic framework underlying the continuous and discrete analysis developed in the sequel. Our goal is twofold: first, to place delay differential equations into the classical theory of $C_{0}$-semigroups and evolution families whenever such a framework is available; second, to formulate the operator splitting and implicit time discretization schemes in a manner that remains meaningful even beyond classical well-posedness.
The section is organized as follows. We first introduce the product-space formulation of delay equations and the associated operators. We then discuss perturbation theory and semigroup generation. Next, we recall the theory of fractional power spaces and interpolation, which plays a central role in the convergence analysis of numerical schemes. Finally, we explain how this continuous framework connects to the discrete resolvent formulations used throughout the paper.
1.1 Product-space formulation
Let $\tau<0$, let $X$ be a Banach space, and fix $1<p<\infty$. For a function $u:[\tau,T]\to X$, we encode the history segment by
$\rho_{t}(\sigma):=u(t+\sigma),\qquad\sigma\in[\tau,0].$
We work on the product space
$\mathcal{E}_{0}:=X\oplus L^{p}([\tau,0];X),\qquad\|(u,\rho)\|_{\mathcal{E}_{0}}:=\|u\|_{X}+\|\rho\|_{L^{p}([\tau,0];X)}.$
The phase variable is $x(t):=(u(t),\rho_{t})\in\mathcal{E}_{0}$. The compatibility relation $\rho_{t}(0)=u(t)$ is encoded in the domain of the transport operator below.
1.2 Transport operator and semigroup generation
Define the transport (left-shift) operator
$C(u,\rho):=(0,\partial_{\sigma}\rho),$
with domain
$\mathcal{D}(C):=\bigl\{(u,\rho)\in X\oplus W^{1,p}([\tau,0];X):\ \rho(0)=u\bigr\}.$
Then $C$ is the generator of a strongly continuous left-shift semigroup $(T_{C}(t))_{t\geq 0}$ on $\mathcal{E}_{0}$ (see, e.g., (Engel and Nagel, 1999, Chap. VI)). In particular, its resolvent
$S_{h}^{C}:=(I-hC)^{-1}\in\mathcal{L}(\mathcal{E}_{0})$
is well defined for every $h>0$ and admits an explicit representation.
1.3 Delay coupling and perturbation theory
The delay interaction is represented by the evaluation map
$\Phi_{\tau}:W^{1,p}([\tau,0];X)\to X,\qquad\Phi_{\tau}(\rho):=\rho(\tau),$
which is well-defined on $W^{1,p}$ but does not extend to a bounded operator on $L^{p}([\tau,0];X)$.
We therefore consider operator families $D(t):\mathcal{D}(C)\to\mathcal{E}_{0}$ of the form
$D(t)(u,\rho):=\bigl(a(t)u+b(t)\Phi_{\tau}(\rho),\,0\bigr),\qquad t\in[0,T],$
where $a(\cdot)$ and $b(\cdot)$ are bounded scalar functions.
1.4 Relative boundedness and semigroup generation
Definition 1.1 (CC–bounded operators).
Let $C:\mathcal{D}(C)\subset X\to X$ be densely defined. An operator $B:\mathcal{D}(C)\to X$ is called CC–bounded with relative bound $\alpha$ if there exists $c_{\alpha}>0$ such that
$\|Bx\|_{X}\leq\alpha\|Cx\|_{X}+c_{\alpha}\|x\|_{X},\qquad x\in\mathcal{D}(C).$
Lemma 1.2 ($C$-boundedness of the delay/reaction block on $\mathcal{E}_{0}$).
Let $\tau<0$ and $1<p<\infty$. Set
$\mathcal{E}_{0}:=\mathbb{R}\oplus L^{p}([\tau,0]),\qquad\|(u,\rho)\|_{\mathcal{E}_{0}}:=|u|+\|\rho\|_{L^{p}}.$
Define
$C(u,\rho):=(0,\partial_{\sigma}\rho),\qquad\mathcal{D}(C):=\{(u,\rho):\rho\in W^{1,p}([\tau,0]),\ \rho(0)=u\},$
and
$D(u,\rho):=(au+b\,\rho(\tau),0),\qquad(u,\rho)\in\mathcal{D}(C),$
with constants $a,b\in\mathbb{R}$. Then $D$ is $C$-bounded with relative bound 0, i.e., for every $\varepsilon>0$ there exists $c_{\varepsilon}>0$ such that
$\|Dx\|_{\mathcal{E}_{0}}\leq\varepsilon\|Cx\|_{\mathcal{E}_{0}}+c_{\varepsilon}\|x\|_{\mathcal{E}_{0}},\qquad x\in\mathcal{D}(C).$
In particular, $D$ is $C$-bounded with relative bound <1<1.
Proof.
Fix $\varepsilon>0$ and let $x=(u,\rho)\in\mathcal{D}(C)$. Since $1<p<\infty$, the Sobolev space $W^{1,p}([\tau,0])$ embeds continuously into $C([\tau,0])$, and the Gagliardo–Nirenberg inequality yields
$\|\rho\|_{L^{\infty}}\leq K\,\|\rho^{\prime}\|_{L^{p}}^{1/p}\,\|\rho\|_{L^{p}}^{1-1/p}$
for some constant $K=K(p,\tau)>0$. By Young’s inequality, for every $\delta>0$ there exists $C_{\delta}>0$ such that
$\|\rho\|_{L^{\infty}}\leq\delta\|\rho^{\prime}\|_{L^{p}}+C_{\delta}\|\rho\|_{L^{p}}.$
Choosing $\delta>0$ sufficiently small (depending on $\varepsilon,a,b$) and using $\rho(0)=u$ and $|\rho(\tau)|\leq\|\rho\|_{L^{\infty}}$, we obtain
$|a||u|+|b||\rho(\tau)|\leq(|a|+|b|)\|\rho\|_{L^{\infty}}\leq\varepsilon\|\rho^{\prime}\|_{L^{p}}+c_{\varepsilon}\|\rho\|_{L^{p}}.$
Since $Cx=(0,\rho^{\prime})$, we have $\|Cx\|_{\mathcal{E}_{0}}=\|\rho^{\prime}\|_{L^{p}}$ and hence
$\|Dx\|_{\mathcal{E}_{0}}=|au+b\rho(\tau)|\leq\varepsilon\|Cx\|_{\mathcal{E}_{0}}+c_{\varepsilon}\|x\|_{\mathcal{E}_{0}},$
which is the desired estimate. The relative bound is therefore 0. ∎
Corollary 1.3 (Semigroup well-posedness for autonomous DDEs).
For $1<p<\infty$, the operator $\mathcal{L}=C+D$ with domain $\mathcal{D}(C)$ generates a strongly continuous semigroup on $\mathcal{E}_{0}$.
Proof.
While $D$ is CC–bounded with relative bound 0, we do not rely on a perturbation theorem here. Instead, semigroup generation for $C+D$ follows from the general delay semigroup framework of (Engel and Nagel, 1999, Sec. VI.6). The proof is ginven in Appendix A.5. ∎
1.5 Well-posedness of the continuous problem
1.5.1 Generation of the semigroup for $C+D$
Proposition 1.4 (Semigroup generation for $C+D$ on $\mathcal{E}_{0}$ ($1<p<\infty$)).
Let $\tau<0$ and $1<p<\infty$. Set
$\mathcal{E}_{0}:=X\oplus L^{p}([\tau,0];X),\qquad\|(u,\rho)\|_{\mathcal{E}_{0}}:=\|u\|_{X}+\|\rho\|_{L^{p}([\tau,0];X)}.$
Define the transport operator
$C(u,\rho):=(0,\partial_{\sigma}\rho),\qquad\mathcal{D}(C):=\bigl\{(u,\rho)\in X\oplus W^{1,p}([\tau,0];X):\ \rho(0)=u\bigr\}.$
Let $a,b\in\mathcal{L}(X)$ be bounded operators and define
$D(u,\rho):=\bigl(au+b\,\rho(\tau),\,0\bigr),\qquad(u,\rho)\in\mathcal{D}(C).$
Then the operator $C+D$ with domain $\mathcal{D}(C)$ generates a $C_{0}$–semigroup on $\mathcal{E}_{0}$.
Proof.
Step 1: $C$ generates the left–shift semigroup on $\mathcal{E}_{0}$. Define $(T(t))_{t\geq 0}$ on $\mathcal{E}_{0}$ by
$T(t)(u,\rho):=(u_{t},\rho_{t}),$
where $\rho_{t}$ is the left shift of $\rho$ by $t$ with inflow $u$ at $\sigma=0$:
$\rho_{t}(\sigma):=\begin{cases}\rho(\sigma+t),&\sigma+t\in[\tau,0],\\[5.69054pt] u,&\sigma+t>0,\end{cases}\qquad\sigma\in[\tau,0],$
and $u_{t}:=\rho_{t}(0)=u$. Then $T(t+s)=T(t)T(s)$, $T(0)=I$, and $\|T(t)\|\leq 1$ on $\mathcal{E}_{0}$. Moreover, $T(t)x\to x$ in $\mathcal{E}_{0}$ as $t\downarrow 0$ for all $x\in\mathcal{E}_{0}$, so $(T(t))_{t\geq 0}$ is a $C_{0}$–semigroup. A standard computation shows that its generator is precisely $C$ with domain $\mathcal{D}(C)$; see, for instance, (Pazy, 1983, Ch. III) or Engel and Nagel (1999).
Step 2: Verification of the Miyadera–Voigt condition in the base norm.
Let $(T(t))_{t\geq 0}$ be the $C_{0}$–semigroup on $\mathcal{E}_{0}$ generated by $C$ from Step 1. For $x=(u,\rho)\in\mathcal{D}(C)$ we have
$DT(s)x=\bigl(au+b\,\rho_{s}(\tau),\,0\bigr),$
where $\rho_{s}$ is given by the shift–with–inflow formula
$\rho_{s}(\sigma)=\begin{cases}\rho(\sigma+s),&\sigma+s\in[\tau,0],\\ u,&\sigma+s>0.\end{cases}$
Fix $t_{0}\in(0,-\tau]$. Then for $0\leq s\leq t_{0}$ we have $\tau+s\leq 0$, hence
$\rho_{s}(\tau)=\rho(\tau+s),$
i.e. no inflow contribution appears at the left endpoint. Therefore,
$\|DT(s)(u,\rho)\|_{\mathcal{E}_{0}}=\|au+b\,\rho(\tau+s)\|_{X}\leq\|a\|\,\|u\|_{X}+\|b\|\,\|\rho(\tau+s)\|_{X}.$
Integrating over $s\in[0,t_{0}]$ and using Hölder’s inequality ($1/p+1/p^{\prime}=1$) yields
$\int_{0}^{t_{0}}\|DT(s)(u,\rho)\|_{\mathcal{E}_{0}}\,ds\leq\|a\|\,t_{0}\,\|u\|_{X}+\|b\|\int_{0}^{t_{0}}\|\rho(\tau+s)\|_{X}\,ds\leq\|a\|\,t_{0}\,\|u\|_{X}+\|b\|\,t_{0}^{1/p^{\prime}}\|\rho\|_{L^{p}}.$
Hence
$\int_{0}^{t_{0}}\|DT(s)(u,\rho)\|_{\mathcal{E}_{0}}\,ds\leq q\,\|(u,\rho)\|_{\mathcal{E}_{0}},\qquad q:=\max\{\|a\|t_{0},\ \|b\|t_{0}^{1/p^{\prime}}\}.$
Choosing $t_{0}>0$ sufficiently small so that $q<1$, the Miyadera–Voigt perturbation condition is satisfied. Consequently, by the Miyadera–Voigt perturbation theorem (see, e.g., (Engel and Nagel, 1999, Thm. III.3.14)), the operator $C+D$ with domain $\mathcal{D}(C)$ generates a $C_{0}$–semigroup on $\mathcal{E}_{0}$.
Proposition 1.5 (Semigroup generation for $A_{0}+C+D$ on $\mathcal{E}_{0}$).
Let $\tau<0$ and $1<p<\infty$. Set
$\mathcal{E}_{0}:=X\oplus L^{p}([\tau,0];X),\qquad\|(u,\rho)\|_{\mathcal{E}_{0}}:=\|u\|_{X}+\|\rho\|_{L^{p}([\tau,0];X)}.$
Let $A:\mathcal{D}(A)\subset X\to X$ generate a $C_{0}$–semigroup $(S(t))_{t\geq 0}$ on $X$ (and in particular, $A$ is densely defined and closed). Define the block operator
$A_{0}:=\mathrm{diag}(A,0),\qquad\mathcal{D}(A_{0}):=\mathcal{D}(A)\oplus L^{p}([\tau,0];X).$
Define the transport operator
$C(u,\rho):=(0,\partial_{\sigma}\rho),\qquad\mathcal{D}(C):=\bigl\{(u,\rho)\in X\oplus W^{1,p}([\tau,0];X):\ \rho(0)=u\bigr\},$
and let $a,b\in\mathcal{L}(X)$ be bounded operators. Define
$D:\mathcal{D}(C)\to\mathcal{E}_{0},\qquad D(u,\rho):=\bigl(au+b\,\rho(\tau),\,0\bigr).$
Set
$G:=A_{0}+C,\qquad\mathcal{D}(G):=\mathcal{D}(A_{0})\cap\mathcal{D}(C),\qquad L:=A_{0}+C+D,\qquad\mathcal{D}(L):=\mathcal{D}(G).$
Then the following hold.
  1. 1.
    $G=A_{0}+C$ generates a $C_{0}$–semigroup $(T(t))_{t\geq 0}$ on $\mathcal{E}_{0}$.
  2. 2.
    $D$ is GG–bounded with relative bound 0, hence in particular with relative bound <1<1.
  3. 3.
    Consequently, $L=A_{0}+C+D$ with domain $\mathcal{D}(G)$ generates a $C_{0}$–semigroup on $\mathcal{E}_{0}$.
Proof.
Step 1: $A_{0}+C$ generates a $C_{0}$–semigroup. Define for $t\geq 0$ an operator $T(t)$ on $\mathcal{E}_{0}$ by
$T(t)(u,\rho):=\bigl(S(t)u,\ \rho_{t}\bigr),$
where $\rho_{t}:[\tau,0]\to X$ is given by the “shift with inflow” rule
$\rho_{t}(\sigma):=\begin{cases}\rho(\sigma+t),&\sigma+t\in[\tau,0],\\[2.84526pt] S(\sigma+t)u,&\sigma+t\in(0,t].\end{cases}$ (1.1)
(So the inflow at $\sigma=0$ is $\rho_{t}(0)=S(t)u$, and the old history is shifted left as long as it stays inside $\tau,0]$.)
We verify that $(T(t))_{t\geq 0}$ is a $C_{0}$–semigroup:
  • Semigroup property. Fix $s,t\geq 0$. Using $S(t+s)=S(t)S(s)$ and the definition (1.1), a direct case distinction on $\sigma+t+s\in[\tau,0]$ or $\sigma+t+s>0$ yields $\rho_{t+s}=(\rho_{s})_{t}$ and hence $T(t+s)=T(t)T(s)$.
  • Strong continuity. Since $S(t)u\to u$ in $X$ as $t\downarrow 0$, it remains to show $\rho_{t}\to\rho$ in $L^{p}$. On the region where $\sigma+t\in[\tau,0]$ we have $\rho_{t}(\sigma)=\rho(\sigma+t)$, which converges to $\rho(\sigma)$ in $L^{p}$ by translation-continuity on a finite interval. The remaining piece where $\sigma+t>0$ has measure $t$ and satisfies $\|\rho_{t}(\sigma)\|_{X}\leq\sup_{0\leq r\leq t}\|S(r)\|\,\|u\|_{X}$; hence its $L^{p}$-norm vanishes as $t\downarrow 0$. Therefore $\rho_{t}\to\rho$ in $L^{p}$ and thus $T(t)x\to x$ in $\mathcal{E}_{0}$.
Hence $(T(t))_{t\geq 0}$ is a $C_{0}$–semigroup on $\mathcal{E}_{0}$.
We now identify its generator. Let $x=(u,\rho)\in\mathcal{D}(G)=\mathcal{D}(A_{0})\cap\mathcal{D}(C)$, so $u\in\mathcal{D}(A)$, $\rho\in W^{1,p}$ and $\rho(0)=u$. A standard differentiation of (1.1) at $t=0$ gives
$\frac{d}{dt}\Big|_{t=0}S(t)u=Au,\qquad\frac{d}{dt}\Big|_{t=0}\rho_{t}(\sigma)=\partial_{\sigma}\rho(\sigma)\ \ \text{in }L^{p},$
and therefore the generator equals $A_{0}+C$ on $\mathcal{D}(G)$. Thus $G=A_{0}+C$ generates $(T(t))_{t\geq 0}$.
Step 2: $D$ is GG–bounded with relative bound 0.
Fix $\varepsilon>0$ and let $x=(u,\rho)\in\mathcal{D}(G)$. By the $W^{1,p}\hookrightarrow C$ embedding and the same Gagliardo–Nirenberg estimate as in Lemma 1.2, there exists (for every $\delta>0$) a constant $C_{\delta}>0$ such that
$\|\rho(\tau)\|_{X}\leq\|\rho\|_{L^{\infty}}\leq\delta\,\|\rho^{\prime}\|_{L^{p}}+C_{\delta}\,\|\rho\|_{L^{p}}.$ (1.2)
Hence
$\|Dx\|_{\mathcal{E}_{0}}=\|au+b\rho(\tau)\|_{X}\leq\|a\|\,\|u\|_{X}+\|b\|\,\|\rho(\tau)\|_{X}$
$\leq\|b\|\delta\,\|\rho^{\prime}\|_{L^{p}}+\|a\|\,\|u\|_{X}+\|b\|C_{\delta}\,\|\rho\|_{L^{p}}.$
We next relate $\|\rho^{\prime}\|_{L^{p}}$ to $\|Gx\|_{\mathcal{E}_{0}}$. Since
$Gx=(Au,\rho^{\prime}),\qquad\|Gx\|_{\mathcal{E}_{0}}=\|Au\|_{X}+\|\rho^{\prime}\|_{L^{p}},$
we have the trivial bound $\|\rho^{\prime}\|_{L^{p}}\leq\|Gx\|_{\mathcal{E}_{0}}$. Also $\|u\|_{X}+\|\rho\|_{L^{p}}=\|x\|_{\mathcal{E}_{0}}$. Therefore,
$\|Dx\|_{\mathcal{E}_{0}}\leq\|b\|\delta\,\|Gx\|_{\mathcal{E}_{0}}+\max\{\|a\|,\|b\|C_{\delta}\}\,\|x\|_{\mathcal{E}_{0}}.$
Choose $\delta>0$ so small that $\|b\|\delta\leq\varepsilon$ and set $c_{\varepsilon}:=\max\{\|a\|,\|b\|C_{\delta}\}$. Then
$\|Dx\|_{\mathcal{E}_{0}}\leq\varepsilon\,\|Gx\|_{\mathcal{E}_{0}}+c_{\varepsilon}\,\|x\|_{\mathcal{E}_{0}},\qquad x\in\mathcal{D}(G).$
Since $\varepsilon>0$ was arbitrary, $D$ is GG–bounded with relative bound 0.
Step 3: Semigroup generation for $A_{0}+C+D$.
By Step 1, $G=A_{0}+C$ generates a $C_{0}$–semigroup $(T(t))_{t\geq 0}$ on $\mathcal{E}_{0}$. By Step 2, $D$ is GG–bounded with relative bound 0, i.e., for every $\varepsilon>0$ there exists $c_{\varepsilon}>0$ such that
$\|Dx\|_{\mathcal{E}_{0}}\leq\varepsilon\,\|Gx\|_{\mathcal{E}_{0}}+c_{\varepsilon}\,\|x\|_{\mathcal{E}_{0}},\qquad x\in\mathcal{D}(G).$
Since $(T(t))$ leaves $\mathcal{D}(G)$ invariant and is strongly continuous on $\mathcal{D}(G)$ endowed with the graph norm, there exists for every $t_{0}>0$ a constant $M_{t_{0}}>0$ such that
$\sup_{0\leq s\leq t_{0}}\|T(s)\|_{\mathcal{L}(\mathcal{D}(G))}\leq M_{t_{0}}.$
Hence, for $x\in\mathcal{D}(G)$,
$\int_{0}^{t_{0}}\|DT(s)x\|_{\mathcal{E}_{0}}\,ds\leq t_{0}\,M_{t_{0}}\,(\varepsilon+c_{\varepsilon})\,\|x\|_{\mathcal{D}(G)}.$
Choosing $\varepsilon>0$ and $t_{0}>0$ sufficiently small, we obtain the Miyadera–Voigt perturbation condition. Therefore, by the Miyadera–Voigt perturbation theorem (see, e.g., (Engel and Nagel, 1999, Thm. III.2.7) (which is formulated for contraction semigroups rather than general $C_{0}$–semigroups) , Voigt (1977), Miyadera (1966), (Bátkai and Piazzera, 2005, Thm. 1.37)), the operator $G+D$ with domain $\mathcal{D}(G)$ generates a $C_{0}$–semigroup on $\mathcal{E}_{0}$. See also Appendix A.5. Since $G+D=A_{0}+C+D=L$, the claim follows. ∎
Remark 1.6 (Why no additional “sum of generators” condition is needed).
Statement (1) is not proved via an abstract “sum of generators” theorem. Instead, it follows from the standard product-space construction for delay equations: one defines a $C_{0}$–semigroup $(T(t))_{t\geq 0}$ on $\mathcal{E}_{0}$ by
$T(t)(u,\rho)=(S(t)u,\rho_{t}),\qquad\rho_{t}(\sigma)=\begin{cases}S(t+\sigma)u,&t+\sigma\geq 0,\\ \rho(t+\sigma),&t+\sigma<0,\end{cases}$
and verifies that its generator coincides with $G=A_{0}+C$ on $\mathcal{D}(G)=\mathcal{D}(A_{0})\cap\mathcal{D}(C)$. We refer to (Bátkai and Piazzera, 2005, Thm. 1.37) (and also Engel and Nagel (1999)) for this classical construction.
The generation of $A_{0}+C$ does not rely on an abstract sum theorem, but follows from the classical product-space semigroup construction for delay equations.
1.6 Non-autonomous case and evolution families
When the coefficients $a(t),b(t)$ depend on time, the associated operator family
$\mathcal{L}(t):=C+D(t)$
defines a non-autonomous evolution equation
$x^{\prime}(t)=\mathcal{L}(t)x(t).$
Lipschitz continuity of $t\mapsto D(t)$ in the CC–graph norm is sufficient for the discrete analysis developed later. However, the existence of an evolution family generally requires stronger time regularity. Accordingly, whenever we appeal to the continuous theory of evolution families, we impose the classical Kato–Tanabe assumption that for each $x\in\mathcal{D}(C)$, the mapping $t\mapsto\mathcal{L}(t)x$ belongs to $C^{1}([0,T];\mathcal{E}_{0})$ (cf. (Pazy, 1983, Thm. 5.3), (Tanabe, 1979, Ch. 7)).
Under this assumption, the family $\mathcal{L}(t)$ generates a unique evolution family $U(t,s)$ satisfying
$\frac{d}{dt}U(t,s)x=\mathcal{L}(t)U(t,s)x,\qquad U(s,s)=I,$
with the growth bound
$\|U(t,s)\|\leq Me^{\omega(t-s)}.$
Remark 1.7.
Time regularity and evolution families. The Lipschitz continuity introduced above is sufficient for the discrete analysis developed in this paper. However, it is in general not sufficient to guarantee the existence of an evolution family for the continuous non-autonomous problem.
Whenever the existence of an evolution family is required, we therefore impose the following additional regularity assumption, in accordance with the classical Kato–Tanabe theory of non-autonomous evolution equations.
For every $x\in\mathcal{D}(C)$, the mapping $t\mapsto\mathcal{L}(t)x$ belongs to $C^{1}([0,T];\mathcal{E}_{0})$.
Under this assumption, the operator family
$\mathcal{L}(t)=C+B(t)\qquad(\text{or more generally }\mathcal{L}(t)=A_{0}+C+B(t))$
with common domain $\mathcal{D}(C)$ generates a unique evolution family $\{U(t,s):0\leq s\leq t\leq T\}\subset\mathcal{L}(\mathcal{E}_{0})$ satisfying
$\frac{d}{dt}U(t,s)x=\mathcal{L}(t)U(t,s)x,\qquad U(s,s)=I.$
Moreover, there exist constants $M,\omega>0$ such that
$\|U(t,s)\|\leq Me^{\omega(t-s)},\qquad 0\leq s\leq t\leq T.$
The combination of sectoriality, perturbation theory, interpolation scales, and sufficient time regularity of $\mathcal{L}(t)$ provides the continuous well-posedness framework when needed. Importantly, the core discrete resolvent analysis developed in this paper does not rely on the existence of such an evolution family, and remains meaningful even beyond the classical well-posedness regime.
1.7 Evolution family for the non-autonomous problem
We record a convenient generation result for the non-autonomous operator family
$L(t)=A_{0}+C+D(t),\qquad t\in[0,T],$
under the standard Kato–Tanabe (“hyperbolic”) assumptions. A systematic treatment can be found in Tanabe’s monograph on evolution equations (and also in Pazy’s book); see, e.g., (Pazy, 1983, Ch. 5), (Tanabe, 1979, Ch. 4).
1.7.1 Kato stability and $C^{1}$-regularity in the graph norm
Let $X$ be a Banach space and $D\subset X$ a dense subspace. For each $t\in[0,T]$ let $A(t):D\to X$ be a closed operator. Write $\|\cdot\|_{D}$ for a Banach norm on $D$ equivalent to the graph norm $\|x\|+\|A(t)x\|$ (uniformly in $t$).
Definition 1.8 (Kato stability on a common domain).
A family $\{A(t)\}_{t\in[0,T]}$ on a common domain $D$ is called (uniformly) Kato stable if there exist constants $M\geq 1$ and $\omega\in\mathbb{R}$ such that for every $m\in\mathbb{N}$ and every partition $0\leq t_{1}\leq t_{2}\leq\cdots\leq t_{m}\leq T$ one has
$\bigl\|(\lambda-A(t_{m}))^{-1}\cdots(\lambda-A(t_{1}))^{-1}\bigr\|\leq M(\lambda-\omega)^{-m},\qquad\lambda>\omega.$
Definition 1.9 ($C^{1}$-regularity in the graph norm).
We say that $A(\cdot)$ is $C^{1}$ as a map $0,T]\to\mathcal{L}(D,X)$ (with respect to $\|\cdot\|_{D}$) if for every $x\in D$ the map $t\mapsto A(t)x$ is continuously differentiable in $X$ and
$\sup_{t\in[0,T]}\|A(t)\|_{\mathcal{L}(D,X)}<\infty,\qquad\sup_{t\in[0,T]}\|A^{\prime}(t)\|_{\mathcal{L}(D,X)}<\infty.$
1.7.2 Evolution family generation theorem (Kato–Tanabe)
Theorem 1.10 (Kato–Tanabe evolution family on a common domain).
Let $\{A(t)\}_{t\in[0,T]}$ be a family of closed operators on $X$ with a common dense domain $D$. Assume:
  1. (H1)
    (uniform Kato stability) $\{A(t)\}$ is Kato stable in the sense of Definition 1.8;
  2. (H2)
    (uniform graph norms) there exists a Banach norm $\|\cdot\|_{D}$ on $D$ such that for all $t\in[0,T]$,
    $\|x\|_{D}\simeq\|x\|+\|A(t)x\|\qquad(x\in D),$
    with constants independent of $t$;
  3. (H3)
    ($C^{1}$-regularity) $A(\cdot)\in C^{1}([0,T];\mathcal{L}(D,X))$ in the sense of Definition 1.9.
Then there exists a unique evolution family $\{U(t,s)\}_{0\leq s\leq t\leq T}\subset\mathcal{L}(X)$ such that:
  1. 1.
    $U(s,s)=I$ and $U(t,r)U(r,s)=U(t,s)$ for $s\leq r\leq t$;
  2. 2.
    for each $x\in D$, the map $t\mapsto U(t,s)x$ is differentiable and
    $\partial_{t}U(t,s)x=A(t)U(t,s)x,\qquad t\in[s,T];$
  3. 3.
    there exist $M\geq 1$ and $\omega\in\mathbb{R}$ such that
    $\|U(t,s)\|_{\mathcal{L}(X)}\leq Me^{\omega(t-s)},\qquad 0\leq s\leq t\leq T.$
Remark 1.11.
Theorem 1.10 is the standard “hyperbolic” non-autonomous generation result attributed to Kato and presented in Tanabe’s monograph. See, e.g., (Pazy, 1983, Ch. 5), (Tanabe, 1979, Ch. 4).
Remark 1.12.
In applications to delay equations, verifying the Kato stability condition is delicate. In the present work we assume (H1)–(H3) when invoking the continuous non-autonomous theory.
1.7.3 Application to $L(t)=A_{0}+C+D(t)$
We now verify the Kato–Tanabe hypotheses for our delay equation framework.
Proposition 1.13 (Evolution family for the non-autonomous delay operator).
Let $1<p<\infty$ and $\tau<0$, and set
$\mathcal{E}_{0}:=X\oplus L^{p}([\tau,0];X),\qquad\mathcal{D}(C):=\{(u,\rho)\in X\oplus W^{1,p}([\tau,0];X):\rho(0)=u\}.$
Let $C(u,\rho):=(0,\partial_{\sigma}\rho)$ on $\mathcal{D}(C)$. Let $A$ be sectorial on $X$ and let $A_{0}:=\mathrm{diag}(A,0)$ on
$\mathcal{D}(A_{0}):=\mathcal{D}(A)\oplus L^{p}([\tau,0];X).$
Assume that the coefficient functions $a(\cdot),b(\cdot)\in C^{1}([0,T])$ are bounded together with their derivatives, and define $D(t):\mathcal{D}(C)\to\mathcal{E}_{0}$ by
$D(t)(u,\rho)=\bigl(a(t)u+b(t)\rho(\tau),\,0\bigr).$
Assume moreover that $D(t)$ is $(A_{0}+C)$–bounded with relative bound <1<1 uniformly in $t$, i.e., there exist $\alpha\in[0,1)$ and $c\geq 0$ such that for all $t\in[0,T]$,
$\|D(t)x\|_{\mathcal{E}_{0}}\leq\alpha\|(A_{0}+C)x\|_{\mathcal{E}_{0}}+c\|x\|_{\mathcal{E}_{0}},\qquad x\in\mathcal{D}(A_{0})\cap\mathcal{D}(C).$
(For the scalar/non-sectorial case take $A_{0}=0$.)
Then the family
$L(t):=A_{0}+C+D(t),\qquad\mathcal{D}(L(t))=\mathcal{D}(A_{0})\cap\mathcal{D}(C),$
generates a unique evolution family $\{U(t,s)\}_{0\leq s\leq t\leq T}$ on $\mathcal{E}_{0}$. Moreover, there exist constants $M\geq 1$ and $\omega\in\mathbb{R}$ such that
$\|U(t,s)\|_{\mathcal{L}(\mathcal{E}_{0})}\leq Me^{\omega(t-s)},\qquad 0\leq s\leq t\leq T.$
Proof.
Set $X_{0}:=\mathcal{E}_{0}$ and $D_{0}:=\mathcal{D}(A_{0})\cap\mathcal{D}(C)$.
Step 1: Each $L(t)$ generates a $C_{0}$-semigroup with uniform bounds. Since $A_{0}$ and $C$ act on different components, $A_{0}+C$ is closed on $D_{0}$. In the sectorial case, $A_{0}$ generates a bounded analytic semigroup on $\mathcal{E}_{0}$ (on the first component and trivially on the second), and $C$ generates the left-shift semigroup on the history component; hence $A_{0}+C$ generates a $C_{0}$-semigroup on $\mathcal{E}_{0}$. (For the non-sectorial case simply take $A_{0}=0$.)
By assumption, $D(t)$ is $(A_{0}+C)$–bounded with relative bound $\alpha<1$ uniformly in $t$. Since $G:=A_{0}+C$ generates an analytic $C_{0}$–semigroup on $\mathcal{E}_{0}$, the perturbation theorem for analytic semigroups (see, e.g., Engel–Nagel, Thm. III.2.10) implies that, for each fixed $t\in[0,T]$, the operator
$L(t)=G+D(t)=A_{0}+C+D(t)$
with domain $\mathcal{D}(G)$ generates a $C_{0}$–semigroup on $\mathcal{E}_{0}$. This can be proven using the proof previously employed in Proposition 1.4. Moreover, the constants in the corresponding Hille–Yosida estimates can be chosen uniformly in $t$ because the relative bound $\alpha$ and the constant $c$ are uniform. Consequently, there exist $M_{0}\geq 1$ and $\omega_{0}\in\mathbb{R}$ such that
$\|(\lambda-L(t))^{-1}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq\frac{M_{0}}{\lambda-\omega_{0}},\qquad\lambda>\omega_{0},\ \ t\in[0,T].$ (1.3)
Step 2: Kato stability (H1). Uniform resolvent bounds (1.3) imply Kato stability (Definition 1.8) with the same constants $M_{0},\omega_{0}$: for any $0\leq t_{1}\leq\cdots\leq t_{m}\leq T$ and $\lambda>\omega_{0}$,
$\bigl\|(\lambda-L(t_{m}))^{-1}\cdots(\lambda-L(t_{1}))^{-1}\bigr\|\leq\prod_{j=1}^{m}\frac{M_{0}}{\lambda-\omega_{0}}=M_{0}^{m}(\lambda-\omega_{0})^{-m}.$
Absorbing $M_{0}^{m}$ into the standard definition (or using the equivalent formulation adopted in Tanabe (1979)/Pazy (1983)) yields (H1). (This is the standard route in Kato–Tanabe theory: uniform Hille–Yosida bounds imply stability.)
Step 3: Uniform equivalence of graph norms (H2). Fix $\lambda>\omega_{0}$. Since $\lambda\in\rho(L(t))$ for all $t$ and the resolvents are uniformly bounded, the norms
$\|x\|_{D_{0},t}:=\|x\|+\|L(t)x\|\quad\text{and}\quad\|x\|_{D_{0}}:=\|x\|+\|(A_{0}+C)x\|$
are uniformly equivalent on $D_{0}$. Indeed, from $L(t)=(A_{0}+C)+D(t)$ and $(A_{0}+C)$–boundedness of $D(t)$ with relative bound $\alpha<1$, we obtain
$\|L(t)x\|\leq\|(A_{0}+C)x\|+\|D(t)x\|\leq(1+\alpha)\|(A_{0}+C)x\|+c\|x\|,$
and conversely,
$\|(A_{0}+C)x\|\leq\|L(t)x\|+\|D(t)x\|\leq\|L(t)x\|+\alpha\|(A_{0}+C)x\|+c\|x\|.$
Hence
$(1-\alpha)\|(A_{0}+C)x\|\leq\|L(t)x\|+c\|x\|,$
which yields the reverse inequality. Collecting both estimates shows that the graph norms are equivalent with constants independent of $t$. Taking $\|\cdot\|_{D_{0}}$ as the reference norm proves (H2).
Step 4: $C^{1}$-regularity in $\mathcal{L}(D_{0},\mathcal{E}_{0})$ (H3). Since $a,b\in C^{1}([0,T])$ and the trace map $\rho\mapsto\rho(\tau)$ is bounded on $W^{1,p}([\tau,0];X)$, the operator $D(t)$ defines a bounded map from $D_{0}$ (with the graph norm of $A_{0}+C$) into $\mathcal{E}_{0}$. Moreover, for each fixed $x=(u,\rho)\in D_{0}$,
$t\mapsto D(t)x=(a(t)u+b(t)\rho(\tau),0)$
is $C^{1}$ in $\mathcal{E}_{0}$, with
$\frac{d}{dt}D(t)x=(a^{\prime}(t)u+b^{\prime}(t)\rho(\tau),0).$
Because $a^{\prime},b^{\prime}$ are bounded, we obtain
$\sup_{t\in[0,T]}\|D(t)\|_{\mathcal{L}(D_{0},\mathcal{E}_{0})}<\infty,\qquad\sup_{t\in[0,T]}\|D^{\prime}(t)\|_{\mathcal{L}(D_{0},\mathcal{E}_{0})}<\infty,$
which yields $D(\cdot)\in C^{1}([0,T];\mathcal{L}(D_{0},\mathcal{E}_{0}))$. Since $A_{0}+C$ is independent of $t$, it follows that $L(\cdot)=A_{0}+C+D(\cdot)\in C^{1}([0,T];\mathcal{L}(D_{0},\mathcal{E}_{0}))$. Thus (H3) holds.
Step 5: Apply the Kato–Tanabe theorem. By Steps 2–4, the family $\{L(t)\}$ satisfies (H1)–(H3) of Theorem 1.10 on $X_{0}=\mathcal{E}_{0}$ with common domain $D_{0}$. Therefore, there exists a unique evolution family $\{U(t,s)\}_{0\leq s\leq t\leq T}$ on $\mathcal{E}_{0}$ solving $\partial_{t}U(t,s)x=L(t)U(t,s)x$ for $x\in D_{0}$, and satisfying the growth bound $\|U(t,s)\|\leq Me^{\omega(t-s)}$. ∎
Remark 1.14 (Why Lipschitz is not enough here).
In Kato–Tanabe hyperbolic theory, Lipschitz continuity of $t\mapsto L(t)$ in the graph norm is often insufficient to ensure differentiability of $t\mapsto U(t,s)x$ for $x$ in the common domain, and the standard convenient criterion is $C^{1}$-regularity in $\mathcal{L}(D,X)$ (as used above); see Tanabe (1979); Pazy (1983). This is the reason why we impose $a,b\in C^{1}([0,T])$ when we need evolution-family generation.
1.8 Fractional power and interpolation spaces
In order to measure the regularity of solutions and to formulate sharp convergence estimates for numerical schemes, we work with fractional domain spaces and interpolation spaces associated with sectorial operators. We follow the standard conventions of Lunardi (1995).
Definition 1.15 (Sectorial operator (Lunardi)).
Let $\mathcal{E}$ be a Banach space. A linear operator $A:\mathcal{D}(A)\subset\mathcal{E}\to\mathcal{E}$ is called sectorial if there exist constants $\omega\in\mathbb{R}$, $\vartheta\in(\pi/2,\pi)$, and $M>0$ such that
$\rho(A)\supset S_{\vartheta,\omega}=\{\lambda\in\mathbb{C}:\lambda\neq\omega,|\arg(\lambda-\omega)|<\vartheta\}$
and
$\|R(\lambda,A)\|_{\mathcal{L}(\mathcal{E})}\leq\frac{M}{|\lambda-\omega|},\quad\forall\lambda\in S_{\vartheta,\omega}.$
(See (Lunardi, 1995, Definition 2.0.1).)
Definition 1.16 (Fractional power spaces).
Let $A$ be a sectorial operator on a Banach space $X$ with $0\in\rho(A)$. For $0<\theta<1$, the fractional power $A^{\theta}$ is defined via the Balakrishnan integral. We define the fractional power space by
$\mathcal{D}(A^{\theta}):=\{x\in X:A^{\theta}x\in X\},$
endowed with the graph norm
$\|x\|_{\mathcal{D}(A^{\theta})}:=\|x\|+\|A^{\theta}x\|.$
Since $0\in\rho(A)$, this norm is equivalent to $\|A^{\theta}x\|$. See, e.g., (Lunardi, 1995, Section 1.14), Engel and Nagel (1999); Haase (2006) for details.
Definition 1.17 (Real interpolation spaces).
Let $(X,\mathcal{D}(A))$ be the Banach couple associated with $A$. For $0<\theta<1$ and $1\leq p\leq\infty$, we define the real interpolation space
$(X,\mathcal{D}(A))_{\theta,p}$
via the KK–method. Throughout this paper we fix $p=1$ and define
$X_{\theta}:=(X,\mathcal{D}(A))_{\theta,1}.$
See, e.g., Lunardi (1995) for details.
Theorem 1.18 (Fractional powers and real interpolation).
Let $A$ be a sectorial operator on a Banach space $X$ with $0\in\rho(A)$. Then for every $0<\theta<1$ one has continuous embeddings
$(X,\mathcal{D}(A))_{\theta,1}\hookrightarrow\mathcal{D}(A^{\theta})\hookrightarrow(X,\mathcal{D}(A))_{\theta,\infty}.$
Moreover, there exist constants $C_{\theta}>0$ such that
$\|A^{\theta}x\|\leq C_{\theta}\|x\|_{(X,\mathcal{D}(A))_{\theta,1}},\qquad\|x\|_{(X,\mathcal{D}(A))_{\theta,\infty}}\leq C_{\theta}\|A^{\theta}x\|.$
See, e.g., (Lunardi, 1995, Theorem 1.15.3) for details.
Remark 1.19.
Throughout the paper we set
$X_{\theta}:=(X,\mathcal{D}(A))_{\theta,1}.$
By Theorem 1.18, we have the continuous embedding $X_{\theta}\hookrightarrow\mathcal{D}(A^{\theta})$, and we shall often identify elements of $X_{\theta}$ with elements of $\mathcal{D}(A^{\theta})$ via this embedding. See, e.g., (Lunardi, 1995, Theorem 1.15.3) for details.
1.9 Connection to numerical schemes
The abstract formulation developed above serves as the continuous reference framework for the numerical analysis carried out in the subsequent chapters. In the autonomous case, implicit Euler discretization approximates the semigroup generated by $\mathcal{L}=C+D$. In the non-autonomous case, it approximates the evolution family generated by $\mathcal{L}(t)$ when such a family exists.
More importantly, the operator-theoretic formulation allows us to define the implicit Euler and Lie–Trotter splitting schemes directly at the level of resolvents, which remains meaningful even in situations where continuous well-posedness fails. This perspective underlies the entire discrete resolvent framework developed in Chapter 2.
Remark 1.20 (How the discrete resolvent framework interfaces with semigroups).
The main results of this thesis (Theorems A and C) compare two discrete propagators: the Lie–Trotter splitting step $P_{h}(t)$ and the (frozen) implicit Euler step $R_{h}(t)$. No continuous evolution object is required for their formulation.
Whenever the continuous operator family admits a well-posed evolution (either a $C_{0}$-semigroup in the autonomous case, or an evolution family $U(t,s)$ in the non-autonomous case), one can recover error bounds with respect to the continuous solution by a simple triangle argument. For instance, in the autonomous case,
$\|e^{t_{n}L}-P_{h}^{n}\|\leq\|e^{t_{n}L}-R_{h}^{n}\|+\|R_{h}^{n}-P_{h}^{n}\|.$
Here the second term is controlled by Theorem A, while the first term is the classical implicit Euler approximation error for the underlying semigroup (see, e.g., Pazy (1983); Engel and Nagel (1999)).
An analogous decomposition applies in the non-autonomous case:
$\|U(t_{n},0)-P_{h}(t_{n-1})\cdots P_{h}(t_{0})\|\leq\|U(t_{n},0)-R_{h}(t_{n-1})\cdots R_{h}(t_{0})\|+\Bigl\|\prod_{k=0}^{n-1}R_{h}(t_{k})-\prod_{k=0}^{n-1}P_{h}(t_{k})\Bigr\|,$
where the last term is precisely the quantity treated by the discrete telescoping analysis in this thesis, and the first term is covered by classical frozen-step/backward-Euler approximation results for evolution families under Kato/Acquistapace–Terreni type hypotheses. See, e.g., Kobayashi (2003, 2005), Tanaka (1994).
Remark 1.21 (Link to classical well-posedness).
Assume that $L(t)=A_{0}+C+D(t)$ generates an evolution family $U(t,s)$ on $\mathcal{E}_{0}$ (e.g. under Kato stability / Acquistapace–Terreni type hypotheses Acquistapace and Terreni (1987)). Then the frozen implicit Euler products $R_{h}(t_{n-1})\cdots R_{h}(t_{0})$ converge to $U(t_{n},0)$ as $h\to 0$. See, e.g., Kobayashi (2003, 2005), Tanaka (1994). Combining this with Theorem C yields
$\|U(t_{n},0)-P_{h}(t_{n-1})\cdots P_{h}(t_{0})\|\leq\|U(t_{n},0)-R_{h}(t_{n-1})\cdots R_{h}(t_{0})\|+\|R\text{-prod}-P\text{-prod}\|.$
2 Discrete resolvent framework and abstract convergence theory
Organization of this section.
We state two main abstract convergence principles: one for the autonomous case and one for the non-autonomous case. First-order variants are formulated as propositions, since they are strengthening statements under additional regularity.
Throughout this section we work on the product space
$\mathcal{E}_{0}:=X\oplus L^{p}([\tau,0];X),\qquad 1<p<\infty,\ \tau<0,$
and use the transport operator
$C(u,\rho):=(0,\partial_{\sigma}\rho),\qquad\mathcal{D}(C):=\{(u,\rho)\in X\oplus W^{1,p}([\tau,0];X):\ \rho(0)=u\}.$
The restriction $1<p<\infty$ is essential in our setting: it ensures that the delay trace becomes bounded after resolvent preconditioning, and in particular avoids pathologies that occur on $L^{1}$.
2.1 Discrete resolvent operators
2.1.1 Resolvent preconditioners
Transport resolvent.
We first introduce the resolvent of the transport operator:
$S_{h}^{C}:=(I-hC)^{-1}\in\mathcal{L}(\mathcal{E}_{0}),\qquad h>0.$ (2.1)
It is well-defined for every $h>0$ and satisfies the standard bounds
$\|S_{h}^{C}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq 1,\qquad\|CS_{h}^{C}\|_{\mathcal{L}(\mathcal{E}_{0})}\lesssim h^{-1}.$ (2.2)
Moreover, for the real interpolation scale
$X_{\theta}:=(\mathcal{E}_{0},\mathcal{D}(C))_{\theta,1},\qquad 0<\theta\leq 1,$
we have the discrete smoothing estimate
$\|S_{h}^{C}\|_{\mathcal{L}(X_{\theta},\mathcal{D}(C))}\lesssim h^{-\theta}.$ (2.3)
Sectorial preconditioner.
In the sectorial block setting we also use the resolvent
$S_{h}^{A}:=(I-hA_{0})^{-1}\in\mathcal{L}(\mathcal{E}_{0}),\qquad A_{0}=\mathrm{diag}(A,0),$
and define the combined preconditioner
$S_{h}^{AC}:=S_{h}^{C}S_{h}^{A}.$ (2.4)
2.1.2 Right-preconditioned delay operator and operational smallness
The delay/reaction part $D(t)$ is typically defined only on $\mathcal{D}(C)$ and does not extend to a bounded operator on $\mathcal{E}_{0}$. Consequently, the splitting step must be formulated through the preconditioned operator $D(t)S_{h}^{C}$ (or $D(t)S_{h}^{AC}$ in the sectorial case).
Standing assumption (bounded preconditioned delay). For each $t\in[0,T]$ and sufficiently small $h>0$, we assume that the composition $D(t)S_{h}^{C}$ extends to a bounded operator on $\mathcal{E}_{0}$ (and similarly $D(t)S_{h}^{AC}$ in the sectorial case). This is the case in all applications in this thesis.
Right-preconditioned operator. We define
$H_{h}(t):=h\,D(t)S_{h}^{C}\qquad(\text{or }H_{h}(t):=h\,D(t)S_{h}^{AC}\text{ in the sectorial case}).$
Assumption 2.1 (Operational smallness).
There exist $h_{0}>0$ and $q\in(0,1)$ such that
$\sup_{t\in[0,T]}\ \sup_{0<h\leq h_{0}}\|H_{h}(t)\|_{\mathcal{L}(\mathcal{E}_{0})}\leq q.$
Although Assumption 2.1 is formulated purely at the discrete level, it is automatically satisfied in most classical delay models under the usual relative boundedness assumption on $D$, as shown in the following proposition.
Proposition 2.2 (Verification of (OS) via relative boundedness).
Assume that the delay operator $D:\mathcal{D}(C)\to\mathcal{E}_{0}$ is CC–bounded with relative bound strictly less than $1$, i.e., there exist $\alpha\in(0,1)$ and $c\geq 0$ such that
$\|Dx\|_{\mathcal{E}_{0}}\leq\alpha\|Cx\|_{\mathcal{E}_{0}}+c\|x\|_{\mathcal{E}_{0}},\qquad x\in\mathcal{D}(C).$
Then there exist $h_{0}>0$ and $q\in(0,1)$ such that
$\sup_{0<h\leq h_{0}}\|hDS_{h}^{C}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq q.$
In particular, the operational smallness condition (OS) is satisfied.
Proof.
Let $y\in\mathcal{E}_{0}$ and set $x=S_{h}^{C}y\in\mathcal{D}(C)$. By the relative boundedness assumption and the standard resolvent bounds
$\|CS_{h}^{C}\|_{\mathcal{L}(\mathcal{E}_{0})}\lesssim h^{-1},\qquad\|S_{h}^{C}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq 1,$
we obtain
$\|hDS_{h}^{C}y\|\leq h\alpha\|CS_{h}^{C}y\|+hc\|S_{h}^{C}y\|\leq\alpha\|y\|+hc\|y\|.$
Choosing $h_{0}>0$ sufficiently small such that $\alpha+h_{0}c<1$ yields the claim. ∎
Proposition 2.3 (Verification of (OS) in the sectorial case).
Assume that $D:\mathcal{D}(C)\to\mathcal{E}_{0}$ is CC–bounded with relative bound $\alpha\in(0,1)$, i.e., there exists $c\geq 0$ such that
$\|Dx\|_{\mathcal{E}_{0}}\leq\alpha\|Cx\|_{\mathcal{E}_{0}}+c\|x\|_{\mathcal{E}_{0}},\qquad x\in\mathcal{D}(C).$
Let $A_{0}=\mathrm{diag}(A,0)$ and set
$S_{h}^{A}:=(I-hA_{0})^{-1},\qquad S_{h}^{AC}:=S_{h}^{C}S_{h}^{A}.$
Assume moreover that $\sup_{0<h\leq h_{0}}\|S_{h}^{A}\|_{\mathcal{L}(\mathcal{E}_{0})}<\infty$ (for example, this holds if $A_{0}$ generates a bounded analytic semigroup). Then there exist $h_{1}>0$ and $q\in(0,1)$ such that
$\sup_{0<h\leq h_{1}}\|hDS_{h}^{AC}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq q.$
In particular, the operational smallness condition (OS) holds with $H_{h}=hDS_{h}^{AC}$.
Proof.
Let $y\in\mathcal{E}_{0}$ and put $x=S_{h}^{AC}y=S_{h}^{C}(S_{h}^{A}y)\in\mathcal{D}(C)$. By the relative boundedness assumption and the resolvent bound $\|CS_{h}^{C}\|\lesssim h^{-1}$,
$\|hDS_{h}^{AC}y\|\leq h\alpha\|CS_{h}^{AC}y\|+hc\|S_{h}^{AC}y\|\leq\alpha\|S_{h}^{A}y\|+hc\,\|S_{h}^{C}\|\,\|S_{h}^{A}y\|.$
Using $\|S_{h}^{C}\|\leq 1$ and $\|S_{h}^{A}\|\leq M_{A}$ (uniformly for $0<h\leq h_{0}$), we obtain
$\|hDS_{h}^{AC}\|\leq M_{A}\bigl(\alpha+hc\bigr).$
Choose $h_{1}\in(0,h_{0}]$ such that $M_{A}(\alpha+h_{1}c)<1$ and set $q:=M_{A}(\alpha+h_{1}c)\in(0,1)$. ∎
Proposition 2.4 (Uniform (OS) for non-autonomous delay operators).
Let $\{D(t)\}_{t\in[0,T]}$ be a family of operators $D(t):\mathcal{D}(C)\to\mathcal{E}_{0}$. Assume that $D(t)$ is CC–bounded with a uniform relative bound $\alpha\in(0,1)$, i.e., there exist $\alpha\in(0,1)$ and $c\geq 0$ such that
$\|D(t)x\|_{\mathcal{E}_{0}}\leq\alpha\|Cx\|_{\mathcal{E}_{0}}+c\|x\|_{\mathcal{E}_{0}},\qquad x\in\mathcal{D}(C),\quad t\in[0,T].$
Then there exist $h_{0}>0$ and $q\in(0,1)$ such that
$\sup_{t\in[0,T]}\ \sup_{0<h\leq h_{0}}\ \|hD(t)S_{h}^{C}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq q.$
Hence Assumption 2.1 holds.
Proof.
Fix $t\in[0,T]$ and let $y\in\mathcal{E}_{0}$, $x=S_{h}^{C}y\in\mathcal{D}(C)$. Using the uniform relative boundedness and $\|CS_{h}^{C}\|\lesssim h^{-1}$, $\|S_{h}^{C}\|\leq 1$,
$\|hD(t)S_{h}^{C}y\|\leq h\alpha\|CS_{h}^{C}y\|+hc\|S_{h}^{C}y\|\leq\alpha\|y\|+hc\|y\|.$
Choose $h_{0}>0$ such that $\alpha+h_{0}c<1$ and take $q:=\alpha+h_{0}c$. The estimate is uniform in $t$. ∎
Proposition 2.5 (Uniform (OS) in the non-autonomous sectorial case).
Let $\{D(t)\}_{t\in[0,T]}$ be operators $D(t):\mathcal{D}(C)\to\mathcal{E}_{0}$. Assume the uniform relative boundedness
$\|D(t)x\|\leq\alpha\|Cx\|+c\|x\|,\qquad x\in\mathcal{D}(C),\ t\in[0,T],$
for some $\alpha\in(0,1)$ and $c\geq 0$. Let $S_{h}^{AC}=S_{h}^{C}S_{h}^{A}$ and assume $\sup_{0<h\leq h_{1}}\|S_{h}^{A}\|\leq M_{A}<\infty$. Then there exist $h_{0}>0$ and $q\in(0,1)$ such that
$\sup_{t\in[0,T]}\ \sup_{0<h\leq h_{0}}\ \|hD(t)S_{h}^{AC}\|\leq q.$
Proof.
As in Proposition 2.3, for $y\in\mathcal{E}_{0}$ we have
$\|hD(t)S_{h}^{AC}y\|\leq\alpha\|S_{h}^{A}y\|+hc\,\|S_{h}^{A}y\|\leq M_{A}(\alpha+hc)\|y\|.$
Choose $h_{0}>0$ such that $M_{A}(\alpha+h_{0}c)<1$ and set $q:=M_{A}(\alpha+h_{0}c)$. ∎
Corollary 2.6 (Relative bound 0 implies (OS)).
Assume that $D:\mathcal{D}(C)\to\mathcal{E}_{0}$ is CC–bounded with relative bound 0, i.e., for every $\varepsilon>0$ there exists $c_{\varepsilon}>0$ such that
$\|Dx\|_{\mathcal{E}_{0}}\leq\varepsilon\|Cx\|_{\mathcal{E}_{0}}+c_{\varepsilon}\|x\|_{\mathcal{E}_{0}},\qquad x\in\mathcal{D}(C).$
Then Assumption 2.1 holds for $H_{h}=hDS_{h}^{C}$. Moreover, if $\sup_{0<h\leq h_{0}}\|S_{h}^{A}\|\leq M_{A}<\infty$, then (OS) also holds for $H_{h}=hDS_{h}^{AC}$.
Proof.
Fix $\varepsilon\in(0,1)$. By relative bound 0 there exists $c_{\varepsilon}\geq 0$ such that $\|Dx\|\leq\varepsilon\|Cx\|+c_{\varepsilon}\|x\|$. Apply Proposition 2.2 with $\alpha=\varepsilon$ to obtain (OS) for $S_{h}^{C}$. For $S_{h}^{AC}=S_{h}^{C}S_{h}^{A}$, the same argument as in Proposition 2.3 yields $\|hDS_{h}^{AC}\|\leq M_{A}(\varepsilon+hc_{\varepsilon})$, so choose $h$ small to make this <1<1. ∎
Corollary 2.7 (Uniform relative bound 0 implies uniform (OS)).
Let $\{D(t)\}_{t\in[0,T]}$ satisfy: for every $\varepsilon>0$ there exists $c_{\varepsilon}>0$ such that
$\|D(t)x\|\leq\varepsilon\|Cx\|+c_{\varepsilon}\|x\|,\qquad x\in\mathcal{D}(C),\ t\in[0,T].$
Then Assumption 2.1 holds uniformly in $t$.
2.1.3 Connection with continuous well-posedness
Assumption 2.1 is purely discrete: it guarantees that the splitting step $P_{h}(t)$ is well-defined and uniformly bounded for small $h$, without invoking any continuous well-posedness of the underlying delay operator. Nevertheless, whenever the classical semigroup framework applies, the discrete assumptions become automatic consequences of standard perturbation theory. We record this compatibility in the following proposition.
Remark 2.8 (Continuous vs. discrete requirements).
In the classical semigroup approach, one typically proves well-posedness of $A_{0}+C+D$ on $\mathcal{E}_{0}=X\oplus L^{p}([\tau,0];X)$ by showing that $D$ is CC–bounded with relative bound $\alpha<1$ (or even 0), which is most conveniently formulated for $1<p<\infty$.
In contrast, the present thesis develops a purely discrete resolvent theory: the definition of the splitting step and all local/global error bounds depend only on the operational smallness condition (OS) for the preconditioned operator $H_{h}(t)=hD(t)S_{h}^{C}$ (and on discrete stability of products of $P_{h}(t)$). Therefore the discrete convergence analysis can be carried out even in settings where the semigroup perturbation argument is not available, including $p=1$, provided that (OS) can be verified.
Proposition 2.9 (Relative bound 0 and compatibility of discrete and continuous frameworks).
Let $X$ be a Banach space, $\tau<0$, and $1<p<\infty$, and set
$\mathcal{E}_{0}:=X\oplus L^{p}([\tau,0];X).$
Let $C$ be the transport operator on $\mathcal{E}_{0}$ with domain $\mathcal{D}(C)$, and let
$D:\mathcal{D}(C)\to\mathcal{E}_{0}$
be CC–bounded with relative bound 0.
Then the following hold.
  1. (i)
    Operational smallness for the splitting scheme. The operational smallness condition (OS) holds for the right–preconditioned operators
    $H_{h}(t)=hD(t)S_{h}^{C}\quad\text{(or }H_{h}=hDS_{h}^{C}\text{ in the autonomous case)}.$
    In particular, the splitting propagator
    $P_{h}(t)=(I-H_{h}(t))^{-1}S_{h}^{C}$
    is well defined and uniformly bounded for all sufficiently small $h>0$. Hence the discrete Lie–Trotter framework is well posed on $\mathcal{E}_{0}$.
  2. (ii)
    Compatibility with the continuous semigroup theory. Remember Propositions 1.4 and 1.5. The well–posedness of the continuous delay equation associated with $C+D$ (or $A_{0}+C+D$ in the sectorial block case) is not deduced from relative boundedness alone, but follows from the general product–space semigroup generation theorems for delay equations; direct proof is given in Appendix A.5. In more detail: see, e.g., Miyadera (1966), Voigt (1977), (Engel and Nagel, 1999, Sec. VI.6), (Bátkai and Piazzera, 2005, Thm. 1.37).
Consequently, the single structural condition that $D$ is CC–bounded with relative bound 0 guarantees the well posedness of the discrete splitting framework, while the well posedness of the continuous problem is ensured independently by the classical product–space semigroup theory. In this sense, the discrete and continuous theories are structurally compatible, but logically distinct.
Remark 2.10.
For completeness we note that the block operator $A_{0}+C$ generates a $C_{0}$–semigroup on $\mathcal{E}_{0}$ whenever $A_{0}$ generates a $C_{0}$–semigroup on $X$. This follows from the standard product–space construction of delay semigroups (see, e.g., Engel and Nagel (1999); Bátkai and Piazzera (2005)), and does not rely on relative boundedness of $D$.
[…]
Figure 2: Logical structure of the operational smallness (OS) condition and its verification mechanisms.
2.1.4 Implicit Euler and Lie–Trotter propagators
For the (possibly time-dependent) full operator
$L(t)=A_{0}+C+D(t),\qquad\mathcal{D}(L(t))=\mathcal{D}(A_{0})\cap\mathcal{D}(C),$
the implicit Euler one-step map is given by the full resolvent
$R_{h}(t):=(I-hL(t))^{-1}=(I-h(A_{0}+C+D(t)))^{-1},$ (2.5)
whenever this inverse exists.
Under Assumption 2.1, the Lie–Trotter splitting step is defined by
$P_{h}(t):=(I-H_{h}(t))^{-1}S_{h}^{C}\qquad(\text{or }P_{h}(t):=(I-H_{h}(t))^{-1}S_{h}^{AC}).$ (2.6)
Note that the correction factor $(I-H_{h}(t))^{-1}$ acts after the transport preconditioner. This order is essential: the operator $D(t)$ is controlled only on $\mathcal{D}(C)$ and must therefore be paired with the regularization provided by $S_{h}^{C}$.
2.1.5 Local resolvent defect
We quantify the one-step discrepancy between implicit Euler and splitting by the local resolvent defect
$E_{h}(t):=R_{h}(t)-P_{h}(t),$
whenever both operators are defined.
The global error analysis proceeds in two conceptually distinct steps:
  • (i)
    estimate $\|E_{h}(t)\|$ on an appropriate regularity space;
  • (ii)
    propagate this local defect under products of $P_{h}(t)$ using discrete stability assumptions.
2.2 Discrete stability and propagation of local defects
2.2.1 Telescoping identity (autonomous case)
In the autonomous case we write $R_{h}(t)=R_{h}$ and $P_{h}(t)=P_{h}$. Then
$R_{h}^{n}-P_{h}^{n}=\sum_{k=0}^{n-1}R_{h}^{n-1-k}\,(R_{h}-P_{h})\,P_{h}^{k}=\sum_{k=0}^{n-1}R_{h}^{n-1-k}\,E_{h}\,P_{h}^{k}.$ (2.7)
This identity is the starting point for all global error estimates.
2.2.2 Stability notions (autonomous case)
We introduce the following standard hierarchy of stability notions on $\mathcal{L}(\mathcal{E}_{0})$.
Definition 2.11 (Power boundedness).
The family $\{P_{h}\}_{0<h\leq h_{0}}$ is uniformly power bounded if
$\sup_{0<h\leq h_{0}}\sup_{n\geq 0}\|P_{h}^{n}\|_{\mathcal{L}(\mathcal{E}_{0})}<\infty.$
Definition 2.12 (Ritt stability).
The family $\{P_{h}\}_{0<h\leq h_{0}}$ is uniformly Ritt stable if
$\sup_{0<h\leq h_{0}}\sup_{n\geq 1}n\|P_{h}^{n}-P_{h}^{n+1}\|_{\mathcal{L}(\mathcal{E}_{0})}<\infty.$
Definition 2.13 (Summability / bounded variation of powers).
We say that $\{P_{h}\}_{0<h\leq h_{0}}$ satisfies summability stability if
$\sup_{0<h\leq h_{0}}\sup_{n\geq 1}\left\|\sum_{k=0}^{n-1}P_{h}^{k}\right\|_{\mathcal{L}(\mathcal{E}_{0})}<\infty.$
These notions form a strict hierarchy
$\text{summability}\Longrightarrow\text{Ritt stability}\Longrightarrow\text{power boundedness},$
corresponding to increasingly weaker control of error accumulation; see, e.g., Gröger (1987); Lubich and Nevanlinna (1991); Nagy and Zemánek (1999); Le Merdy (2002); Gomilko and Zemánek (2011); Schweizer (2012).
2.3 Two abstract convergence principles
We now state two abstract principles that serve as the backbone for the instantiations in later sections. The first covers the autonomous case; the second covers the non-autonomous (time-dependent) case.
Theorem 2.14 (Abstract convergence principle: autonomous case).
Let $X$ be a Banach space and let $\{R_{h}\}_{0<h\leq h_{0}}$ and $\{P_{h}\}_{0<h\leq h_{0}}$ be families in $\mathcal{L}(X)$. Assume that:
  • (A1)
    (Uniform boundedness of the implicit scheme)
    $\sup_{0<h\leq h_{0}}\sup_{n\geq 0}\|R_{h}^{n}\|_{\mathcal{L}(X)}<\infty.$
  • (A2)
    (Local defect on a regularity space) There exist a Banach space $X_{\theta}\hookrightarrow X$ and constants $C>0$, $\theta\in(0,1]$ such that
    $\|R_{h}-P_{h}\|_{\mathcal{L}(X_{\theta},X)}\leq Ch^{\theta}.$
  • (A3)
    (Discrete stability of $P_{h}$) One of the following holds uniformly in $h$:
    • (i)
      summability stability,
    • (ii)
      Ritt stability,
    • (iii)
      power boundedness.
Then for every $T>0$ there exists $C_{T}>0$ such that
$\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(X_{\theta},X)}\leq C_{T}h^{\theta},\qquad nh\leq T,$
under assumptions (i) or (ii). Under assumption (iii), one obtains the weaker estimate with linear growth in $T$, as in the classical Lax-type accumulation mechanism.
Proposition 2.15 (First-order variant on a stronger space).
In addition to the assumptions of Theorem 2.14, assume that for some Banach space $X_{1}\hookrightarrow X$ the local defect satisfies
$\|R_{h}-P_{h}\|_{\mathcal{L}(X_{1},X)}\leq Ch.$
Then the corresponding global estimate holds on $X_{1}$ with order $h$ under the same discrete stability assumptions as in Theorem 2.14.
Theorem 2.16 (Abstract convergence principle: non-autonomous case).
Let $t_{k}=kh$ and consider families $\{R_{h}(t_{k})\}$ and $\{P_{h}(t_{k})\}$ of bounded operators. Define the time-ordered products
$R_{h}^{(n)}:=R_{h}(t_{n-1})\cdots R_{h}(t_{0}),\qquad P_{h}^{(n)}:=P_{h}(t_{n-1})\cdots P_{h}(t_{0}).$
Assume that:
  • (C1)
    (Uniform boundedness of the implicit products)
    $\sup_{nh\leq T}\|R_{h}^{(n)}\|_{\mathcal{L}(X)}<\infty.$
  • (C2)
    (Local defect) There exist $X_{\theta}\hookrightarrow X$ and $C>0$, $\theta\in(0,1]$ such that
    $\|R_{h}(t_{k})-P_{h}(t_{k})\|_{\mathcal{L}(X_{\theta},X)}\leq Ch^{\theta},\qquad\text{uniformly in }k.$
  • (C3)
    (Product stability of $\{P_{h}(t_{k})\}$) The family of products $P_{h}^{(n)}$} satisfies a stability condition strong enough to control the time-ordered telescoping sum.
Then there exists $C_{T}>0$ such that
$\|R_{h}^{(n)}-P_{h}^{(n)}\|_{\mathcal{L}(X_{\theta},X)}\leq C_{T}h^{\theta},\qquad nh\leq T.$
Remark 2.17 (Concrete sufficient stability conditions).
Typical sufficient conditions for (C3) include:
  • uniform summability of products,
  • bounded variation of time-ordered powers,
  • a suitable non-autonomous Ritt-type condition,
  • uniform exponential stability.
Each of these conditions allows control of the time-ordered telescoping sum and hence yields global convergence.
3 Non-sectorial case
In this section we show how the abstract discrete–resolvent framework developed in section 2 is concretely realized for delay equations in the non-sectorial setting. This case exhibits the strongest dependence on discrete stability and therefore provides the clearest illustration of the stability hierarchy underlying the convergence mechanism.
3.1 Autonomous non-sectorial case
We begin with the autonomous non-sectorial setting. The purpose of this section is twofold:
  • to derive sharp local defect estimates for Lie–Trotter splitting on fractional regularity spaces;
  • to demonstrate explicitly how the resulting global comparison depends on discrete stability properties of the splitting propagator.
Throughout this section we assume that the delay operator is autonomous, i.e. $D(t)\equiv D$.
3.1.1 Setting and one-step maps
Recall that the one-step operators are given by
$R_{h}:=(I-h(C+D))^{-1},\qquad S_{h}^{C}:=(I-hC)^{-1},\qquad H_{h}:=hDS_{h}^{C},\qquad P_{h}:=(I-H_{h})^{-1}S_{h}^{C},$
whenever these operators are well defined. The associated local resolvent defect is
$E_{h}:=R_{h}-P_{h}.$
We work on the fractional interpolation spaces
$X_{\theta}:=(\mathcal{E}_{0},\mathcal{D}(C))_{\theta,1},\qquad 0<\theta\leq 1.$
3.1.2 Technical toolbox: transport resolvent, trace control, and fractional smoothing
Lemma 3.1 (Transport resolvent bounds).
Let $C$ be the left-shift generator on $\mathcal{E}_{0}$. Then $S_{h}^{C}:=(I-hC)^{-1}\in\mathcal{L}(\mathcal{E}_{0})$ exists for every $h>0$ and
$\|S_{h}^{C}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq 1,\qquad\|CS_{h}^{C}\|_{\mathcal{L}(\mathcal{E}_{0})}\lesssim h^{-1}.$
Moreover, for every $0<\theta\leq 1$,
$\|S_{h}^{C}\|_{\mathcal{L}(X_{\theta},\mathcal{D}(C))}\lesssim h^{-\theta},\qquad X_{\theta}:=(\mathcal{E}_{0},\mathcal{D}(C))_{\theta,1}.$
Sectorial resolvent bounds.
This lemma is only needed later when the present-state component contains a sectorial block $A_{0}=\mathrm{diag}(A,0)$ and one uses the combined preconditioner $S_{h}^{AC}$ instead of $S_{h}^{C}$.
Lemma 3.2 (Sectorial resolvent bounds (when $A_{0}$ is present)).
Assume that $A_{0}=\mathrm{diag}(A,0)$ is sectorial on $\mathcal{E}_{0}$ with $0\in\rho(A_{0})$. Then $S_{h}^{A}:=(I-hA_{0})^{-1}$ exists for all $h>0$ and
$\|S_{h}^{A}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M_{A},\qquad\|A_{0}S_{h}^{A}\|_{\mathcal{L}(\mathcal{E}_{0})}\lesssim h^{-1}.$
Consequently, the combined preconditioner $S_{h}^{AC}:=S_{h}^{C}S_{h}^{A}$ satisfies
$\|S_{h}^{AC}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M_{A},\qquad\|S_{h}^{AC}\|_{\mathcal{L}(X_{\theta},\mathcal{D}(C))}\lesssim h^{-\theta},\quad 0<\theta\leq 1.$
Trace control on the fractional scale.
A central difficulty in delay equations on product spaces is that the point evaluation $\Phi_{\tau}$ is not bounded on $L^{p}([\tau,0];X)$. The next lemma shows that applying the transport resolvent $S_{h}^{C}$ regularizes the history component and yields a usable trace bound.
Lemma 3.3 (Trace bound through the transport resolvent $S_{h}^{C}$).
Let $\tau<0$, let $1<p<\infty$, and let $X$ be a Banach space. Set
$\mathcal{E}_{0}:=X\oplus L^{p}([\tau,0];X),\qquad\|(f,g)\|_{\mathcal{E}_{0}}:=\|f\|_{X}+\|g\|_{L^{p}([\tau,0];X)}.$
Let $C$ be the transport (left-shift) operator on $\mathcal{E}_{0}$,
$C(u,\rho):=(0,\partial_{\sigma}\rho),\qquad\mathcal{D}(C):=\{(u,\rho)\in X\oplus W^{1,p}([\tau,0];X):\rho(0)=u\}.$
For $h>0$ denote $S_{h}^{C}:=(I-hC)^{-1}\in\mathcal{L}(\mathcal{E}_{0})$ and the trace $\Phi_{\tau}(u,\rho):=\rho(\tau)\in X$.
Then for every $y=(f,g)\in\mathcal{E}_{0}$ the element $x=(u,\rho):=S_{h}^{C}y\in\mathcal{D}(C)$ is given by
$u=f,\qquad\rho(\sigma)=e^{\sigma/h}f+\frac{1}{h}\int_{\sigma}^{0}e^{(\sigma-s)/h}\,g(s)\,ds,\qquad\sigma\in[\tau,0].$ (3.1)
In particular, there exists a constant $c=c(\tau,p)\geq 1$ such that for all $h>0$,
$\|\Phi_{\tau}S_{h}^{C}(f,g)\|_{X}=\|\rho(\tau)\|_{X}\leq\|f\|_{X}+\frac{c}{h}\,\|g\|_{L^{p}([\tau,0];X)}.$ (3.2)
Equivalently,
$\|\Phi_{\tau}S_{h}^{C}\|_{\mathcal{L}(\mathcal{E}_{0},X)}\leq 1+\frac{c}{h}.$ (3.3)
Proof.
Let $y=(f,g)\in\mathcal{E}_{0}$ and write $x=(u,\rho)=S_{h}^{C}y\in\mathcal{D}(C)$, i.e. $(I-hC)(u,\rho)=(f,g)$. Then $u=f$ and $\rho-h\rho^{\prime}=g$ on $\tau,0]$ with boundary condition $\rho(0)=u=f$. Solving this first-order equation yields (3.1).
Evaluating (3.1) at $\sigma=\tau$ gives
$\rho(\tau)=e^{\tau/h}f+\frac{1}{h}\int_{\tau}^{0}e^{(\tau-s)/h}\,g(s)\,ds.$
Since $e^{\tau/h}\leq 1$, we have $\|e^{\tau/h}f\|_{X}\leq\|f\|_{X}$. For the integral term, use Hölder’s inequality:
$\left\|\int_{\tau}^{0}e^{(\tau-s)/h}g(s)\,ds\right\|_{X}\leq\|e^{(\tau-\cdot)/h}\|_{L^{p^{\prime}}([\tau,0])}\,\|g\|_{L^{p}([\tau,0];X)},$
where $p^{\prime}$ is the conjugate exponent. A direct computation shows that $\|e^{(\tau-\cdot)/h}\|_{L^{p^{\prime}}([\tau,0])}\leq c(\tau,p)\,h^{1/p^{\prime}}$ for some constant $c(\tau,p)$ independent of $h$. Multiplying by $h^{-1}$ yields (3.2), and (3.3) follows immediately. ∎
Corollary 3.4 (Concrete bound for delay operators after preconditioning).
Let $D:\mathcal{D}(C)\to\mathcal{E}_{0}$ be a delay operator of the form
$D(u,\rho)=(au+b\,\rho(\tau),\,0),\qquad a,b\in\mathbb{R}.$
Then there exists $c=c(\tau,p)\geq 1$ such that for all $h>0$ and all $(f,g)\in\mathcal{E}_{0}$,
$\|DS_{h}^{C}(f,g)\|_{\mathcal{E}_{0}}\leq|a|\,\|f\|_{X}+|b|\left(\|f\|_{X}+\frac{c}{h}\|g\|_{L^{p}([\tau,0];X)}\right).$
3.1.3 Operational smallness and well-definedness of the splitting step
Assumption 3.5 (Operational smallness (OS)).
There exist $h_{0}>0$ and $q\in(0,1)$ such that
$\sup_{0<h\leq h_{0}}\|H_{h}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq q.$
Under (OS) the inverse $(I-H_{h})^{-1}$ exists and satisfies
$\sup_{0<h\leq h_{0}}\|(I-H_{h})^{-1}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq\frac{1}{1-q},$
so that $P_{h}$ is well defined and uniformly bounded on $\mathcal{E}_{0}$.
Remark 3.6 (Verification of (OS) under relative bound 0).
If $D:\mathcal{D}(C)\to\mathcal{E}_{0}$ is CC–bounded with relative bound 0, then (OS) is not an extra hypothesis. Indeed, for every $\varepsilon>0$ there exists $b(\varepsilon)\geq 0$ such that
$\|Dx\|\leq\varepsilon\|Cx\|+b(\varepsilon)\|x\|,\qquad x\in\mathcal{D}(C),$
and therefore, using Lemma 3.1,
$\|hD(I-hC)^{-1}\|\leq h\varepsilon\|C(I-hC)^{-1}\|+h\,b(\varepsilon)\|(I-hC)^{-1}\|\leq\varepsilon+O(h)\quad(h\downarrow 0).$
Choosing first $\varepsilon>0$ and then $h>0$ small yields $\|H_{h}\|\leq q<1$.
3.1.4 Local resolvent defect (autonomous)
Lemma 3.7 (Right-preconditioned resolvent identity).
Assume (OS) and that $R_{h}\in\mathcal{L}(\mathcal{E}_{0})$ exists. Then
$R_{h}-P_{h}=R_{h}H_{h}-(I-H_{h})^{-1}H_{h}S_{h}^{C}.$
Proof.
Rewrite $R_{h}=(I-hC-hD)^{-1}$ and compare with $P_{h}=(I-hDS_{h}^{C})^{-1}S_{h}^{C}$, where $S_{h}^{C}=(I-hC)^{-1}$. ∎
Proposition 3.8 (Local defect on $\mathcal{D}(C)$ (first order)).
Assume (OS) and that $R_{h}$ exists with $\sup_{0<h\leq h_{0}}\|R_{h}\|\leq M_{R}$. If in addition $D:\mathcal{D}(C)\to\mathcal{E}_{0}$ is CC–bounded with relative bound 0, then there exists $C>0$ such that
$\|R_{h}-P_{h}\|_{\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})}\leq Ch,\qquad 0<h\leq h_{0}.$
Proof.
By Lemma 3.7,
$\|R_{h}-P_{h}\|_{\mathcal{D}(C)\to\mathcal{E}_{0}}\leq\Bigl(\|R_{h}\|+\|(I-H_{h})^{-1}\|\,\|S_{h}^{C}\|\Bigr)\,\|H_{h}\|_{\mathcal{D}(C)\to\mathcal{E}_{0}}.$
Under (OS) we have $\|(I-H_{h})^{-1}\|\leq(1-q)^{-1}$ and $\|S_{h}^{C}\|\leq 1$. Moreover, relative bound 0 implies $\|hDS_{h}^{C}\|_{\mathcal{D}(C)\to\mathcal{E}_{0}}\lesssim h$. ∎
Proposition 3.9 (Local defect on $X_{\theta}$ (fractional scale)).
Assume (OS) and that $R_{h}$ exists with $\sup_{0<h\leq h_{0}}\|R_{h}\|\leq M_{R}$. Let $0<\theta<1$ and set $X_{\theta}=(\mathcal{E}_{0},\mathcal{D}(C))_{\theta,1}$. Then there exists $C_{\theta}>0$ such that
$\|R_{h}-P_{h}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{\theta}\,h^{1-\theta},\qquad 0<h\leq h_{0}.$
Proof.
From Lemma 3.7 it suffices to bound $\|H_{h}\|_{X_{\theta}\to\mathcal{E}_{0}}$. Using a graph-norm control $\|Dy\|\lesssim\|y\|+\|Cy\|$ on $\mathcal{D}(C)$ and applying it to $y=S_{h}^{C}x$ gives
$\|H_{h}x\|=h\|DS_{h}^{C}x\|\lesssim h\|S_{h}^{C}x\|+h\|CS_{h}^{C}x\|.$
Now $\|S_{h}^{C}\|_{X_{\theta}\to\mathcal{E}_{0}}\lesssim 1$ and $\|CS_{h}^{C}\|_{X_{\theta}\to\mathcal{E}_{0}}\lesssim h^{-\theta}$ by Lemma 3.1. Hence $\|H_{h}\|_{X_{\theta}\to\mathcal{E}_{0}}\lesssim h+h\cdot h^{-\theta}\lesssim h^{1-\theta}$. ∎
3.1.5 Global comparison via telescoping (autonomous)
For $n\in\mathbb{N}$ we use the telescoping identity
$R_{h}^{n}-P_{h}^{n}=\sum_{k=0}^{n-1}R_{h}^{\,n-1-k}\,(R_{h}-P_{h})\,P_{h}^{k}.$ (3.4)
To avoid a factor $n\sim T/h$ we require discrete stability information on $\{P_{h}\}$ that controls defect propagation.
Assumption 3.10 (Summability / bounded partial sums).
There exists $M_{\mathrm{sum}}\geq 1$ such that for all $0<h\leq h_{0}$ and $N\geq 1$,
$\Bigl\|\sum_{k=0}^{N-1}P_{h}^{k}\Bigr\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M_{\mathrm{sum}}.$
Proposition 3.11 (Global error under summability (autonomous)).
Assume (OS), $\sup_{0<h\leq h_{0}}\|R_{h}\|\leq M_{R}$, and Assumption 3.10. Then for every $T>0$ there exists $C_{T}>0$ such that
$\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})}\leq C_{T}\,h,\qquad nh\leq T,$
and for every $0<\theta<1$,
$\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{T}\,h^{1-\theta},\qquad nh\leq T.$
Proof.
Let $E_{h}:=R_{h}-P_{h}$. By (3.4),
$R_{h}^{n}-P_{h}^{n}=\sum_{k=0}^{n-1}R_{h}^{n-1-k}E_{h}P_{h}^{k}.$
Using boundedness of $R_{h}$ and summability,
$\Bigl\|\sum_{k=0}^{n-1}E_{h}P_{h}^{k}\Bigr\|=\|E_{h}\sum_{k=0}^{n-1}P_{h}^{k}\|\leq M_{\mathrm{sum}}\|E_{h}\|.$
Insert Proposition 3.8 or Proposition 3.9. ∎
Proposition 3.12 (Global bound under mere power-boundedness (autonomous)).
Assume (OS), $\sup_{0<h\leq h_{0}}\|R_{h}\|\leq M_{R}$, and $\sup_{0<h\leq h_{0}}\sup_{k\geq 0}\|P_{h}^{k}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M_{P}$. Then for $nh\leq T$,
$\|R_{h}^{n}-P_{h}^{n}\|\leq C\,n\,\|R_{h}-P_{h}\|,$
and hence one generally loses uniform convergence on fractional scales.
Proof.
Estimate (3.4) by $\|R_{h}^{n-1-k}\|\leq M_{R}$ and $\|P_{h}^{k}\|\leq M_{P}$. ∎
3.1.6 Main theorem (autonomous non-sectorial)
Theorem 3.13 (Main Theorem A (autonomous non-sectorial case)).
Consider the autonomous delay equation on $\mathcal{E}_{0}$ with operators $C$ and $D$. For $0<h\leq h_{0}$, define
$P_{h}:=\bigl(I-hD(I-hC)^{-1}\bigr)^{-1}(I-hC)^{-1},\qquad R_{h}:=(I-h(C+D))^{-1},$
whenever they exist. Assume:
  1. (A1)
    (OS) holds (Assumption 3.5).
  2. (A2)
    For every $T>0$ there exists $M_{T}$ such that $\sup_{0<h\leq h_{0}}\sup_{n:nh\leq T}\|R_{h}^{n}\|\leq M_{T}$.
  3. (A3)
    $D:\mathcal{D}(C)\to\mathcal{E}_{0}$ is CC–bounded with relative bound 0.
  4. (A4)
    A strong discrete stability condition holds for $\{P_{h}\}$ (e.g. summability of powers as in Assumption 3.10).
Then for every $T>0$ there exists $C_{T}>0$ such that, for $nh\leq T$,
$\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{D}(C)\to\mathcal{E}_{0}}\leq C_{T}h,\qquad\|R_{h}^{n}-P_{h}^{n}\|_{X_{\theta}\to\mathcal{E}_{0}}\leq C_{T}h^{1-\theta}\quad(0<\theta<1).$
3.2 Non-autonomous non-sectorial case
In this section we extend the discrete resolvent comparison to the non-autonomous setting. The key new difficulty is that nn–step propagators are time-ordered products, so global control requires product stability assumptions as in the abstract non-autonomous theory.
3.2.1 Frozen one-step maps and goal
Fix $T>0$ and a stepsize $h>0$ with grid $t_{k}:=kh$. We compare the frozen implicit Euler resolvents
$R_{h}(t):=(I-h(C+D(t)))^{-1}$
with the frozen Lie–Trotter splitting steps
$P_{h}(t):=(I-H_{h}(t))^{-1}S_{h}^{C},\qquad H_{h}(t):=hD(t)S_{h}^{C},\qquad S_{h}^{C}:=(I-hC)^{-1}.$
Our aim is to verify the hypotheses of the abstract non-autonomous convergence principle (Theorem 2.16 in Section 2) for this concrete non-sectorial model.
3.2.2 Uniform frozen-step defect
Let $0<\theta<1$ and set $\alpha:=1-\theta$. Write $X_{\alpha}=(\mathcal{E}_{0},\mathcal{D}(C))_{\alpha,1}$.
Lemma 3.14 (Fractional bound for $H_{h}(t)=hD(t)S_{h}^{C}$).
Assume $D(t)\in\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})$ uniformly in $t\in[0,T]$. Then there exists $C_{\theta}>0$ such that for all sufficiently small $h$,
$\sup_{t\in[0,T]}\|H_{h}(t)\|_{\mathcal{L}(X_{1-\theta},\mathcal{E}_{0})}\leq C_{\theta}\,h^{\theta}.$
Proof.
Use $\|S_{h}^{C}\|_{X_{1-\theta}\to\mathcal{D}(C)}\lesssim h^{-(1-\theta)}$ from Lemma 3.1. ∎
Proposition 3.15 (Uniform frozen-step local defect on $X_{1-\theta}$).
Fix $0<\theta<1$. Assume a uniform operational smallness condition on $[0,T]$: there exist $h_{0}>0$ and $q\in(0,1)$ such that
$\sup_{t\in[0,T]}\sup_{0<h\leq h_{0}}\|H_{h}(t)\|\leq q,$
and assume the frozen resolvents exist with $\sup_{t\in[0,T]}\sup_{0<h\leq h_{0}}\|R_{h}(t)\|<\infty$. Then there exists $C_{T,\theta}>0$ such that
$\|R_{h}(t)-P_{h}(t)\|_{\mathcal{L}(X_{1-\theta},\mathcal{E}_{0})}\leq C_{T,\theta}\,h^{\theta},\qquad t\in[0,T],\ 0<h\leq h_{0}.$
Proof.
As in the autonomous case, use the right-preconditioned identity $R_{h}(t)-P_{h}(t)=R_{h}(t)H_{h}(t)-(I-H_{h}(t))^{-1}H_{h}(t)S_{h}^{C}$ and Lemma 3.14. ∎
3.2.3 Why frozen analysis is insufficient and the main non-autonomous result
Even though Proposition 3.15 is uniform in $t$, it does not by itself imply global convergence: the nn–step propagators are time-ordered products. The global difference expands into a non-commutative telescoping sum, so one needs product-stability assumptions as in the abstract theory.
Theorem 3.16 (Main Theorem C (non-sectorial, non-autonomous case)).
Let $D(\cdot)$ be time-dependent and consider the frozen one-step maps
$R_{h}(t):=(I-h(C+D(t)))^{-1},\qquad P_{h}(t):=(I-hD(t)S_{h}^{C})^{-1}S_{h}^{C},\qquad S_{h}^{C}:=(I-hC)^{-1}.$
Fix $T>0$ and set $t_{k}:=kh$. Assume the hypotheses of the abstract non-autonomous non-sectorial principle Theorem 2.16, i.e.
  • (C1)
    uniform boundedness of the implicit products $\|R_{h}(t_{n-1})\cdots R_{h}(t_{0})\|$ for $nh\leq T$;
  • (C2)
    a uniform frozen-step defect bound $\|R_{h}(t_{k})-P_{h}(t_{k})\|_{X_{\theta}\to\mathcal{E}_{0}}\leq Ch^{\theta}$ on $[0,T]$;
  • (C3)
    a uniform non-autonomous product stability condition for $\{P_{h}(t_{k})\}$ as required in Theorem 2.16.
Then there exists $C_{T}>0$ such that for all $x\in X_{\theta}$ and $nh\leq T$,
$\left\|R_{h}(t_{n-1})\cdots R_{h}(t_{0})x-P_{h}(t_{n-1})\cdots P_{h}(t_{0})x\right\|_{\mathcal{E}_{0}}\leq C_{T}h^{\theta}\|x\|_{X_{\theta}}.$
3.3 Connection to continuous theory
This section is discrete by design and does not assume continuous well-posedness. When a $C_{0}$-semigroup or evolution family exists, the discrete results are compatible with the classical product approximation theory (e.g. implicit Euler/Chernoff). One may keep the corresponding propositions in Section 2 or place them here if one wants a self-contained narrative.
3.3.1 Connection to evolution families for frozen implicit Euler products
Under classical Kato–Tanabe or Acquistapace–Terreni type assumptions, existence of an evolution family is guaranteed; see, e.g., (Pazy, 1983, Ch. 5), Tanabe (1979), Acquistapace and Terreni (1990).
Proposition 3.17 (Evolution family connection for frozen implicit Euler products).
Assume that $L(t):=C+D(t)$ satisfies classical non-autonomous generation hypotheses (e.g. of Kato–Tanabe or Acquistapace–Terreni type), so that there exists a strongly continuous evolution family $U(t,s)$ on $\mathcal{E}_{0}$ generated by $L(t)$. Existence of the evolution family under such assumptions is classical; see, e.g., (Pazy, 1983, Ch. 5), Tanabe (1979), Acquistapace and Terreni (1990).
Let $t_{k}:=kh$, and define the frozen implicit Euler resolvents
$R_{h}(t_{k}):=(I-hL(t_{k}))^{-1},\qquad 0\leq k\leq n-1,\quad t_{n}\leq T.$
Then, under the standard regularity assumptions ensuring first-order convergence of the implicit Euler product approximation, there exists $C_{T}>0$ such that
$\Bigl\|U(t_{n},0)x-\prod_{j=0}^{n-1}R_{h}(t_{j})x\Bigr\|\leq C_{T}\,h\,\|x\|_{Y},\qquad x\in Y,$
for a suitable regularity space $Y$ (for instance $Y=\mathcal{D}(L(0))$ or a graph/interpolation space, depending on the chosen non-autonomous framework).
Proof.
The convergence of frozen implicit Euler products toward the evolution family is well known; see, e.g., Tanaka (1994); Kobayashi (2003, 2005). ∎
4 Sectorial case
4.1 Discrete resolvents and splitting propagators for the three-factor model
We consider the abstract delay system formulated on the product space $\mathcal{E}_{0}$ with operator decomposition
$L(t)=A_{0}+C+D(t),$
where
$A_{0}=\begin{pmatrix}A&0\\ 0&0\end{pmatrix}\quad\text{(sectorial diffusion part)},\qquad C=\begin{pmatrix}0&0\\ 0&\partial_{\sigma}\end{pmatrix}\quad\text{(transport/shift)},$
and $D(t)$ denotes the bounded delay/reaction operator acting on $\mathcal{D}(C)$. In the autonomous case, $D(t)\equiv D$ is independent of time.
Implicit Euler resolvent.
For each time $t$ and stepsize $h>0$, we define the frozen implicit Euler resolvent
$R_{h}(t):=(I-hL(t))^{-1}=(I-h(A_{0}+C+D(t)))^{-1},$
whenever it exists as a bounded operator on $\mathcal{E}_{0}$. In the autonomous case, we simply write
$R_{h}:=(I-h(A_{0}+C+D))^{-1}.$
Two-factor preconditioner.
We introduce the resolvents associated with the diffusion and transport parts:
$S_{h}^{A_{0}}:=(I-hA_{0})^{-1},\qquad S_{h}^{C}:=(I-hC)^{-1},$
and define the combined (right) preconditioner
$S_{h}^{AC}:=S_{h}^{C}S_{h}^{A_{0}}=(I-hC)^{-1}(I-hA_{0})^{-1}.$
Right-preconditioned delay operator.
Since $D(t)$ is not bounded on $\mathcal{E}_{0}$ but only defined on $\mathcal{D}(C)$, we introduce the right-preconditioned operator
$H_{h}(t):=h\,D(t)S_{h}^{AC}.$
In the autonomous case, this reduces to
$H_{h}:=h\,DS_{h}^{AC}.$
Lie–Trotter splitting propagator.
The one-step splitting map is defined by
$P_{h}(t):=(I-H_{h}(t))^{-1}S_{h}^{AC}=(I-hD(t)S_{h}^{AC})^{-1}S_{h}^{AC},$
whenever $(I-H_{h}(t))^{-1}$ exists as a bounded operator on $\mathcal{E}_{0}$. In the autonomous case, we simply write
$P_{h}:=(I-hDS_{h}^{AC})^{-1}S_{h}^{AC}.$
Local resolvent defect.
The fundamental quantity governing convergence is the local defect between implicit Euler and splitting:
$E_{h}(t):=R_{h}(t)-P_{h}(t)\qquad\text{(non-autonomous case)},$
and in the autonomous case
$E_{h}:=R_{h}-P_{h}.$
The subsequent analysis is devoted to estimating the size of $E_{h}(t)$ in appropriate operator norms and to propagating this local defect to global error bounds under suitable discrete stability assumptions.
Proposition 4.1 (Local defect for the three-factor model: non-sectorial loss $h^{1-\theta}$).
Let $\mathcal{E}_{0}$ and $C$ be as in Section 1, and let $A_{0}$ be sectorial on the uu–component. For $t\in[0,T]$ and $h>0$, define
$G:=A_{0}+C,\qquad S_{h}^{AC}:=(I-hC)^{-1}(I-hA_{0})^{-1},\qquad L(t):=G+D(t).$
Assume that the exact two-factor resolvent identity holds:
$(I-hG)^{-1}=S_{h}^{AC}\qquad\text{on }\mathcal{E}_{0}.$ (4.1)
Define the frozen implicit Euler resolvent and the Lie–Trotter splitting step by
$R_{h}(t):=(I-hL(t))^{-1}=(I-h(G+D(t)))^{-1},\qquad P_{h}(t):=(I-hD(t)S_{h}^{AC})^{-1}S_{h}^{AC}.$
Let $X_{\theta}:=(\mathcal{E}_{0},\mathcal{D}(C))_{\theta,1}$ for $0<\theta<1$.
Assume moreover:
  1. (OS)
    (Operational smallness) There exist $h_{0}>0$ and $q\in(0,1)$ such that
    $\sup_{t\in[0,T]}\sup_{0<h\leq h_{0}}\|hD(t)S_{h}^{AC}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq q.$
  2. (D-bdd)
    $D(t)\in\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})$ uniformly in $t$:
    $\sup_{t\in[0,T]}\|D(t)\|_{\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})}\leq M_{D}.$
  3. (C-smooth)
    (transport resolvent smoothing) for $0<h\leq h_{0}$,
    $\|S_{h}^{C}\|_{\mathcal{L}(X_{\theta},\mathcal{D}(C))}\lesssim h^{-\theta}.$
    (Equivalently, $\|S_{h}^{AC}\|_{\mathcal{L}(X_{\theta},\mathcal{D}(C))}\lesssim h^{-\theta}$, since $S_{h}^{A_{0}}$ is bounded on $\mathcal{E}_{0}$.)
Then there exists $C_{T,\theta}>0$ such that for all $t\in[0,T]$ and $0<h\leq h_{0}$,
$\|E_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}:=\|R_{h}(t)-P_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{T,\theta}\,h^{1-\theta}.$
In particular, the local defect exhibits the non-sectorial loss $h^{1-\theta}$ coming from the transport regularization.
Proof.
Fix $t\in[0,T]$ and abbreviate $D:=D(t)$. By (4.1), we have the right-preconditioned factorization
$I-h(G+D)=(I-hG)\,(I-hD(I-hG)^{-1})=(I-hG)\,(I-hDS_{h}^{AC}).$ (4.2)
Since $(I-hG)^{-1}=S_{h}^{AC}$ is bounded on $\mathcal{E}_{0}$, the identity (4.2) implies
$R_{h}(t)=(I-h(G+D))^{-1}=(I-hDS_{h}^{AC})^{-1}S_{h}^{AC}.$ (4.3)
But the right-hand side is exactly the definition of $P_{h}(t)$. Hence,
$E_{h}(t)=R_{h}(t)-P_{h}(t)=0\qquad\text{as operators on }\mathcal{E}_{0}.$
To obtain a meaningful defect estimate (and to recover the non-sectorial order $h^{1-\theta}$), one must distinguish implemented three-factor preconditioners from the exact two-factor resolvent. In the present proposition, however, we are assuming the exact identity (4.1); therefore the splitting step coincides with the implicit Euler resolvent and the defect vanishes. ∎
4.2 Autonomous sectorial case
This section instantiates the discrete-resolvent framework of Section 1 in the sectorial setting and aligns it with the right-preconditioned three-factor splitting introduced there. The key point is that the delay/reaction operator $D$ need not be bounded on $\mathcal{E}_{0}$, so the naive resolvent $(I-hD)^{-1}$ is not available. All discrete propagators are therefore defined via right-preconditioned resolvents, which ensures well-posedness of the splitting step and is compatible with the smoothing mechanism generated by the sectorial part; see, e.g., Pazy (1983); Engel and Nagel (1999); Bátkai and Piazzera (2005).
Throughout this section we assume that $A_{0}$ generates a bounded analytic semigroup on $\mathcal{E}_{0}$. We write
$S_{h}^{A_{0}}:=(I-hA_{0})^{-1},\qquad S_{h}^{C}:=(I-hC)^{-1},\qquad S_{h}^{AC}:=S_{h}^{C}S_{h}^{A_{0}}.$
4.2.1 Right-preconditioned three-factor splitting
Let $D:\mathcal{D}(C)\to\mathcal{E}_{0}$ denote the delay/reaction part. We impose the following discrete smallness condition, which guarantees that the splitting step is well-defined.
Assumption 4.2 (Sectorial operational smallness).
There exist $h_{0}>0$ and $q\in(0,1)$ such that
$\sup_{0<h\leq h_{0}}\ \bigl\|\,h\,D\,S_{h}^{AC}\bigr\|_{\mathcal{L}(\mathcal{E}_{0})}\leq q.$
Under Assumption 4.2, the Lie–Trotter one-step map is defined by
$P_{h}:=(I-hDS_{h}^{AC})^{-1}S_{h}^{AC}\ \in\ \mathcal{L}(\mathcal{E}_{0}).$
Lemma 4.3 (Well-posedness and uniform boundedness of the three-factor step).
Assume (OS-sectorial). Then for every $0<h\leq h_{0}$ the inverse $(I-hDS_{h}^{AC})^{-1}$ exists in $\mathcal{L}(\mathcal{E}_{0})$ and admits the Neumann series representation
$(I-hDS_{h}^{AC})^{-1}=\sum_{m=0}^{\infty}(hDS_{h}^{AC})^{m},$
with convergence in operator norm. Moreover,
$\sup_{0<h\leq h_{0}}\ \|(I-hDS_{h}^{AC})^{-1}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq\frac{1}{1-q},\qquad\sup_{0<h\leq h_{0}}\ \|P_{h}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq\frac{\sup_{0<h\leq h_{0}}\|S_{h}^{AC}\|_{\mathcal{L}(\mathcal{E}_{0})}}{1-q}.$
Proof.
Fix $0<h\leq h_{0}$. By (OS-sectorial), $\|hDS_{h}^{AC}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq q<1$, hence the Neumann series $\sum_{m\geq 0}(hDS_{h}^{AC})^{m}$ converges in operator norm to $(I-hDS_{h}^{AC})^{-1}$ and
$\|(I-hDS_{h}^{AC})^{-1}\|\leq\sum_{m=0}^{\infty}\|hDS_{h}^{AC}\|^{m}\leq\sum_{m=0}^{\infty}q^{m}=\frac{1}{1-q}.$
The bound for $P_{h}=(I-hDS_{h}^{AC})^{-1}S_{h}^{AC}$ follows immediately. ∎
4.2.2 Implicit Euler resolvent and reference propagator
The one-step implicit Euler map for the full operator $A_{0}+C+D$ is
$R_{h}:=(I-h(A_{0}+C+D))^{-1},$
whenever this inverse exists. We view $R_{h}$ as the reference discrete propagator and measure the splitting error through the local resolvent defect
$E_{h}:=R_{h}-P_{h}.$
4.2.3 Smoothing-compatible local defect
The three-factor structure is designed to respect the block smoothing induced by the sectorial part. In the block-decoupled product-space formulation of delay equations, the resolvent of $A_{0}+C$ factorizes exactly, so that no additional consistency error is introduced at the $A_{0}+C$ level.
Proposition 4.4 (Exact one–sided factorization of the $(A_{0}+C)$–resolvent).
Let $\mathcal{E}_{0}:=X\oplus Y$ be a Banach product space. Assume that
$A_{0}=\begin{pmatrix}A&0\\ 0&0\end{pmatrix},\qquad\mathcal{D}(A_{0})=\mathcal{D}(A)\oplus Y,$
where $A:\mathcal{D}(A)\subset X\to X$ is closed. Let $C:\mathcal{D}(C)\subset\mathcal{E}_{0}\to\mathcal{E}_{0}$ be a closed operator of the form
$C(u,y)=(0,C_{Y}y),$
with some closed $C_{Y}:\mathcal{D}(C_{Y})\subset Y\to Y$, and assume that
$\mathcal{D}(C)\subset X\oplus\mathcal{D}(C_{Y})\quad\text{and}\quad C(\mathcal{D}(C))\subset\mathcal{D}(A_{0}).$ (4.4)
(For instance, (4.4) holds in the delay product–space setting, where $C$ is the transport operator on the history component.)
Fix $h>0$ and assume that the resolvents
$(I-hA_{0})^{-1}\in\mathcal{L}(\mathcal{E}_{0}),\qquad(I-hC)^{-1}\in\mathcal{L}(\mathcal{E}_{0})$
exist. Then the resolvent of $G:=A_{0}+C$ exists and satisfies the exact one–sided factorization
$(I-h(A_{0}+C))^{-1}=(I-hC)^{-1}(I-hA_{0})^{-1}.$ (4.5)
In general, the reversed order need not hold unless one additionally has $(Au,0)\in\mathcal{D}(C)$ for all $u\in\mathcal{D}(A)$ (e.g. $\mathcal{D}(C)=X\oplus\mathcal{D}(C_{Y})$).
Proof.
Set $G:=A_{0}+C$ and let $x\in\mathcal{D}(C)$. By (4.4) we have $Cx\in\mathcal{D}(A_{0})$ and hence the composition $A_{0}C$ is well-defined on $\mathcal{D}(C)$. Moreover, since $Cx=(0,C_{Y}y)$ has vanishing XX–component, we obtain
$A_{0}Cx=(A0,0)=(0,0),\qquad x\in\mathcal{D}(C).$
Consequently, for all $x\in\mathcal{D}(A_{0})\cap\mathcal{D}(C)$,
$(I-hA_{0})(I-hC)x=x-h(A_{0}+C)x=(I-h(A_{0}+C))x.$ (4.6)
Now fix $y\in\mathcal{E}_{0}$ and set
$x:=(I-hC)^{-1}(I-hA_{0})^{-1}y.$
Then $(I-hA_{0})^{-1}y\in\mathcal{D}(A_{0})$ and $x\in\mathcal{D}(C)$ by construction, so $x\in\mathcal{D}(A_{0})\cap\mathcal{D}(C)=\mathcal{D}(G)$. Using (4.6),
$(I-h(A_{0}+C))x=(I-hA_{0})(I-hC)x=(I-hA_{0})(I-hA_{0})^{-1}y=y.$
Thus $x=(I-h(A_{0}+C))^{-1}y$, proving (4.5). ∎
Remark 4.5 (Why the $A_{0}$$C$ splitting error vanishes in the block model).
Proposition 4.4 is an algebraic consequence of the block structure: $A_{0}$ acts only on the present-state component, while $C$ acts only on the history component. Hence the resolvent of $A_{0}+C$ factorizes exactly on $\mathcal{E}_{0}$, and any two-factor Lie–Trotter resolvent splitting between $A_{0}$ and $C$ produces no one-step defect.
Consequently, in the three-factor delay operator
$L(t)=A_{0}+C+D(t),$
the only source of splitting error in the right-preconditioned step
$P_{h}(t)=(I-hD(t)S_{h}^{AC})^{-1}S_{h}^{AC}$
is the coupling/delay operator $D(t)$ through its interaction with the preconditioner $S_{h}^{AC}$; there is no separate $A_{0}$$C$ commutator contribution.
Remark 4.6 (Why analytic perturbation does not apply to $A_{0}+C$).
One might wonder whether the analytic perturbation theorem (for instance, (Pazy, 1983, Thm. 2.1)) could be applied to conclude that $A_{0}+C$ generates an analytic semigroup whenever $A_{0}$ does. However, the assumptions of such results are not satisfied in the present product–space setting.
Indeed, analytic perturbation theorems typically require that the perturbation operator $C$ be relatively bounded with respect to $A_{0}$ on $D(A_{0})$ with sufficiently small relative bound. In our framework,
$A_{0}=\mathrm{diag}(A,0),\qquad C(u,\rho)=(0,\partial_{\sigma}\rho),$
the operator $A_{0}$ acts only on the present component $u$, whereas $C$ differentiates exclusively in the history component. Consequently, $C$ cannot be controlled in norm by $A_{0}$ on $D(A_{0})$, and in general $D(C)\not\supset D(A_{0})$. Hence $C$ fails to be $A_{0}$–bounded in the sense required by analytic perturbation theory.
This structural mismatch reflects the intrinsic anisotropy of the product space: analytic smoothing produced by $A_{0}$ on the present component does not propagate to the history component governed by the non–sectorial transport operator $C$. As a result, even when $A_{0}$ generates an analytic semigroup on $X$, the sum $A_{0}+C$ typically generates only a $C_{0}$–semigroup on the product space, and no global analytic smoothing can be expected.
Lemma 4.7 (Fractional size of the preconditioned delay operator (three-factor model)).
Let $0<\theta<1$ and $X_{\theta}:=(\mathcal{E}_{0},\mathcal{D}(C))_{\theta,1}$. Assume:
  • (i)
    $D(t)\in\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})$ uniformly for $t\in[0,T]$, i.e.
    $\sup_{t\in[0,T]}\|D(t)\|_{\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})}\leq M_{D};$
  • (ii)
    the transport resolvent satisfies
    $\|S_{h}^{C}\|_{\mathcal{L}(X_{\theta},\mathcal{D}(C))}\leq c_{\theta}\,h^{-\theta},\qquad S_{h}^{C}:=(I-hC)^{-1};$
  • (iii)
    $S_{h}^{A_{0}}:=(I-hA_{0})^{-1}$ is uniformly bounded on $\mathcal{E}_{0}$:
    $\sup_{0<h\leq h_{0}}\|S_{h}^{A_{0}}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq M_{A}.$
Define $S_{h}^{AC}:=S_{h}^{C}S_{h}^{A_{0}}$ and $H_{h}(t):=hD(t)S_{h}^{AC}$. Then there exists $C_{\theta}>0$ such that
$\sup_{t\in[0,T]}\|H_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{\theta}\,h^{1-\theta},\qquad 0<h\leq h_{0}.$
Proof.
Since $A_{0}$ acts only on the present-state block, $S_{h}^{A_{0}}$ preserves $X_{\theta}$ and is bounded there with the same bound as on $\mathcal{E}_{0}$. Hence, for $x\in X_{\theta}$ and $t\in[0,T]$,
$\|H_{h}(t)x\|_{\mathcal{E}_{0}}\leq h\,\|D(t)\|_{\mathcal{D}(C)\to\mathcal{E}_{0}}\,\|S_{h}^{C}\|_{X_{\theta}\to\mathcal{D}(C)}\,\|S_{h}^{A_{0}}\|_{\mathcal{E}_{0}\to\mathcal{E}_{0}}\,\|x\|_{X_{\theta}}\leq h\,M_{D}\,c_{\theta}h^{-\theta}\,M_{A}\,\|x\|_{X_{\theta}}.$
This yields the claim with $C_{\theta}:=M_{D}c_{\theta}M_{A}$. ∎
Proposition 4.8 (Local defect in the sectorial block three-factor model).
Let $0<\theta<1$ and $X_{\theta}=(\mathcal{E}_{0},\mathcal{D}(C))_{\theta,1}$. Assume the block setting of Proposition LABEL:prop:AC_resolvent_exact_factorization and define
$S_{h}^{AC}:=(I-hC)^{-1}(I-hA_{0})^{-1},\qquad H_{h}(t):=hD(t)S_{h}^{AC},\qquad P_{h}(t):=(I-H_{h}(t))^{-1}S_{h}^{AC}.$
Assume:
  • (i)
    (OS-sectorial) there exist $h_{0}>0$ and $q\in(0,1)$ such that
    $\sup_{t\in[0,T]}\sup_{0<h\leq h_{0}}\|H_{h}(t)\|_{\mathcal{L}(\mathcal{E}_{0})}\leq q;$
  • (ii)
    the implicit Euler resolvent exists and is uniformly bounded,
    $R_{h}(t):=(I-h(A_{0}+C+D(t)))^{-1}\in\mathcal{L}(\mathcal{E}_{0}),\qquad\sup_{t\in[0,T]}\sup_{0<h\leq h_{0}}\|R_{h}(t)\|\leq M_{R};$
  • (iii)
    the fractional bound of Lemma 4.7 holds: $\sup_{t\in[0,T]}\|H_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{\theta}h^{1-\theta}$.
Then there exists $C_{T,\theta}>0$ such that
$\|R_{h}(t)-P_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{T,\theta}\,h^{1-\theta},\qquad 0<h\leq h_{0},\ t\in[0,T].$
In particular, the local defect order is governed by the transport smoothing and does not improve automatically by the presence of the sectorial block $A_{0}$.
Proof.
Fix $t\in[0,T]$ and set $G:=A_{0}+C$. Abbreviate
$S:=(I-hG)^{-1}=S_{h}^{AC},\qquad H:=hD(t)S.$
By the (purely algebraic) right-preconditioned factorization, whenever the inverses exist,
$(I-h(G+D(t)))^{-1}=(I-hG)^{-1}\bigl(I-hD(t)(I-hG)^{-1}\bigr)^{-1}=S(I-H)^{-1}.$
Hence
$R_{h}(t)=S(I-H)^{-1},\qquad P_{h}(t)=(I-H)^{-1}S.$
Therefore,
$R_{h}(t)-P_{h}(t)=S(I-H)^{-1}-(I-H)^{-1}S=R_{h}(t)\,H\,S-(I-H)^{-1}H\,S,$
and taking norms on $\mathcal{L}(X_{\theta},\mathcal{E}_{0})$ yields
$\|R_{h}(t)-P_{h}(t)\|_{X_{\theta}\to\mathcal{E}_{0}}\leq\Bigl(\|R_{h}(t)\|+\|(I-H)^{-1}\|\Bigr)\,\|H\|_{X_{\theta}\to\mathcal{E}_{0}}\,\|S\|.$
Under (OS-sectorial), $\|(I-H)^{-1}\|\leq(1-q)^{-1}$ uniformly in $t$. Moreover $\|S\|\leq\|(I-hC)^{-1}\|\|(I-hA_{0})^{-1}\|\leq M_{A}$. Using $\|R_{h}(t)\|\leq M_{R}$ and $\|H\|_{X_{\theta}\to\mathcal{E}_{0}}\leq C_{\theta}h^{1-\theta}$ proves the claim. ∎
Proposition 4.9 (Sectorial-block three-factor model reduces to the non-sectorial local defect).
Let $\mathcal{E}_{0}$ be the product space and let $C:\mathcal{D}(C)\subset\mathcal{E}_{0}\to\mathcal{E}_{0}$ be the transport generator. Let
$A_{0}:=\begin{pmatrix}A&0\\ 0&0\end{pmatrix},\qquad L(t):=A_{0}+C+D(t),\qquad t\in[0,T],$
where $D(t):\mathcal{D}(C)\to\mathcal{E}_{0}$ is the delay/reaction operator (possibly time-dependent). For $h>0$ define
$S_{h}^{C}:=(I-hC)^{-1},\qquad S_{h}^{A_{0}}:=(I-hA_{0})^{-1},\qquad S_{h}^{AC}:=S_{h}^{C}S_{h}^{A_{0}}.$
Assume the block commutation $A_{0}C=CA_{0}$ and that $A_{0}$ acts only on the $u$-block while $C$ acts only on the history block (as in the standard product-space formulation). Then
$(I-h(A_{0}+C))^{-1}=S_{h}^{AC}=(I-hC)^{-1}(I-hA_{0})^{-1}.$ (4.7)
Fix $0<\theta<1$ and let $X_{\theta}:=(\mathcal{E}_{0},\mathcal{D}(C))_{\theta,1}$ be the real interpolation space. For $t\in[0,T]$ define the (frozen) implicit Euler resolvent and the Lie–Trotter splitting step by
$R_{h}(t):=(I-hL(t))^{-1}=(I-h(A_{0}+C+D(t)))^{-1},\qquad P_{h}(t):=(I-H_{h}(t))^{-1}S_{h}^{AC},$
with the right-preconditioned operator
$H_{h}(t):=hD(t)S_{h}^{AC}.$
Assume operational smallness (OS) (in the form of Assumption 2.1 with $S_{h}^{AC}$): there exist $h_{0}>0$ and $q\in(0,1)$ such that
$\sup_{t\in[0,T]}\sup_{0<h\leq h_{0}}\|H_{h}(t)\|_{\mathcal{L}(\mathcal{E}_{0})}\leq q.$
Assume furthermore that $D(t)$ is uniformly bounded on $\mathcal{D}(C)$:
$\sup_{t\in[0,T]}\|D(t)\|_{\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})}\leq M_{D},$
and that the transport smoothing estimate holds:
$\|S_{h}^{C}\|_{\mathcal{L}(X_{\theta},\mathcal{D}(C))}\lesssim h^{-\theta},\qquad 0<h\leq h_{0}.$
Then there exists $C_{\theta}>0$ such that for all $t\in[0,T]$ and $0<h\leq h_{0}$,
$\|R_{h}(t)-P_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{\theta}\,h^{1-\theta}.$ (4.8)
In particular, the local defect mechanism is identical to the non-sectorial two-factor theory: the rate $h^{1-\theta}$ is governed by the transport smoothing on $X_{\theta}$ and is not improved merely by the presence of the sectorial block $A_{0}$ in the three-factor model.
Proof.
Step 1 (Exact one-sided resolvent factorization for $A_{0}+C$). Let $y=(f,g)\in\mathcal{E}_{0}$ and set $x=(u,\rho):=(I-h(A_{0}+C))^{-1}y$. Then $(I-h(A_{0}+C))x=y$ expands to
$(I-hA)u=f,\qquad\rho-h\partial_{\sigma}\rho=g,\qquad\rho(0)=u.$
Hence $u=(I-hA)^{-1}f$ and $\rho$ is given by the transport resolvent with inflow $u$, i.e. $(u,\rho)=(I-hC)^{-1}(u,g)$. Therefore,
$(I-h(A_{0}+C))^{-1}(f,g)=(I-hC)^{-1}\bigl((I-hA)^{-1}f,\ g\bigr)=(I-hC)^{-1}(I-hA_{0})^{-1}(f,g),$
which proves (4.7).
Step 2 (Fractional bound for $H_{h}(t)$). By definition $S_{h}^{AC}=(I-h(A_{0}+C))^{-1}$ and hence
$\|S_{h}^{AC}\|_{\mathcal{L}(X_{\theta},\mathcal{D}(C))}\leq\|S_{h}^{C}\|_{\mathcal{L}(X_{\theta},\mathcal{D}(C))}\,\|S_{h}^{A_{0}}\|_{\mathcal{L}(\mathcal{E}_{0})}\lesssim h^{-\theta},$
where $\|S_{h}^{A_{0}}\|$ is uniformly bounded since $A_{0}=\mathrm{diag}(A,0)$. Therefore,
$\|H_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq h\,\|D(t)\|_{\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})}\,\|S_{h}^{AC}\|_{\mathcal{L}(X_{\theta},\mathcal{D}(C))}\lesssim h\cdot M_{D}\cdot h^{-\theta}\lesssim h^{1-\theta},$
uniformly in $t$.
Step 3 (Right-preconditioned defect identity and conclusion). Fix $t$ and abbreviate $S:=S_{h}^{AC}$ and $H:=H_{h}(t)=hD(t)S$. Whenever the inverses exist, the right-preconditioned factorization gives
$R_{h}(t)=(I-h(A_{0}+C+D(t)))^{-1}=S(I-H)^{-1},\qquad P_{h}(t)=(I-H)^{-1}S.$
Hence
$R_{h}(t)-P_{h}(t)=S(I-H)^{-1}-(I-H)^{-1}S=R_{h}(t)\,H\,S-(I-H)^{-1}H\,S.$
Taking norms on $\mathcal{L}(X_{\theta},\mathcal{E}_{0})$ yields
$\|R_{h}(t)-P_{h}(t)\|_{X_{\theta}\to\mathcal{E}_{0}}\leq\bigl(\|R_{h}(t)\|+\|(I-H)^{-1}\|\bigr)\,\|H\|_{X_{\theta}\to\mathcal{E}_{0}}\,\|S\|.$
Under (OS), $\|(I-H)^{-1}\|\leq(1-q)^{-1}$ uniformly in $t$, and consequently $\|R_{h}(t)\|=\|S(I-H)^{-1}\|\leq\|S\|(1-q)^{-1}$. Moreover $\|S\|$ is uniformly bounded on $\mathcal{E}_{0}$. Combining these uniform bounds with Step 2 gives
$\|R_{h}(t)-P_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\lesssim h^{1-\theta},$
which is (4.8). ∎
Proposition 4.10 (Global consequence via the abstract principles).
Assume the setting of Proposition 4.9 and let
$E_{h}(t):=R_{h}(t)-P_{h}(t).$
Then the local defect satisfies
$\|E_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\lesssim h^{1-\theta}$
uniformly on $[0,T]$.
Consequently,
  • in the autonomous case ($D(t)\equiv D$), global convergence on finite time intervals follows directly from Theorem 2.14;
  • in the non-autonomous case, global convergence follows from Theorem 2.16 under the corresponding product stability assumption.
4.2.4 Interpretation
The autonomous sectorial block model clarifies the respective roles of analytic smoothing and discrete stability in delay equations on product spaces. The sectorial part $A_{0}$ indeed provides analytic smoothing on the present-state component, in accordance with the classical theory of analytic semigroups and sectorial operators; see, e.g., Lunardi (1995); Haase (2006). However, in the three-factor formulation $\mathcal{L}=A_{0}+C+D$ the dominant contribution to the Lie–Trotter defect is typically governed by the non-sectorial transport dynamics generated by $C$ on the history component. As a consequence, analytic smoothing of $A_{0}$ alone does not automatically improve the local splitting order on fractional spaces: the characteristic transport-driven loss $h^{1-\theta}$ persists unless further structure is imposed on the coupling and the splitting design. Therefore, in this thesis, even for sectorial block-structured delay models, the abstract convergence principles of Main Theorems A and C are applied, rather than the sectorial-specific hierarchy.
Remark 4.11 (Finite-time convergence versus long-time behavior).
It is important to distinguish between two conceptually different questions.
(i) Finite-time convergence. Assume that the continuous operator $L$ generates a $C_{0}$-semigroup $(U(t))_{t\geq 0}$ on $\mathcal{E}_{0}$ and that the implicit Euler resolvent satisfies the classical convergence property
$R_{h}^{n}x\longrightarrow U(T)x\qquad\text{as }h\to 0,\ nh=T\ \text{fixed},$
which is well known for sectorial generators; see, e.g., (Pazy, 1983, Ch. 5), Kobayashi (2003, 2005). If, in addition, the splitting scheme satisfies a finite-time error estimate
$\|R_{h}^{n}-P_{h}^{n}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq C_{T}\,\varepsilon(h),\qquad nh\leq T,$
with $\varepsilon(h)\to 0$ as $h\to 0$ for each fixed $T>0$, then a direct application of the triangle inequality yields
$P_{h}^{n}x\longrightarrow U(T)x,\qquad h\to 0,\ nh=T,$
for every fixed $T>0$. Thus, convergence of Lie–Trotter splitting to the continuous solution is guaranteed on every finite time interval whenever both approximations are consistent with the same continuous limit.
(ii) Long-time behavior. A fundamentally different question is whether the numerical approximation remains accurate as $T\to\infty$. If the constant $C_{T}$ in the above estimate grows with $T$ (for instance $C_{T}\sim T^{\alpha}$), then the bound does not provide any uniform control in time, and no conclusion on the long-time behavior can be drawn. Uniform-in-time convergence or preservation of asymptotic properties (e.g. exponential stability, invariant sets, attractors) typically requires additional assumptions, such as strong stability, spectral gap conditions, or structure-preserving properties of the numerical scheme; see, e.g., Hundsdorfer and Verwer (2003); Blanes and Casas (2016).
In this thesis, we focus primarily on the first aspect: finite-time convergence of discrete propagators toward the implicit Euler scheme (and, whenever available, toward the underlying semigroup or evolution family). The investigation of long-time qualitative behavior constitutes a separate and substantially more delicate problem.
4.3 Non-autonomous sectorial case
We briefly discuss the non-autonomous extension of the sectorial block model. Although the present-state component is governed by a sectorial operator, the overall delay system retains a genuinely non-autonomous product structure, which plays a decisive role in the convergence analysis.
Throughout this section, we assume that the sectorial operator $A_{0}$ generates a bounded analytic semigroup on $\mathcal{E}_{0}$, and that $D(\cdot)$ satisfies the regularity assumptions imposed in Section 3 (e.g. boundedness and Lipschitz continuity in time with values in $\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})$).
4.3.1 Frozen discrete propagators
Fix a time step $h>0$ and let $t_{k}:=kh$. At each time level we define the frozen one-step splitting propagator by
$P_{h}(t_{k}):=(I-hD(t_{k})S_{h}^{AC})^{-1}S_{h}^{AC},$
and the corresponding frozen implicit Euler resolvent by
$R_{h}(t_{k}):=(I-h(A_{0}+C+D(t_{k})))^{-1},$
whenever these inverses exist. The associated frozen local defect is defined by
$E_{h}(t_{k}):=R_{h}(t_{k})-P_{h}(t_{k}).$
The $n$-step discrete propagators are given by the time-ordered products
$P_{h}^{(n)}:=P_{h}(t_{n-1})\cdots P_{h}(t_{0}),\qquad R_{h}^{(n)}:=R_{h}(t_{n-1})\cdots R_{h}(t_{0}).$
4.3.2 Uniform local defect estimate
The local defect estimate established in the autonomous case extends uniformly to the non-autonomous setting.
Proposition 4.12 (Uniform frozen local defect).
Assume that $D(\cdot)$ is bounded and Lipschitz continuous in time with values in $\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})$. Then for every $0<\theta<1$ there exist constants $C_{\theta}>0$ and $h_{1}>0$ such that
$\|R_{h}(t_{k})-P_{h}(t_{k})\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{\theta}h^{1-\theta},\qquad 0<h\leq h_{1},\quad t_{k}\in[0,T].$
Proof.
The proof follows the same right-preconditioned resolvent argument as in the autonomous case. The Lipschitz continuity of $D(t)$ ensures that all constants can be chosen uniformly in $t\in[0,T]$. ∎
4.3.3 Time-ordered products and loss of autonomous structure
The global error in the non-autonomous setting is governed by time-ordered products of frozen one-step propagators,
$R_{h}(t_{n-1})\cdots R_{h}(t_{0}),\qquad P_{h}(t_{n-1})\cdots P_{h}(t_{0}),$
which admit the non-commutative telescoping representation
$\sum_{j=0}^{n-1}R_{h}(t_{n-1})\cdots R_{h}(t_{j+1})\bigl(R_{h}(t_{j})-P_{h}(t_{j})\bigr)P_{h}(t_{j-1})\cdots P_{h}(t_{0}).$
In contrast to the autonomous case, these products cannot be reduced to powers of a single operator. In the autonomous setting one has, for example,
$P_{h}^{n}=(P_{h})^{n},$
so that discrete smoothing or stability properties of $P_{h}$ (such as sectorial resolvent bounds, Ritt conditions, or summability of powers) can be applied directly to the entire $n$-step propagator. This permits a global control mechanism in which local defects are suppressed uniformly by analytic or discrete smoothing.
In the non-autonomous situation, however, each factor $P_{h}(t_{k})$ depends on the time level $t_{k}$. The $n$-step propagator is therefore a genuinely non-commutative product of different operators, and no functional calculus or power estimate for a single operator is available. Even if each frozen operator $P_{h}(t_{k})$ individually contains a sectorial component and thus enjoys analytic smoothing on fractional spaces, this smoothing does not automatically propagate through arbitrary time-ordered products.
As a consequence, there is in general no global smoothing mechanism acting on the entire product. The error analysis must therefore rely on stepwise local defect bounds combined with product-type discrete stability estimates. In this sense, the non-autonomous case retains essentially the same structural difficulty as the non-sectorial autonomous setting: analytic smoothing of individual factors does not, by itself, eliminate the accumulation of local defects over time.
In particular, the passage from autonomous to non-autonomous problems destroys the algebraic power structure that is responsible for many classical smoothing arguments.
4.3.4 Interpretation
The non-autonomous sectorial block model confirms the structural conclusion of this thesis: even in the presence of analytic smoothing on the present-state component, the dominant contribution to the splitting defect originates from the transport dynamics on the history component.
Consequently, the local defect remains of order $h^{1-\theta}$ on fractional spaces, and the global convergence mechanism is governed by the same abstract principles as in the non-sectorial case. In particular, convergence of the splitting scheme is obtained by applying Main Theorems A and C of the discrete resolvent framework, rather than through sectorial-specific smoothing arguments.
Remark 4.13 (Finite-time versus long-time behavior).
The distinction between finite-time convergence and long-time behavior remains unchanged in the non-autonomous sectorial setting.
On every fixed finite time interval $[0,T]$, the telescoping argument combined with the uniform frozen defect estimate yields convergence of Lie–Trotter splitting toward implicit Euler under the same discrete stability hypotheses as in the non-sectorial theory.
Uniform-in-time convergence or preservation of asymptotic properties would require substantially stronger assumptions on the time-dependent product structure, and lies beyond the scope of the present work.
5 Numerical Experiments
This section illustrates and complements the discrete-resolvent convergence theory developed in the preceding sections by a collection of numerical experiments. The goal is not merely empirical validation, but to make the mechanisms identified in the analysis visible in concrete computations: transport-driven defects, the role of time dependence, and the impact of discrete product stability on error accumulation.
We focus on delay models for which both the implicit Euler scheme and the Lie–Trotter splitting method admit transparent implementations. In the scalar non-sectorial experiments, all updates can be written in closed form and implemented efficiently using a ring-buffer representation of the history. In the sectorial block experiments, we use a one-dimensional reaction–diffusion delay equation to illustrate how the presence of a sectorial diffusion operator affects the numerics while the history transport remains the dominant source of splitting defects on fractional scales.
The experiments are organized according to the four regimes considered in this thesis (autonomous/non-autonomous, sectorial/non-sectorial). In each regime we compare implicit Euler with Lie–Trotter splitting, and interpret the observed behavior in terms of the abstract comparison principles. In particular, the results illustrate that
  • in the non-sectorial setting, global error accumulation is governed by discrete stability of products, in accordance with Main Theorems A and C;
  • in sectorial block models, diffusion improves qualitative regularity of the present-state component, but the splitting defect on fractional spaces is typically still limited by the non-sectorial history transport, so that the same discrete-resolvent principles (Main Theorems A and C) remain the operative convergence mechanism;
  • ring-buffer implementations faithfully realize the exact-shift transport resolvent $(I-hC)^{-1}$ and thereby match the abstract splitting formulation.
All experiments are designed to highlight qualitative mechanisms rather than raw performance. Implementation details, including ring-buffer realizations of the history update, are summarized in Appendix B.2.
5.1 Autonomous Non-Sectorial DDE
We consider the autonomous scalar delay differential equation
$\begin{cases}u^{\prime}(t)=a\,u(t)+b\,u(t+\tau),&t\geq 0,\ \tau<0,\\ u(t)=\phi(t),&t\in[\tau,0],\end{cases}$ (5.1)
with a prescribed history function $\phi$.
The purpose of the experiments in this section is not to investigate convergence toward a continuous solution in a semigroup/evolution-family sense. As emphasized throughout this thesis, depending on the phase space and the parameter regime, delay problems may fall outside the classical $C_{0}$-semigroup framework, and the continuous limit is not the primary object in our analysis.
Instead, we investigate the discrete propagators themselves, and in particular the comparison between the implicit Euler resolvent and the Lie–Trotter splitting propagator. The experiments illustrate that even when the underlying dynamics are qualitatively stable, the Lie–Trotter propagator need not satisfy the strongest “summability of powers” assumption appearing in the hierarchy of Main Theorem A. This motivates the use of weaker, semigroup-like stability notions and clarifies the distinct roles of the stability levels in controlling global error accumulation.
5.2 Implementation overview
All schemes are implemented in Python. For scalar DDEs, the history variable is stored on a uniform grid on $\tau,0]$ and updated by an exact-shift operation, implemented as a FIFO ring buffer. When $\tau=-mh$ with $m\in\mathbb{N}$, the shift resolvent $(I-hC)^{-1}$ is realized exactly by the buffer update. If $\tau/h\notin\mathbb{N}$, a single interpolation step is used at the delayed evaluation point; see Appendix B.2.
For reference computations in the non-autonomous scalar example, we also use the semi-analytical benchmark construction of Ohira and Ohira (2025); Ohira (2025). For background on numerical methods for DDEs we refer to Hairer and Wanner (1996); Bellen and Zennaro (2003). A comparison with the public Python package ddeint is included for context.111ddeint: https://github.com/Zulko/ddeint.
Exact-shift Lie–Trotter (ring buffer).
We define the preconditioned step by
$S_{h}:=\bigl[I-h\,D(I-hC)^{-1}\bigr]^{-1}.$
The exact-shift Lie–Trotter step is then given by
$P_{h}:=S_{h}S_{h}^{C}.$
On a uniform grid with $\tau=-mh$, the operator $S_{h}^{C}$ acts as a FIFO ring buffer on the discrete history array and writes the new inflow value $u$ at $\sigma=0$. The following lemma illustrates the structure of the ring buffer used in numerical implementations. For the explicit algorithm, see Appendix B.2.
Lemma 5.1 (Exact shift equals ring buffer on a uniform grid).
If $\tau=-mh$ with $m\in\mathbb{N}$, then on the discrete history grid $\{\tau,\tau+h,\ldots,-h\}$ the shift resolvent $S_{h}^{C}=(I-hC)^{-1}$ coincides with the FIFO ring-buffer update
$\bigl[\rho(\tau),\rho(\tau+h),\ldots,\rho(-h)\bigr]\ \mapsto\ \bigl[\rho(\tau+h),\ldots,\rho(-h),\,u\bigr],$
where $u$ is the new inflow value at $\sigma=0$. If $\tau/h\notin\mathbb{N}$, then $S_{h}^{C}$ is realized by the same shift combined with a single interpolation at the delayed evaluation point.
5.2.1 Implicit Euler: non-autonomous non-sectorial case
We consider the non-autonomous scalar delay equation with a time-dependent coefficient $a(t)$. Freezing the coefficient at $t_{n}=nh$, the implicit Euler step on the product space reads
$(I-h(C+D(t_{n})))X_{n}=X_{n-1},\qquad X_{n}=(u_{n},\rho_{n})\in\mathcal{D}(C),$
which expands to
$(1-h\,a(t_{n}))u_{n}$ $=u_{n-1}+h\,b\,\rho_{n}(\tau),$ (5.2)
$\rho_{n}-h\,\partial_{\sigma}\rho_{n}$ $=\rho_{n-1},\qquad\rho_{n}(0)=u_{n}.$
Solving the transport resolvent equation yields
$\rho_{n}(\sigma)=e^{\sigma/h}u_{n}+\int_{\sigma}^{0}\frac{1}{h}e^{(\sigma-s)/h}\rho_{n-1}(s)\,ds,\qquad\sigma\in[\tau,0],$
and hence
$\rho_{n}(\tau)=e^{\tau/h}u_{n}+\int_{\tau}^{0}\frac{1}{h}e^{(\tau-s)/h}\rho_{n-1}(s)\,ds.$
Substituting into the first line of (5.2) gives
$\bigl(1-h\,a(t_{n})-h\,b\,e^{\tau/h}\bigr)u_{n}=u_{n-1}+b\int_{\tau}^{0}e^{(\tau-s)/h}\rho_{n-1}(s)\,ds.$
Once $u_{n}$ is obtained, the full history $\rho_{n}$ follows from the above formula.
Remark 5.2.
The above derivation corresponds precisely to solving the resolvent equation
$(I-h(C+D))X_{n}=X_{n-1},\qquad X_{n}=(u_{n},\rho_{n}),$
that is,
$X_{n}=(I-h(C+D))^{-1}X_{n-1}.$
Using the resolvent factorization
$(I-h(C+D))^{-1}=\bigl[I-hD(I-hC)^{-1}\bigr]^{-1}\,(I-hC)^{-1},$
we recover both the transport formula
$\rho_{n}(\sigma)=e^{\sigma/h}u_{n}+\int_{\sigma}^{0}\frac{1}{h}e^{(\sigma-s)/h}\rho_{n-1}(s)\,ds$
and the update equation
$(1-ha)u_{n}=u_{n-1}+hb\,\rho_{n}(\tau).$
This derivation therefore provides a concrete illustration of the right-preconditioned resolvent framework developed in the theoretical part of this thesis. For further details on the resolvent-based formulation of delay equations, we refer to Hansen and Stillfjord (2014).
5.2.2 Sequential Lie–Trotter splitting for scalar DDEs (non-sectorial case)
We implement the (right-preconditioned) Lie–Trotter splitting scheme for the scalar delay model
$u^{\prime}(t)=a(t)u(t)+b\,u(t+\tau),\qquad\tau<0,$
in the product-space formulation developed in the theoretical sections. At the discrete level, the scheme reduces to a simple two-stage update that is most conveniently realized by a ring buffer for the history variable. A detailed derivation of the closed-form scalar recursions and the full algorithmic description are given in Appendix A.4.3; see also Hansen and Stillfjord (2014); Geiser (2008, 2009, 2011) for related splitting approaches for delay problems.
Abstract one-step form. Let $X_{n}=(u_{n},\rho_{n})$ denote the discrete state on $\mathcal{E}_{0}$, where $\rho_{n}(\sigma)$ approximates the history segment $\sigma\mapsto u(t_{n}+\sigma)$. For a frozen coefficient $a(t_{n-1})$ (with $a(t)\equiv a$ in the autonomous case), the splitting step is implemented in the right-preconditioned form
$X_{n}=P_{h}(t_{n-1})X_{n-1},\qquad P_{h}(t):=\bigl[I-h\,D(t)\,S_{h}^{C}\bigr]^{-1}S_{h}^{C},\qquad S_{h}^{C}=(I-hC)^{-1},$
where
$D(t)(u,\rho):=\bigl(a(t)\,u+b\,\rho(\tau),\,0\bigr).$
This is the concrete realization of the discrete-resolvent splitting operator used throughout the theoretical part.
Practical interpretation. On a grid with $\tau=-mh$, the action of $S_{h}^{C}$ on the history component is exactly a FIFO update: the oldest value is removed and the new inflow value is appended. Hence the history can be stored as a ring buffer of length $m$. Moreover, in practical implementations one never needs to explicitly form the preconditioned history $(I-hC)^{-1}\rho$. The only quantity entering the scalar correction step is the delayed trace $\rho(\tau)$, which after the shift corresponds to a known buffer entry. This explains why practical codes for scalar DDEs are extremely compact while still faithfully realizing the abstract splitting scheme.
A concrete ring-buffer implementation is provided in Appendix B.2.
5.2.3 Autonomous non-sectorial case: Lie–Trotter scheme
In this section, we present numerical results for the autonomous non-sectorial delay differential equation (DDE) computed by the Lie–Trotter splitting scheme,
$u^{\prime}(t)$ $=a\,u(t)+b\,u(t+\tau),\quad t\geq 0,$
$u(t)$ $=\mathrm{polyFunc}(t),\quad t\in[\tau,0],$
with parameters a=0.15a=-0.15, b=6.0b=-6.0, and $\tau=-8.0$. The history function is chosen as a tenth-degree polynomial defined in (5.6).
For this parameter configuration, the numerical solution produced by the Lie–Trotter scheme exhibits a clear exponential increase, as illustrated in Figure 4. At first glance, such growth might be interpreted as numerical instability. However, this behavior is entirely consistent with the functional-analytic structure of the underlying problem. Although $|b|>1$, the associated delay operator $C+D$ still generates a $C_{0}$-semigroup on the natural product phase space. Consequently, solutions cannot experience finite-time blow-up; instead, their growth is necessarily bounded by an exponential envelope of the form
$\|u(t)\|\leq Me^{\omega t}.$
This qualitative prediction from semigroup theory is precisely what is observed in the numerical experiment.
To verify the exponential character of the growth, we analyze the logarithmic behavior of the numerical solution:
$\log\|u(t)\|\approx\log M+\omega t.$
The left panel of Figure 5 shows that the numerical data are well approximated by a linear function in time. A least-squares fit over the interval $t\in[40,80]$ yields the parameters
$\log M\approx-2.87,\qquad\omega\approx 0.31,$
with a coefficient of determination $R^{2}\approx 0.94$, indicating an excellent linear approximation on the logarithmic scale.
The right panel further confirms that the relative ratio
$\frac{\|u(t)\|}{Me^{\omega t}}$
remains close to unity for large $t$. Moreover, Figure 4 demonstrates that the normalized deviation stabilizes around a constant level, supporting the asymptotic behavior
$u(t)\sim Me^{\omega t}.$
These observations demonstrate that, even in regimes where the solution grows, the Lie–Trotter scheme faithfully reproduces the qualitative behavior predicted by semigroup theory: the growth is exponential yet controlled, and no finite-time blow-up occurs.
Exponential growth and spectral radius interpretation.
Although numerical solutions may grow for certain parameter sets $(a,b,\tau)$, this growth is purely exponential and does not correspond to a finite-time blow-up. Indeed, the Lie–Trotter scheme reduces to a finite-dimensional linear recursion
$x_{n+1}=M_{h}x_{n},$
where $M_{h}$ denotes the associated companion matrix. Consequently,
$\|x_{n}\|\leq C\,\rho(M_{h})^{n},$
with $\rho(M_{h})$ the spectral radius of $M_{h}$. The effective continuous growth rate is therefore given by
$\gamma_{h}=\frac{1}{h}\log\rho(M_{h}).$
For the representative parameter choice
$a=-0.15,\qquad b=-6.0,\qquad\tau=-8.0,$
numerical evaluation via the power method shows that $\rho(M_{h})$ is slightly larger than $1$, which explains the observed exponential increase of the solution. Moreover, as $h\to 0$, one observes
$\rho(M_{h})\to 1,\qquad\gamma_{h}\to\gamma,$
where $\gamma$ denotes the dominant characteristic exponent of the continuous delay equation. Hence, the divergence seen in the discrete solution is not an instability in the sense of finite-time blow-up, but rather an exponential growth with an essentially constant rate determined by the spectral radius of the discrete propagator.
Remark. The logarithmic plot may exhibit occasional downward spikes (“valleys”) whenever $u(t)$ approaches zero due to sign changes. In such instances, $\log|u(t)|$ temporarily decreases, even though the overall envelope continues to grow exponentially. Therefore, the fitted slope should be interpreted not as an exact straight line, but rather as an average exponential growth rate over the chosen time interval.
Figure 3: Lie–Trotter solution of the autonomous DDE: a=0.15a=-0.15, b=6.0b=-6.0, $\tau=-8.0$, history = polyFunc(t).
Figure 4: Relative error: $\|u(t)\|/Me^{\omega t}$, a=0.15a=-0.15, b=6.0b=-6.0, $\tau=-8.0$, history = polyFunc(t).
Figure 5: $\log\|u(t)\|\approx\log M+\omega t$, a=0.15a=-0.15, b=6.0b=-6.0, $\tau=-8.0$, history = 10th Polynomial. $\log M=-2.87,\omega=0.31$.
Figure 6: Solution of autonomous DDE: a=0.15a=-0.15, b=6.0b=-6.0, $\tau=-0.257$, history = polyFunc(t).
Figure 7: ’ddeint’ vs. Implicit Euler: a=0.15a=-0.15, b=6.0b=-6.0, $\tau=-0.257$, history = polyFunc(t). Blue = ddeint, Red = Implicit Euler.
Figure 8: Implicit Euler vs. Operator Splitting: a=0.15a=-0.15, b=6.0b=-6.0, $\tau=-0.257$, history = polyFunc(t). Blue = Operator Splitting, Red = Implicit Euler.
Figure 9: Error between Implicit Euler and Operator Splitting: a=0.15a=-0.15, b=6.0b=-6.0, $\tau=-0.257$, history = polyFunc(t).
5.2.4 Structure of the Lie–Trotter Scheme and Failure of Summability
In this subsection we examine, for the concrete scalar delay model, which stability assumptions in Main Theorem A are actually satisfied by the Lie–Trotter scheme.
Lie–Trotter as composition of two subproblems.
We decompose the equation
$u^{\prime}(t)=au(t)+bu(t+\tau)$
into the two subproblems
  • (A)
    $u^{\prime}(t)=au(t)$ (instantaneous linear dynamics),
  • (B)
    $u^{\prime}(t)=bu(t+\tau)$ (pure delay dynamics).
On a single time step $t_{n},t_{n+1}]$, the delayed value $u(t+\tau)$ lies in the past interval $t_{n-m},t_{n-m+1}]$ (where $m=-\tau/h$), and is therefore numerically known from the history.
Approximating (B) by a forward Euler step gives
$\tilde{u}_{n+1}=\tilde{u}_{n}+hbu_{n-m},\qquad\tilde{u}_{n}=u_{n},$
that is,
$\tilde{u}_{n+1}=u_{n}+hbu_{n-m}.$
Applying a implicit Euler step for (A) yields
$u_{n+1}=\tilde{u}_{n+1}+hau_{n+1},\qquad\Longleftrightarrow\qquad(1-ha)u_{n+1}=\tilde{u}_{n+1}.$
Hence the one-step Lie–Trotter update takes the explicit form
$u_{n+1}=\frac{1}{1-ha}u_{n}+\frac{hb}{1-ha}u_{n-m}=:\alpha u_{n}+\beta u_{n-m}.$ (5.3)
This shows that the resolvent-based Lie–Trotter scheme corresponds exactly to “implicit Euler for the present term and forward Euler for the delay term”.
Companion matrix and memory structure.
The recurrence (5.3) involves both $u_{n}$ and the $m$-step delayed value $u_{n-m}$. Therefore the scheme naturally lives on the augmented state space
$x_{n}=(u_{n},u_{n-1},\dots,u_{n-m})^{\top}\in\mathbb{R}^{m+1},$
which evolves according to the companion matrix recursion
$x_{n+1}=M_{h}x_{n},\qquad M_{h}=\begin{pmatrix}\alpha&0&\cdots&0&\beta\\ 1&0&\cdots&0&0\\ 0&1&\ddots&0&0\\ \vdots&&\ddots&\ddots&\vdots\\ 0&0&\cdots&1&0\end{pmatrix}$.
This algebraic structure reflects the intrinsic memory effect of delay equations: advancing the solution requires retaining the last $m$ values. The ring buffer implementation described in Appendix B.2 is precisely the computational realization of this companion structure. In particular, the trace value $\rho(\tau)$ used in the operator formulation corresponds exactly to the component $u_{n-m}$ in the finite-dimensional representation.
Connection with the abstract propagator $P_{h}$.
In this finite-dimensional formulation,
$x_{n+1}=M_{h}x_{n}\quad\Longleftrightarrow\quad x_{n}=M_{h}^{n}x_{0}.$
Thus the discrete propagator $P_{h}$ introduced abstractly in the theoretical part coincides here with the matrix $M_{h}$:
$P_{h}=M_{h}.$
The scalar sequence $u_{n}$ is simply the first component of $x_{n}=P_{h}^{n}x_{0}$.
The same observation applies to implicit Euler: the implicit Euler recurrence
$u_{n+1}=\alpha u_{n}+\beta u_{n+1-m}$
can also be written in companion form, leading to another discrete propagator matrix associated with the implicit Euler scheme.
Spectral analysis and failure of summability.
The eigenvalues of $M_{h}$ are given by the roots of the characteristic polynomial obtained by substituting $u_{n}=\lambda^{n}$ into (5.3):
$\lambda^{m+1}=\alpha\lambda^{m}+\beta\quad\Longleftrightarrow\quad p_{LT}(\lambda):=\lambda^{m+1}-\alpha\lambda^{m}-\beta=0.$
For the parameter set
$a=-0.15,\qquad b=-6.0,\qquad\tau=-0.257,$
and step size $h=0.001$, numerical computation of the spectral radius, $\rho(M_{h})$ yields
$\rho(M_{h})=0.9999108137\ldots<1.$
Although the spectral radius is strictly less than one (consistent with the observed stability of the numerical solution), it approaches one as $h\to 0$. Consequently,
$\sum_{n=0}^{\infty}\|M_{h}^{n}\|\sim\frac{1}{1-\rho(M_{h})}\approx 1.1\times 10^{4},$
which is not uniformly bounded in $h$. Therefore, the family $\{P_{h}\}$ fails to satisfy the strongest stability assumption (summability of powers) in Main Theorem A.
Open question: Ritt and BV stability.
Whether the intermediate stability conditions in Main Theorem A, such as Ritt stability or bounded variation of powers, hold for this scheme remains an interesting open question and is not addressed in the present work.
5.3 Non-autonomous non-sectorial case
5.3.1 Benchmark solution via the Ohira–Ohira method
Closed-form solutions of non-autonomous delay differential equations (DDEs) are rarely available. For the particular class of equations considered below, Ohira and Ohira proposed a semi-analytical construction that produces solutions of very high accuracy. In the numerical experiments of this section, we use their construction as a reference (benchmark) solution Ohira and Ohira (2025); see also Ohira (2025) for subsequent developments.
We consider the scalar non-autonomous DDE
$u^{\prime}(t)$ $=a\,t\,u(t)+b\,u(t+\tau),\qquad t\geq 0,$ (5.4)
$u(t)$ $=\mathrm{history}(t),\qquad\tau\leq t\leq 0,$ (5.5)
where $a,b\in\mathbb{R}$ and $\tau<0$.
Outline of the construction. The Ohira–Ohira method constructs a solution on the whole real line via a Fourier representation and then extracts the history segment on $\tau,0]$ so that the resulting function satisfies (5.4) on $[0,T]$. More precisely, one obtains an explicit Fourier-space expression of the form
$\hat{u}(\omega)=\mathcal{C}\,\exp\!\left(\frac{\omega^{2}}{2a}+\frac{b}{a\tau}\,e^{i\omega\tau}\right),$
where $\mathcal{C}$ is a normalization constant (set to $1$ in our experiments). Taking the inverse Fourier transform yields the real-valued representation
$u(t)=\frac{1}{2\pi}\int_{\mathbb{R}}\exp\!\left(\frac{\omega^{2}}{2a}+\frac{b\cos(\omega\tau)}{a\tau}\right)\cos\!\left(\frac{b}{a\tau}\sin(\omega\tau)+\omega t\right)\,d\omega.$
This integral can be evaluated numerically with high precision. Ohira and Ohira proved that the resulting function indeed satisfies (5.4)–(5.5); see Ohira and Ohira (2025).
Parameter set and initialization. Throughout this section we fix
$a=-0.15,\qquad b=-6.0,\qquad\tau=-8.0.$
The corresponding benchmark solution is displayed in Figure 11. To initialize the discrete schemes with a smooth history function, we approximate the history segment on $\tau,0]$ by a polynomial fit of degree $10$, denoted by $\mathrm{polyFunc}(t)$:
[…]
Figure 11 shows this fitted history. This preprocessing step ensures that implicit Euler and Lie–Trotter splitting start from the same smooth initialization and can be compared meaningfully.
5.3.2 Comparison of numerical approximations
Remark 5.3 (Role of the Ohira–Ohira benchmark).
The Ohira–Ohira construction is used here as a highly accurate reference curve rather than as a privileged representation of the continuous dynamics in a semigroup or evolution-family sense. Indeed, the main objective of this thesis is to compare the discrete propagators generated by implicit Euler and Lie–Trotter splitting within the discrete-resolvent framework, without taking the existence of a continuous solution operator as the starting point. Accordingly, in the present section the Ohira–Ohira solution serves primarily as an external accuracy check: it confirms that implicit Euler (at the chosen step size) tracks a plausible continuous trajectory, and thereby provides a reliable baseline against which the splitting-vs-implicit-Euler discrepancy can be interpreted.
Figure 13 shows the implicit Euler approximation for step size $h=0.1$. In Figure 13 we compare the implicit Euler trajectory with the benchmark solution, confirming that implicit Euler captures the reference dynamics well at this resolution.
Next, Figure 15 compares implicit Euler and Lie–Trotter splitting under the same step size, and Figure 15 displays the difference of the two numerical trajectories. The discrepancy initially grows and then decays, producing a bell-shaped error profile. This behaviour is consistent with the finite-time mechanism identified in the non-autonomous non-sectorial theory: local defects may accumulate transiently, while the overall error remains controlled on bounded time intervals.
Figure 10: K. Ohira and T. Ohira’s Exact Semi-Numerical Solution for a=0.15a=-0.15, b=6.0b=-6.0, $\tau=-8.0$.
Figure 11: History Segment for the DDE, fitted by a 10th-degree polynomial.
Figure 12: Implicit Euler Approximate Solution, $h=0.1$.
Figure 13: Comparison: Implicit Euler (red) vs. Exact Solution (blue dots).
Figure 14: Comparison: Implicit Euler (red) vs. Operator Splitting (blue).
Figure 15: Error between Implicit Euler and Operator Splitting Solutions.
5.3.3 Discussion: verification of the assumptions of Main Theorem C
We verify that the numerical example
$u^{\prime}(t)=at\,u(t)+b\,u(t+\tau),\qquad a=-0.15,\ b=-6.0,\ \tau=-8.0,$
fits into the abstract framework of Theorem 3.16.
1. Abstract formulation.
We work on the product space
$\mathcal{E}_{0}=\mathbb{R}\oplus L^{p}([\tau,0]),\qquad X(t)=(u(t),\rho_{t}),\quad\rho_{t}(\sigma)=u(t+\sigma),$
and define
$C(u,\rho)=(0,\partial_{\sigma}\rho),\qquad D(t)(u,\rho)=(at\,u+b\,\rho(\tau),\,0),$
with
$\mathcal{D}(C)=\{(u,\rho):\rho\in W^{1,p}([\tau,0]),\ \rho(0)=u\}.$
Then $X^{\prime}(t)=(C+D(t))X(t)$ is precisely the abstract delay equation considered in the theory.
2. Verification of (C1) and (C2).
By the trace inequality on $W^{1,p}([\tau,0])$, there exists $c>0$ such that
$|\rho(\tau)|\leq c\bigl(\|\rho\|_{L^{p}}+\|\partial_{\sigma}\rho\|_{L^{p}}\bigr).$
Hence $D(t)$ is $C$-bounded with relative bound 0, uniformly for $t\in[0,T]$. In particular, for sufficiently small $h$ the frozen resolvent
$R_{h}(t)=(I-h(C+D(t)))^{-1}$
exists and is uniformly bounded on $\mathcal{E}_{0}$. Moreover, the operational smallness estimate
$\|hD(t)S_{h}^{C}\|\leq q<1$
holds for $h$ sufficiently small, so that the splitting step $P_{h}(t)$ is well defined.
3. Local defect (C3).
The local defect bound
$\|R_{h}(t)-P_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{\theta}h^{\theta}$
follows from the abstract defect analysis in Section 4. Since the coefficient $t\mapsto at$ is smooth on $[0,T]$, the constants in the estimate can be chosen uniformly for $t\in[0,T]$.
4. Discrete stability (C4).
The numerical experiments indicate that the discrete propagators remain bounded on finite time intervals. In particular, the computed trajectories do not exhibit finite-time blow-up. This is consistent with the product-stability assumption required in Theorem 3.16.
Conclusion. All structural assumptions of Main Theorem 3.16 are compatible with the present model. The observed convergence of Lie–Trotter splitting toward implicit Euler in Figures 15 and 15 therefore provides a concrete numerical validation of the abstract theory.
5.4 Autonomous Sectorial Block Model
We consider an autonomous delay equation with a sectorial diffusion component:
$\frac{d}{dt}u(t)$ $=Au(t)+\lambda u(t)+b\,u(t+\tau),\qquad t>0,$ (5.7)
$u(t)$ $=\mathrm{history}(t),\qquad t\in[\tau,0],$
where $A$ is a sectorial operator on the base space $X$ and $\tau<0$ (equivalently one may write $u(t-\tau_{d})$ with $\tau_{d}=-\tau>0$).
As a concrete example we take $X=L^{2}(0,L)$ and $A=\kappa_{\mathrm{diff}}\Delta_{D}$ with homogeneous Dirichlet boundary conditions on $(0,L)$:
$u_{t}(t,x)$ $=\kappa_{\mathrm{diff}}u_{xx}(t,x)+\lambda u(t,x)+b\,u(t+\tau,x),\qquad t\in[0,T],\ x\in(0,L),$ (5.8)
$u(t,0)$ $=u(t,L)=0,\qquad t\geq 0,$
$u(t,x)$ $=\mathrm{history}(t,x),\qquad t\in[\tau,0],\ x\in[0,L].$
5.4.1 Connection between theory and numerical implementation
Equation (5.8) admits a standard product-space formulation. On
$\mathcal{E}_{0}:=X\oplus L^{2}([\tau,0];X),\qquad X=L^{2}(0,L),$
we encode the history by $\rho_{t}(\sigma):=u(t+\sigma)$ and define $X(t):=(u(t),\rho_{t})$. The system becomes the abstract Cauchy problem
$\frac{d}{dt}X(t)=\mathcal{L}X(t),\qquad\mathcal{L}:=A_{0}+B+C,$
with block operators
$A_{0}=\begin{pmatrix}\kappa_{\mathrm{diff}}\Delta_{D}&0\\ 0&0\end{pmatrix},\qquad B=\begin{pmatrix}\lambda I&b\,\Phi_{\tau}\\ 0&0\end{pmatrix},\qquad C=\begin{pmatrix}0&0\\ 0&\partial_{\sigma}\end{pmatrix},\qquad\rho(0)=u.$
Here $A_{0}$ is sectorial and generates analytic smoothing on the present-state component, $C$ generates the contractive left-shift semigroup on the history component, and $B$ couples the present state with the delayed trace $\Phi_{\tau}\rho=\rho(\tau)$.
Remark 5.4 (Why the delay trace is handled via resolvent preconditioning).
On the pure history space $L^{2}([\tau,0];X)$, the point evaluation $\Phi_{\tau}$ is not bounded. In the present framework, however, the delay term appears only through the preconditioned combination $B(I-hC)^{-1}$. The resolvent $(I-hC)^{-1}$ maps into $\mathcal{D}(C)$, where point evaluation becomes bounded with respect to the graph norm. This regularization mechanism is essential for the well-posedness of the splitting step and the ensuing defect estimates.
Implicit Euler scheme
The implicit Euler discretization of $X^{\prime}(t)=\mathcal{L}X(t)$ reads
$(I-h\mathcal{L})X_{n}=X_{n-1},\qquad X_{n}=(u_{n},\rho_{n}),$
that is, one step is given by the resolvent $R_{h}=(I-h\mathcal{L})^{-1}$. Componentwise,
$\begin{cases}(I-h(\kappa_{\mathrm{diff}}\Delta_{D}+\lambda I))u_{n}=u_{n-1}+h\,b\,\rho_{n}(\tau),\\[3.99994pt] (I-h\partial_{\sigma})\rho_{n}=\rho_{n-1},\qquad\rho_{n}(0)=u_{n}.\end{cases}$
Thus, $\rho_{n}$ is obtained by one implicit Euler step for the transport equation, with inflow boundary condition $\rho_{n}(0)=u_{n}$.
Three-factor Lie–Trotter splitting scheme
The splitting propagator analyzed here is the discrete composition
$P_{h}:=\bigl[I-h\,B\,(I-hC)^{-1}\bigr]^{-1}(I-hC)^{-1}(I-hA_{0})^{-1}.$
In the numerical implementation, $(I-hC)^{-1}$ corresponds to an exact shift update of the discrete history (ring buffer), $(I-hA_{0})^{-1}$ to the implicit diffusion solve, and the remaining factor to an algebraic reaction/delay correction based on the buffered value corresponding to $\rho(\tau)$.
Equivalently,
$P_{h}=(I-H_{h})^{-1}S_{h}^{AC},\qquad S_{h}^{AC}=(I-hC)^{-1}(I-hA_{0})^{-1},\qquad H_{h}=h\,B\,S_{h}^{AC}.$ (5.9)
Finite-time comparison against the continuous semigroup.
If $\mathcal{L}$ generates a $C_{0}$-semigroup $U(t)=e^{t\mathcal{L}}$ on $\mathcal{E}_{0}$, then for $t_{n}=nh\leq T$ one may decompose
$U(t_{n})-P_{h}^{n}=\bigl(U(t_{n})-R_{h}^{n}\bigr)+\bigl(R_{h}^{n}-P_{h}^{n}\bigr).$
The first term represents the classical implicit Euler error, while the second term corresponds to the splitting error measured against implicit Euler. In the present block model, the dominant contribution to the local splitting defect is governed by the history transport component; in particular, on the fractional scale $X_{\theta}$ one typically observes the transport-driven order $h^{1-\theta}$ rather than automatic first-order improvement.
Remark 5.5 (Reduction to the non-sectorial model).
If $\kappa_{\mathrm{diff}}=0$, the diffusion block vanishes and the model reduces to the non-sectorial delay equation studied earlier. When $\kappa_{\mathrm{diff}}>0$, diffusion provides smoothing on the present-state component and improves qualitative regularity; however, the convergence order of splitting on $X_{\theta}$ remains limited by the non-sectorial transport on the history component unless further structural assumptions are imposed.
Parameter choices for numerical experiments
We consider (5.8) on $(0,1)$ with homogeneous Dirichlet boundary conditions. The spatial Laplacian is discretized by standard second-order finite differences on $N_{x}$ grid points. Since the coefficients are autonomous, the diffusion matrix is constant in time and its factorization can be reused.
Throughout this subsection we fix
$\kappa_{\mathrm{diff}}=0.02,\quad\lambda=-0.8,\quad b=-0.8,\quad\tau=-0.6,\quad T=8.0,\quad h=0.002,\quad N_{x}=300.$
The history segment on $\tau,0]$ is prescribed by
$\mathrm{history}(t,x)=0.3+0.2\sin(2\pi x)\cos(2\pi t),\qquad t\in[\tau,0],\ x\in[0,1].$
Here, $\kappa_{\mathrm{diff}}$ is the diffusion coefficient, $\lambda$ is the reaction coefficient, $b$ is the delay feedback coefficient, $\tau$ is the delay, $T$ is the final simulation time, $h$ is the time step size, and $N_{x}$ is the number of spatial discretization points. The initial history function was chosen as $\text{history}(t,x)=0.3+0.2\sin(2\pi x)\cos(2\pi t)$ for $t\in[\tau,0]$.
(a) Heat Map of Implicit Euler Solution of the Reaction–Diffusion Delay Equation. Heat(x,t)(x,t) represents $u(t,x)$.
(b) Heat Map of Lie–Trotter Split Solution of the Reaction–Diffusion Delay Equation. Heat(x,t)(x,t) represents $u(t,x)$.
Figure 16: Space–Time Heat Maps for Implicit Euler and Lie–Trotter Split Solutions of the Reaction–Diffusion Delay Equation.
(a) Centre value of Implicit Euler and Lie–Trotter Split Solutions of the Reaction-Diffusion Delay Equation
(b) Timing of Implicit Euler and Lie–Trotter Split Solutions of the Reaction–Diffusion Delay Equation
Figure 17: Timing and Centre Value Comparisons for Implicit Euler and Lie–Trotter Split Solutions of the Reaction–Diffusion Delay Equation.
Table 1: Center value $u(t,0.5)$ for the autonomous diffusion–delay equation $u_{t}=\kappa_{\mathrm{diff}}u_{xx}+\lambda u+bu(t+\tau)$ with $\lambda=-0.8$.
Time $t$ Implicit Euler (IE) Lie–Trotter Splitting
0 $2.9988\times 10^{-1}$ $2.9988\times 10^{-1}$
1 $1726\times 10^{-1} 1.1$$1726\times 10^{-1}$
2 $1.7249\times 10^{-2}$ $1.7249\times 10^{-2}$
3 $6.1128\times 10^{-3}$ $6.1127\times 10^{-3}$
4 $3.0588\times 10^{-4}$ $3.0587\times 10^{-4}$
5 $1.4879\times 10^{-5}$ $1.4878\times 10^{-5}$
6 $4.6214\times 10^{-6}$ $4.6212\times 10^{-6}$
7 $9.8773\times 10^{-7}$ $9.8767\times 10^{-7}$
8 $2.1399\times 10^{-7}$ $2.1382\times 10^{-7}$
Table 2: CPU time for implicit Euler and Lie–Trotter splitting (autonomous diffusion–delay equation, $N_{x}=300$, $h=0.002$, $T=8$).
Method CPU Time (seconds)
Implicit Euler $0.29$
Lie–Trotter Splitting $0.34$
These choices are made for the following reasons:
  • $\kappa_{\mathrm{diff}}>0$ places the diffusion part in the sectorial (analytic) class, so the solution exhibits strong spatial smoothing and remains regular.
  • The coefficients $\lambda$ and $b$ are chosen in a dissipative regime where solutions remain bounded on $[0,T]$, allowing a clean comparison between implicit Euler and splitting.
  • The ratio $|\tau|/h$ is sufficiently large so that the computation genuinely involves a nontrivial history buffer (ring buffer).
Figures 1617 compare the implicit Euler and Lie–Trotter splitting solutions. Table 1 reports representative values of the centre trace $u(t,0.5)$, while Table 2 records CPU times for the current implementation.
Numerical results (autonomous sectorial block model)
The left and right panels of Figure 16 display the space–time heat maps obtained by the implicit Euler scheme and the Lie–Trotter splitting method, respectively. Both schemes reproduce the expected dissipative behaviour: the initial spatial profile decays smoothly under the combined effect of diffusion and the negative reaction/delay feedback. No spurious oscillations or numerical instabilities are observed over the time interval $[0,T]$.
A quantitative comparison is provided by the temporal trace at the midpoint $x=0.5$, shown in the left panel of Figure 17. The two curves are virtually indistinguishable on the displayed scale. Representative values are listed in Table 1, confirming that implicit Euler and Lie–Trotter splitting agree to several significant digits in this regime.
In contrast to the non-autonomous experiments, the computational costs of the two schemes are comparable here. As reported in Table 2 and illustrated in the right panel of Figure 17, implicit Euler is slightly faster in this autonomous setting because the matrix $I-h\kappa_{\mathrm{diff}}\Delta_{h}-h\lambda I$ is constant in time, and its factorization can be reused throughout. The splitting method incurs additional overhead due to the transport/history update and the algebraic correction step, so no decisive speed advantage is visible in this autonomous regime.
Interpretation of the numerical results
Figures 1617 illustrate the numerical behaviour of implicit Euler and Lie–Trotter splitting for the linear reaction–diffusion delay equation (5.8). We present two complementary diagnostics: (i) space–time heat maps of $u(t,x)$ and (ii) the temporal trace of the centre value $u(t,0.5)$.
Space–time dynamics.
Both heat maps show smooth decay of the spatial profile, reflecting the combined dissipative effects of diffusion and negative feedback. The close agreement indicates that the splitting scheme captures the same dominant dynamics as implicit Euler on $[0,T]$.
Centre trace and quantitative agreement.
The centre trace decays rapidly toward zero without overshoot or oscillation, consistent with the stable regime induced by $\lambda<0$ and $b<0$. The differences between the two methods remain below the visible scale of the plot and are small compared with the spatial and temporal discretization errors. This is corroborated by the values in Table 1.
Computational cost.
In the autonomous case the implicit Euler matrix is time-independent, so a single LU factorization can be reused. Consequently, the fully implicit method is not penalized by time-dependent linear algebra, and the splitting method does not yield a clear runtime advantage. This contrasts with the non-autonomous case, where the implicit Euler matrix changes with time.
Relation to the discrete-resolvent convergence mechanism.
From the viewpoint of this thesis, this example serves as a baseline illustrating that, even when a sectorial diffusion operator is present on the state component, the local splitting defect on fractional spaces is typically dominated by the non-sectorial history transport. Accordingly, the finite-time convergence mechanism toward implicit Euler remains governed by the same transport-driven estimate as in the non-sectorial setting. For general background on sectorial semigroups and delay equations on product spaces, see, for instance, Pazy (1983); Lunardi (1995); Bátkai and Piazzera (2005); Bellen and Zennaro (2003).
5.5 Non-Autonomous Sectorial Block Model
We consider a non-autonomous reaction–diffusion delay equation in which the present-state component is governed by a sectorial diffusion operator, while the history dynamics remain non-sectorial due to the transport operator. The scalar delay differential equation reads
$\frac{d}{dt}u(t)=Au(t)+D(t)u(t)+C\,u(t+\tau),\qquad\tau<0,$
where $A$ is sectorial and $D(t)$ is a time-dependent perturbation.
As a concrete model we study the heat equation with time-dependent reaction and delayed feedback:
$u_{t}(t,x)$ $=\kappa_{\mathrm{diff}}u_{xx}(t,x)+\lambda(t)u(t,x)+b\,u(t+\tau,x),\quad t\in[0,T],\ x\in(0,L),$ (5.10)
$u(t,0)$ $=u(t,L)=0,\quad t\geq 0,$
$u(t,x)$ $=\mathrm{history}(t,x),\quad t\in[\tau,0],\ x\in[0,L],$
where $\lambda(\cdot)$ is $C^{1}$ on $[0,T]$. This model coincides with the autonomous sectorial case studied in the previous section, except that the constant coefficient $\lambda$ is replaced by a time-dependent coefficient $\lambda(t)$.
5.5.1 Connection between theory and numerical implementation
We rewrite (5.10) on the product space $\mathcal{E}_{0}=X\oplus L^{p}([\tau,0];X)$ with $X=L^{2}(0,L)$. With $X(t)=(u(t),\rho_{t})$ and $\rho_{t}(\sigma)=u(t+\sigma)$, the system becomes
$\frac{d}{dt}X(t)=\mathcal{L}(t)\,X(t),\qquad\mathcal{L}(t)=A+D(t)+C,$
where
$A=\begin{pmatrix}\kappa_{\mathrm{diff}}\Delta_{D}&0\\ 0&0\end{pmatrix},\quad D(t)=\begin{pmatrix}\lambda(t)I&b\,\Phi_{\tau}\\ 0&0\end{pmatrix},\quad C=\begin{pmatrix}0&0\\ 0&\partial_{\sigma}\end{pmatrix}$.
Here $A$ is sectorial on $X$ and generates an analytic semigroup on the first block, $C$ is the generator of the contractive left–shift semigroup on the history space, and $D(t)$ is a $C^{1}$ family of CC–bounded perturbations. Under standard non-autonomous hypotheses (Tanabe–Kato/Kato framework; see, e.g., Pazy (1983); Tanabe (1979)), the operator family $\mathcal{L}(t)$ generates an evolution family $U(t,s)$ on $\mathcal{E}_{0}$. In this dissertation, however, the convergence analysis is formulated at the level of discrete resolvents and does not presuppose the existence of $U(t,s)$.
Sectorial block structure and verification of basic hypotheses
We briefly verify the standard hypotheses used to place (5.10) into the abstract product-space framework.
  • (S1)
    $A=\kappa_{\mathrm{diff}}\Delta_{D}$ with homogeneous Dirichlet boundary conditions is sectorial on $X=L^{2}(0,L)$ and generates an analytic semigroup.
  • (S2)
    $C(u,\rho)=(0,\partial_{\sigma}\rho)$ with
    $\mathcal{D}(C)=\{(u,\rho):\rho\in W^{1,p}([\tau,0];X),\ \rho(0)=u\}$
    is the standard left–shift generator on the history space.
  • (S3)
    The trace operator $\Phi_{\tau}:\rho\mapsto\rho(\tau)$ is bounded $\mathcal{D}(C)\to X$ with respect to the graph norm $\|x\|+\|Cx\|$. Hence each $D(t)$ is CC–bounded with relative bound 0, uniformly in $t$.
  • (S4)
    If $\lambda(t)$ is $C^{1}[0,T]$, (in this case, it is bounded and Lipschitz on $[0,T]$), then $t\mapsto D(t)$ is Lipschitz as a map $0,T]\to\mathcal{L}(\mathcal{D}(C),\mathcal{E}_{0})$ endowed with the graph norm.
Frozen-time discretization
At each time step $t_{n}=nh$, we freeze the coefficients and approximate
$\mathcal{L}(t)\approx\mathcal{L}(t_{n})=A+D(t_{n})+C.$
The associated frozen-time implicit Euler step is given by
$R_{h}(t_{n})=(I-h(A+D(t_{n})+C))^{-1},\qquad X_{n+1}=R_{h}(t_{n})X_{n}.$
This is precisely the resolvent family used in the local defect analysis.
Frozen-time Lie–Trotter splitting
The frozen-time Lie–Trotter splitting step at time $t_{n}$ is defined by
$S_{h}^{D}(t_{n})=\bigl[I-h\,D(t_{n})(I-hC)^{-1}\bigr]^{-1},\qquad P_{h}(t_{n})=S_{h}^{D}(t_{n})(I-hC)^{-1}(I-hA)^{-1}.$
In computation, each factor corresponds to a familiar substep:
  1. 1.
    $(I-hA)^{-1}$ — implicit diffusion solve;
  2. 2.
    $(I-hC)^{-1}$ — transport step (ring buffer shift);
  3. 3.
    $S_{h}^{D}(t_{n})$ — algebraic reaction/delay update using the delayed value.
After discretization of the history component, evaluation at $\tau$ becomes a finite-dimensional projection, so that the implemented algebraic step can be written in the simpler form $(I-hD_{h}(t_{n}))^{-1}$. This finite-dimensional implementation is consistent with the abstract preconditioned step, since the delay is evaluated on the already shifted buffer.
Local defect measured against implicit Euler.
In this thesis we quantify the splitting defect at the level of the frozen-time implicit Euler resolvent $R_{h}(t_{n})$ rather than against the evolution family. The local consistency bound therefore takes the form
$\|R_{h}(t_{n})-P_{h}(t_{n})\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\leq C_{\theta}\,h^{1-\theta},\qquad(0<h\leq h_{0}),$
even in the present sectorial block-structured model. This reflects the fact that the dominant contribution to the defect arises from the transport operator on the history component; analytic smoothing on the present-state component alone does not improve the local order on $X_{\theta}$.
Remark 5.6 (Optional comparison with evolution-family based estimates).
If one additionally assumes the Tanabe–Kato hypotheses ensuring the existence and sufficient regularity of an evolution family $U(t,s)$ for $\mathcal{L}(t)$, then the one-step error against $U$ can be decomposed as
$U(t_{n+1},t_{n})-P_{h}(t_{n})=\bigl(U(t_{n+1},t_{n})-R_{h}(t_{n})\bigr)+\bigl(R_{h}(t_{n})-P_{h}(t_{n})\bigr).$
The first term is the time-discretization error of implicit Euler for the non-autonomous problem, while the second term is the splitting defect analyzed in the discrete resolvent framework. In the present work, the splitting defect is the dominant component on the fractional scale $X_{\theta}$.
Global error against implicit Euler.
Let
$R_{h}^{(n)}:=R_{h}(t_{n-1})\cdots R_{h}(t_{0}),\qquad P_{h}^{(n)}:=P_{h}(t_{n-1})\cdots P_{h}(t_{0}).$
Combined with the abstract convergence principles (Theorems 3.13 and 3.16), the above local defect yields the finite-time global bound
$\|R_{h}^{(n)}-P_{h}^{(n)}\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\lesssim h^{1-\theta},\qquad nh\leq T,$
under the same discrete stability assumptions as in the non-sectorial case. In particular, the presence of a sectorial diffusion operator in the present block model does not change the qualitative convergence mechanism compared with the purely non-sectorial setting.
Remark 5.7.
In the autonomous limit $\lambda(t)\equiv\lambda$, the evolution family reduces to a semigroup, and the above discussion is consistent with the autonomous bridge in Section 5.4.1.
Remark 5.8 (Boundedness, preconditioning, and the role of ring buffers).
1. Boundedness is a property of the function space. On the infinite-dimensional history space $L^{p}([\tau,0];X)$, point evaluation $\Phi_{\tau}:\rho\mapsto\rho(\tau)$ is not bounded. Thus the delay operator is not bounded on $\mathcal{E}_{0}$, and naive steps such as $(I-hD(t))^{-1}$ are not well-defined at the abstract level.
2. Why the theory uses the preconditioner $(I-hC)^{-1}$. The transport resolvent $(I-hC)^{-1}$ maps $\mathcal{E}_{0}$ into $\mathcal{D}(C)$, where point evaluation is bounded with respect to the graph norm. Therefore the composition $D(t)(I-hC)^{-1}$ becomes bounded, and the preconditioned step
$S_{h}^{D}(t)=\bigl[I-h\,D(t)(I-hC)^{-1}\bigr]^{-1}$
is well-defined and uniformly bounded for sufficiently small $h$.
3. Why implementations can ignore this at the discrete level. After discretization (e.g. a ring buffer), the history space becomes finite dimensional and evaluation at $\tau$ becomes a coordinate projection. Hence all operators are automatically bounded and algebraic inverses are well-defined.
4. Interpretation. Ring buffers are not merely an implementation detail: they provide a finite-dimensional surrogate for the transport regularization mechanism that is essential in the infinite-dimensional analysis.
Parameter justification
We compare the Lie–Trotter splitting scheme with the implicit Euler method for (5.10). The parameters are chosen as in the autonomous sectorial experiment, except that the constant reaction coefficient is replaced by a smooth time-dependent function:
$\lambda(t)=\lambda_{0}+\lambda_{1}\sin\!\Bigl(\frac{2\pi t}{T_{\lambda}}\Bigr),\qquad\lambda_{0}=-0.8,\quad\lambda_{1}=0.2,\quad T_{\lambda}=4.0,$
so that $\lambda(t)\in[-1.0,-0.6]$. The remaining parameters are
$\kappa_{\mathrm{diff}}=0.02,\qquad b=-0.8,\qquad\tau=-0.6,\qquad h=0.002,\qquad T=8.0,\qquad N_{x}=300.$
The history function is prescribed as
$u(t,x)=0.3+0.2\sin(2\pi x)\cos(2\pi t),\qquad t\in[\tau,0],$
which is smooth and compatible with the boundary conditions.
Numerical results
Table 3: Center value $u(t,0.5)$ for the non-autonomous diffusion–delay equation with $\lambda(t)=-0.8+0.2\sin(\tfrac{2\pi t}{4})$.
Time $t$ Implicit Euler (IE) Lie–Trotter Splitting
0 $2.9988\times 10^{-1}$ $2.9988\times 10^{-1}$
1 $1.2276\times 10^{-1}$ $1.2276\times 10^{-1}$
2 $814\times 10^{-2} 1.881$$814\times 10^{-2}$
3 $5.6897\times 10^{-3}$ $5.6898\times 10^{-3}$
4 $3.9143\times 10^{-4}$ $3.9141\times 10^{-4}$
5 $2.0877\times 10^{-5}$ $2.0878\times 10^{-5}$
6 $6.3167\times 10^{-6}$ $6.3166\times 10^{-6}$
7 $1.4738\times 10^{-6}$ $1.4738\times 10^{-6}$
8 $3.9015\times 10^{-7}$ $3.8738\times 10^{-7}$
Table 4: CPU time for implicit Euler and Lie–Trotter splitting (non-autonomous diffusion–delay equation, $N_{x}=300$, $h=0.002$, $T=8$).
Method CPU Time (seconds)
Implicit Euler $3.1516$
Lie–Trotter Splitting $0.1100$
Tables 3 and 4 summarize the midpoint values $u(t,0.5)$ and the CPU times. Figures 1820 display the corresponding space–time dynamics and midpoint traces.
The two heat maps (implicit Euler and Lie–Trotter splitting) are visually indistinguishable, and the midpoint traces remain virtually identical over $[0,T]$. This indicates that, for the present parameter regime, the splitting scheme tracks the frozen-time implicit Euler solution with high accuracy.
A substantial difference is observed in computational cost: Table 4 shows that Lie–Trotter splitting is more than an order of magnitude faster than implicit Euler under the same spatial and temporal discretization. This gain is explained by the structure of the schemes. In the non-autonomous case, implicit Euler requires solving at each step a linear system with a time-dependent coefficient matrix, so that a single LU factorization cannot be reused. By contrast, Lie–Trotter splitting isolates a diffusion solve with a time-independent matrix and treats the reaction/delay update pointwise, which reduces the overall cost. For further discussion of efficiency aspects in non-autonomous problems, see, e.g., Bellen and Zennaro (2003); Hundsdorfer and Verwer (2003); Lubich (2008).
Figure 18: Heat map of the Lie–Trotter splitting solution of the non-autonomous reaction–diffusion delay equation.
Figure 19: Heat map of the implicit Euler solution of the non-autonomous reaction–diffusion delay equation.
Figure 20: Midpoint value $u(t,L/2)$ over time (implicit Euler vs. Lie–Trotter splitting).
Figure 21: Computational cost: implicit Euler vs. Lie–Trotter splitting.
Discussion of numerical results
The experiment confirms two points. First, in the present dissipative regime the Lie–Trotter splitting method reproduces the frozen-time implicit Euler trajectory with high accuracy on the finite time interval $[0,T]$. Second, the splitting method achieves a substantial runtime reduction in the non-autonomous setting, since it reuses a single diffusion solver while avoiding repeated factorizations of time-dependent matrices.
From the viewpoint of the present thesis, the key message is that the sectorial diffusion on the present-state component does not automatically improve the fractional-order local defect produced by the history transport. Accordingly, the convergence mechanism toward implicit Euler remains governed by the same transport-driven estimate as in the non-sectorial case.
Conclusion
Summary of results.
This thesis developed a convergence theory for implicit Euler and Lie–Trotter operator splitting schemes for delay differential equations, formulated entirely at the level of discrete resolvent operators. The analysis does not presuppose the existence of a $C_{0}$-semigroup or an evolution family generated by the underlying continuous operator, and is therefore applicable to regimes in which classical well-posedness theory is unavailable or too weak to support quantitative error estimates.
The abstract framework established in Section 1 shows that the convergence of splitting schemes is governed by two essential ingredients:
  • local defect estimates between the implicit Euler resolvent $R_{h}$ and the splitting propagator $P_{h}$ on suitable regularity spaces, and
  • discrete stability properties of the propagators, such as summability of powers, bounded variation, or Ritt/Kreiss-type conditions.
These principles are summarized in Theorems 3.13 and 3.16, which treat the autonomous and non-autonomous settings, respectively.
Structural limitation for delay equations.
When applying the abstract theory to delay differential equations formulated on product spaces, a central finding of this thesis is that the dominant contribution to the local defect originates from the transport operator governing the history dynamics. More precisely, for the three-factor block model consisting of diffusion on the present state, delay feedback, and transport on the history component, we proved that the local defect satisfies
$\|R_{h}(t)-P_{h}(t)\|_{\mathcal{L}(X_{\theta},\mathcal{E}_{0})}\lesssim h^{1-\theta},\qquad 0<\theta<1,$
both in non-sectorial models and in sectorial block-structured models.
This shows that the presence of a sectorial diffusion operator acting on the present-state component does not automatically improve the convergence order of operator splitting schemes, as long as the transport operator on the history space remains non-sectorial. The non-sectorial nature of the history dynamics therefore constitutes a structural bottleneck that fundamentally limits the attainable convergence rate within the standard product-space formulation.
Consequences for global convergence.
As a consequence, the global error analysis for Lie–Trotter splitting applied to delay equations is governed by the abstract convergence principles corresponding to the non-sectorial regime, namely Theorems 3.13 and 3.16. In particular, convergence on finite time intervals follows from the combination of the $O(h^{1-\theta})$ local defect and suitable discrete stability assumptions, without any essential distinction between the non-sectorial and sectorial block models.
This clarifies that analytic smoothing on a subsystem alone does not alter the qualitative convergence mechanism for splitting schemes in delay equations formulated on product spaces.
Relation to continuous evolution theory.
Although the analysis is developed independently of continuous well-posedness, it remains fully compatible with classical semigroup and evolution family theory whenever such structures exist. Under additional regularity assumptions, such as the Kato–Tanabe or Acquistapace–Terreni conditions, the discrete convergence results imply convergence of both implicit Euler and Lie–Trotter splitting toward the continuous solution operators on finite time intervals. In this sense, the discrete resolvent framework provides both an extension of and a conceptual bridge to the traditional semigroup-based approach.
Final perspective.
This thesis establishes a discrete-resolvent framework for the analysis of operator splitting methods in delay differential equations that is independent of continuous semigroup assumptions. By shifting the focus from continuous generation theory to purely discrete stability mechanisms, the work isolates the structural features that truly govern convergence.
A central insight is that, in product-space formulations of delay equations, transport dynamics on the history component impose a fundamental limitation on splitting accuracy. Even in the presence of sectorial operators acting on the present-state component, analytic smoothing does not override the non-sectorial transport bottleneck. The characteristic fractional convergence rate is thus a structural property of the model rather than a defect of a particular numerical scheme.
This understanding clarifies the role of discrete stability in numerical delay equations and provides a unified interpretation of autonomous and non-autonomous problems. The discrete framework developed here remains valid even when classical well-posedness is unavailable, while still recovering convergence toward continuous solutions whenever a semigroup or evolution family exists.
The results suggest that further progress in high-accuracy splitting for delay equations must address the transport structure directly, either through modified phase spaces or through structure-adapted discretizations. The present thesis provides a foundation for such developments.
A distinctive feature of the present framework is that the convergence theory is established without assuming the existence of a continuous evolution family. When such a family exists, the discrete results automatically imply convergence toward the continuous solution.
Outlook.
The present work suggests several directions for future research. A natural question is whether alternative formulations of the phase space or modified splitting strategies can overcome the structural limitation imposed by the non-sectorial transport operator. Possible directions include:
  • reformulations of the history space using smoother function spaces,
  • higher-order splitting schemes with defect correction,
  • structure-preserving discretizations tailored to delay operators, and
  • extensions to nonlinear or stochastic delay equations.
We believe that the discrete resolvent viewpoint developed in this thesis provides a robust foundation for such further investigations.
Appendix A Auxiliary Functional–Analytic Material
This appendix collects basic functional-analytic properties of the transport operator used throughout the paper. These results ensure that the abstract framework introduced in Section 1 is mathematically well posed and compatible with classical semigroup theory.
A.1 Functional-analytic properties of the transport operator
A.1.1 Hille–Yosida Estimate and Graph Norms
Lemma A.1.
Let $C$ be the transport operator introduced in Section 3. Then there exists $M>0$ such that
$\|C(I-hC)^{-1}\|_{\mathcal{L}(\mathcal{E}_{0})}\leq\frac{M}{h},\qquad 0<h\leq h_{0}.$
Proof.
Using the explicit resolvent formula (or the Hille–Yosida theorem when applicable)
$CR(\lambda,C)=\lambda R(\lambda,C)-I,$
we obtain
$\|CR(\lambda,C)\|\leq\lambda\|R(\lambda,C)\|+1.$
Setting $\lambda=\frac{1}{h}$ and using the Hille–Yosida estimate
$\|R(\lambda,C)\|\leq\frac{M}{\lambda-\omega},$
we obtain
$\|CR(\tfrac{1}{h},C)\|\leq\frac{M}{1-h\omega}+1.$
Since $CR(\tfrac{1}{h},C)=h\,C(I-hC)^{-1}$, we obtain
$\|h\,C(I-hC)^{-1}\|=\|CR(\tfrac{1}{h},C)\|\leq\frac{M}{1-h\omega}+1.$
Hence
$\|C(I-hC)^{-1}\|\leq\frac{1}{h}\left(\frac{M}{1-h\omega}+1\right)\lesssim\frac{1}{h},$
for $0<h\leq h_{0}$. ∎
A.1.2 Closedness and density of the transport operator
Lemma A.2 (closedness and density of the transport/shift generator).
Let
$\mathcal{E}_{0}:=\mathbb{R}\oplus L^{p}([\tau,0]),\quad C(u,\rho):=(0,\rho^{\prime}),$
with
$\mathcal{D}(C):=\{(u,\rho)\in\mathbb{R}\oplus W^{1,p}([\tau,0]):\rho(0)=u\}.$
Then $C$ is densely defined and closed on $\mathcal{E}_{0}$.
Proof.
  • 1.
    Density of $\mathcal{D}(C)$ in $\mathbb{R}\oplus L^{p}(I)$
    Take any $\left(u_{0},\rho_{0}\right)\in\mathbb{R}\oplus L^{p}(I)$. Choose a standard mollification $\tilde{\rho}_{n}\in C^{\infty}(I)\subset W^{1,p}(I)$ with $\tilde{\rho}_{n}\rightarrow\rho_{0}$ in $L^{p}(I)$. Now adjust the trace at 0 without changing $L^{p}$ much: pick a smooth bump $\psi_{n}\in C_{c}^{\infty}([-1/n,0])$ with $\psi_{n}(0)=1$ and $\left\|\psi_{n}\right\|_{L^{p}}\leq 1/n$. Set
    $\rho_{n}:=\tilde{\rho}_{n}+\left(u_{0}-\tilde{\rho}_{n}(0)\right)\psi_{n}.$
    Then $\rho_{n}\in W^{1,p}(I),\rho_{n}(0)=u_{0}$, and $\left\|\rho_{n}-\rho_{0}\right\|_{L^{p}}\leq\left\|\tilde{\rho}_{n}-\rho_{0}\right\|_{L^{p}}+\left|u_{0}-\tilde{\rho}_{n}(0)\right|\left\|\psi_{n}\right\|_{L^{p}}\rightarrow 0$. Hence $\left(u_{0},\rho_{n}\right)\in\mathcal{D}(C)$ and $\left(u_{0},\rho_{n}\right)\rightarrow\left(u_{0},\rho_{0}\right)$ in $\mathcal{E}_{0}$. So $\mathcal{D}(C)$ is dense.
  • 2.
    Closedness of $C$ on $\mathcal{E}_{0}$:
    Suppose $\left(u_{n},\rho_{n}\right)\in\mathcal{D}(C)$ with
    $\left(u_{n},\rho_{n}\right)\rightarrow(u,\rho)\text{ in }\mathcal{E}_{0},\quad A\left(u_{n},\rho_{n}\right)=\left(au_{n},\rho_{n}^{\prime}\right)\rightarrow(v,\eta)\text{ in }\mathcal{E}_{0}.$
    Then $u_{n}\rightarrow u$ in $\mathbb{R}$ and $au_{n}\rightarrow v$ in $\mathbb{R}$, hence $v=au$. For the second component: $\rho_{n}\rightarrow\rho$ in $L^{p}(I)$ and $\rho_{n}^{\prime}\rightarrow\eta$ in $L^{p}(I)$. Therefore $\left(\rho_{n},\rho_{n}^{\prime}\right)\rightarrow(\rho,\eta)$ in $L^{p}\times L^{p}$. The graph of the derivative operator with domain $W^{1,p}(I)$ is closed in $L^{p}\times L^{p}$, so $\rho\in W^{1,p}(I)$ and $\rho^{\prime}=\eta$.
    Finally, the trace map $T:W^{1,p}(I)\rightarrow\mathbb{R},T(\varphi)=\varphi(0)$, is continuous. Since $\rho_{n}\rightarrow\rho$ in $W^{1,p}$ (we have both $\rho_{n}\rightarrow\rho$ and $\rho_{n}^{\prime}\rightarrow\rho^{\prime}$ in $L^{p}$ ), we get $\rho_{n}(0)\rightarrow\rho(0)$. But $\rho_{n}(0)=u_{n}\rightarrow u$. Hence $\rho(0)=u$.
    Thus $(u,\rho)\in\mathcal{D}(C)$ and $C(u,\rho)=\left(au,\rho^{\prime}\right)=(v,\eta)$. So $C$ is closed.
A.1.3 Generator property of $C$ on $L^{p}$
Lemma A.3 (generator property of the transport/shift operator).
Let $I=[\tau,0]$ (finite interval). Define
$C:\mathcal{D}(C)\subset L^{p}(I)\rightarrow L^{p}(I),\quad C\rho=\rho^{\prime},$
with the domain
$\mathcal{D}(C):=\left\{\rho\in W^{1,p}(I):\rho(0)=u\right\}.$
Then $C$ generates the left-shift semigroup $(S(t))_{t\geq 0}$ on $L^{p}(I)$ defined by
$(S(t)\rho)(\sigma):=\begin{cases}\rho(\sigma+t),&\sigma+t\in I\\ 0,&\sigma+t\notin I\end{cases}$
for $\rho\in L^{p}(I),t\geq 0,\sigma\in I$.
Proof.
We verify the Hille–Yosida estimate. For $\lambda>0$ and $\rho\in L^{p}(I)$, the resolvent equation $(\lambda I-C)\varphi=\rho$ reads
$\lambda\varphi(\sigma)-\varphi^{\prime}(\sigma)=\rho(\sigma),\quad\sigma\in I,\quad\varphi(\tau)=0.$
The unique solution is given by
$\varphi(\sigma)=\int_{\tau}^{\sigma}e^{-\lambda(\sigma-s)}\rho(s)ds,\quad\sigma\in I.$
Using Minkowski’s integral inequality, we estimate
$\|\varphi\|_{L^{p}(I)}$ $=\left(\int_{\tau}^{0}\left|\int_{\tau}^{\sigma}e^{-\lambda(\sigma-s)}\rho(s)ds\right|^{p}d\sigma\right)^{1/p}$
$\leq\int_{\tau}^{0}\left(\int_{s}^{0}e^{-\lambda p(\sigma-s)}d\sigma\right)^{1/p}|\rho(s)|ds$
$=\left(\frac{1-e^{-\lambda p|\tau|}}{\lambda p}\right)^{1/p}\int_{\tau}^{0}e^{-\lambda s}|\rho(s)|ds$
$\leq\left(\frac{1-e^{-\lambda p|\tau|}}{\lambda p}\right)^{1/p}\left(\int_{\tau}^{0}e^{-\lambda ps}ds\right)^{1/p}\|\rho\|_{L^{p}(I)}$
$=\frac{1-e^{-\lambda|\tau|}}{\lambda}\|\rho\|_{L^{p}(I)}\leq\frac{|\tau|}{\lambda}\|\rho\|_{L^{p}(I)}.$
Thus the Hille–Yosida estimate holds:
$\|R(\lambda,C)\|_{\mathcal{L}(L^{p}(I))}\leq\frac{1}{\lambda},\quad\lambda>0.$
By the Hille–Yosida theorem, $C$ generates a $C_{0}$-semigroup on $L^{p}(I)$. It is straightforward to verify that this semigroup coincides with the left-shift semigroup defined above. ∎
A.1.4 Relative boundedness of $D$
Lemma A.4 (relative boundedness of the delay operator).
Let $I=[\tau,0]$ (finite interval) and $1<p<\infty$. Define
$D:L^{p}(I)\rightarrow\mathbb{R},\quad D\rho=\int_{\tau}^{0}\rho(\sigma)\mu(d\sigma),$
where $\mu$ is a finite Borel measure on $I$. Then $D$ is relatively bounded with respect to $C$ with relative bound 0, i.e., for every $\varepsilon>0$ there exists $b(\varepsilon)\geq 0$ such that
$\|D\rho\|_{\mathbb{R}}\leq\varepsilon\|C\rho\|_{L^{p}(I)}+b(\varepsilon)\|\rho\|_{L^{p}(I)}$
for all $\rho\in\mathcal{D}(C)$
Proof.
Since $1<p<\infty$, the Sobolev embedding theorem and the Gagliardo–Nirenberg inequality yields
$\|\rho\|_{L^{\infty}}\leq K\,\|\rho^{\prime}\|_{L^{p}}^{1/p}\,\|\rho\|_{L^{p}}^{1-1/p}$
for some constant $K=K(p,\tau)>0$. By Young’s inequality, for every $\delta>0$ there exists $C_{\delta}>0$ such that
$\|\rho\|_{L^{\infty}}\leq\delta\|\rho^{\prime}\|_{L^{p}}+C_{\delta}\|\rho\|_{L^{p}}.$
for all $\rho\in W^{1,p}(I)$. Therefore
$|D\rho|$ $=\left|\int_{\tau}^{0}\rho(\sigma)\mu(d\sigma)\right|\leq\|\rho\|_{\infty}|\mu|(I)$
$\leq(\delta\|\rho^{\prime}\|_{L^{p}}+C_{\delta}\|\rho\|_{L^{p}})|\mu|(I)$
$=|\mu|(I)\delta\|\rho^{\prime}\|_{L^{p}}+|\mu|(I)C_{\delta}\|\rho\|_{L^{p}}$
$=\varepsilon\|\rho^{\prime}\|_{L^{p}}+b(\varepsilon)\|\rho\|_{L^{p}}$
for all $\rho\in\mathcal{D}(C)$. Given $\varepsilon>0$, we can choose $\delta=\varepsilon/|\mu|(I)$ and set $b(\varepsilon)=|\mu|(I)C_{\delta}$. ∎
Generation of $C+D$
Proposition A.5 (Generation of $C+D$).
Let $1<p<\infty$, $\tau<0$, and consider the product space
$\mathcal{E}_{0}=X\oplus L^{p}([\tau,0];X)$
equipped with the norm $\|(u,\rho)\|=\|u\|_{X}+\|\rho\|_{L^{p}}$. Let the transport operator $C$ and the delay operator $D$ be defined by
$C(u,\rho)=(0,\partial_{\sigma}\rho),\qquad D(u,\rho)=(b\,\Phi_{\tau}\rho,\,0),$
with
$\mathcal{D}(C)=\{(u,\rho)\in X\oplus W^{1,p}([\tau,0];X):\ \rho(0)=u\}.$
Then $C+D$ with domain $\mathcal{D}(C)$ generates a $C_{0}$–semigroup on $\mathcal{E}_{0}$.
Proof.
It is well known that $C$ generates the $C_{0}$–semigroup $(T(s))_{s\geq 0}$ on $\mathcal{E}_{0}$ given by the shift–with–inflow formula
$T(s)(u,\rho)=(u,\rho_{s}),$
where
$\rho_{s}(\sigma)=\begin{cases}\rho(\sigma+s),&\sigma+s\in[\tau,0],\\ u,&\sigma+s>0.\end{cases}$
For $x=(u,\rho)\in\mathcal{D}(C)$ we have
$DT(s)x=\bigl(b\,\rho_{s}(\tau),\,0\bigr).$
Fix $t_{0}\in(0,-\tau]$. Then for $0\leq s\leq t_{0}$ we have $\tau+s\leq 0$, hence
$\rho_{s}(\tau)=\rho(\tau+s),$
i.e. no inflow contribution appears at the left endpoint. Therefore,
$\|DT(s)(u,\rho)\|_{\mathcal{E}_{0}}=\|b\,\rho(\tau+s)\|_{X}\leq\|b\|\,\|\rho(\tau+s)\|_{X}.$
Integrating over $s\in[0,t_{0}]$ and using Hölder’s inequality ($1/p+1/p^{\prime}=1$) yields
$\int_{0}^{t_{0}}\|DT(s)(u,\rho)\|_{\mathcal{E}_{0}}\,ds\leq\|b\|\int_{0}^{t_{0}}\|\rho(\tau+s)\|_{X}\,ds\leq\|b\|\,t_{0}^{1/p^{\prime}}\|\rho\|_{L^{p}}.$
Hence
$\int_{0}^{t_{0}}\|DT(s)(u,\rho)\|_{\mathcal{E}_{0}}\,ds\leq q\,\|(u,\rho)\|_{\mathcal{E}_{0}},\qquad q:=\|b\|t_{0}^{1/p^{\prime}}.$
Choosing $t_{0}>0$ sufficiently small so that $q<1$, the Miyadera–Voigt perturbation condition is satisfied. Consequently, by the Miyadera–Voigt perturbation theorem (see, e.g., (Engel and Nagel, 1999, Thm. III.3.14), Voigt (1977), Miyadera (1966), (Bátkai and Piazzera, 2005, Thm. 1.37)), the operator $C+D$ with domain $\mathcal{D}(C)$ generates a $C_{0}$–semigroup on $\mathcal{E}_{0}$. ∎
A.2 Technical lemmas for discrete error analysis
This section collects purely discrete identities and estimates used to control error propagation for (possibly non-autonomous) splitting propagators. No commutativity is assumed.
A.2.1 Non-autonomous telescoping identity (discrete Duhamel formula)
Lemma A.6 (Non-autonomous telescoping identity).
Let $\mathcal{E}_{0}$ be a Banach space and let $R_{k},P_{k}\in\mathcal{L}(\mathcal{E}_{0})$ $(k=0,1,\dots,n-1)$ be bounded operators. Set the (time-ordered) products
$\mathcal{R}_{j:k}:=\begin{cases}R_{j}R_{j-1}\cdots R_{k},&j\geq k,\\ I,&j<k,\end{cases}\qquad\mathcal{P}_{j:k}:=\begin{cases}P_{j}P_{j-1}\cdots P_{k},&j\geq k,\\ I,&j<k.\end{cases}$
Then with $E_{k}:=R_{k}-P_{k}$ one has the exact identity
$\mathcal{R}_{n-1:0}-\mathcal{P}_{n-1:0}=\sum_{k=0}^{n-1}\mathcal{R}_{n-1:k+1}\,E_{k}\,\mathcal{P}_{k-1:0}.$ (A.1)
Proof.
Expand
$\mathcal{R}_{n-1:0}-\mathcal{P}_{n-1:0}=(R_{n-1}-P_{n-1})\mathcal{P}_{n-2:0}+R_{n-1}\bigl(\mathcal{R}_{n-2:0}-\mathcal{P}_{n-2:0}\bigr),$
and iterate. This yields (A.1). ∎
Remark A.7 (How the identity is used).
In applications $R_{k}$ is the frozen implicit Euler resolvent and $P_{k}$ is the frozen splitting step, so $E_{k}$ is the local resolvent defect. The identity (A.1) reduces global error bounds to estimating sums of the form $\sum\mathcal{R}_{n-1:k+1}E_{k}\mathcal{P}_{k-1:0}$ under suitable discrete stability assumptions.
A.2.2 Abel transformation (summation-by-parts) in the non-commutative setting
Lemma A.8 (Abel (summation-by-parts) identity).
Let $X$ be a Banach space, let $T\in\mathcal{L}(X)$, and let $(\tau_{k})_{k\geq 0}\subset X$. Define partial sums $A_{k}:=\sum_{j=0}^{k}\tau_{j}$ (with $A_{-1}:=0$). Then for every $n\geq 1$ one has the exact identity
$\sum_{k=0}^{n-1}T^{\,n-1-k}\tau_{k}=T^{\,n-1}A_{0}+\sum_{m=1}^{n-1}\bigl(T^{\,m-1}-T^{\,m}\bigr)\,A_{n-1-m}+A_{n-1}.$ (A.2)
Proof.
Write $\tau_{k}=A_{k}-A_{k-1}$ and expand:
$\sum_{k=0}^{n-1}T^{n-1-k}(A_{k}-A_{k-1})=\sum_{k=0}^{n-1}T^{n-1-k}A_{k}-\sum_{k=0}^{n-1}T^{n-1-k}A_{k-1}.$
Shift indices in the second sum and simplify to obtain (A.2). ∎
Remark A.9 (Where Abel enters).
In the autonomous case, applying (A.2) to the standard error representation $e_{n}=\sum_{k=0}^{n-1}T^{n-1-k}\tau_{k}$ moves derivatives (finite differences) from the partial sums $A_{k}$ to the powers $T^{m}-T^{m-1}$. This is precisely where Ritt/Kreiss-type assumptions yield improved bounds.
A.2.3 Abel summation gains under discrete stability assumptions
Proposition A.10 (Abel estimate under power-boundedness / Ritt / bounded variation).
Let $T\in\mathcal{L}(X)$ and assume $\sup_{m\geq 0}\|T^{m}\|\leq M$.
(i) Power-bounded case. For any sequence $(\tau_{k})$ with partial sums $A_{k}=\sum_{j=0}^{k}\tau_{j}$,
$\left\|\sum_{k=0}^{n-1}T^{n-1-k}\tau_{k}\right\|\leq(M+1)\,\max_{0\leq k\leq n-1}\|A_{k}\|+\Bigl(\sum_{m=1}^{n-1}\|T^{m-1}-T^{m}\|\Bigr)\max_{0\leq k\leq n-1}\|A_{k}\|.$ (A.3)
(ii) Ritt case (logarithmic accumulation). If in addition
$\sup_{m\geq 1}m\|T^{m}-T^{m-1}\|\leq K,$
then
$\left\|\sum_{k=0}^{n-1}T^{n-1-k}\tau_{k}\right\|\leq C(M,K)\,(1+\log n)\,\max_{0\leq k\leq n-1}\|A_{k}\|.$ (A.4)
(iii) Bounded-variation case (uniform accumulation). If instead
$\sum_{m=0}^{\infty}\|T^{m+1}-T^{m}\|<\infty,$
then
$\left\|\sum_{k=0}^{n-1}T^{n-1-k}\tau_{k}\right\|\leq C_{\mathrm{BV}}\,\max_{0\leq k\leq n-1}\|A_{k}\|,\qquad n\geq 1.$ (A.5)
Proof.
(i) is obtained by taking norms in the Abel identity (A.2). (ii) follows since $\|T^{m-1}-T^{m}\|\leq K/m$ and $\sum_{m=1}^{n-1}1/m\leq 1+\log n$. (iii) follows by replacing the harmonic bound with the summable series. ∎
Remark A.11 (Why this is exactly the “stability hierarchy” in the main text).
The three regimes above correspond to increasingly strong control of $\sum\|T^{m}-T^{m-1}\|$. In the error propagation formula, this quantity measures how local defects accumulate over time. This is the structural reason why the main theorems distinguish power-boundedness, Ritt stability, and bounded-variation assumptions.
Remark A.12.
The identity (A.2) is a discrete analogue of summation-by-parts (Abel transformation) and is a standard tool in the analysis of error propagation for operator iterations and time-discretization schemes; see, e.g., Lubich and Nevanlinna (1991); Gröger (1987). Its role in connection with Ritt-type stability conditions is discussed in Nagy and Zemánek (1999); Schweizer (2012).
A.3 Transport Resolvents
This section collects several local and global error estimates for operator splitting schemes and resolvent approximations that are used throughout the main text Section 2. We begin with the fundamental local error estimate for the implicit Euler method applied to a $C_{0}$-semigroup, and then proceed to splitting-specific bounds.
Remark A.13 (Notation).
Throughout this section, $\mathcal{E}_{0}$ denotes a Banach space, $L:\mathcal{D}(L)\subset\mathcal{E}_{0}\to\mathcal{E}_{0}$ is a densely defined closed operator generating a $C_{0}$-semigroup $(U(t))_{t\geq 0}$, and $R_{h}:=(I-hL)^{-1}$ denotes the resolvent at $\lambda=1/h$. The graph norm on $\mathcal{D}(L)$ is denoted by $\|x\|_{\mathcal{D}(L)}:=\|x\|_{\mathcal{E}_{0}}+\|Lx\|_{\mathcal{E}_{0}}$.
The following lemma estimates the local error of the implicit Euler method for a $C_{0}$-semigroup.
Lemma A.14 (Local Error Per Step of the Implicit Euler Method).
Let $L:\mathcal{D}(L)\subset\mathcal{E}_{0}\to\mathcal{E}_{0}$ be a densely defined closed linear operator such that $(U(t))_{t\geq 0}$ is a $C_{0}$-semigroup defined by $C+D_{\text{bd}}$ on a Banach space $\mathcal{E}_{0}$, and define the implicit Euler step by $R_{h}:=(I-hL)^{-1}$. Then for every $x\in\mathcal{E}_{0}$, the local error satisfies
$\|U(h)x-R_{h}x\|_{\mathcal{E}_{0}}\leq Ch\|Lx\|_{\mathcal{E}_{0}},$
where $C>0$ is a constant independent of $x$ and $h$, for all sufficiently small $h>0$. In particular,
$\|U(h)-R_{h}\|_{\mathcal{L}(\mathcal{D}(L)\to\mathcal{E}_{0})}\leq Ch.$
Proof.
Let $x\in\mathcal{D}(L)$ be arbitrary. First, observe that
$R_{h}x-x=(I-hL)^{-1}x-x=h(I-hL)^{-1}Lx=hR_{h}Lx,$
where we used that $L$ and $(I-hL)^{-1}$ commute on $\mathcal{D}(L)$. On the other hand, by the variation-of-constants formula, the semigroup satisfies
$U(h)x-x=\int_{0}^{h}\frac{d}{dt}U(t)x\,dt=\int_{0}^{h}LU(t)x\,dt.$
Since $LU(t)x=U(t)Lx$ for $x\in\mathcal{D}(L)$, thus:
$U(h)x-x=\int_{0}^{h}U(t)Lx\,dt.$
Subtracting the two expressions gives
$U(h)x-R_{h}x=\int_{0}^{h}U(t)Lx\,dt-hR_{h}Lx.$
Taking norms and applying the triangle inequality yields
$\|U(h)x-R_{h}x\|_{\mathcal{E}_{0}}\leq\int_{0}^{h}\|U(t)Lx\|_{\mathcal{E}_{0}}dt+h\|R_{h}Lx\|_{\mathcal{E}_{0}}.$
Using $\|U(t)\|\leq Me^{\omega t}$ and $\|R_{h}\|\leq M^{\prime}$ for small $h$, we estimate
$\int_{0}^{h}\|U(t)Lx\|_{\mathcal{E}_{0}}dt\leq Me^{\omega h}h\|Lx\|_{\mathcal{E}_{0}},\quad h\|R_{h}Lx\|_{\mathcal{E}_{0}}\leq hM^{\prime}\|Lx\|_{\mathcal{E}_{0}}.$
Thus,
$\|U(h)x-R_{h}x\|_{\mathcal{E}_{0}}\leq\left(Me^{\omega h}+M^{\prime}\right)h\|Lx\|_{\mathcal{E}_{0}}\leq Ch\|Lx\|_{\mathcal{E}_{0}},$
where $C:=Me^{\omega h_{0}}+M^{\prime}$ is independent of $h\in(0,h_{0}]$. ∎
This estimate is classical; see, e.g., (Pazy, 1983, Chap. 9), (Tanabe, 1979, Chap. 5), or Kobayashi (2003).
Lemma A.15 (Transport resolvent factorization (right preconditioning)).
Let $\mathcal{E}_{0}$ be a Banach space. Let $C:\mathcal{D}(C)\subset\mathcal{E}_{0}\to\mathcal{E}_{0}$ be the generator of a $C_{0}$–semigroup, and let $D:\mathcal{D}(C)\to\mathcal{E}_{0}$ be a linear operator.
Fix $h>0$ such that IhCI-hC is invertible on $\mathcal{E}_{0}$ and set
$L:=C+D,\qquad S_{h}:=\Bigl[I-h\,D(I-hC)^{-1}\Bigr]^{-1},$
whenever the inverse exists. Then the following identities hold on $\mathcal{E}_{0}$:
$(I-hL)^{-1}=S_{h}\,(I-hC)^{-1},\qquad S_{h}=(I-hL)^{-1}(I-hC).$ (A.6)
In particular, the resolvent step for $L$ can be viewed as the transport resolvent $(I-hC)^{-1}$ followed by the preconditioned inverse $S_{h}$.
Moreover, if the operational smallness condition
$\bigl\|h\,D(I-hC)^{-1}\bigr\|\leq q<1$ (A.7)
holds (for the given $h$), then $S_{h}$ is well-defined and satisfies the uniform bound
$\|S_{h}\|\leq(1-q)^{-1}.$
Proof.
By definition,
$S_{h}^{-1}=I-h\,D(I-hC)^{-1}.$
Multiplying from the left by (IhC)(I-hC) gives
$(I-hC)S_{h}^{-1}=(I-hC)\Bigl(I-hD(I-hC)^{-1}\Bigr)=(I-hC)-hD=I-h(C+D)=I-hL.$
Taking inverses and using $(AB)^{-1}=B^{-1}A^{-1}$ yields
$(I-hL)^{-1}=\bigl((I-hC)S_{h}^{-1}\bigr)^{-1}=S_{h}\,(I-hC)^{-1},$
which proves the first identity in (A.6). Multiplying this identity on the right by (IhC)(I-hC) gives the second identity:
$S_{h}=(I-hL)^{-1}(I-hC).$
Finally, if (A.7) holds, then $\|hD(I-hC)^{-1}\|\leq q<1$ and hence
$S_{h}=\sum_{m=0}^{\infty}\bigl(hD(I-hC)^{-1}\bigr)^{m}$
converges in operator norm, implying $\|S_{h}\|\leq(1-q)^{-1}$. ∎
Similar resolvent factorizations appear implicitly in perturbation arguments for delay equations and operator splitting; see, e.g., Bátkai and Piazzera (2005); Bátkai et al. (2017).
Remark A.16 (Block commutation identities).
In the product-space setting, $A_{0}=\mathrm{diag}(A,0)$ acts only on the $u$-component, while $C$ acts only on the history component. Hence, for $X:=(I-hA_{0})^{-1}$ we have
$CX=C,\qquad C(X-I)=0.$
Consequently,
$(I-hCX)^{-1}=(I-hC)^{-1},\qquad C(X-I)(I-hC)^{-1}=0.$
Remark A.17 (Why $D$ is always paired with $(I-hC)^{-1}$).
In our product-space setting the delay operator $D$ is typically not bounded on $\mathcal{E}_{0}=X\oplus L^{p}([\tau,0];X)$, because point evaluation $\Phi_{\tau}:\rho\mapsto\rho(\tau)$ is not continuous on $L^{p}$. However, $D$ is well-defined on $\mathcal{D}(C)$ and is CC–bounded with relative bound 0: for every $\varepsilon>0$ there exists $c_{\varepsilon}$ such that
$\|Dx\|\leq\varepsilon\|Cx\|+c_{\varepsilon}\|x\|,\qquad x\in\mathcal{D}(C).$
Since the transport resolvent satisfies $(I-hC)^{-1}:\mathcal{E}_{0}\to\mathcal{D}(C)$ and $\|(I-hC)^{-1}\|\leq 1$, we obtain
$\|h\,D(I-hC)^{-1}\|\leq h\Bigl(\varepsilon\|C(I-hC)^{-1}\|+c_{\varepsilon}\|(I-hC)^{-1}\|\Bigr)\lesssim\varepsilon+c_{\varepsilon}h.$
Choosing $\varepsilon$ and then $h$ sufficiently small yields $\|h\,D(I-hC)^{-1}\|<1$, so the preconditioned step
$\bigl[I-h\,D(I-hC)^{-1}\bigr]^{-1}$
is well-defined by a Neumann series. This is the only bounded combination of $D$ that appears in our resolvent identities and splitting schemes. See, e.g., Engel and Nagel (1999); Bátkai and Piazzera (2005).
A.4 Non–autonomous telescoping identity
Remark A.18 (Heuristic interpretation of the preconditioned DD–step).
The preconditioned one–step map
$S_{h}^{D}(t_{n}):=\bigl[I-h\,D(t_{n})(I-hC)^{-1}\bigr]^{-1}$
is introduced at the resolvent level. It is a bounded operator on $\mathcal{E}_{0}$ for sufficiently small $h>0$.
Formally, recalling that $(I-hK)^{-1}=I+hK+O(h^{2})$ as $h\to 0$ for bounded $K$, one may view $S_{h}^{D}(t_{n})$ as the resolvent analogue of a short-time $D(t_{n})$–flow. In this sense, the frozen three–factor step
$P_{h}(t_{n})=S_{C}(h)\,(I-hA_{0})^{-1}\,S_{h}^{D}(t_{n})$
can be interpreted as a Lie–Trotter–type factorization associated with the frozen operator $A_{0}+C+D(t_{n})$.
We emphasize that this viewpoint is used only as intuition. All rigorous estimates in this thesis are derived from resolvent identities, preconditioning by $(I-hC)^{-1}$, and interpolation–space bounds, without invoking operator exponentials of the non–sectorial part.
A.4.1 From Ritt/Kreiss stability to variation bounds
Definition A.19 (Ritt and Kreiss stability).
Let $X$ be a Banach space and $T\in\mathcal{L}(X)$.
  • (i)
    $T$ is a Ritt operator if
    $\sup_{n\geq 0}\|T^{n}\|\leq M,\qquad\sup_{n\geq 1}n\,\|T^{n}-T^{n-1}\|\leq M_{1}$
    for some constants $M,M_{1}>0$.
  • (ii)
    $T$ satisfies a Kreiss resolvent condition if
    $\sup_{|z|>1}(|z|-1)\,\|(zI-T)^{-1}\|\leq K$
    for some constant $K>0$.
Lemma A.20 (Ritt $\Rightarrow$ logarithmic variation bound).
Let $T\in\mathcal{L}(X)$ be a Ritt operator. Then for every $N\in\mathbb{N}$,
$\sum_{n=0}^{N-1}\|T^{n+1}-T^{n}\|\;\leq\;M_{1}\sum_{n=0}^{N-1}\frac{1}{n+1}\;\leq\;M_{1}\,(1+\log N).$
In particular, Ritt stability yields at most logarithmic growth of the total variation of the powers.
Proof.
By the Ritt bound $n\|T^{n}-T^{n-1}\|\leq M_{1}$ we have $\|T^{n+1}-T^{n}\|\leq M_{1}/(n+1)$ for $n\geq 0$. Summing this estimate gives the claim. ∎
Remark A.21 (On bounded variation assumptions).
The uniform bounded-variation property $\sup_{N}\sum_{n=0}^{N-1}\|T^{n+1}-T^{n}\|<\infty$ is strictly stronger than the Ritt condition in general. In the main text we therefore treat bounded-variation assumptions as separate discrete stability hypotheses when needed.
A.4.2 Non–autonomous telescoping identity
Lemma A.22 (Non–autonomous telescoping identity).
Let $\{U(t,s)\}_{0\leq s\leq t\leq T}$ be an evolution family on a Banach space $\mathcal{E}_{0}$. Let $\{P_{h}(t_{k})\}_{k\geq 0}$ be bounded one–step operators with step size $h>0$ and $t_{k}:=kh$. Define the local defect
$E_{h}(t_{k}):=U(t_{k+1},t_{k})-P_{h}(t_{k}).$
Then for every $n\in\mathbb{N}$ with $t_{n}\leq T$,
$U(t_{n},0)-P_{h}(t_{n-1})\cdots P_{h}(t_{0})=\sum_{k=0}^{n-1}U(t_{n},t_{k+1})\,E_{h}(t_{k})\,P_{h}(t_{k-1})\cdots P_{h}(t_{0}),$ (A.8)
where $P_{h}(t_{-1})\cdots P_{h}(t_{0})$ is interpreted as the identity.
Proof.
Set
$Q_{n}:=P_{h}(t_{n-1})\cdots P_{h}(t_{0}),\qquad Q_{0}:=I.$
For each $k=0,\dots,n-1$ define
$W_{k}:=U(t_{n},t_{k})\,Q_{k}.$
Then $W_{0}=U(t_{n},0)$ and $W_{n}=Q_{n}$. Moreover,
$W_{k+1}-W_{k}$ $=U(t_{n},t_{k+1})Q_{k+1}-U(t_{n},t_{k})Q_{k}$
$=U(t_{n},t_{k+1})\bigl(P_{h}(t_{k})-U(t_{k+1},t_{k})\bigr)Q_{k}$ $=-\,U(t_{n},t_{k+1})E_{h}(t_{k})Q_{k}.$
Summing $W_{k+1}-W_{k}$ from $k=0$ to n1n-1 yields
$Q_{n}-U(t_{n},0)=\sum_{k=0}^{n-1}(W_{k+1}-W_{k})=-\sum_{k=0}^{n-1}U(t_{n},t_{k+1})E_{h}(t_{k})Q_{k},$
which is equivalent to (A.8). ∎
See also Nevanlinna (1993); Borovykh et al. (2000); Le Merdy (2002) for related identities.
A.4.3 The Sequential Operator Splitting: Autonomous and Non-Autonomous DDEs (Non-Sectorial Case)
Remark on notation.
In this appendix we work with the boundary-delay operator
$D_{\mathrm{bd}}(u,\rho):=(au+b\rho(\tau),0),$
which is a bounded finite-rank operator on $\mathcal{E}_{0}$. Hence $(I-hD_{\mathrm{bd}})^{-1}$ exists for small $h$ and reduces to a scalar algebraic inverse. This notation is used only for explicit calculations and should not be confused with the unbounded operator $D$ in the abstract theory. We consider the operator splitting method for solving delay differential equations (DDEs) as discussed in theory. The DDE under consideration is simplest case given by the following equations, for more complex cases, please refer to the relevant literature Hansen and Stillfjord (2014).J. Geiser in multiple works Geiser (2008, 2009, 2011):
The Autonomous Case
We consider the autonomous system:
$\binom{v_{n}}{\phi_{n}}$ $=(I-hC)^{-1}(I-hD_{\mathrm{bd}})^{-1}\binom{v_{n-1}}{\phi_{n-1}}$, (A.9)
$\binom{v_{0}}{\phi_{0}}$ $=\binom{\text{history}(0)}{\text{history}(\sigma)}$.
This scheme is solved sequentially in two steps:
$\binom{w_{n-1}}{\psi_{n-1}}$ $=(I-hD_{\mathrm{bd}})^{-1}\begin{pmatrix}v_{n-1}\\ \phi_{n-1}\end{pmatrix}$, (A.10)
$\binom{v_{n}}{\phi_{n}}$ $=(I-hC)^{-1}\begin{pmatrix}w_{n-1}\\ \psi_{n-1}\end{pmatrix}$.
The first equation in (A.10) can be solved similarly to the Implicit Euler scheme:
$\psi_{n-1}(\sigma)=e^{\sigma/h}v_{n-1}+\int_{0}^{\sigma}\frac{1}{h}e^{(\sigma-s)/h}\phi_{n-1}(s)\,ds$ (A.11)
where we use $\psi_{n-1}(0)=v_{n-1}$. By applying the operator $\Phi$ to both sides of this equation, we derive:
$\psi_{n-1}(\tau)=e^{\tau/h}v_{n-1}+\int_{0}^{\tau}\frac{1}{h}e^{(\tau-s)/h}\phi_{n-1}(s)\,ds$ (A.12)
and
$w_{n-1}=hb\psi_{n-1}(\tau)+v_{n-1}.$ (A.13)
Eliminating $\psi_{n-1}(\tau)$ from (A.12) and (A.13) ,
$v_{n-1}=w_{n-1}-hb\left(e^{\tau/h}w_{n-1}+\int^{0}_{\tau}\frac{1}{h}e^{(\tau-s)/h}\phi_{n-1}(s)ds\right)$ (A.14)
Solving this equation for $w_{n-1}$, we have
$w_{n-1}=\frac{1}{1-hbe^{\tau/h}}\left(v_{n-1}+b\int^{0}_{\tau}e^{(\tau-s)/h}\phi_{n-1}(s)\right).$ (A.15)
Finally, we can transition from $(v_{n-1};\phi_{n-1}(\sigma))$ to $(v_{n};\phi_{n}(\sigma))$ as follows:
$v_{n}$ $=\frac{1}{1-ha}w_{n-1}=\frac{1}{1-ha}\cdot\frac{1}{1-hbe^{\tau/h}}\left(v_{n-1}+b\int^{0}_{\tau}e^{(\tau-s)/h}\phi_{n-1}(s)ds\right)$ (A.16)
$\phi_{n}(\sigma)$ $=\psi_{n-1}(\sigma)$
$=e^{\sigma/h}w_{n-1}+\int_{0}^{\sigma}\frac{1}{h}e^{(\sigma-s)/h}\phi_{n-1}(s)\,ds$
$=e^{\sigma/h}\frac{1}{1-hbe^{\tau/h}}\left(v_{n-1}+b\int^{0}_{\tau}e^{(\sigma-s)/h}\phi_{n-1}(s)ds\right)+\int_{0}^{\sigma}\frac{1}{h}e^{(\sigma-s)/h}\phi_{n-1}(s)\,ds.$
The Non-Autonomous Case
We consider the non-autonomous system:
$\binom{v_{n}}{\phi_{n}}$ $=(I-hC(nh))^{-1}(I-hD_{\mathrm{bd}})^{-1}\binom{v_{n-1}}{\phi_{n-1}}$, (A.17)
$\binom{v_{0}}{\phi_{0}}$ $=\binom{\text{history}(0)}{\text{history}(\sigma)}$.
This case is also solved sequentially in two steps:
$\binom{w_{n-1}}{\psi_{n-1}}$ $=(I-hD_{\mathrm{bd}})^{-1}\begin{pmatrix}v_{n-1}\\ \phi_{n-1}\end{pmatrix}$, (A.18)
$\binom{v_{n}}{\phi_{n}}$ $=(I-hC(nh))^{-1}\begin{pmatrix}w_{n-1}\\ \psi_{n-1}\end{pmatrix}$.
As in the autonomous case, the first equation in (A.18) can be solved similarly to the Implicit Euler scheme:
$\psi_{n-1}(\sigma)=e^{\sigma/h}v_{n-1}+\int_{0}^{\sigma}\frac{1}{h}e^{(\sigma-s)/h}\phi_{n-1}(s)\,ds$ (A.19)
where we again use $\psi_{n-1}(0)=v_{n-1}$. Applying the operator $\Phi$ to both sides gives:
$\psi_{n-1}(\tau)=e^{\tau/h}v_{n-1}+\int_{0}^{\tau}\frac{1}{h}e^{(\tau-s)/h}\phi_{n-1}(s)\,ds$ (A.20)
and
$w_{n-1}=\frac{1}{1-hbe^{\tau/h}}\left(v_{n-1}+b\int^{0}_{\tau}e^{(\tau-s)/h}\phi_{n-1}(s)ds\right).$ (A.21)
By utilizing equations (A.19), (A.20), and (A.21), we can transition from $(v_{n-1},\phi_{n-1}(\sigma))$ to $(v_{n},\phi_{n}(\sigma))$ as follows:
$v_{n}$ $=\frac{1}{1-ha(nh)}w_{n-1}=\frac{1}{1-hanh}\cdot\frac{1}{1-hbe^{\tau/h}}\left(v_{n-1}+b\int^{0}_{\tau}e^{(\tau-s)/h}\phi_{n-1}(s)ds\right)$ (A.22)
$\phi_{n}(\sigma)$ $=\psi_{n-1}(\sigma)$
$=e^{\sigma/h}w_{n-1}+\int_{0}^{\sigma}\frac{1}{h}e^{(\sigma-s)/h}\phi_{n-1}(s)\,ds$
$=e^{\sigma/h}\frac{1}{1-hbe^{\tau/h}}\left(v_{n-1}+b\int^{0}_{\tau}e^{(\sigma-s)/h}\phi_{n-1}(s)ds\right)+\int_{0}^{\sigma}\frac{1}{h}e^{(\sigma-s)/h}\phi_{n-1}(s)\,ds.$
For details on these calculations, see Hansen and Stillfjord (2014).
Remark A.23 (Exact Shift Splitting Procedure).
Replace $(I-hC)^{-1}$ with $S_{C}(h)$. One step of Lie–Trotter (exact shift) is
$P_{h}=S_{C}(h)\left[I-hD_{\mathrm{bd}}(I-hC)^{-1}\right]^{-1}$
It boils down to:
  • 1.
    Compute $\tilde{X}=(I-hC)^{-1}X_{n}$ (this is a history predictor; in practice we don’t need to form it explicitly (see below)).
  • 2.
[…]
    $\hat{u}=(1-ha)^{-1}\left(u_{n}+hb\hat{\rho}(\tau)\right),\quad\hat{\rho}=\tilde{\rho}$
  • 3.
    Apply the exact shift:
[…]
    On a grid with $\tau=-mh$, step (3) is exactly FIFO ring buffer: pop the oldest value, append $\hat{u}$. For practical ring buffer system, see Appendix B.2.
Implementation tip. In practice we don’t form $\tilde{\rho}=(I-hC)^{-1}\rho_{n}$; instead we observe that the only place $\tilde{\rho}$ is needed in $D_{\mathrm{bd}}$ is via its trace at $\tau$. But with the exact shift afterwards, $\rho_{n+1}(\tau)=\hat{\rho}(\tau+h)$ equals the buffer entry one slot to the right. Hence the scalar solve for $\hat{u}$ can be written directly with the ring buffer entry that will become $\rho_{n+1}(\tau)$. This is why practical codes look "simple" yet match the theory.
Appendix B Program Listings: Core Numerical Algorithms
This appendix summarizes the essential Python implementations used for the numerical experiments in Section 5. The purpose is not to provide exhaustive code, but to document the algorithmic structure underlying the implementations and to clarify the computational distinction between the fully implicit Euler method and the Lie–Trotter splitting scheme.
B.1 Algorithmic comparison: Implicit Euler vs. Lie–Trotter
Both methods approximate the same delayed reaction–diffusion equation under identical discretization settings. The difference lies in how the nonlinear and history-coupled system is treated.
Implicit Euler (fully coupled solve)
At each time step, implicit Euler requires solving a nonlinear system
$u^{n+1}=u^{n}+hF(t_{n+1},u^{n+1},u^{n+1-m}),$
which is implemented via Newton iteration. A simplified schematic is:
for n in range(steps):
    u_iter = u_prev.copy()
    for k in range(max_iter):
        F = residual(u_iter, u_prev, u_delay)
        J = jacobian(u_iter)              # depends on current iterate
        du = solve(J, -F)
        u_iter += du
    u = u_iter
The Jacobian depends on the iterate $u^{n+1,(k)}$ and must be recomputed and factorized at each Newton step, resulting in substantial computational cost.
Lie–Trotter splitting (decoupled steps)
The Lie–Trotter method separates the diffusion and reaction–delay components. The linear diffusion step is solved implicitly using a constant matrix, while the nonlinear reaction–delay step is computed explicitly or in closed form.
# Pre-factorization (once)
LU = factorize(I - h * kappa * Laplacian)
for n in range(steps):
    # A-step: diffusion
    uA = LU.solve(u_prev)
    # B-step: reaction + delay
    u = reaction_delay_flow(uA, u_delay, h)
    update_history_buffer(u)
Since the diffusion matrix is constant, its factorization is reused throughout the computation, leading to a substantial reduction in computational cost. This structural difference explains the performance gap observed in Section 5.
The algorithmic distinction between fully implicit schemes and splitting methods follows the classical numerical analysis literature; see, e.g., Hundsdorfer and Verwer (2003); Blanes and Casas (2016).
B.2 Ring buffer for delay handling
The discrete delay term $u(t+\tau)$ is handled using a ring buffer (circular queue). When $|\tau|=mh$ for some integer $m$, the delayed value corresponds exactly to the solution from $m$ time steps earlier.
Let $\{u^{n}\}_{n\geq 0}$ denote the discrete solution at $t_{n}=nh$. The ring buffer stores the most recent $m+1$ values:
$\mathcal{B}_{n}=(u^{n-m},\dots,u^{n}),$
so that
$u^{n-m}=\mathcal{B}_{n}[0],\qquad u^{n}=\mathcal{B}_{n}[m].$
Each update discards the oldest value and appends the newest one, requiring only constant memory and constant-time updates.
Algorithmic structure
Algorithm 1 Ring buffer update for discrete delay
1:step size $h$, delay $\tau=mh$, initial history $\{u^{-m},\dots,u^{0}\}$
2:Initialize buffer $\mathcal{B}\leftarrow[u^{-m},\dots,u^{0}]$
3:for $n=0,1,\dots$ do
4:  $u^{n}\leftarrow\mathcal{B}[m]$
5:  $u^{n-m}\leftarrow\mathcal{B}[0]$
6:  Compute $u^{n+1}=\Psi_{h}(u^{n},u^{n-m})$
7:  Update buffer: $\mathcal{B}\leftarrow[u^{n-m+1},\dots,u^{n+1}]$
8:end for
Minimal Python implementation
The following short routine implements one time step of a scalar delayed equation using a ring buffer:
def step(u_n, hist, h, a, b):
    u_delay = hist[0]
    u_next  = (u_n + h*b*u_delay) / (1 - h*a)
    hist = np.roll(hist, -1)
    hist[-1] = u_next
    return u_next, hist
This implementation mirrors exactly the discrete history transport structure used throughout the numerical experiments.
The ring buffer thus provides a computational realization of the abstract history formulation
$u_{t}(\sigma)=u(t+\sigma),\qquad\partial_{t}u_{t}=\partial_{\sigma}u_{t},$
used in the operator-theoretic analysis of delay equations Hale and Verduyn Lunel (1993); Bátkai and Piazzera (2005). From a numerical viewpoint, this corresponds to the standard discrete history management in DDE solvers; see, e.g., Bellen and Zennaro (2003); Lubich (1984).
Acknowledgments
First, I would like to express my sincere gratitude to my supervisor, Professor Toru Ohira of the Graduate School of Mathematics at Nagoya University, for his invaluable guidance, encouragement, and insightful discussions throughout this research project. I would also like to thank Professor Toshiaki Hishida of the Graduate School of Mathematics at Nagoya University for enthusiastically answering my questions and for his invaluable advice on generating semigroups and evolution families while preparing this paper. Thanks also go to: Professor Mitsuru Sugimoto, who allowed me to attend his seminar and provided guidance. Thanks also go to Yukihiko Nakata of Aoyama Gakuin University and Junya Nishiguchi of Tohoku University. Lastly, I would like to thank Professor Jong Son Shin at Shizuoka University for his continuous support.